A Complete Bibliography of Publications in the Journal of Cell Biology: 1965–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

20 June 2023 Version 1.01

Title word cross-reference

19 [WS68]. 5 [Com66b, PS65b, RCV69]. 6 [OE66b]. ^{++} [GC66b]. ^{131} [MBM67]. ^{14} [BT68]. ^{3} [CL66, Com66b, MO69, NL66a, NL66b, Rez69, SP67a, Sco67, Sim67, SS65c, SK67b, Sti67, SMP65, TVP67, VM66b, WD68b, WHL69]. ^{55} [Orl68]. _{1} [MSP66, RS67b, TAP67]. _{2} [GG69, MSP66]. _{3} [GB69b]. α [KY66, LH67, Mil69]. β [Boq69, HSC68].

-Aminouracil-Induced [PS65b]. -amylase [KY66]. -cells [Boq69]. -Day [WS68]. -glucuronidase [HSC68]. -H [Com66b]. -hydroxy [LH67]. -Hydroxytryptamine [RCV69]. -keratin [Mil69]. -Phosphatase [OE66b]. -Thymidine [SS65c, SMP65, VM66b].

22S [BB69b].

4-Dichlorobenzyl [BP68a].

5 [TSACBB69]. 5’-Nucleotidase [TSACBB69]. 50S [Nan67].

ÂGes [Bea68].

A5969 [MMB65]. Abdominal [CKKW66, TP68a, TP68b]. Abnormal [ML66b]. Absence [CR66, KEW67b, NLS67, RS67b, SF69b]. Absolute [BS67]. Absorbed [BT68, SWR68]. Absorbing [HB67, SE67]. Absorption [CBP67, Cho65, FF67, Jer66, KW65, KC69, OYHM65]. Absorptive [OYHM65]. Acanthamoeba [BK68, BK69, KW67b, USK69, WK69]. Accentuated [MO69]. Accessory [All69]. Accumulation [BF65, CP68, GC66b, LT69, NM67, SK67b]. Accumulations [BN65a]. Acer [WN65a]. Acetate [HF68]. Acetobacter [BHO65]. Acetyl [Pog69]. Acetylation [PAM67]. Acetylcholine [HH67]. Acetylcholinesterase [BB66, BTD66, DK67, Kar65c, RB67]. Achalarus [RC67a]. Achnanthes [Hol68]. Acid [BSH68, Ber67b, BH67, BMBW67, Blu65, BT68, DK67, Dem68, FW65, FC68, JEZ69, JH65, KW65, Kra67, LCW65, LS68a, Lym67, Maa67, MC65, MK65b, PAM67, RAH69, SSSR67, Sed69, SP67b, SG69b, SB65c, WK66, WSW65, WWS66, WARS65, Wil65a, LH67]. Acid-1- [BT68]. Acid-Induced [RAH69]. Acid-Secreting [Sed69]. Acids [DSDD65, Hof65b, KSB65, PA67]. Acinar [AOS69]. Acridine [CB69a, MU67]. Acriflavin [Sim68]. Acriflavine [HA69, SvA67]. Acrolein [SC68]. Acrosome [BBFA67, ND65a, ND65b]. Acrosomes [DMEA67]. Across [SM65c]. Acth [Kah68]. Acth-Induced [Kah68]. Actidione [CBR65]. Actin [Aro65]. Acting [Ros69a]. Actinomycin [CP67a, DK68, Ebs67, EKSG69, Jag69, Joh69, Kah68, Lan69a, MS69c, PK69, Sim67, Sus65, WARS65, dMN67]. Action [BP68a, BT67a, BT67b, BFDP68, DDGL66, DC67, FJH65, MR68, ZZ67]. Activated [VHR65, MP67a]. Activation [vdD69]. Active [AH66b, BKS69, Com67, GW67a, GKS^{+}67, GC66b, VU68, WB69, WHT67]. Actively [HS65a]. Activities [DSA65, ENQ65, KHSR66, NQVF66, Sti68, UAH68, ZOE67]. Activity [BDRN65, BD67, CP67b, DP67, DSDD65, Fah69, Fis68, GP65, GTLE65, Hac66, Hac68, HT66, HK65, KP68, LS68a, LT68a, Maa67, Mat67, Nag65, OVF69, PM65, SDV68, SF69a, SPW^{+}67, SKWH68, ST65, TS65a, TMZ68, Too65, WN65b, HSC68, LH67]. Actomyosin [IKNG68]. Acylation [SWS68]. Adenine [BF65, Lan66, SGF68]. Adenohypophysis [Din69]. Adenosine [FP66, GP65, Gib65b, Nag65, RB66, SDV68, SF69a, MP67a]. Adepidermal [Kel66]. Adipose [LDRA67, LC66, OVF69]. Administration [CPR67, CL66, WSW65, ZOE67]. Adn [CA68]. Adrenal [CvD67, DS69a, GWS65, PVM69, TVM69, VM65, VM66a, VM66b]. Adrenals [Kah68]. Adrenocortical [NC69]. Adult [DBN68, HTBM65, Hir68, MCP69, MP68, MYS^{+}67, SOZC68]. Aedes [TSW67]. Aegypti [TSW67]. Aerobic [MB67]. Affecting [WRJ67]. Affinity [ML66a]. After [Cam65, CPR67, CL66, EKG66, Gou68a, GTLE65, HF68, MP68, Nan68, PVM69, RSL69, See67, SF69b, SR65, Tro65, Tro67]. Age [SOK65]. Aged [SOK65]. Agent [SA66]. Agents [GKS^{+}67]. Aggregates [CB66]. Aggregation [GL68, MWGM65]. Aging [GB69a]. Agranular [Lin67]. Aid [Kol65, KA65]. Albino [Par67]. Album [GS67a]. Albumin [MBM67, RTL69, SH67b]. Alcohol [DDGL66]. Aldehyde [Gun65, SD69a, Sus67]. Alga [BB67c, Bro69, Ram69, Urb69]. Algae [SS65c]. Alignment [HC68]. Alkaline [GC66c, GCG67, GB69b, MG67, SC67b, TSACBB69]. Allium [Fus68, Laf65]. Alpha [Hol67, OVF69]. Alpha-One [Hol67]. Alteration [LS68b]. Alterations [CT67, GWS65, KLW68, Lan65b, LSL68, MR68, SS69c, TMZ68, WB69, WLKS68, vdD67]. Altered [RBS69, SGA67, SAR69]. Alveolar [KCW67, LL66, SKK68]. Alveolar-Capillary [SKK68]. Amantadine [Pog69]. Ameba [CM65, OK67]. Ameba-To-Flagellate [OK67]. Amebae [Fli69, Rot67]. Americana [ST65, Smi66b, Sri66]. Amine [CY69, CYD69]. Amine-Storing [CYD69]. Amino [KW65, PA67]. Aminouracil [PS65b]. Amitosis [MB65]. Amnion [SW67]. Amoeba [CG69, Fli68, Lan65a, ZR65]. Among [PR68, Tay65b, TVM69]. Amount [BTVM69, NM67, SE67]. Amounts [Ebs67]. Amphibia [Sre65, WD68c]. Amphibian [FP65a, FP66, GW66, Kar65a, Lan67b, MK65b, Rog68, TD65, Too65]. Ampullate [BP69a]. Amylase [SP66, KY66]. Amyloid [EKG66, SC67a, SC67b]. Anaerobic [MB67]. Anaerobically [DNC69]. Anaerobically-Grown [DNC69]. Analogy [MCP69]. Analysis [Cog66, Fuj67, Him67a, HD67, Koe67, LHS^{+}65, MD69, She67, SC67a, Ste65a, TM69, VM66a]. Anaphase [RJ69]. Anatomic [DCL69]. Anatomy [BW66]. Anchor [KEW67a]. Anchoring [LB68]. Animal [Gar69, KM65, ML66a, TGB69, WG68a]. Animals [OJP^{+}68]. Annelides [Bas66, Bas66]. Annulate [Eve68a, Eve68b, Kes65, Kes68, KB69, KHH68, ML66b, SF69a]. Annuli [Fel65]. Anomalous [Faw68, Ste65b]. Anopheles [Sch69]. Ant [LR66]. Anterior [CM69, MSH69, SF66, TKH69, WST68]. Antibiotic [DSDD65]. Antibodies [BMH68, KC69, NP67]. Antibody [CP67b, MCB68, Pep66, SMW67, SFG67]. Antibody-Producing [SMW67]. Antibody-Staining [Pep66]. Antibody-Treated [MCB68]. Anticellular [CP67b]. Antigen [BA65a, BA65b, BA67a, BA67b]. Antigen-Induced [BA65a, BA65b, BA67a, BA67b]. Antigens [NP67]. Anuran [WD68c]. Apex [GS67a]. Aphid [Smi65]. Apical [SS65a, WG68b]. Aplysia [CKKW66]. Apparatus [Ais69, BH67, BB69b, Bro69, CMM66, DF66, EZ68, GSW65, Hir65, Kan65, KF65, Kan67, Mol65, MM66a, Nic68, OBJ^{+}68, OBJ69, RHL69, RS67a, Rin67, RJ69, RM66, SWDK67, Ste67, Wil69]. Apparent [Ech65]. Appearance [Man69]. Appearances [RSL69]. Apposed [BR69]. Araneus [BP69a]. Arbacia [And68, GTP69, Kar65b, LA68, TM69, TG69]. Area [BDRN65]. Arenarum [EP68]. Arginine [LN65a]. Argon [BOR69]. Armored [Rob66]. Arms [Wil69]. Arrangement [Ech65, Hag66, LL65, Met67, WRJ67].

Arrangements [BW67, MMB65]. Array [BE68, RK67b]. Arrays [Fla68a]. Arrest [KS66]. Arrowhead [IBH69]. Artefact [PA67]. Artifactual [Tor65]. Ascaris [Foo68, Ros65a, Ros65b, Ros69b]. Ascidian [DD65, Eve68b]. Ascites [BS66b, GK68, HBT69]. Ascorbic [SSSR67]. Ascospore [Car67]. Ase [TS65a]. Asexual [AHS67, Sti68]. Aspects [EF69, Fla68b, KS66, Pep66, SS68c, VU68]. Assay [RB69a, RC66, Wen66]. Associated [And69, BH67, DFW66, EZ68, GC66a, GG69, Lan68, LL66, MR68, MK68, SS65a, SKEB67, SBK67, ZOE67]. Association [HS67a, Kar65c, KH69, LC66, TSACBB69]. Assortment [KK67]. Astrocytes [BRH65, MK65a, PV67]. Asynchrony [KRC69]. ATP [Lan67a, RE69, TS65a]. Atpase [Mat67]. Attached [Fal69b]. Attachment [Eve68a, FLV68]. Atypical [HOZ68, NSHW68, SB65d]. Augusta [Nic68]. Aurelia [Huf69]. Aureus [Sug66]. Autonomic [ESL68, RB67]. Autophagic [FHC68a, FHC68b, FC68]. Autophagy [DdD67, DBdD67]. Autoradiographic [KW65, Sus65]. Autoradiography [Stu65, SH65, VHR65]. Avian [AHHS66, EE69, HS66, LC66]. Axon [ESL68, PSPO68]. Axonal [MG68, SR65]. Axons [DS68, PV67]. Axoplasm [Met69, MI69]. Axostyles [GC65c]. Azo [Kar69, HSC68]. Azo-Dye [Kar69]. Azoindoxyl [HH66]. Azurophil [BF68a, BF68b].

B [FW67]. Bacillus [FJH65, LvI65a, Nan67, Nan68, Nan69, SB65a]. Bäckerhefe [Moo66]. Backwards [DB68a]. Bacteria [PO65, SS66a]. Bacterial [HvIvZ66, SKHR68, vIHvZ66, vI66]. Bacterium [vIHvZ67]. baker [Moo66, HJ68]. Band [Kel67]. Bands [Faw68, SKN65]. Barrier [RK67a]. Barriers [DCL69]. Basal [All69, Arg65, DB66, HP66b, HvIvZ66, Hof65b, KEW67b, Las69, Par69b, Ter67, vIHvZ66]. Basalar [RC67a]. Base [DE69]. Based [Kel67]. Basement [KEW67b, SS65d, SBK67, Ter67]. Bases [Hac66, Hac68]. Basic [DMEA67, DD65, KA65]. Basique [Che67]. Basis [Kar67, KK67, SKK68]. Basket [KK69]. Basophils [FH65]. Beadings [ESL68]. Beads [KW67b, WK69]. Bean [New67a, New67b, NSHW68, WR68]. Beating [MS69c]. Beef [TKH69, VF69]. Beet [CHE66, ECH67]. Before [DS69b]. Behavior [BF68c, KEW67b]. Beltian [Ric68]. Berghei [VR67]. Beta [CHE66, HKL69]. Between [BLRG67, BR69, Din69, GW66, GTLE65, Las67b, MCP69, MM66a, Ove69, Pen66, PM65, She66, SE66, TRH66, WKG66, WD68c]. Bicarbonate [SC68]. Bicarbonate-Containing [SC68]. Binaire [PVT68]. Binding [Him67b, PA67, RS69, Sim67, HA68]. Binucleate [KS67]. Biochemical [BBdD65, CG65, CG66, DMEA67, DS69a, HSS67, LT68a, LHS^{+}65, MO68, MYS^{+}67, SS68a, SHW69, THNB69, dBVSH65, WSGH69]. Biochemistry [RLV65]. Biogenesis [CWL67, DSP66a, DSP66b, JKL68, KBL69, OSP67a, OSP67b, WHL68]. Biogenic [CY69]. Biological [CM69, SE69]. Biosynthesis [BLD^{+}66, CHB65, DS69a, FM67, SVD68]. Birds [Sla65]. Birefringence [For65, NA65, RS67a]. Birefringent [Aro66, MS69a]. Bladder [Cho65, DCL69, HSM65, HB66]. Blast [PCS69]. Blastula [LT67]. Blattarian [HSS67]. Blepharisma [DH66, Dem68, Ken66]. Blood [Beh67, BMH68, BP68b, BP68c, Las67b, LS68b, MD69, RK67a, SK68b, SS65d]. Blood-Brain [RK67a]. Blowfly [GB66b, TS65a]. Blue [LvI65a, LvI65b, SB65a]. Bodies [All69, Arg65, DB66, Hes65a, HvIvZ66, Hof65b, Laf65, Lan69a, LL66, LC68, MP66, RA65, Ric68, SM67, Tro65, Tro67, Wil67, vIHvZ66]. Body [AG68, BS66a, CY69, CYD69, Com66b, Fri69, Jas67, Lin67, LC65, LC68, MM65]. Bombyx [MMNT68a, MMNT68b, TMMN68a, TMMN68b]. Bone [AC67, CPR67, HN66, Vae68, VHR65, WN65b]. Border [EC65, OEC65, Ove68]. Borders [BP69b]. Born [KCR67b]. Bound [KHSR66, PM69]. Bovine [CSJ67, FFO69, SKN65, SS69c, SGR67, WCJS67]. Bowman [Die67]. Brain [BRH65, Bri65a, BR69, HTBM65, Kir69, LRM66, MP68, Ram65, RK67a]. Brains [Sre65]. Brasiliensis [Jam66]. Breakdown [Amm69]. Breathing [SF65]. Brevipes [Hol68]. Brewer [DNC69]. Bridges [Wil69, ZG68]. Brown [BB67c, LDRA67, OVF69, THNB69]. Brunner [Fri65]. Brush [BP69b, EC65, OEC65, Ove68]. Bryopsis [Urb69]. Buds [BS65]. Buffer [Gar69]. Bufo [EP68]. Bundles [KEW67a, Met67, RR69].

C [BT68, Fra67, DK68, Les67, MC69]. Calcification [And67b, And69, MK68]. Calcium [BD69, Bor68, HSM65, HS66, KS69, LL69a, SH67a, TG68, Win67]. Calcium-Insoluble [KS69]. Calculation [SE67]. Calf [KT68a]. Californica [CKKW66]. Calliphora [GB66b, TS65a]. Cambium [Sri66]. Canaliculate [Lin67]. Canard [TVP67]. Canary [Sla65]. Cancer [Lan69b]. Canine [CB66, Ich65]. Capillaries [BP68b, BP68c, CP69a, CP69b, KSR67]. Capillary [Kar67, KCW67, SKK68]. Capra [KO65]. Capsule [Die67]. Carausius [SS65b]. Carbohydrate [Hof68, NL66b, RL67]. Carbohydrate-Rich [RL67]. Carbohydrates [RHL69]. Carcinogenesis [Kar69]. Carcinoma [GK68]. Cardiac [CB66, ENQ65, Faw68, JS67, Man68, Man69, Sla65, SJ68a, SB68a, Tha66b]. Carotid [BS66a, CY69, CYD69]. Cartilage [And67b, And69, CP68, HD68, MPO67, MK68, SSW67]. Castellanii [BK68, BK69, USK69, WK69]. Castrate [Ber68]. Cat [FM69, HTBM65, Ito65, MF69, MP66, MP68, NOP69]. Catalase [Fah69, FN69, VF69]. Cation [DN66]. Caudatum [RP67]. Caupo [Gou68a]. Cecropia [AH66b, Sta65]. Cell [AOS69, Asb67, AB65a, BKS69, BA65a, BA65b, BA67a, BA67b, Bon69, BN65b, BW66, BR69, BL68, Cam65, Die66, DFW66, EKSG69, FP65a, Fel66, FLV68, Fle65, Fli65, FCLT69, Fuj67, GS67b, GC66c, GCG67, GB69b, HTBM65, Ham66, HT66, HB66, HN66, HJ68, Hes65a, Hey66, Hir65, Hof65a, HS65a, Hof68, IBH69, JEE69, JP67a, JP67b, JP68a, JP68b, JP68c, KEW67b, KW67a, KK67, Koe67, KH69, KHSR66, Las67a, LT68b, LS68b, Les67, LT69, Løv69, MT67, ML66a, MD66a, MD66b, MBMB68, MO69, MCK65, MW66, NKKW69, Ove65, Phi65, PS65b, Pre66, RL67, Ray67, RC66, Rey65, Riz67, RS67b, RM69, RS69, RM66, ROJ65, Rub69b, SR68, SP67a, Sch69, Sed69, SR69, SGF68, SS69a, SCP67, SH68b, SHW69, Ste65c, SK68c, Sto68, SE66]. Cell [SH67b, TESW69, TCL65, TL67, TT69, TMF66, VF67, WD68b, WSGH69, Wei65, Wei66, Wei67b, Wei67a, Wil65a, Wol67, WN65a, You67, ZG68, dMN67, MP67a, Moo66]. Cell-Plate [HJ68]. Cells [AH66a, AC67, AS67, ABK66, BLT69, BSH68, BS65, BM69, Ber67b, BF69, BH67, BH69a, Bla66a, Bla66b, BS66b, BN65a, BHB68, BFB^{+}69, Bor68, BP69b, Bri65b, BMR67, CPR67, CP67b, Cha65, Chr65, CKKW66, CL66, CLB65, Com66a, CP68, DFM65, DSDD65, Die67, DK68, Ech65, EC65, EBH67, EKSG69, ETS68, ET69, EC67, EB66, Fal69a, FH65, FC67, FB66, Fla68a, Fla68b, FD65, Flo69, FOP^{+}69, FC68, FS69, Gar68, GK68, GMGFLS65, GL67, GAPRC68, GMB65, GW67b, GB69b, GPB68, HL66, HP66a, HP66b, HD66, HRS69, Hol67, HBT69, HCGP67, HAKB65, HA66, HB67, IP65, ISS68, JEE69, JT67, JH65, JLL68, Joh69, Kah68, Kar69, KS67, KB65a, KB67, KL65, KCW67, Kla68, Kle69, KZvW65, KHH68, KRC69, KM65, KCR67a, KCR68, Laf65]. Cells [Lan65b, LN65a, Lan67a, Las69, LF69, LL65, Len66, Les67, LL69b, Lou68, LF65, LC66, MD69, MS69c, Mer68, Mis66, Mou65, NL66a, NL66b, New67b, NC69, NdM66, Oka67, OYHM65, OJP^{+}68, Orl68, OEC65, Ove66, Ove69, Pad66, Pad69, Par67, PVM69, Pen66, PP69, PK69, Pla69, RL67, RHL69, RP66, Ray67, RBS69, Rez69, RTL69, Ric68, RJM68, Ros65b, Rub69a, Rys67, SOZC68, SMW67, Sco67, See66a, See66b, See66c, See66d, She66, SFM67, SFM69, SCP67, SWP69, Sil69, SWDK67, SW67, Sla65, SF66, SVWD68, SAW^{+}66, SE66, Sus67, SW66, SGA67, SAR69, SH65, Tan69, TMMN68a, Tay66, Tha66a, TAP67, TM67, TRH66, TL67, WG68a, WO67a, WO67b, WSW65, WLKS68, Wil65b, WST68, WG68b, WN65b, vBPB66, Boq69, HA68]. Cellular [CT67, DdD67, DBdD67, Fab68, Fri68, Gre65, Koe68, Pad66, SS66a, SB67, War69]. Cellules [Bea68, TVP67]. Cellulose [BHO65, Hey66, Mil69]. Center [SH65]. Centers [Wen66]. Central [HZL66, HD67, ST68, ST65, VCM65]. Central-Pair [ST68]. Centrifugal [Zal69]. Centrifugation [BHB68, Gil65]. Centriole [MG66, MW66, Phi67, RJM68]. Cepa [Fus68, Laf65]. Ceratocystis [Ais69]. Cercopides [Gou68b]. Cerebellum [Fuj67, Ros66, SOZC68]. Cerebral [KK67, MK65a, PPKA68, Ram65, dBVSH65]. Cerebrospinal [Bri65a]. Cerevisiae [CWL67, Wil65a]. Certain [GS67a, GC65c, Phi69]. Cervical [NDRFK66]. Cestode [RC68]. Chambers [Ros67a, RKC68]. Chang [EBH67, JEE69]. Change [Kra67, Sto69]. Changes [AOS69, BP69a, BA65a, BA65b, BA67a, BA67b, CG65, CPR67, CSJ67, GGF68, Hes65a, Ich65, JEE69, KI67, KT68b, Maa67, MK68, OYHM65, Ove65, PaSN65, Par69b, RJ69, RK67c, SF65, SK68a, THNB69, VM65, WSW65, WCJS67, dWA65]. Changing [NA65]. Channel [Ter65]. Chaos [CM65, Nac68]. Chara [She67].

Character [DGT65, GWT65]. Characteristic [SR68]. Characteristics [HTBM65, JB68, Men69, SS68a]. Characterization [ADC69, BTG69, DDB69, FFO69, HSS67, KM65, MRH^{+}67, MO68, RAH69, SRZB66, SK68c, WR68]. Chard [KSB65]. Charge [MD69, SD69b, WK66]. Chelation [CT67]. Chemical [CSJ67, GCG67, KMK66, Nan69, RCV69]. Chemically [DB68b, HP66b, Rob69a]. Chenopodium [GS67a]. Chez [CA68, PVT68, VP69]. Chick [AR67, Bye66, CHB65, DB66, EI67, Fis67, HR67, KGO69, Les67, ML66b, OS69a, OS69b, Ove69, PK69, She66, She68, SFM67, TRH66, Wei68, KY66]. Chicken [Hes65a]. Chinese [Bri65b, EKSG69, ETS68, ET69, PP69, RS67b, SB66, TAP67]. Chironomus [DE69, DL69, SS66b]. Chlamydomonas [JP68c, Rin67, RMR69]. Chloramphenicol [SMW67]. Chloride [PCS69, SK68a]. Chlorophenoxyisobutyrate [SA66]. Chlorophyll [LP66]. Chloroplast [ADC69, BP68a, BB67c, Car69a, DC67, Fal69b, GC66a, Gun65, LP66, Lym67, OSP67a, OSP67b, Par65a, SSS68, SS65c]. Chloroplasts [ADC69, KSB65, NM67, PaSN65, PNGM66, SSS68, SL66, SSWG68, TGB69]. Cholesterol [DS69a, FCLT69, HHM69, SD69a, TSACBB69]. Cholesterol-Fed [FCLT69]. Cholesteryl [SSGF69]. Choline [HB69a, Pla69]. Cholinergic [ESL68, Rob69b]. Cholinesterase [BDRN65, BD67, DK67, NQVF66, ST65]. Chondrioids [Nan68]. Chondrococcus [PO67a, PJO67, PO67b]. Chondromucoprotein [HD68]. Chorioallantoic [Oll68, SB68b]. Choroid [DB66]. Chromatin [Com66b, Com67, DDB69]. Chromatographic [Kor67]. Chromatolytic [HNV67]. Chromatophile [Vau66]. Chromatophore [BTH68]. Chromosomal [BTG69, GW67a, HP68, SR67a]. Chromosome [AM69, DE69, Hes65c, HB69b, KS67, Moe69a, NAL^{+}67, Nic65, NK69, Wol68]. Chromosomes [Amm69, Bar67, BH69b, Cav66, DE69, Him67b, HA66, MU67, MO69, MB65, PHP67, Stu65, Tay65b, TW65, WH66]. Chrysophycean [Bro69]. Chylomicron [SSGF69]. Cilia [All68, Cob69, Din69, DB66, Gib65b, Gib65a, HOZ68, MP68, RB66, Ree65b, RRG68, RC69, Sat65, Sat68, SB66, Win67]. Ciliary [Arg65, Hes65a, Mat67, RB69a, Tor65]. Ciliata [Amm69]. Ciliate [DH66, Jen67, LK67b]. Ciliated [All69]. Ciliumlike [Act66]. Cinereum [AS67]. Circadian [SP67a]. Circulating [BMH68, SK67a, SK68b]. Circumfusion [Ros67a, RKC68]. Cisternae [Buc66, Wei68]. Cisternal [BN65a]. Citrate [VC65]. Classes [SH68a, SS69a]. Cleavage [Arn69, Lan66, LT67, TM69]. Cleaved [DB68b]. Clonal [BLT69, RTL69]. Cluster [CKKW66]. Cmrl [HP66b]. Cmrl-1415 [HP66b]. Coat [GBW69, GB66b, Ito65, RL67]. Coated [FF67]. Coating [MP65]. Cochlea [SR65]. Cochlear [DFW66]. Cockroach [Hag66, HS68]. Coenocytic [Urb69]. Cohnii [KR69b]. Colcemid [KS67, KS66, SB66, ZZ67]. Colchicine [BT67a, BT67b, Kla68, Pad66, RS69, RC69, Tay65a, War68]. Colchicine-Binding [RS69]. Cold [CvD67, PVM69, THNB69, TVM69, VM65, VM66a, VM66b]. Cold-Exposed [THNB69]. Cold-Treated [CvD67]. coli [FW67, BR67, Non68, HBS69, SB65b]. Collagen [Bac67, CHB65, CP68, FM67, HSS67, Par69a, RB65]. Collagenolytic [WN65b]. Collecting [GGM^{+}68, GGBO69]. Colloid [PW67]. Colloidal [MD69, Pad69]. Colon [WWS66]. Columnar [Die67]. Columnaris [PO67a, PJO67, PO67b]. Combined [BBdD65, HSC68, HH66, LHS^{+}65, SB65a, SB65b]. Common [Car69b]. Communication [JL68, LK67a, LP67, Pen66]. Companion [WN65a]. Comparative [BP66, BD69, Fel69a, KS69, MW67, RC67a]. Comparison [ADC69, NL66a, Pag65]. Compartment [PWFP68]. Compartmentation [CG68]. Compartments [Bri65a]. Completion [Gou68a]. Complex [Aik67, Ber67b, Buc66, Hic66a, KB65a, Man69, NL66b, RHL69, RI67, SS65b, Som65, WB69, WG68b]. Complexe [VP69]. Complexes [Fli69, Gar68, HH66, IBH69, Moe69b, SBK67]. Component [Aro66, Phi65, Rob69b]. Components [CE67, HG69, MG68]. Composition [CSJ67, DE69, Gou68b, KMK66, LSL68, Luc65b]. Compound [WG65]. Compounds [RB66]. Computing [Oka67]. Concentration [RD69, RE69, SG69a, SR67b, WRJ67]. Concentric [NC69]. Concerning [Har65]. ConcrÉTions [Gou68b]. Conditions [WRJ67]. Conducting [BDRN65]. Cones [Coh68, Eak65, YD68]. Configuration [Kel67, KB65a, Kra67, Lin67, Met69, PS66]. Confirmation [Hir68]. Conformation [LS68b]. Confronting [Wei68]. Connections [MM66a, Ram69, She66]. Consequences [DB68a]. Conservation [SW67]. Constancy [CvD67]. Constituents [Die66, SM65a]. Constrictions [MO69]. Contact [YF69]. Contacts [YB67]. Containing [FE69, SD69a, SC68, SWH66, Sus65, Tro65, WWS66]. contamination [Mil69]. Content [AMH69, BF68a, BF68b, BB67b, CvD67, EH68, GW67a, TVM69, VM65]. Contiguities [Tha66b]. Continuity [WS65]. Contractile [DS69b, GC65c, KR69a]. Contractility [CFK65a, CFK65b, RB69a]. Contraction [Aro67, Lan65b, MSL69, Pod68, SB68a, Ste65b, Ste65a]. Contrast [Fel69b, HH66, SWH66]. Contribution [WK66]. Contributions [Fle65]. Control [BLT69, Hof68, Sil69, SH67b]. Convenient [RC66]. Coordinated [SM69]. Copepod [WF69]. Copilia [WF69]. Coprophila [GGK66, Phi66b, Phi66a]. Cord [BBP65, MS69b, SFM69]. Cordycepin [SWP69]. Core [Moe69b]. Cores [Bie66]. Cornea [CHB65]. Corneal [SP67a]. Corpus [AH69a, AH69b, LR66]. Corpuscles [NOP69]. Correlated [BRGG65, BRG68, SHW69, WSGH69]. Correlates [dBVSH65]. Correlation [RM69]. Correspondence [PM65]. Corresponding [Koe68]. Cortex [All69, Fuj67, GGF68, HTBM65, KS69, KK67, MK65a, PPKA68, Ram65]. Cortical [And68, Bon69, Huf69, Kah68, KI67, Ric68, Tay66]. Cortisone [AMH69, KLW68, WLKS68]. Cortisone-Induced [KLW68, WLKS68]. Cotton [Hey66, Jen68]. Cotyledonary [Yat65]. Counts [SP67a]. Coupled [BL68]. Coupling [SB68a, HSC68]. Cours [Bea68, PVT68]. Cpib [SA66]. Crab [HC69b]. Crabs [Lan69b]. Crassa [DGT65, DDB69, GWT65, HW67, Luc65a, Luc65b, MT67, Sti68]. Crayfish [KB68]. Cricket [KMK66]. Crista [NSHW68]. Cristae [Lin65]. Crithidia [HA69]. Critical [Wol68]. Cross [All68]. Cross-Sections [All68]. Cruentum [GC65a, GC66a]. Crush [HN65]. Crustacean [Fah67, Ros69a, WG65, WF69]. Crustaces [Che67]. Crypt [EB66]. Crypts [TT69, VF67]. Crystal [Kar65b]. Crystalline [Ber69, Hol68, MS69b, SB69b, Wil65b]. Crystallization [Bye66]. Crystalloid [Bie66, FH65, Oll68]. Crystals [BM69, KH69, MS69a]. Cucurbita [CE68a, CE68b]. Culicidae [Sch69]. Cultivation [HP66a]. Culture [Amm69, DFM65, GMB65, JT67, Kah68, PM65, Pla69, Ros67a, RKC68, SWDK67, SW66, TVP67, UAH68]. Cultured [AC67, BMR67, BBP65, Cav66, EBH67, Ish68, ISS68, Joh69, KL65, Kla68, LN65a, MS69c, Oka67, PP69, PK69, RP66, RS67b, SSSR67, SW67, WO67a, WO67b, WB69]. Cultures [BN65b, BBPM67, EF69, Gar69, GC66c, HN66, KWM69, LG65, MBB67a, MBB67b, MBMB68, RC66, SFM69]. Current [SE69]. Cycle [AOS69, AB65a, DK68, EKSG69, Fli65, GB69b, Hof68, Oka67, Par69b, Pre66, RJM68, RM69, RS69, SCP67, SM65b, Sti68, TP68a, TP68b, VF67, Wil65a]. Cycloheximide [VF67]. Cysts [MSW68]. Cytochalasin [KRC69]. Cytochalasin-Induced [KRC69]. Cytochemical [BA69, DMEA67, DD65, Ebs67, FK66, Fah69, FN69, Fri69, GP65, Gol67, LF65, Mat65, SB65c, Sus65, HSC68]. Cytochemistry [ET66, HH66, dFM66, Rub69a, Sal67, Sal69, SB65b, TB65, TE66]. Cytochimique [Che67]. Cytochrome [MC69, SPW^{+}67, SKWH68]. Cytochromes [TCL65]. Cytodifferentiation [AWE67, KGO69, LT67]. Cytologic [Lan65b]. Cytological [Aro65, DH66, GP66, GWS65, LF65, MBB67a, MBB67b]. Cytology [WKS66]. Cytolysomes [TD65]. Cytophotometrique [PVT68]. Cytoplasm [Nac68, SS66b, TMMN68a, Tay66, vI66]. Cytoplasmic [AF68, Blo67, Coo66, DD65, DGT65, FH66, Fli68, Fus68, GWT65, GB66b, HBS65, Kes65, KB68, LC66, Mou65, NB65, PS65a, Ric68, SD68b, TM67, VR67]. Cytotoxic [DS66, DSDD65].

D [CP67a, DK68, Ebs67, EKSG69, GAPRC68, Jag69, Joh69, Kah68, Lan69a, MS69c, PK69, Pog69, Sim67, Sus65, dMN67]. D- [Sim67]. D-Treated [Lan69a]. Dab [SKWH68]. Damage [HB69b, KCW67]. Dans [Che67, Gou68b, Bas66]. Daphnia [SRZB66]. Darkness [Eak65]. Daunomycin [DSDD65]. Day [AB65a, WS68]. Death [Bon69]. Decapodes [Che67]. Decrease [PVM69]. Deferens [FF67, Mer68, Rob69b]. Deficiency [SSSR67]. Deficient [EBH67, JEE69]. Defined [HP66a, HP66b, Rob69a]. Deformability [Wei65, Wei66, Wei67b, Wei67a]. Degeneration [Fri68, SR65]. Degradation [EBH67, WARS65]. Dehydration [Mit69, SWR68]. Dehydroascorbic [MC65]. Dehydrogenase [OVF69, SB65a, SB65b]. Dehydrogenases [FK66]. Delayed [PHP67]. Delimitation [Car67]. Demonstration [Aro65, Can68, JEZ69, SB65a, SPW^{+}67, SKWH68, HSC68]. Dendraster [Wen66]. Denervation [MMA65]. Dense [PRGC69]. Density [BLRG67, Gil65, RAH69, SM65a]. Denucleation [SK67a]. Deoxyribonucleic [BMBW67, Wil65a]. Deoxythymidine [Tak68]. Dependence [Fli69]. Dependent [RI67, SL66]. Depleted [FFS67]. Depletion [CT67]. Depolarization [MU67]. Deposit [Pod68]. Deposition [Ray67]. Deprivation [LN65a, NDRFK66, dWA65]. Derivative [SMP65]. Dermal [BTH68, DDGL66]. Dermis [PF65]. Derniers [Bea68]. Description [Lou68, WLKS68]. Desmosomes [Kel66, Loc65, RSL69]. Despite [Gou68a]. Destruction [Amm69]. Detail [Lan65b]. Detectable [RS67b]. Detection [AP69, CP67a, RHL69]. Determination [BB67b, BTVM69, RD69, VHR65]. Determining [Koe68]. Deuteron [SOZC68]. Deux [Bea68]. Developing [Asb67, Ech65, EI67, IBH68, Kel66, Man68, MK65b, Orl68]. Development [And67b, Car69a, CSJ67, CHB65, DFM65, DF66, EE69, Fli69, Fra67, Fri68, GTP69, Gol68b, KCW67, KT68b, KRC69, Len69, LA68, OS69a, OS69b, PM68, SM69, SS66a, SFM69, Sho66, SS65d, TG69, TMK68, VR67, War69, WCJS67, WN65a]. Developmental [Hes65a, LN65b]. Diabetes [OJP^{+}68]. Diameter [Wol68]. Diaminobenzidine [SKWH68]. Diatom [Hol68]. Diatoms [RLV65]. Diazophthalocyanins [TB65]. Dichlorobenzyl [BP68a]. Dictyosome [MEK68]. Dictyosomes [MM66a]. Dietary [vdD67]. Difference [WD68c]. Differences [BF68a, BF68b, TVM69, WKG66]. Different [DK68, MSP66, SS69a, TVM69]. Differential [MI69, MMA65, PK69, WR68]. Differentialis [Lan68]. Differentiated [AH66a, SFM67]. Differentiating [BC65, CB65, CE67, Rub69b]. Differentiation [And67a, And68, BH69a, CLB65, Eak65, EI67, Fla68b, Fuj67, GL67, MB67, MR65, MG67, PM65, SM69, SMW67, VR67, ZG68]. Difficulty [Car69b]. Diffraction [Bar69, Ber69, EV65, SB69b, TG68]. Diffusible [YF69]. Digestion [Dem68, EKG66, TW65, WR68]. Digestive [GMB65, Len66, MBM67, SLAR69]. Digitonin [ADC69, SD69a]. Digitonin-Containing [SD69a]. Digitonin-Treated [ADC69]. Dimensional [RR69]. Dimensions [Tha66a]. Diminuta [RC68]. Dimorphic [Ram69]. Dinoflagellate [KR69b]. Dioxide [Pad69]. Diploid [HHM69, Sch69]. Diptera [Phi65, Sch69]. Dipteran [DL69, LN65b]. Direct [HN66]. Directional [DFW66]. Disc [RSL69]. Discharge [MPD65]. Disclosed [MI69]. Discs [FM65, Rob69a]. Disease [WA65]. Dish [EF69]. Disorganization [SGF68]. Dispersion [ABK66]. Displaced [Vau66]. Disruption [EE68, Ste65a]. Distinct [BHdD69, SH68a, TSACBB69]. Distribution [BN69, Bri65a, BS69, DDBE69, Dem68, HS66, KT68b, Lan68, LT69, MPO67, PR68, RD69, SP66, Tay65b]. Dividing [HEL67, PR68]. Division [AB68, Cam65, DK68, Fel66, FCLT69, Gou68a, HW67, Hir65, Jen67, JP68c, KR69b, PVT68, PS65b, RE69, SP67a, Sto68, SE66].

Division-Synchronizing [RE69]. Divisions [Bri65b, HAKB65, PVT68]. Dl [TVP67]. Dl-Leucine- [TVP67]. DNA [Amm69, BP66, BTVM69, BB67c, CP67a, CvD67, Coo66, CR66, DE69, Ebs67, FCLT69, GK68, GS67b, GW67b, GG68, GG69, HEL67, Him67b, HG69, HBS65, HP68, KL65, dFM66, MU68, MSH69, Mou65, MM66b, NP65, PR68, PVM69, Pre66, RP66, RC66, Rod69, RI67, SR67a, SSS68, See66a, See66b, See66c, See66d, SL66, She67, SVWD68, SS65c, SvA67, SMP65, Sto69, Tak65, Tak68, Tay65b, TVM69, VM65, VM66a, WD68c, vBPB66]. DNA-Dependent [RI67, SL66]. DNA-Induced [GS67b]. DNA-Like [SVWD68]. DNA-Protein [Him67b]. Dog [PW67]. Dollar [Wen66]. Domestic [MP67b]. Domestica [PR69]. Dormant [HS65a]. Dorsal [BTD66, BBPM67, KT68b, MBB67a, MBB67b, NQVF66]. Double [FP65b]. Double-Staining [FP65b]. Doublet [ST68]. Dragonfly [KB69]. Drone [PB69]. Droplet [KCR67b, PW67]. Droplets [ALEW66, LC66, SWH66, Tro67]. Drosophila [TMK67, BP66, BB69a, BBB68, FM65, Fri68, Hes65c, HP68, MM66b, NP65, Riz67, Rob69a, Sho66, Sho67, TWE68]. Drug [Orr65a, Orr65b, OEE65, OE66a]. Drug-Hydroxylating [Orr65a]. Drug-Metabolizing [OEE65, OE66a]. Ductal [Mor68]. Due [PA67]. D’Une [Che67]. Duodenal [HB67, OYHM65, TT69]. Duration [TAP67]. Durations [MSP66]. During [AR67, Amm69, AS66, AOS69, AWE67, And67a, BF68c, Beh67, BP69a, Bon69, BA69, BT68, CG65, CSJ67, CHB65, DSA65, Die66, DK68, DR66, Fel66, Fel68, FF67, Gol68a, Gol68b, GG69, Jer66, Kar65a, Kar69, KMK66, KB67, KZvW65, KT68b, Kra67, LS68a, Lea68, LT67, Løv69, Maa67, McC69, Mit69, MH67, Nic65, Ove65, Par69b, Pre66, RCE69, RM69, RS69, SWR68, SM69, See67, SCP67, SK68a, Ste65c, Sti68, Sto69, Tak65, Tak68, Tay65b, TM67, TP68a, TP68b, TL67, TMK68, VR67, War69, WPP69, WCJS67, ZG68, dWA65]. Dyads [Ros69b]. Dye [Kar69, HSC68]. Dynamic [AOS69, GSS69].

Ear [FD65]. Earle [KHH68, KH69]. Early [Fel68, HJ68, KC69, MCP69, PV67, RCE69, SOZC68, She66, She68, TRH66, Tro65]. Earthworm [Cog66]. Earthworms [Bac67]. Eccrine [Ell65]. Edema [HZL66]. EDTA [HL66]. Edulis [ND65a, ND65b]. Eel [BB66, Ham65]. Effect [AMH69, BH68, BHHW69, CBR65, DN66, DSDD65, Fel65, GGM^{+}68, Gib65b, GKS^{+}67, HSM65, Hes65c, Ken66, Kla68, Kue69, MC65, MCB68, NDRFK66, PAM67, Pog69, SK68a, Sim68, SVD68, SH67a, SGR67, SHMR69, Sus67, Tan69, TAP67, VF67, War68, Wol68]. Effects [BN65b, Bor68, CG68, CP67b, CYD69, DP67, DS69a, DK68, Fli68, GSS69, HL66, HA69, HS65a, Kah68, Lan65a, Lan66, LN65a, LG65, MS69c, RB66, RE69, SOZC68, WO67a, WST68, ZR65]. Efferent [SR65]. Egg [Arn69, BF68c, KS69]. Eggs [Gou68a, Har65, Hir65, Jag69, KI67, Lan66, MCB68, RS67a, TCV69, TM69, Wen66, ZZ67]. Egregia [BB67c]. Ehrlich [GK68, HBT69]. Ehrlich-LettrÉ [GK68]. Einzelelementen [Gir65].

Elaborate [Ish68]. Elaboration [VP69]. Elastic [GRH66, RB69c]. Electric [BB66, Kar65c, She65]. Electrical [LF69, Men69]. Electron [And67b, AP69, BS67, BEB67, BB67a, BE65, Ber69, BA69, BS69, BA66, BBP65, BBPM67, BT68, CB66, CY69, Cho65, CL66, DFM65, DK67, DS68, DB68b, DR66, EV65, Fel65, Fis67, FC67, GC66b, GBW69, GB66b, HSS67, HK65, HH66, IKNG68, JT67, Joh69, Kar65a, KA65, KCR67a, KCR68, Laf65, LR66, LH65, LT68b, LS68b, MD69, MPD65, MI69, Mit69, MH67, Nac68, Nag65, NP67, NL66b, NM67, NdM66, RL67, Ree65a, RD69, Rot67, SS68a, Sal67, Sal69, SBS69, SD69a, Sco67, SB65a, SB65b, SC67a, SFG67, SG69a, SSG69, SD67, SKHR68, ST65, SKEB67, SB69b, SAW^{+}66, SH65, SB68b, TG68, Too65, TRH66, Tro65, Tro67, VC65, VF69, WSB68, WA65, dMN67, HSC68]. Electron-Opaque [Fel65, Tro65, Tro67]. Electron-Probe [RD69]. Electron-Transport [SKEB67]. Electronique [TVP67]. Electrophoresis [PNGM66]. Electrophorus [Kar65c]. Electrophysiological [She66, She68]. Electroplaque [BB66]. Electrostatic [Kol65]. Element [Oll68]. Elements [BC65, CB65, CE67, Gir65, Koe67, Koe68, MI69]. Eleven [AB65a]. Eleven-Day [AB65a]. Elimination [SD67]. Elongation [RMR69]. Elongatum [AHS67]. Emarginata [HC69b]. Embedding [BMR67, LH65]. Embryo [AR67, AB65a, Bye66, DB66, GL68, HP66b, MK65b, She66, She68, TRH66, Wei68]. Embryogenesis [Jen68, Kar65a]. Embryonic [CLB65, CP68, Fis67, KEW67a, Les67, MCP69, Man69, Ove69, PM65, SFM67, SFM69, KY66]. Embryos [Eve68b, ML66b, TCV69]. Emptying [SH68b]. Endings [Ber67a, BLS67, HR67, LS65, Las69, SR65]. Endocrine [FOP^{+}69, GAPRC68, OJP^{+}68]. Endocytic [WG68b]. Endodermis [Bon68]. Endogenous [Fis68, Lan66, MC69]. Endometrium [HOZ68, Ter65]. Endoplasmatischen [Gir65]. Endoplasmic [BF65, Bas66, DSP66a, DSP66b, Gir65, KHH68, MS69b, NC69, OEE65, OE66a, OE66b, WK66, WKG66]. Endoplasmique [Bas66]. Endosperm [HJ68]. Endothelial [KSR67, MSL69]. Endothelium [SM65c]. Endplate [KZ69a]. Energy [LDRA67, War69]. Energy-Transfer [LDRA67]. Enhancing [SWH66]. Entamoeba [MSW68]. Enteric [Ito65]. Entorhinal [GGF68]. Entry [BF68c, Fra65]. Enucleation [Fli68]. Envelope [And67a, Eve68a, HBS69, MS69b]. Envelopes [SK68c, WF68]. Environment [Mer68]. Enzymatic [BTG69, HCGP67, SG68]. Enzymatically [BKS69]. Enzyme [BF68a, BF68b, HK65, HH66, McC69, NP67, Orr65a, OEE65, OE66a, Rub69a, Sal67, Sal69, Sti68, TW65, UAH68, WR68]. Enzyme-Labeled [NP67]. Enzyme-Membrane [OE66a]. Enzymes [DDBE69, HB67, MCB68, RDS^{+}69, RB69b, SPT65, VRN^{+}68, ZOE67]. Enzymically [WHT67]. Eosinophil [MdHP66]. Eosinophils [SP67b]. Ephestia [Ove66]. Epidermal [FE69, KEW67b, NKKW69, SS65b, SS68c]. Epidermis [Cog66, FP66, Kel66]. Epiphyseal [And69, SSW67]. Epithelia [DB68a, MP65]. Epithelial [AS67, BP69b, EC65, HB66, JT67, KW67a, OEC65, Ter67, UAH68, WG68b].

Epithelioid [SW66]. Epithelium [BT68, Cho65, DB66, EE69, FOP^{+}69, Hic65, Hic66a, Hic66b, KKHE65, Lav69, MW67, SP67a, Tor65, YB67, YB69]. Equine [PAM67, Pog69]. error [Mil69]. Erythroblast [SD67]. Erythroblasts [SK67a]. Erythrocephala [GB66b, TS65a]. Erythrocyte [Mit69, See67]. Erythrocytes [BMH68, CCIC69, SKN65, TD65, Too65]. Erythrophores [Mat65, MO68]. Erythropoiesis [GW66, RCE69]. Erythropoietic [GW67b]. Escherichia [FW67, HBS69, SB65b]. Establishment [RTL69, RCE69, Sch69]. Ester [SSGF69]. Estrogen [Ber68, KGO69, OS69a, OS69b, RK67c]. Estrogen-Induced [KGO69]. Estrogen-Treated [Ber68]. Etch [ADC69]. Etching [Nan68, Nan69]. Ethanol [Win67]. Ethanol-Calcium-Isolated [Win67]. Ethione [SF69b]. Ethione-Induced [SF69b]. Ethionine [BF65, SGF68]. Ethionine-Induced [BF65]. Ethyl [SA66]. Etiolated [Jac68, SK68a]. Etude [Che67, TVP67]. Euglena [Blu65, Car69a, Coo66, Lym67, MEK68, SL66, SB65c, Som65]. Euplotes [Pre66]. Eurystomus [Pre66]. Evaluation [MD69, MW66]. Events [Cam65, GB69b, RCE69]. Everted [KW65]. Evidence [BH69b, Che67, Coh68, MK68, Mou65, NM67, PB69, She68, Sus65, WD68b, vdD67, Mil69]. Evolution [VP69]. Examination [BEB67, BB67a, Kar65a]. Examined [SJ68b]. Excentricus [Wen66]. Exceptions [Phi69]. Excess [DDGL66]. Exchange [KB68]. Exchanges [Fel65, Fel66, Fel68]. Excised [Hof65a]. Excitation [SB68a]. Excitation-Contraction [SB68a]. Exconjugants [Amm69]. Excretion [Mor68]. Exhibiting [Faw68]. Exocrine [JP67a, JP67b, JP68a, JP68b]. Exoerythrocytic [HHS66]. Exogenous [MC69, RK67a, HA68]. Expansion [PWFP68]. Expectation [MW66]. Expelled [SD69b]. Experimental [OJP^{+}68, Mil69]. Experimentally [HZL66, RBS69, SGA67, SAR69]. Explanation [SD69b]. Explanted [HP66b]. Exponential [Oka67]. Exposed [KHH68, THNB69, VM65, VM66b]. Exposure [PVM69, TVM69]. Expression [SMW67]. Extensor [RC67a]. External [Les67]. Extracellular [Bou69, Bru69, Coh68, PB69, VCM65]. Extract [CPR67]. Extracted [CFK65a, CFK65b]. Extraction [BT68, SSWG68]. Extracts [SC67b]. Extranucleaire [Che67]. Extraocular [CMDLB68]. Extrusion [CP68, SK68b, Szo65, TMMN68a]. Eye [PB69, WG65, WF69]. Eyes [Sho66].

Faba [Laf65, PS65b, TW65, WD68b]. Fabricius [Hag66]. Factor [YF69]. Facultatively [MB67]. Fagacearum [Ais69]. Fallax [Aik67, HHS66]. Fasciculata [HA69]. Fast [Fah67, MG68, Men69, Ros69a]. Fast-Acting [Ros69a]. Fasting [Tro67]. Fat [LC65, LC68, THNB69, Tro67]. Fate [EC68, Vau66]. Fats [Jer66]. Fatty [RAH69]. Fe [Orl68]. Feather [KEW67a]. Feature [BBFA67]. Features [Blo67, CHE66, EKG66, Nan68, Ros66]. Fed [FCLT69]. Feeding [AHHS66]. Felderstruktur [HMW66]. Femoral [Hag66, PR69]. Fenestrations [KSR67]. Ferritin [Bri65a, Tor65]. Fertilization [Gou68a]. Fertilized [KI67]. Fetal [Bel69, EF69, MYS^{+}67, RCE69, Sco67, WSB68, WS68, War69].

Feulgen [BS66b]. Fiber [CF66a, For65, GP66, Ham66, Met69, MI69, RM66, RB69c, SDV68]. Fibers [Ber67a, Ber68, BRGG65, BRG68, BD67, CFK65a, CFK65b, EP68, GRH66, Hes65b, HD66, Hey66, HMW66, MR68, Pod68, RRG68, Ros65a, Ros69b, Smi66a, Smi66b, SJ68b, WS65, WSB68, WS68, Wol68]. Fibres [Pag65]. Fibril [Bru69, PF65, SC67a]. Fibrillar [CE67, Lan69a, MR65, Nac68, Ove69, SBK67]. Fibrils [SC67b]. Fibroblasts [DDGL66, EF69, HHM69, KB65b, KH69, MCP69, RK67c, SSSR67, SD68a, SB66, WARS65]. Fibroin [DFM67]. Field [Kol65]. Figures [HK65]. Filament [HS68, KEW67a, Kel67, PH67]. Filamentous [Met69, Phi65, SM67, TWE68]. Filaments [IBH68, KR68, KH69, LB68, LC66, PV67]. Films [DB68b]. Fils [VP69]. Filters [FBS68, MBMB68]. Fine [AHS67, Bak65, BP69a, BB66, Bon68, BC65, BK68, BK69, BP69b, Bri65b, BD68, CPR67, CF66a, CMDLB68, Chr65, Cog66, CB65, DDGL66, DFM67, Ell65, Fah67, FB66, FFS67, Fli65, FBS68, Fri65, Fri66, Gal66, GGF68, GAPRC68, GRH66, GC65c, Gun65, Ham65, Ham66, HHS66, HD66, Hic65, HB67, Ich65, Jen67, JP68c, Kar69, KMK66, Kel66, Kel67, KZ69a, Kes68, KCW67, Lav69, LT67, LA68, MR68, Mat65, MK65a, Moe69a, MMP65, Nan67, New67a, NOP69, Ove65, Pag65, PO65, PO67a, PJO67, PO67b, PS65a, Phi66a, Pic65, Ram65, RK67a, RC67a, RLV65, Rin67, RK67c, Rub69a, Rud65, SOK65, SVR67, She65, SFM67, SK68a, SP68, Sri66, Ste65c, SBK67, TB65, TM69, TMZ68, TP68a, TP68b, Urb69, VR67, WSW65, Wol67, WN65a, YTST65, ZMBB66]. Fine [dWA65, vI66]. First [Gou68a, Lan66]. Fish [MS69b]. Fishes [LWS^{+}68]. Fixation [AS66, EVD65, Gun65, HF68, Mit69, MH67, Nan69, SWR68, SD69a, Sus67]. Fixatives [LS68b, PA67, SC68]. Fixed [CCIC69, LT68a, SWH66, SHG68, vBPB66]. Flagella [HvIvZ66, Phi69, ST68, vIHvZ66]. Flagellar [Bar69, DF66, Phi66c, Rin67, RC67b, RMR69]. Flagellate [Nic68, OK67]. Flagellates [GC65c, RC67b]. Flame [MD66a]. Flight [Smi65, Smi66a, TS65a]. Flow [GGBO69, PNGM66]. Fluid [Bri65a, DB68a, KWWL66]. Fluid-Transporting [DB68a]. Fluorescein [Aro65]. Fluorescein-Labeled [Aro65]. Fluorescence [GL65, KA65, LP66, MU67, MU68]. Focal [MK65a]. Follicle [HA67, KP68, Par67]. Follicles [RD69, UAH68]. Follicular [BW66]. Following [CY69, Fus68, Gun65, HN65, HBT69, JEE69, KS67, MBB67a, MBB67b, Pod68, Sto68, SB68b, TVM69, Vau66, WSW65]. Form [Bas66, SB65d]. Formaldehyde [MBM67, SH67a]. Formaldehyde-Treated [MBM67]. Formation [AC67, And67a, Arn69, Bon69, Bro69, DR66, FH66, Fis67, FW67, Fla68a, Foo68, GTP69, Gou68b, HJ68, HFC68, HKL69, Ish68, IBH69, LC65, Luc65a, MP68, PW67, Phi67, Pla69, Pod68, RLV65, RB65, RI67, SB66, TG69]. Forme [Bas66]. Formica [LR66]. Forming [MM66a, RRG68]. Found [EKG66]. Fowl [MP67b]. Fraction [SH68a]. Fractionation [LRM66, PNGM66]. Fractions [ADC69, BBdD65, BEB67, BB67a, BHB68, KKHE65, Kar65c, SWS68, SK68c, SHMR69]. Fracture [BA66]. Fragmentation [SF69b]. Fragments [BFB^{+}69, DDBE69, HBS69, McC65, WK66, WKG66]. Fraxinus [Sri66]. Free [KHSR66, PA67]. Freeze [ADC69, BA66, Nan68, Nan69, VCM65]. Freeze-Etch [ADC69]. Freeze-Etching [Nan68, Nan69]. Freeze-Fracture [BA66]. Freeze-Substitution [VCM65]. French [Bas66]. Frog [BLRG67, BTD66, BD67, Coh68, DP67, EE68, MMA65, Pag65, Pea65, Ree65b, SB68a]. Frogs [Pic65]. Frozen [BL67, Fis68, LBHT67]. Fruit [GC65b]. Function [Bon68, GW69, GC65c, HSM65, Hic66a, HS65b, IKNG68, KLW68, Løv69, Nic65, NDRFK66, Rud65, SD68a, SB67, SF66, TESW69, UR66a, UR66b]. Functional [BB69a, CKKW66, DB68a, Hes65c, SMW67, SS68c, TGB69]. Functions [FF67]. Fundulus [LT67, TL67]. Fungi [Ais69]. Fungus [Phi66b]. Furrow [Arn69]. Furrowing [TM69]. Further [Hir65, Nac68, Orr65a, SK68c, Sug66]. Fusarium [Ais69, RA65]. Fusiform [Fal69a]. Fusion [Fra65, HFC68, LA68, RSBG66, Urb69].

G [GG69, MSP66, RS67b, TAP67]. Galactosamine [Pog69]. Galactose [NL66a, SK67b, WHL69]. Galactose- [SK67b, WHL69]. Galactose-H [NL66a]. Gallbladder [KWWL66]. Gallinaceum [TSW67]. Gamete [BF68c, Fra65]. Gametophyte [Kel69]. Ganglia [BTD66, BBPM67, MBB67a, MBB67b, NQVF66, Wei68]. Ganglion [CKKW66, Hes65a, HNV67, KT68b, NDRFK66]. Ganglionic [Lan68]. Garter [Hes65b, HMW66]. Gas [SK68c]. Gastric [Rub69a, Sed69]. Gastrointestinal [FOP^{+}69]. Gastrula [Bak65, LT67]. Gastrulation [TL67]. Generation [AB65a]. Genetic [SD68a]. Genetically [Com67]. Geometry [DB68a]. Gerbil [NC69]. Gerl [HNV67]. Germ [ZG68]. German [Gir65, Moo66]. Germinal [SH65]. Germs [KEW67a]. ghost [MP67a]. Giant [MR68, Met69, MI69, Phi67, SW66]. Giardia [Fri66]. Gl-C [Fra67]. Gland [AOS69, BP69a, BW66, DS69a, DL69, EE69, GWS65, HS66, KCR67b, LS65, LN65b, MMNT68a, MMNT68b, NP65, Phi65, Riz67, SS65a, SF66, SS68c, TMMN68a, TMMN68b, YB67]. Glande [Bea68]. Glands [BB69a, BBB68, CM69, Ell65, Fri65, KGO69, PS65a, SS66b, TR69, TMK67]. Glandular [YB67]. Glial [HD66]. Gliosomes [Sre65]. Glis [SOK66]. Globular [Kel66]. Glomerulosa [GWS65]. Glucagon [DdD67, DBdD67]. Glucose [CL66, Jer66, NL66a, NL66b, NDRFK66, OE66b, Tro67, dWA65]. Glucose- [OE66b]. Glucose-H [CL66, NL66a, NL66b]. glucuronidase [HSC68]. Glucuronylating [ZOE67]. Glutamic [LT68a, LT68b]. Glutamine [EBH67, JEE69, PM65]. Glutamine-Deficient [EBH67, JEE69]. Glutaraldehyde [AS66, HF68, SH67a]. Glyceride [Jer66]. Glycerinated [Gib65a, RB66]. Glycerol [CFK65a, CFK65b]. Glycerol-Extracted [CFK65a, CFK65b]. Glycerophosphate [OVF69]. Glycogen [BHP65, CL66, Fla68a, Gar68, IP65, Len66, RC68, WD68a].

Glycogen-Membrane [Gar68]. Glycogenosis [HS67b]. Glycolysis [Cha65]. Glycolytic [FK66]. Glycoproteins [NL66a]. Gnat [Phi66b]. Goat [KO65]. Goblet [Ber67b, NL66b, TT69]. Gold [DK67]. Gold-Thiocholine [DK67]. Gold-Thiolacetic [DK67]. Goldfish [MO68]. Golgi [AG68, Ber67b, BH67, Bro69, Buc66, CMM66, EZ68, FFO69, Fli69, Fri69, Hic66a, KB65a, Lan68, Man69, Mol65, MM66a, NL66a, NL66b, RHL69, WB69]. Gondii [VP69]. Gossypium [Yat65]. Gracilis [Blu65, Lym67, MEK68, SL66, SB65c, Som65]. Gradient [SM65a]. Gradients [MK65b]. Grain [SP67a]. Grains [Tak65, Tak68]. Granular [KHH68, MS69b]. Granule [AS66, Can68, FH66, HKL69, SR68, Wol67]. Granules [And68, AF68, BHdD69, BF66, BF68a, BF68b, Bel69, CYD69, CM69, FH65, FC68, GC66a, GB69a, GHL69, HFL69, HYL69, Lav69, Len66, LC65, LC68, Man69, MP65, Mat65, MdHP66, PRGC69, Sho66, TKH69]. Granulolysis [OJP^{+}68]. Graphic [Oka67]. Graphite [DB68b]. Grasshopper [Lan68, Lea68, Moe69b, Ste65c]. Green [LF65]. Greening [SK68a]. Ground [Nac68]. Group [TSACBB69]. Grouping [BMH68]. Growing [HS65a, KI67, Ray67, SSW67]. Grown [DNC69, Pla69]. Growth [BLT69, CG65, EKSG69, Gol68a, HS65a, Hol67, JL68, LS68a, LK67a, LP67, Løv69, MCP69, Oka67, Par69b, Sto69, TMZ68, TMK68, WO67a, WF68]. Guinea [Chr65, FH65, FC68, HH67, Mer68, Non68, PCS69, Rob69b, SP66]. Guppyi [DR66]. Gut [AS67, Cob69]. Gyrodinium [KR69b].

H [TVP67, CL66, Com66b, MO69, NL66a, NL66b, Rez69, SP67a, Sco67, Sim67, SS65c, SK67b, Sti67, SMP65, VM66b, WD68b, WHL69]. H-Labeled [WD68b]. Habenula [MP66]. Haemanthus [HJ68]. Hagen [SS65a]. Hair [DFW66, Par67, Par69b, UAH68]. Hairs [DFW66, NB65]. Halobacterium [SR67b, SK68c]. Halobium [SR67b, SK68c]. Hamster [BRH65, Bri65b, CYD69, EKSG69, ETS68, ET69, GS67b, HT65, KW65, PP69, RS67b, SK67b, Sti67, SB66, TAP67]. Healing [RB65]. Heart [BD69, MS69c, RK67b, SJ68b, SRZB66, SS69c, WHT67]. Heat [SB67]. Heat-Sensitive [SB67]. Heavy [Aro65, IBH69, MK65a]. HeLa [BKS69, BFB^{+}69, BN65b, Bor68, BTG69, CP67b, DSDD65, DK68, FC67, GCG67, GB69b, HRS69, HA66, JH65, RJM68, RM69, RS69, SWP69, SVWD68, Sus67]. Helical [Ech65, KH69, Met67]. Heliozoa [THM66, TP67, TB69]. Helleri [Mat65]. Heme [GL65]. Hemidesmosomes [Kel66]. Hemiptera [War68]. Hemoglobin [EC68, GW66]. Hemolysis [See67]. Henle [SM67]. Hensel [EP68]. Hepatectomy [Tro65]. Hepatic [ALEW66, Bie66, Fah69, KO65, LT68a, NdM66, RCE69, SF65, SGF68, SA66, TESW69, dMN67]. Hepaticae [CK67, CK68a]. Hepatitis [DFM65]. Hepatocytes [TMK68]. Hepatoma [Fla68a, Fla68b, Pla69, RTL69]. Heterochromatin [Him67a, Sim67]. Heteroclitus [LT67]. Heterogeneity [DBN68]. Heterogeneous [DDBE69]. Heterophil [BHdD69]. Heterotrichous [DH66]. Hexagonal [BE68, RK67b]. High [BF69, Car69b, HH66, Lan65a, Lan66, OVF69, Orl68, SC67a, Sim67, SK67b, Sti67, ZR65].

High-Resolution [Car69b, Orl68, SC67a, SK67b, Sti67]. High-Yield [BF69]. Higher [HBS65]. Hillock [PSPO68]. Hippocampus [Ber67a]. Hircus [KO65]. Hirsutum [Yat65]. Hirudo [LF69]. Histamine [MSL69]. Histamine-Type [MSL69]. Histochemical [LT68a, MHA65]. Histochemistry [LWS^{+}68]. Histogenesis [KZ69b]. Histological [WCJS67]. Histone [BMBW67, GW67a, HEL67, KT68b, PAM67, Pre66, Vau66, Sto69]. Histone/ [Sto69]. Histones [BA65a, BA65b, BA67a, BA67b, Com67, Vau66, HA68]. Holes [See67]. Holographic [Fel69b]. Homogenates [Kir69]. Homogeneous [BHB68]. Homologous [EV65]. Homoptera [Smi65]. Homopteres [Gou68b]. Honeybee [PB69]. Hormonal [GC66c, Ich65]. Hormone [BLT69, BN65b, Bor68, GH65, GTLE65, TKH69]. Horn [WST68]. Horseradish [CK68b, MHA65]. Host [EC67]. House [KMK66]. Housefly [PR69]. Hsm [BHP65, MSP66]. Hudsonius [Ber67a]. Human [AM69, AH69a, AH69b, AC67, AB65b, BH67, Bie66, Cav66, Com66b, EF69, FH65, GB69a, GC66c, HOZ68, HA67, HF68, Hol67, HHM69, KS67, MCP69, Mat67, Mit69, OR68, PHP67, Ram65, RC66, RO68, SSSR67, SW67, SBK67, TR69, Ter65, Wil65b, ZMBB66]. Hyaloid [Las67b]. Hyaluronic [BSH68]. Hydei [BB69a, BBB68, Hes65c]. Hydra [Len66, MPD65]. Hydrochloride [Pog69]. Hydrocortisone [GB69b]. Hydrodynamic [RB66]. Hydrolases [Maa67]. Hydrolysis [BS66b, Lan67a]. Hydrostatic [Lan65a]. hydroxy [LH67]. Hydroxylating [Orr65a]. Hydroxylation [Orr65b]. Hydroxypropyl [LH65]. Hydroxytryptamine [RCV69]. Hyla [Bak65]. Hymenolepis [RC68]. Hyperthermic [RE69]. Hypertrophy [KHH68]. Hypnoides [Urb69]. Hypolipidemic [SA66]. Hypophyse [TVP67]. Hypothermia [Bye66]. Hypotonic [See67]. Hypoxia [CYD69]. Hypoxic [RE69].

I-Albumin [MBM67]. Icosahedral [dZG69]. Identifiable [CKKW66]. Identification [JT67, MPO67, RCV69]. Identified [RCV69]. Ii [HS67b, TVM69]. III [Oll68, PVM69]. Ileal [WG68b]. im [Moo66]. Images [Fel69b]. Imaginal [FM65, Rob69a]. Immature [GC65b]. Immediate [Hol67]. Immunochemical [BB69b]. Importance [HH66]. Improved [Gil65, HP66b, SdZ67]. Inactive [Com67, GW67a]. Inclusions [AS67, MS69b, Riz67, WR68]. Incomplete [Beh67]. Incorporation [Bry66, Cav66, GK68, HB69a, HM67, ISS68, Par65b, SS65c, WHL69, vBPB66]. Increase [Løv69, ZOE67]. Increased [HH67, ML66b]. Incubated [GAPRC68, ML66b]. Incubation [JEE69]. Indentation [DS69b]. Indirect [Wen66]. Individual [DE69, Mer68]. Individus [VP69]. Individus-Fils [VP69]. Induced [AC67, And67b, BF65, BA65a, BA65b, BA67a, BA67b, Bye66, DBdD67, EB66, Fel69b, For65, GGF68, GS67b, Gol68a, HC69a, HK65, HZL66, HC68, HB69b, Kah68, KLW68, KGO69, KRC69, KH69, MSL69, NAL^{+}67, OYHM65, OEE65, PaSN65, PCS69, PS65b, RAH69, SGF68, SF69b, SK68b, SD68b, SS69c, WR68, WLKS68]. Inducer [DdD67]. Induces [Hol67].

Inducibility [HS65b]. Induction [GC66c, Orr65a, OE66a, OE66b, Pad66, TM69]. Infected [CHE66, LF65]. Infection [SB68b]. Influence [DdD67, Luc65b, SLAR69]. Influenza [SFG67]. Infusionum [Rud65]. Inhibited [Sus65]. Inhibiting [DSDD65]. Inhibition [Gou68a, Jag69, JLL68, Lym67, RC69, Sto68, Tay65a, WPP69, YF69]. Inhibitor [RDS^{+}69, RB69b, WB69]. Inhibitors [BH68, DS69a, Kue69, WST68]. Initial [FJH65, PSPO68]. Initiation [SD68a]. Injected [Bri65a, LPB^{+}68, NL66b]. Injection [Tro67]. Injections [CY69]. Injury [MK65a, Rey65, SR69]. Inner [DFW66, FD65, PWFP68, SG68]. Innervation [BR67, Mun65, Rob69b]. Inorganic [LT69, SB69a]. Insect [HS67a, LC65, LC68, Phi69, Rob66, ST65, SS65b, Smi66b, SS68c, TM67, TP68a, TP68b]. Insects [HSS67]. Insignis [LK67b]. Insoluble [KS69]. Insulin [BLD^{+}66]. Intact [Flo69, HCGP67]. Inter [dZG69]. Inter- [dZG69]. Interaction [OS69a, OS69b]. Interactions [Fel69a, GP67, GP68, PG68]. Intercalated [RSL69]. Intercellular [JL68, LK67a, LP67, Ove69, RK67b, She68, ZG68]. Intercisternal [AG68, Mol65]. Intercostal [KZ69b]. Interference [BTVM69, MI69]. Interferon [JLL68]. Intermediate [IBH68]. Intermediate-Sized [IBH68]. Intermittent [PVM69, TVM69]. Intermittently [VM65, VM66b]. Interneuronal [LR66]. Internodal [BD67]. Interpeduncular [MP66]. Interphase [Fel68, Him67b, HS65b, KB67, TMZ68]. Interpretation [PVT68, Mil69]. Interrelationships [Tha66a]. Interruption [SWP69]. Intersegmental [Smi66b, TP68a, TP68b]. Interstitial [Chr65, FC68]. Intervals [Tro65]. Intestin [Gou68b]. Intestinal [Ber67b, BP69b, BT68, CBP67, CT67, CP69a, CP69b, EC65, Ito65, JB68, Jer66, KC69, Lan65b, MG67, OEC65, Ove68]. Intestine [DS68, EB66, KW65, MW67, SK67b, Sti67, VF67]. Intimately [BR69]. Intoxicated [GK69]. Intracellular [AS67, ALEW66, CP68, GMB65, GL65, GH65, HS66, JP67a, JP67b, JP68a, JP68b, RE69, SS67b, SS67a, SS68b, SS69b, UR66a, UR66b, WRJ67, WR68, Zal69, dZG69]. Intraepidermal [Mun65]. Intramitochondrial [IP65, Len66, Oll68, SM67, Wil67]. Intranuclear [Boq69, Buc66, Col65, Eve68a, Eve68b, Kar65b, Kes65, Lan69a, SVD68]. Intranucleolar [KB68, Rod69]. Intrapopulation [SM65b]. Invadens [MSW68]. Invaginating [HMW66]. Invertebrate [Ste65b]. Investigation [Bar69, Too65]. Investigations [IKNG68]. Involution [BA69]. Involvement [Par69a]. Iodide [NKKW69]. Iodine [RD69]. Ion [HC69a, LL69a, NM67, PWFP68]. Ion-Induced [HC69a]. Ionic [IKNG68, LSL68, LG65, Pen66]. Ionizing [MK65a]. Ions [Bul69, LT69, TS69]. Ipomoea [Ech65]. Iron [MD69, OYHM65, WWS66]. Iron-Containing [WWS66]. Irradiated [FB66, Pic65, SAW^{+}66]. Irradiation [BOR69, For65, GSS69, MBB67a, MBB67b, MCK65, SOZC68]. Irreversible [dWA65]. islet [Boq69]. Islets [GHL69, HKL69, HFL69, HYL69, MC65]. Isolated [BHP65, BP65, BE65, BE68, BB69a, BF69, BB69b, BLS67, BP69b, BRGG65, BRG68, CMM66, Fra66, GGM^{+}68, GK69, GGBO69, HC69a, Hof65b, HCGP67, HKL69, KT68a, Koe67, Koe68, Met67, MYS^{+}67, PS66, RS67a, Rys67, SS68a, SSS68, SSWG68, SS69c, Tay66, Win67, Wol65, Wol68]. Isolation [Arg65, BKS69, BFB^{+}69, CM69, DDB69, EEKS65, FFO69, FM65, Gil65, GC65b, GHL69, GCG67, HTBM65, HB66, HK65, HBT69, HFL69, HYL69, HA66, Jac68, KS69, KK67, LS65, LRM66, MRH^{+}67, McC65, MSW68, RAH69, SH68a, ST67, SWDK67, SK68c, TCV69, WHT67]. Isoptera [SS65a].

Japanese [Ham65]. Japonicus [Ham66]. Jejunal [SWR68]. Junction [DK67, Len69, NSHW68, RB67]. Junctions [BE68, BR69, BL68, FP65a, LR66, Las67a, Las69, RK67b, She68, SFM69, TRH66]. Jute [Hey66].

K-activated [MP67a]. Katherinae [HJ68]. Kb [LN65a, Tan69]. keratin [Mil69]. Keratinization [FE69]. Keratinizing [MP65, Par67]. Keratinocytes [Bel69]. Kerverni [Car67]. Kidney [Bul69, LF65, MHA65, PM68, PM69, SPT65, SCP67, SM67]. Kidneys [GK69]. Killing [EKSG69]. Kinase [HS65b]. Kindtii [WG65]. Kinetic [KS66]. Kinetics [Fab68, Hof65a, HS65a, SM65b]. Kinetoplast [BD68, HA69, Sim68]. Kinetoplastic [SvA67]. Kinetosomes [Arg65]. Kinocilium [FD65]. Knowledge [Fle65]. Kühniella [Ove66].

L [BW67, Ech65, SD68a, Yat65, Cog66, HA68, HC69b, JLL68, KB65b, KHH68, KH69, LH67]. L- [LH67]. L-Cell [KH69]. L-Strain [SD68a]. L. [BP68a, Fus68, Met69, MI69, ST65, Smi66b, Sri66]. L5178Y [WO67a, WO67b]. Label [Ebs67, Fus68]. Labeled [Aro65, BS69, Cav66, CY69, Jer66, NP67, PVM69, RDS^{+}69, RB69b, SP67a, Tay65b, WD68b]. Labeling [GMGFLS65, MO69]. Labial [TR69]. Lacrimal [YB67]. laevis [Ove65]. Lamellae [Eve68a, Eve68b, GC66a, Kes65, Kes68, KB69, KHH68, LP66, ML66b, Par65a, SF69a]. Lamellar [Lin67, NLS67]. Lamellated [LL66]. Lamellibranch [Cob69]. Lamina [Par69b, Ter67]. Lamouroux [Urb69]. Lamprey [Jas67]. Langerhans [Bel69, Can68, GHL69, HKL69, HFL69, HYL69, MC65, Mis66, NKKW69, SR68, Wol67]. Langmuir [Wol68]. Lanthanum [Can68, Les67, Ove68, RSL69]. Lanthanum-Staining [Les67]. Lapse [CP67b]. Large [GC65b, LPB^{+}68, MW67, SKHR68]. Large-Scale [GC65b, LPB^{+}68]. Larvae [FM65]. Larvaires [Bea68]. Larval [BBB68, DL69, NP65]. Laser [BOR69, GSS69, ROJ65, SAW^{+}66, Tan69]. Late [MCP69, SD67]. Lateral [Bon69, FD65, Ham65, SP68]. Latex [KW67b, WK69]. Lattice [HS68]. Layer [Die67, Kel66, Les67]. Layered [KM65, KWM69]. Layers [FW67]. Lead [FP65b, GK69, Pod68, VC65]. Lead-Intoxicated [GK69]. Leading [Cam65]. Leaf [FN69, Ric68]. Leaves [CHE66, GPB68, Jac68, SK68a]. Lebendnachweis [Gir65]. Lebistes [DR66].

Lecithin [SS69b]. Leech [LF69, Met67]. Leg [RC67a]. Leishmania [Sim68]. Length [BLRG67]. Lens [Ove65]. Lepidopteran [RC67a]. Leptodora [WG65]. Lesion [FJH65]. Lesions [WRJ67]. Lethal [DK68]. LettrÉ [GK68]. Leucine [TVP67]. Leucophaea [Hag66]. Leukocyte [MdHP66]. Leukocytes [BHdD69, BF66, BF68a, BF68b, Cav66, GKS^{+}67, HF68, PAM67]. Level [Rub69a, WSB68, vdD67]. Levels [Lan66, ZOE67]. Lewisi [BD68]. Libellula [KB69]. Libinia [HC69b]. Life [Oka67, RM69, RS69, SCP67]. Ligament [CSJ67, WCJS67]. Light [Aro67, BBPM67, Gib65b, Ken66, Laf65, NP67, NdM66, PaSN65, SdZ67, dMN67]. Light-Induced [PaSN65]. Lignification [LG65]. Like [FW67, MMB65, SVWD68]. Lilium [Moe69a]. Limb [SM67]. Limited [CR66, DNC69]. Limulus [Las67a]. Line [FD65, GS67b, Ham65, KC68, MO69, SHR67, SHMR69, WSB68]. Lines [Sch69]. Lining [KCW67]. Linkage [Coh68]. Linked [Hac66, Hac68]. Linoleic [BT68]. Lipid [BF65, BT68, DNC69, FFS67, Gol68b, LC66, MH67, NLS67, SWR68, SWH66, SS67b, SS67a, SS68b, SM65c, Tro65]. Lipid-Containing [SWH66, Tro65]. Lipid-Depleted [FFS67]. Lipid-Limited [DNC69]. Lipids [Kor67, Mit69, USK69]. Lipofuscin [FC68, GB69a, SOK65]. Liquid [TCL65]. Liver [AMH69, AP69, BF65, BBdD65, BE65, BE68, BF69, BH67, CG66, CG68, CL66, DBN68, DB65, DdD67, DBdD67, Fab68, FFO69, Flo69, Gil65, GB66a, GB69a, Gol68b, GC66b, HL66, HT66, HK65, HCGP67, IS69, JEE69, Kue69, LT68b, LPLD69, Lou68, MBM67, MYS^{+}67, Orr65a, Orr65b, OE66b, Pen66, PLD69, RK67b, Rey65, SS68a, SH68a, SD69a, SB69a, SG68, SR69, SRRK69, SKEB67, SHW69, SWS68, SS69b, SSGF69, SB65d, SS69c, SH67b, Tro65, Tro67, TMF66, VG68, VF69, WSGH69, WLKS68, Wil65b, vdD69, HSC68]. Livers [Kar69, LPB^{+}68, RAH69]. Living [Baj67, For65, MCK65]. Loaded [TG68]. Lobster [Men69]. Local [For65, WRJ67]. Localization [BSH68, Ber67b, BB66, Bul69, DK67, DS68, FK66, Fah69, FP66, Fis68, FC68, FN69, Gol67, GL65, GH65, GCG67, HH67, HB67, KR68, KA65, Lan67a, LL69a, LP66, MT67, McC69, Nag65, NP67, RK67a, RC68, RB67, SEG67, SB65b, SP67b, ST65, SB65c, Sus67, Tak68, TS65a, Tor65, WSW65, WWS66, MP67a]. Localized [Ove68]. Located [SB67]. Location [NM67, Zad66]. Loci [GW67a]. Locust [DSA65]. Loew [For65]. Loligo [Arn69, Met69, MI69]. Long [BBPM67, HN66, SOZC68]. Long-Term [BBPM67, SOZC68]. Longitudinal [Rob69b]. Looping [Kel67]. Lors [VP69]. Loss [AH66a, MH67, SWR68, SC68, SvA67, VM66a]. Low [RAH69, She68, SR67b, HA68]. Low-Resistance [She68]. Lugubris [LR66]. Lumbricoides [Foo68, Ros65a, Ros65b, Ros69b]. Lumbricus [AWE67, Cog66]. Lung [EF69, LL66, Rei66, RS67b, VRN^{+}68]. Lungs [KCW67]. Lutein [FB66]. Luteum [AH69a, AH69b]. Lyciades [RC67a]. Lymphatic [LB68, WA65]. Lymphocyte [BHHW69]. Lymphocytes [GL68, MMP65, PCS69, Pog69, RBLH68, SR67a, TMZ68]. Lymphoid [BA65a, BA65b, BA67a, BA67b, BFdD67, BdD67a, BdD67b, SS69a].

Lyncozymba [Ham65]. Lysed [CG66, SK68c, VG68]. Lysine [Cav66, Sus65]. Lysine-Containing [Sus65]. Lysis [CG68, See67, SR67b]. Lysolecithin [See67]. Lysophosphatides [SWS68]. Lysosomal [CB69a, HB67, KKHE65, SPT65]. Lysosome [SF66]. Lysosomes [BP65, BFdD67, BdD67a, BdD67b, BHDW68, BHHW69, DdD67, DBdD67, HBHW68, HN65, HNV67, HBT69, HS67b, LPB^{+}68, Mor68, SPT65]. Lysozyme [Maa67].

M [KC68, SHMR69]. M-Line [SHMR69]. Macromolecular [Cam65, RM69]. Macronuclear [Pre66, SMP65, WF68]. Macrophages [BA69, EC68, HFC68, Par69a, SD69b, SW66]. Maculosus [KP68]. Maderae [Hag66]. Magnesium [DP67]. Maintenance [Fra67, HT66, Rob69a]. Maize [Him67a, Jac68]. Malarial [AHHS66, AB68]. Male [Sho67, TWE68]. Malhamensis [SLAR69]. Mammalian [BBFA67, BF68c, Beh67, BM69, BHB68, GMB65, HP66a, HN66, HP69, HAKB65, KK67, KB67, KL65, Kle69, Lan67a, LSL68, LL69a, Oka67, RP66, RSL69, SWDK67, Stu65, SE66, Too65, WO67a, WO67b]. Man [Ell65, Gol67, MdHP66]. Mannose [WHL69]. Mannose- [WHL69]. Marine [Bro69, RS67a, SS65c, Urb69]. Marker [CB69a]. Marmorata [She65]. Marrow [HN66]. Marsupial [BH69b]. Mass [BBB68]. Mast [Cha65, CLB65, Com66a, FH65, GL67, Pad66, Pad69]. Material [AG68, Fel65, GGK66, LL66, Les67, Ove69, SHR67]. Materials [Cho65, HN66, TMMN68a]. Mating [EZ68]. Matrix [And69, Hol68, MPO67, PRGC69]. Maturation [Com66a]. Mature [CE67, MU68]. Mealybug [RR69]. Measles [NAL^{+}67]. Measurement [MPNK69, SFG67, SG69a, SSG69]. Mechanical [Hac66, Hac68, Men69]. Mechanism [AHHS66, BT67a, BT67b, BFDP68, CT67, DN66, GC66c, HMS65, KR69a, OBJ^{+}68, OBJ69, Orr65b, PS65b, SK67a, SK68b, SH68b, Tay65a]. Mechanisms [SB68a, Vae68]. Media [HP66a, Rob69a, SR67b]. Mediators [MSL69]. Medicinalis [LF69]. Medium [Amm69, EBH67, HP66b, JEE69, LH65]. Medulla [CvD67, PVM69, TVM69, VM65, VM66a, VM66b]. Megaterium [FJH65]. Megoura [Smi65]. Meig [GB66b]. Meiosis [Jag69, Tay65b]. Meiotic [Gou68a, Moe69a, RI67, WH66]. Melanogaster [FM65, Fri68, HP68, MM66b, NP65, Rob69a, Sho66, Sho67]. Melanogenesis [McC69]. Melanoma [Sil69]. Melanoplus [Lan68]. Melanosomes [Mis66]. Melting [Aro66]. Membranaire [VP69]. Membrane [ADC69, BBFA67, BFB^{+}69, CP67b, DDBE69, FW67, Fla68a, Fra65, Gar68, GB66b, HB66, HBS69, JT67, KEW67b, Kar65c, KHSR66, Les67, MP65, OE66a, PWFP68, Pla69, PM69, SKK68, See67, SKEB67, SS65d, SWS68, Sug66, SBK67, SB68b, Ter67, WK66, WKG66, WG68b]. Membrane-Bound [KHSR66, PM69]. Membrane-Coating [MP65]. Membrane-Envelope [HBS69]. Membrane-Glycogen [Fla68a]. Membrane-Like [FW67].

Membranes [BF68c, BE65, BE68, BR69, BDRN65, DSP66a, DSP66b, DBN68, DC67, Fal69b, FFO69, Fli68, Fra66, KB68, KL65, LS68b, MCK65, OSP67a, OSP67b, OEE65, OE66a, SDV68, SG68, SWH66, SD68b, SRRK69, SK68c, SE69, WG68a, WSL65, MP67a]. Membranous [Nic68]. Menziesii [BB67c]. Mercuric [PCS69]. Meristem [BMBW67]. Meristematic [Laf65, LL69b]. Meristems [PS65b, WD68b]. Meromyosin [Aro65, IBH69]. Mesenchymal [LL65]. Mesenchyme [GTP69, TG69]. Mesentery [DS68, NOP69]. Mesodermal [PK69]. Mesosomes [Nan68]. Mesothelium [CK68b]. Messenger [HRS69]. Metabolic [GKS^{+}67, dFM66, SK68a]. Metabolically [Hac66, Hac68]. Metabolism [Bor68, Pla69, Rei66, SSGF69, War69]. Metabolizing [OEE65, OE66a]. Metachromasy [FW65, LCW65]. Metakinesis [Har65]. Metanephric [PM68]. Metaphase [HA66, RJ69]. Metaphase-To-Anaphase [RJ69]. Methacrylate [LH65]. Method [BKS69, BA66, Gil65, GMGFLS65, GCG67, Koe68, LHS^{+}65, MU68, Oka67, RC66, SWH66, SE67, Wen66, Wol68]. Methods [DK67, Pep66, RDS^{+}69, RB69b, WKS66]. Methylcarbamate [BP68a]. Mg [MP67a]. Mice [HBT69, KSR67, KCR67b, MW67, SOK65, SOZC68, dMN67]. Micro [BOR69]. Micro-Irradiation [BOR69]. Microanalysis [RD69]. Microbeam [For65, MCK65, Ste65a, SAW^{+}66, Tan69]. Microbodies [Afz65, AF68, BN69, Bie66, Fah69, FN69, RBS69, SA66, SGA67, SAR69, TMF66, TMK68]. Microcrystalline [Hey66]. Microcylinders [Lin65]. Microdissection [IS69]. Microextracts [SE67]. Microfibril [NB65]. Microfibrils [Ove66]. Micrographs [Ber69, SB69b]. Microincineration [TG68]. Micromanipulation [NK69]. Micronucleaires [PVT68]. Micronuclear [RP67]. Microscope [BS67, BS69, BBP65, BBPM67, CB66, CL66, DFM65, DS68, DR66, EV65, Fis67, FC67, GC66b, GSS69, GBW69, HK65, Laf65, LR66, LT68b, MD69, MPD65, Nac68, NL66b, NdM66, RL67, Sal67, Sal69, SBS69, SdZ67, Sco67, SAW^{+}66, TVP67, TRH66, WSB68, WA65]. Microscopic [And67b, AP69, BEB67, BB67a, BE65, BA69, CY69, DK67, HSS67, IKNG68, JT67, Kar65a, NP67, NM67, SS68a, SC67a, SD67, SKHR68, ST65, SH65, SB68b, VHR65, dMN67, HSC68]. Microscopy [BTVM69, BA66, BT68, Cho65, DB68b, GB66b, HH66, Joh69, KA65, KCR67a, KCR68, LH65, LS68b, LP66, MU68, MI69, Mit69, MH67, Nag65, Ree65a, Rot67, SD69a, SB65a, SB65b, SFG67, SG69a, SSG69, TG68, Too65, VC65, VF69]. Microsomal [DBN68, OEE65, SP66, TSACBB69]. Microsomes [BD69, DN66, IS69, Orr65a, Orr65b, dBVSH65]. Microspectrophotometric [She67]. Microspectrophotometry [TCL65, Wol69]. Microspikes [Tay66]. Microtubulaire [Bas66]. Microtubular [Bas66, BM69, RR69, SS65b]. Microtubule [Beh67, Gal66, NB65, SS65a, SC68, vIHvZ67]. Microtubule-Associated [SS65a]. Microtubules [Bar69, Beh67, Bou69, EP68, GTP69, HT65, HJ68, HC68, Kir69, MP67b, Ove66, PV67, Rob66, ST67, ST68, Tay66, THM66, TP67, TG69, TB69, Wil69, Wol65]. Microvilli [BP69b, FC67, Ito65, MW67]. Midgut [AH66b, SVR67]. Migration [AP69]. Millipore [FBS68, MBMB68]. MinÉRales [Gou68b]. Minor [GPB68, HG69]. Mirabilis [HvIvZ66, vIHvZ66]. Miscible [LH65]. Mise [Che67]. Mite [Aro67]. Mitochondria [BB67b, BRH65, CG66, CG68, CWL67, DNC69, Die66, FFS67, GB66a, Gol68b, GK69, GC66b, GTLE65, Hac66, Hac68, HC69a, HL66, HA69, JEE69, JKL68, Kah68, KBL69, KSB65, LT68a, LPB^{+}68, Luc65a, MYS^{+}67, MC69, NSHW68, PR68, PRGC69, SB69a, SG68, Sla65, SKEB67, SS69c, Tan69, Ted65, TG68, WHL68, Wil65b]. Mitochondrial [Ash65, BB67c, CA68, EH68, HW67, KLW68, Lin65, LDRA67, Luc65b, OVF69, OYHM65, PWFP68, Par65b, SK68b, SB65d, WD68c, ZOE67]. Mitocondria [VG68]. Mitogen [BHHW69]. Mitomycin [DK68]. Mitoses [Din69]. Mitosis [AR67, BH69a, CBR65, GW67b, HRS69, IBH68, KB67, Kra67, Lea68, Man68, MMP65, Nic65, Rot67, RSBG66, SW67, Ste65c, Tay65a, TM67]. Mitosis-Related [SW67]. Mitotic [Ais69, Baj67, Bar67, BB69b, BH68, Bri65b, DSDD65, Hir65, Hof68, HAKB65, JH65, Kan65, KF65, Kan67, KS66, KB65b, KWM69, RS67a, RJ69, SM65b, SWDK67, Ste67, Wen66, Wil69, WST68]. Mixtures [BHB68]. Mobility [CP67b]. Mode [Gou68b, SLAR69]. Model [RB69a, Rub69b]. Models [Kel67, SE69]. Modifications [Bea68, WG68b]. Modified [Cob69]. Modulations [PB68]. Mold [CBR65, GG68, NA65]. Molecules [WD68c]. Molischianum [GSW65, HF65a]. Mollusc [Cob69]. Molting [TP68a, TP68b]. Mongolian [NC69]. Monkey [LF65, SCP67]. Monoamine [SEG67]. Monocytes [SW66]. Monolayer [SFM69]. Monolayers [GL68]. Mori [MMNT68a, MMNT68b, TMMN68a, TMMN68b]. Morosus [SS65b]. Morphogenesis [All69, Gre65, KZ69a, Ove66, TMK68]. Morphogenetic [EF69, RCE69]. Morphogenic [Bak65]. Morphological [BBdD65, BRGG65, BRG68, CKKW66, GKS^{+}67, KK67, KS66, MO68, MYS^{+}67, SR67b, SS68c, dBVSH65, VU68]. Morphology [BP66, Fra65, Rod69, SGR67, SE66]. Morphometric [SHW69, WKS66, WSGH69]. Mosaic [EC67, Kla68, KZvW65, WRJ67, WR68]. Mosquito [SVR67]. Mosquitoes [TSW67]. Moth [Sta65]. Mother [Ech65]. Motility [Rob66]. Motion [Fel69b, Rin67]. Motion-Induced-Contrast [Fel69b]. Motor [Ber67a, Hes65b, KZ69a]. Mounting [Kol65]. Mouse [AMH69, AC67, Asb67, AB65a, BS66b, Bry66, DFM65, Din69, EB66, Flo69, Fri65, Fuj67, GGF68, Gar68, HL66, HP66b, HB67, Jag69, KHSR66, MBM67, MG67, Mor68, NdM66, RK67b, SD69a, Sil69, SH65, TESW69, Tro65, Tro67, TT69, WWS66, WARS65]. Movement [ECH67, HC68, RR69]. Movements [Bak65, Pad66, TL67]. Moyen [Gou68b]. Mucocysts [TS65b]. Mucopolysaccharides [Ber67b, NL66a]. Mucosa [FOP^{+}69, HS66, HB67, JB68, SWR68, SBK67]. Mucosubstances [WWS66]. Mucous [JT67]. Mucus [Lav69, NL66b]. Mul [OBJ^{+}68]. Multilayering [EF69]. Multimicronucleatum [OBJ69]. Multinucleate [KRC69]. Multinucleated [SW66]. Multiperforate [RA65]. Multiple [FW67, KM65, KWM69, Moe69b]. Multiple-Layered [KM65, KWM69]. Multipurpose [Ros67a, RKC68]. Multivesicular [Fri69, LC68]. Murine [DS69a]. Muris [Fri66]. Musca [PR69]. Muscle [Aro66, Aro67, BD69, Ber67a, Ber68, BRGG65, BLRG67, BRG68, CB69a, CFK65a, CFK65b, CMDLB68, CB69b, EEKS65, EE68, ENQ65, EI67, FK66, Faw68, Fis67, FS69, Gar68, GP66, Gol68a, Gol67, GTLE65, Hag66, HS68, Hes65b, HMS65, HMW66, Ish68, IBH68, JS67, Kel67, KR68, KZ69b, KR69a, KC68, Lan65b, Lan67a, Man68, Men69, Mer68, MMA65, Non68, Pag65, PH67, PR69, Pod68, RC67a, Rez69, Rob69b, Ros65a, Ros65b, Ros67b, Ros68, Ros69a, Ros69b, RK67c, SVR67, SM69, SM65a, SFM67, SFM69, Sla65, Smi65, Smi66a, Smi66b, SJ68a, SB68a, SH68b, SRZB66, Ste65b, Ste65a, SGR67, SHR67, Tha66b, Tha66a, WS65, WSB68, WS68, WD68a, Zad66]. Muscles [Fah67, Jas67, RDS^{+}69, RB69b, TS65a, TP68a, TP68b]. Mutant [Fri68, Sho66, Sho67]. Myelin [HB69a, HK65, HZL66, HD67, Hir68, KA65, NLS67, Ros66]. Myelinated [BD67, Hes65a]. Myelocytes [FH66]. Myocardium [FM69, LSL68, LL69a, MF69, ML66b]. Myocytes [Man69]. Myoepithelium [Ell65, YB67]. Myofibril [Fis67, Pep66]. Myofibrillar [TS65a]. Myofibrils [SGR67]. Myofilament [BLRG67, Hag66]. Myofilaments [Non68]. Myogenesis [BH68, War68]. Myogenic [BH69a]. Myoglobin [Gol67]. Myoplasm [MMA65]. Myosin [CB66, KR68]. Mytilus [Amm69, ND65a, ND65b]. Myxomycete [MR65].

N [Pog69]. N-Acetyl-D-Galactosamine [Pog69]. Nadi [SPW^{+}67]. Naegleria [DF66]. Nalidixic [Lym67]. Native [BHP65]. Nature [Col65, HMW66, MM65, PRGC69, SKN65]. Necrosis [EB66]. Necturus [KP68]. Need [SH67b]. Neff [BK69, WK69]. Negative [ADC69, Bar67, CMM66, Hey66, SF69a]. Negatively [Bar69, BE65]. Negri [MM65]. Nematoda [Foo68]. Nematode [Jam66]. Neonatal [MYS^{+}67]. Neoplastic [MYS^{+}67]. Nephridial [OBJ^{+}68, OBJ69]. Nephron [BN69, GBW69]. Nephrotoma [For65]. Nerve [Asb67, Ber67a, BLS67, BD67, BS69, DH67, EP68, Ham66, HB69a, HD66, HN65, LS65, MR68, Met69, MI69, NQVF66, SDV68, SFM69, SR65, TR69, Tha66b, Tha66a, YB67]. Nerve-Glandular [YB67]. Nerve-Muscle [SFM69]. Nerve-Myoepithelium [YB67]. Nerves [ESL68, PV67]. Nervous [HZL66, HD67, Mer68, Met67, SOK65, ST65, VCM65, dWA65]. Network [Met69]. Networks [Ish68]. Neural [LHS^{+}65]. Neuroblast [Ste65c]. Neuroblasts [Lea68]. Neurofibrillar [Met67]. Neurofilaments [WST68]. Neuromuscular [DK67, Len69, RB67]. Neuron [PPKA68, RCV69]. Neurons [EH68, HNV67, HP69, KT68b, Lan68, Las67b, LWS^{+}68, Pic65, WB69]. Neurosecretory [CKKW66]. Neurosecretosomes [BLS67]. Neurospora [DGT65, DDB69, GWT65, HW67, Luc65a, Luc65b, MT67, Sti68]. Neurotubules [WST68]. Neutral [Mor68]. Neutron [VHR65]. Neutron-Activated [VHR65]. Nevadensis [SS65a]. Newborn [DBN68, GB69a, KC69, SM65c, WSB68]. Newly [HP66b, SP66]. Newt [Bar67, Kel66, SF69a]. Nexuses [CB69b]. Nicotiana [Mou65]. Nippostrongylus [Jam66]. Nitrogen [TCL65]. Nodal [BD67]. Nodosum [HNV67]. Non [CHB65, Tak65]. Non-Collagen [CHB65]. Non-Random [Tak65]. Nonaqueous [SSWG68]. Nondroplet [SKWH68]. Noninfected [JLL68]. Nonmitotic [KWM69]. Nonprotein [HP66a]. Nonspecific [DK67]. Norepinephrine [BS69, DS68]. Normal [BS66b, GK69, HNV67, KEW67b, Lou68, ML66a, PRGC69, See66a, See66b, See66c, See66d, SB65d, Wil65b]. Normoblasts [SD69b]. Notes [Baj67]. Nuchal [CSJ67, WCJS67]. Nuclear [AB68, Eve68a, Fel65, Fra66, FS69, Gal67, GW69, GCG67, HA68, HBS65, Kar65a, KB68, KL65, KRC69, Kue69, Laf65, MS69b, Ric68, SH68a, See66a, See66b, See66c, See66d, She67, SCP67, SD67, SVWD68, SHG68, Tak65, TMMN68a, TVM69, VR67, WSL65]. Nuclear-Cytoplasmic [KB68]. Nuclei [AMH69, BB69a, BKS69, BBB68, CvD67, EEKS65, Fli65, Gil65, GCG67, GG68, HTBM65, HC68, KK67, KMK66, KB69, KT68a, LRM66, MP66, Phi65, Rod69, SS68a, SP67a, SD69b, TCV69, TMZ68, VM65, VM66a, VM66b, WHT67, vBPB66]. Nucleic [DSDD65, FW65, Hof65b, KSB65, LCW65, MK65b]. Nucleocytoplasmic [Fel66, Fel68, Fel69a, GP67, GP68, PG68]. Nucleoids [TMF66]. Nucleolar [BP66, CG65, GGK66, Joh69, SCP67, SWP69, TS69, Ter65]. Nucleoli [BOR69, HN66, HAKB65, Lan67b, LP69, ML66a, NP65, SS68a, SGF68, SF69b, Sus65, Sus67, Szo65, dMN67]. Nucleolus [Bri65b, GG69, LL69b, PP69, SB67, Ste65c, WPP69]. Nucleolus-Associated [GG69]. Nucleoside [DP67, ENQ65, NQVF66, Too65]. Nucleotidase [TSACBB69]. Nucleotide [Lan66]. Nucleotides [vBPB66]. Nucleus [BTVM69, Blo67, EH68, Fli69, HR67, LT69, RM66, Rog68, SS65b, SP68, SS66b, Moo66]. Numbers [ML66b]. Nutrition [SLAR69]. Nystromi [Ham65].

Obligately [MB67]. Obliquely [Ros65a, Ros67b, Ros68]. Observations [Aro67, BE65, BRH65, Blu65, Bro69, DH66, FB66, GKS^{+}67, Ham65, Har65, HF65a, HF65b, HC69b, Koe67, KH69, MPD65, Phi66b, RL67, TR69, Ted65]. Observed [Beh67, HZL66, Nan67]. Obtained [ADC69]. Occurrence [Afz65, AF68, Faw68, MRH^{+}67]. Occurring [Die67]. Ochromonas [SLAR69]. Octagonal [Gal67]. Octopus [YTST65]. Odonata [Smi66a]. Old [MD69]. Olfactory [ABK66, Ree65b]. Oligodendroglial [Hir68]. Ommatidia [Las67a]. Oncopeltus [WH66]. One [Hol67]. Onion [BMBW67]. Onset [MK68, PHP67]. Ontogenesis [AB65b]. Oocyst [TSW67]. Oocyte [And67a, And68, DD65, HA67, Rog68]. Oocytes [Fra65, KI67, Kes65, KB68, KB69, Lan67b, Lan69a, MS69a, SF69a, Sta65]. Oogenesis [DR66, ZG68, Die66]. Ooplasm [MM66b]. Opaque [Fel65, Tro65, Tro67]. Operation [Tro65]. Opossum [Mun65, SR67a]. Optic [PV67]. Optical [Bar69, Ber69, SB69b, WF69]. Oral [Fra67, JT67, KCR67a, KCR68, SBK67]. Orange [CB69a, MU67]. Order [Kle69]. Ordered [Phi65]. Organ [Ham65, She65]. Organelle [Løv69, SS65a]. Organelles [DFW66, TGB69, Zal69]. Organism [MMB65]. Organization [BH69b, DB66, EC65, HD66, KP68, LL69b, MU68, Nac68, OEC65, PH67, PR69, Pep66, Ros65a, SOK66, Smi65, Smi66a]. Organized [BBPM67]. Organotypic [MBB67a, MBB67b]. Organotypiqe [TVP67]. Organs [FD65, SS69a]. Orientation [NB65, Tak65]. Orientations [MW66]. Origin [And68, BF66, CG69, Hir68, LF69]. Orotic [RAH69]. Orthophosphate [LT69, TS69]. Orthoptera [MU68]. Osmiophilia [HH66]. Osmiophilic [ALEW66, LL66, SWH66, SPW^{+}67, SKWH68]. Osmium [AS66, CCIC69, Gun65, HF68, Kor67, NKKW69, SWH66, SHG68]. Osmium-Tetroxide [Gun65]. Osmotic [CG68, EE69]. Osmotically [CG66, SS69c, VG68]. Osmotically-Induced [SS69c]. Osteogenic [Sco67]. Other [Die66, GS67a, Ros66, SM65a, Sla65]. Oto [SWH66]. Outer [Coh68, McC65, RRG68, SG68, ST68, SKEB67, You67, YB69]. Outer-Doublet [ST68]. Outgrowth [UAH68]. Ovalbumin [KGO69]. Ovalbumin-Secreting [KGO69]. Ovarian [Bla66a, Bla66b]. Ovary [FB66, Kar65b, KB65a]. Oviduct [KGO69, OS69a, OS69b]. Ovulation [Jag69]. Ovum [ZMBB66]. Oxalacetic [LT68a, LT68b]. Oxidase [SEG67, SPW^{+}67, SKWH68, LH67]. Oxidation [MC69]. Oxidative [CG68, ZOE67]. Oxygen [KCW67, RM69, SF65, dWA65]. Oxygen-Poisoned [KCW67]. Oxysporum [Ais69]. Oxytalan [CF66a].

P [CE68a, CE68b]. Pacinian [NOP69]. Packing [BLRG67]. Pair [ST68]. Pairing [Moe69a]. Palaemonetes [HD66]. Paludosa [HS65a]. Pancreas [GAPRC68, HH67, Ich65, MC65, Mor68, OJP^{+}68, SP66, KY66]. Pancreatic [JP67a, JP67b, JP68a, JP68b, Boq69]. Pancreozymin [HH67]. Paneth [TT69]. Papillae [GB66b]. Paracristalline [Bas66]. Paracrystalline [Bas66]. Paragonia [TWE68]. Paramecium [Huf69, OBJ^{+}68, OBJ69, PVT68, RP67]. Parameters [Oka67]. Parasite [AB68]. Parasites [AHHS66]. Parasitized [SKN65]. Parasitizing [LK67b]. Parathyroid [BN65b, Bor68, CPR67]. Parenchymal [BF69, HCGP67, Lou68, Rey65, SR69, WLKS68, Wil65b]. Pargyline [MP68]. Parietal [Die67, Rub69a]. Parotid [AOS69]. Partial [Tro65]. Partially [JLL68]. Participation [And68, DBdD67, YB69]. Particle [BW67]. Particles [Ash65, IP65, LP69, MK65a, MBM67, Rog68, TSACBB69, WRJ67]. Particular [And68]. Particulate [BBdD65, SM65a, SK68c]. Parvum [DS66]. Passage [MCP69, SM65c]. Paths [GGBO69]. Pathway [Las67b]. Pathways [Gre65]. Pattern [NA65, Phi69]. Patterns [Kla68, MI69, MO69, RR69, SR67a]. Pealii [Met69, MI69, Arn69]. Peculiar [Lin67]. Pederine [BFDP68]. Pedunculatum [LR66]. Pellicle [DS69b, Som65, TS65b]. Pellicular [Aik67]. Penaeus [Ham66].

Penetration [SD68a]. Penicillin [FJH65]. Perfusate [LSL68]. Perfused [GGM^{+}68, Ich65]. Perfusion [KWM69]. Period [KL65]. Periodic [Faw68, KW67a]. Peripheral [HB69a, NQVF66]. Periplaneta [ST65, Smi66b]. Peritoneal [Cha65]. Permanent [MO69]. Permeability [CT67, CK68b, DCL69, Hic66b, Kar67, KT68a, SKK68, Ted65, WSL65]. Peroxidase [CK68b, Kar67, LG65, MHA65, RK67a, SKK68, Sed69, SJ68b]. Peroxidatic [Fah69]. Peroxisomes [BN69, Fah69, FN69, LPB^{+}68, LPLD69, PLD69]. Persistence [HN66, HRS69]. Petri [EF69]. pH [HA68]. Phagocytes [Kar65b]. Phagocytic [CP67b, Mis66]. Phagocytosis [CM65, KW67b, MPNK69, MWGM65, WK69]. Phalacrocleptes [LK67b]. Phase [GG69, Oka67, RS67b, SP67a, WO67b, WF68]. Phenobarbital [DBN68, OEE65, OE66a, OE66b, ZOE67]. Phenobarbital-Induced [OEE65]. Phenobarbital-Treated [DBN68]. Phenomenes [PVT68]. Phenotypic [AH66a]. Phenylhydrazine [SK68b]. Phenylhydrazine-Induced [SK68b]. Phloem [CE68a, CE68b, ECH67, GPB68, WN65a]. Phlorizin [Sti67]. Phlorizin- [Sti67]. Phosphatase [BH67, DP67, Dem68, ENQ65, FC68, GC66c, GCG67, GB69b, LS68a, MG67, NQVF66, OE66b, SP67b, Too65, WSW65]. Phosphatases [Blu65, ET66, Lan68, SB65c, TE66]. Phosphate [TG68]. Phosphate-Loaded [TG68]. Phosphatidylcholines [EV65]. Phosphatidylinositol [HH67]. Phosphodiesterase [LS68a, TSACBB69]. Phospholipase [Les67]. Phospholipid [MK65b]. Phospholipids [CF66b, EVD65]. Phosphorus [VHR65]. Phosphorylation [CG68]. Phosphotungstic [JEZ69, SG69b]. Photocytes [Bas66]. Photoreceptor [Wol69, You67]. Photosynthetic [GSW65]. Phycomyces [Wol69, Zal69]. Physarum [CBR65, CR66, GSS69, GG68, GG69, HG69, NA65, RSBG66]. Physical [CF66b, DdD67]. Physicochemical [WKG66]. Physiological [BRGG65, BRG68, JB68]. Phytoflagellate [DS66]. Phytohemagglutinin [PAM67, Pog69]. Piesma [AS67]. Pig [FH65, FC68, HH67, Mer68, Non68, Rob69b, SCP67, SP66]. Pigment [GS67b, GB69a, KB65a, Mat65, SOK65, Sho66, Sil69, YB69]. Pigment-Producing [KB65a]. Pigmentation [Ken66]. Pigs [Chr65, PCS69]. Pinacyanol [Tan69]. Pincytosed [EC68]. Pinealocyte [Lin65, Lin67]. Pinocytic [HFC68]. Pinocytosis [ZR65]. Pinocytotic [BMH68]. Pipiens [KI67, War68]. Pisum [Hof65a]. Pit [Ram69]. Pituitary [BLT69, BLS67, CM69, GH65, LS65, MSH69, SF66, TKH69]. Placenta [Oll68]. Plant [Bar69, CMM66, Die66, Fal69a, Fis68, LL69b, MM66a, Ray67, Tur68, dZG69]. Plants [GS67a, HBS65]. Plasma [BE65, BE68, BFB^{+}69, GB66b, KHSR66, SRRK69, SWS68, Sug66, WK66, WKG66]. Plasma-Cell [KHSR66]. Plasmalemmal [PB68]. Plasmodia [MR65, NA65, RSBG66]. Plasmodium [AHS67, Aik67, HHS66, TSW67, VR67]. Plastids [Die66, New67a]. Plate [HJ68, SSW67]. Platelet [MWGM65]. Platelets [Beh67, HT65].

Pleurochrysis [Bro69]. Pleuropneumonia [MMB65]. Pleuropneumonia-Like [MMB65]. Plexus [DB66]. Plugs [RA65]. Point [Aro66, Wol68]. Poisoned [KCW67]. Pokeweed [BHHW69]. Polarized [Aro67, MU68]. Pollen [Ech65, Tak65, Tak68]. Polycephalum [CBR65, CR66, GSS69, GG68, GG69, HG69, NA65, RSBG66]. Polychaete [LK67b]. Polymerase [Fis68, vBPB66]. Polymerization [Beh67]. Polymerizing [SKWH68]. Polymers [MT67]. Polymorphonuclear [BF66, BF68a, BF68b]. Polynoinæ [Bas66, Bas66]. Polyribosomes [BN65a, EBH67, KH69, MRH^{+}67, SKHR68, vdD69]. Polysaccharide [MPO67, MK68]. Polysomes [Fal69b, PS66, SH68a]. Polytene [Amm69, BB69a, MU67, Rod69]. Pool [Lea68, Luc65b]. Pools [SMP65]. Population [Hof65a, HS65a, Sto69]. Populations [BHdD69, KM65, RBLH68, WD68b]. Pore [DS69b, Fel65]. Pores [Gal67]. Porphyridium [GC65a, GC66a]. Position [Nic65]. Possibilities [dZG69]. Possible [FC68, GPB68, Mil69]. Posterior [BLS67, LS65, MMNT68a, MMNT68b, TMMN68a, TMMN68b]. Postfixation [HF68]. Postnatal [Fuj67, PV67, SM69, TMK68]. Postovulatory [KP68]. Postrigor [SGR67]. Postsynaptic [MP66]. Potassium [Bul69]. Potential [ROJ65, SMW67]. Practical [WKS66]. Prawn [DH67, HD66]. Pre [KP68, SGR67, TMK68]. Pre- [KP68, SGR67, TMK68]. Preceding [KL65]. Precipitates [SHG68]. Precursor [LP69, Luc65b]. Precursors [CY69]. Prefixation [Wol68]. Pregamiques [PVT68]. Premelanosomes [Par67]. Preneoplastic [Kar69]. Preparation [BF69, BBB68, Hol67, HCGP67]. Preparations [DB69, HN66, JLL68]. Preparative [FM65]. Prepared [SD69a]. Presence [Hol68, KSR67, NP65, Pod68, SMW67]. Present [RL67]. Preservation [Mit69, Nan69, NLS67]. Pressure [Lan65a, Lan66, ZR65]. Pretreatment [HBT69]. Prevailing [Phi69]. previous [Mil69]. Pricked [KI67]. Primary [And67a, GTP69, HHM69, TG69]. Primate [SD68b]. Principle [DS66]. Pringsheim [Bro69]. Prior [Lan66]. Prismatic [BRH65]. Prismatic-Type [BRH65]. Probe [RD69]. Procedure [BMR67, Jac68]. Process [Act66, GMB65, SF66, YB69]. Processes [BH69a, Fle65, Tha66a]. Processing [BT68, WPP69]. Produced [CT67, FHC68a, FHC68b, LSL68, MK65a, SF65]. Produces [MS69a]. Producing [KB65a, New67b, SMW67]. Production [BLT69, BP69a, GS67b, HH66, KM65, PP69, Sil69]. Products [Amm69, Kor67, Lan67a]. Profiles [MW66]. Progesterone [OS69a, OS69b]. Progression [Hof68]. Projections [Kel69]. Proliferating [Kar69]. Proliferation [Asb67, CLB65, Fab68, Fuj67, OEE65, OE66a, OE66b]. Proline [ISS68]. Prolixus [TP68a, TP68b, War68]. Pronuclear [LA68, Urb69, ZMBB66]. Pronuclei [Szo65]. Proof [Gir65]. Properties [Baj67, BB69a, CG66, CM69, DMEA67, DdD67, GHL69, HFL69, HYL69, Mat65, MBM67, RB66, RB69a, RRG68, SG68, ST68]. Prophase [RI67]. Proplastids [BW67, Jac68]. Protein [AMH69, AS66, AP69, BP69a, BB69b, CHB65, CE68a, CE68b, CR66, FF67, Gol68a, Gol68b, Gou68a, HP66a, HBS69, Him67b, Hol67, HP69, JH65, JLL68, KS69, Koe68, KA65, Kue69, KHSR66, LC65, LC68, MPO67, MK68, MS69c, MG68, New67a, Pre66, RRG68, RS69, Rys67, SR67a, See66a, See66b, See66c, See66d, ST67, ST68, SF69b, SG69a, Sta65, SHMR69, VF69, WPP69, YD68, vdD67]. Protein-Polysaccharide [MPO67]. Protein-Storing [New67a]. Proteine [Che67]. Proteins [Bry66, DMEA67, DD65, FE69, GP67, GP68, JP67a, JP67b, JP68a, JP68b, LS68b, PG68, SG69b, SW67]. Proteolytic [IS69, MCB68]. Proteus [CG69, Fli68, HvIvZ66, Lan65a, LvI65b, ZR65, vIHvZ66]. Prothoracique [Bea68]. Protocollagen [CP68]. protofibril [Mil69]. Protoplasts [GC65b]. Protozoa [Amm69, Jen67]. Protozoan [All69, RC67b]. Protrusions [TM67]. Provision [Hof68]. Prymnesium [DS66]. Pseudogloiophloea [Ram69]. Pseudoplatanus [WN65a]. Pterinosomes [Mat65, MO68]. Puffer [LWS^{+}68]. Pulchella [KB69]. Pulverization [KS67, MO69]. Pulverizations [NAL^{+}67]. Pump [SH67a]. Punctulata [And68, GTP69, Kar65b, LA68, TM69, TG69]. Pure [SF65]. Purified [JLL68]. Purkinje [SJ68b]. Puromycin [EB66, GGF68, Kah68, SVD68]. Puromycin-Induced [EB66, GGF68]. Purpurea [Ech65]. Pwm [BHHW69]. Pyramidal [PPKA68]. Pyrenoids [Hol68]. Pyriformis [All69, Arg65, BHP65, Cam65, CG65, CA68, EZ68, Fli65, Fra67, Gib65a, GW69, HEL67, Hof65b, LS68a, MSP66, Par65b, PR68, SMP65, Sto68, TS65b, Win67, WF68]. Pyrimethamine [AB68]. Pyrimidines [Cam65]. Pyroantimonate [Bul69, SHG68]. Pyroantimonate-Osmium [SHG68]. Pyrophosphatase [SB69a].

Quantitative [BTVM69, Fuj67, GB66a, LWS^{+}68, Lou68, RDS^{+}69, RB69b, Sal67, Sal69, SD69a, SFG67, SG69a, SSG69, Sim67, Stu65, TVP67, VHR65, WLKS68].

Rabbit [ABK66, BHdD69, CMDLB68, GGM^{+}68, GAPRC68, HM67, KWWL66, Koe67, SP67b, SM65a, WD68a, ZG68]. Rabbits [PCS69]. Radiation [HS65a, HB69b, LN65a, RP66]. Radiation-Induced [HB69b]. Radiation-Stimulated [RP66]. Radioactive [ISS68]. Radioactivity [Koe68]. Radioautographic [AP69, CP67a, CY69, HB69a, HH67, Jer66, NL66a, Ray67, Sus67, Tak68, VM66b, WHL69]. Radioautographie [TVP67]. Radioautography [BS67, BSH68, BS69, Car69b, CL66, Com66b, DS68, Joh69, NL66b, NAL^{+}67, NdM66, Orl68, PA67, Sal67, Sal69, SBS69, SdZ67, Sco67, Sim67, SK67b, Sti67, dMN67]. Radish [BN65a, NB65]. Ramie [Hey66]. Rana [KI67, War68]. Random [Tak65]. Raphanus [Laf65]. Rapid [AS66, BKS69, SE67]. Rat [AOS69, BF65, BLT69, BBdD65, BN69, BE65, BE68, BF69, Bul69, BBP65, BBPM67, BT68, CB69a, CG66, CG68, CBP67, Cha65, CL66, CLB65, Com66a, CM69, DDGL66, DB65, DdD67, DBdD67, DS68, ENQ65, Fah69, FB66, Fla68a, Fla68b, FOP^{+}69, FF67, Gil65, GL68, GB66a, GC66b, GHL69, GTLE65, HT66, HK65, Hic65, Hic66b, HZL66, Hir68, HN65, HNV67, HCGP67, HFL69, HYL69, IP65, JB68, Kah68, Kar69, KZ69b, KCW67, Kue69, LT68a, LT68b, LPLD69, Lin65, Lin67, Lou68, LRM66, MSH69, MBB67a, MBB67b, MS69c, MC65, MHA65, MMP65, Nag65, NDRFK66, NdM66, OVF69, Oll68, Orr65b, PV67, PPKA68, Pla69, PLD69, RL67, RHL69, RD69, SS68a, SH68a, SP67a, SM69, SB69a, SG68, Sco67, SRRK69, SP68, SWS68, SS69b, SSGF69, SB65d, SS69c]. Rat [SM67, Szo65, TMF66, VG68, Vau66, VF67, VM66a, WS65, WSW65, WHL69, WHT67, WLKS68, WN65b, ZOE67, vdD69, HSC68]. Rat-Liver [CG66, GC66b]. Rate [HS65a, SP67a, WO67a]. Rates [MPNK69]. Ratio [Sto69]. Ratios [Gol68b]. Rats [Ber68, CvD67, DBN68, GK69, LPB^{+}68, NL66b, OYHM65, PCS69, SF65, SSW67, SM65c, THNB69, TVM69, TMK68, VM65, VM66b, WSB68, WS68]. Ray [EV65]. Rays [Hes65c]. Reabsorption [Maa67]. Reaction [And68, BBFA67, Kor67, ND65a, ND65b, SPW^{+}67, SFG67, SJ68b]. Reactivation [Gib65a, Win67]. Reactive [NKKW69]. Reactivity [SG69b]. Reagent [SKWH68]. Reagents [SPW^{+}67, TB65]. Reaggregating [Ove69]. Receptor [IP65]. Recognition [SD69b]. Reconstitution [SC67b, SHR67, SHMR69]. Reconstruction [SR68]. Recovery [KS66]. Rectal [GB66b]. Red [LS68b, MD69, Mor68, Orl68, Ram69, MP67a]. Reduction [For65, LvI65a, LvI65b, SD69b]. Redundant [Ros66]. Reference [And68, DFW66, Ish68, Mat65]. Reformation [SF69b]. Regenerating [Fab68, Rez69]. Regeneration [Ove65, RC67b, RC69]. Regilla [Bak65]. Region [Lan68, NL66a]. Regions [BD67]. Regulation [GB69b, GG68, SF66, WPP69, vdD67]. Reinhardi [JP68c]. Reinvestigation [All68]. Related [GWS65, RB66, SW67]. Relation [Din69, EKSG69, ECH67, EC67, PWFP68, Rob66, SMP65, TL67, WHL69]. Relationship [BLRG67, Bro69, GW66, GTLE65, OEE65, OE66a, OE66b, SB68a, SE66, Vau66]. Relationships [SP67a]. Release [Amm69, AS66, SPT65]. Remark [BB67b]. Removal [Cam65, Les67, SHR67]. Renal [GGM^{+}68, GGBO69, Maa67]. Renewal [BS65, TT69, You67, YD68, YB69]. Reovirus [SD68a]. Repair [OR68, RO68]. Repeat [KW67a]. Repeating [GB66b]. Replacement [Cam65]. Replication [BA66, BD68, Car69a, GG68, GG69, Lym67, MG66, PHP67, Rod69, SR67a, SH67b, YF69]. Reproduction [KZvW65]. Required [Cam65]. Requirement [HHM69]. Research [ROJ65]. Resemble [AF68]. Resembling [Buc66, NSHW68]. Reserpine [CYD69]. Resistance [She68]. Resolution [BHdD69, BHB68, Car69b, Orl68, SBS69, SdZ67, SC67a, Sim67, SK67b, Sti67]. Resorption [Par69a, SSW67, Vae68]. Respiration [EH68, HBS69, RE69]. Respiratory [HT66]. Response [BP65, BMH68, EE69, HH67, Ich65, MK65a, MMA65, RK67c, SB66, SA66, vdD67]. Responses [LF69]. Retention [AS66, SD69a]. Reticulatus [DR66]. Reticulocytes [SK68b]. Reticulum [BF65, Bas66, DSP66a, DSP66b, DB69, ET66, EI67, GP65, Gir65, HS67a, Hes65b, KHH68, Lin67, MS69b, NC69, OE66b, PR69, Pea65, Ros69a, SM69, SH67a, TE66, WSB68, WK66, WKG66]. Retikulums [Gir65]. Retina [HM67, PM65, SOK66, YTST65]. Retinal [IP65, Koe67, Las67b, Mat67, McC65, YD68, YB69]. Revealed [CMM66, Hey66, Nag65, Pep66, TW65]. Reversal [BF65, SGF68]. Reversible [EBH67, JEE69, WB69, dWA65]. Rhabdome [PB69]. Rhodnius [TP68a, TP68b, War68]. Rhodospirillum [GSW65, HF65a, HF65b]. Riboflavin [TESW69]. Ribonuclease [LS68a]. Ribonucleases [UR66a, UR66b]. Ribonucleic [JH65, PAM67, WARS65]. Ribonucleoprotein [KHSR66, Rog68, Sus65, Sus67, vI66]. Ribosomal [MMB65, Nan67]. Ribosome [Bye66, Flo69, LP69, SVD68]. Ribosomes [BP68a, CG69, DB65, Ech65, LL65, MSW68, PM69]. Rich [RL67]. Rigor [SGR67]. Rigor-Shortened [SGR67]. Ring [Lan67b]. Rings [SK67b, Sti67]. Rise [ZOE67]. RNA [ETS68, ET69, Fis68, GW67a, GW66, HRS69, Kar65a, KB67, Koe67, NdM66, RP67, See66a, See66b, See66c, See66d, SL66, SCP67, SWP69, SD68a, SVWD68, SS66b, SE66, WPP69]. Rod [Koe67, YB69]. Rodents [MdHP66]. Rods [Coh68, Col65, Eak65, Mat67, YD68, Boq69]. Role [HC68, KS69]. Room [TCL65]. Root [BMBW67, BN65a, Bon68, Bon69, BTD66, BBPM67, Him67a, KT68b, MBB67a, MBB67b, Mou65, NB65, New67a, New67b, NSHW68, NQVF66, PS65b]. Rootlet [Mat67]. Roots [Fus68, Hof65a, HS65a]. Rostratus [OK67]. Roth [Ech65]. Rotifer [MD66a, MD66b]. Rough [DBN68, Fal69b, NC69]. Rough-Surfaced [DBN68]. Rous [SD68b, SB68b]. Routine [FP65b]. Rubrum [HF65b]. Ruby [SAW^{+}66, Tan69]. Rumen [Lav69]. Runt [WA65].

S [GB69b, MSP66]. Saccharomyces [CWL67, Wil65a]. Saccobolus [Car67]. Saccules [Coh68]. Sacs [KW65]. Salivary [BB69a, BBB68, DL69, LN65b, NP65, PS65a, Phi65, Riz67, SS66b, TR69, TMK67]. Salt [EE69, SR67b]. Salts [FP65b]. Sand [Wen66]. Saponin [See67]. Sarcolemma [WS65]. Sarcoma [SD68b, SB68b]. Sarcomere [BLRG67, Ste65a]. Sarcoplasm [MMA65]. Sarcoplasmic [DB69, ET66, EI67, GP65, HS67a, Hes65b, PR69, Pea65, Ros69a, SM69, SH67a, TE66, WSB68]. Sarcotubular [EE68, SM65a]. Sartorius [EE68, Pea65]. Satellite [Rez69]. Sativum [Hof65a]. Sativus [Laf65]. Saturated [EV65]. Say [AS67]. Scale [GC65b, LPB^{+}68, Ove66, Rob66]. Scattered [Gib65b]. Scherffelii [Bro69]. Schiller [KR69b]. Schizobranchia [LK67b]. Schwann [Asb67, Ham66]. Sciara [GGK66, PS65a, Phi65, Phi66b, Phi67, Phi66a]. Sciatic [Asb67, HN65]. Sea [And68, Fra65, Ham65, Har65, Hir65, KS69, Lan66, LA68, MCB68, ST68, TCV69, ZZ67]. Secondary [MO69]. Secrete [RTL69]. Secreting [KGO69, NL66a, Sed69]. Secretion [BW66, DL69, KY66, SS67b, SS67a, SS69b]. Secretory [AOS69, CM69, EE69, GHL69, HFL69, HYL69, JP67a, JP67b, JP68a, JP68b, MM66a, Riz67, SF66]. SÉCrÉTrices [Bea68]. Section [Ree65a, SR65]. Sections [All68, Ash65, BL67, Fis68, FP65b, Hey66, Kol65, LBHT67, SSG69, VHR65]. Sedimentation [BHdD69, PNGM66]. Sedimentation-Electrophoresis [PNGM66]. Seed [GS67a]. Seedlings [BW67, BP68a]. Seeds [Yat65]. Seen [Ash65]. Segment [Coh68, McC65, PSPO68, YB69]. Segments [DE69, SH68b, You67]. Segregation [MB65]. Selected [BMR67, Koe67]. Selective [EE68, JEZ69, SWP69]. Sendai [NAL^{+}67]. Sensitive [SB67]. Sensitivity [BS67, DFW66, PK69]. Sensory [DFW66, FD65]. Separated [BE68]. Separation [LPB^{+}68, SM65a, SS69a, Zal69]. Septal [RA65]. Septate [Loc65]. Septations [RA65]. Serial [SR68]. Sericatus [BP69a]. Series [EV65]. Serum [FCLT69, Hol67, ML66a]. Several [KS69, MU68]. Sex [Com66b]. Sham [Tro65]. Shape [FS69]. Sheath [Ham66, HD67]. Sheaths [DH67, HD66, Ros66]. Sheep [YB67]. Shell [HS66, RLV65]. Shocks [RE69]. Shoot [GS67a]. Short [HN66]. Shortened [SGR67]. Shortening [RMR69]. Showing [MO69]. Shown [CL66, NL66b]. Shrew [SOK66]. Shrimp [Ham66]. Sialic [Kra67, WK66]. Sieve [BC65, CE67, ECH67]. Significance [GPB68]. Significant [SF69b]. Silica [RLV65]. Silk [BP69a, DFM67, MMNT68a, MMNT68b, TMMN68a, TMMN68b]. Silkworm [MMNT68a, MMNT68b, TMMN68a, TMMN68b]. Simian [LF65]. Simple [BKS69, BA66, Oka67, Rub69b]. Simplified [VC65]. Simultaneous [HEL67, HF68, HSC68]. Single [BRGG65, BRG68, EH68, Gir65, Løv69, SAW^{+}66, TCL65]. Sinusoid [KO65]. Sirmate [BP68a]. Site [DCL69, LF69]. Sites [AP69, CL66, HD68, Kar65a, NKKW69, PP69]. Situ [BMR67, GB66b, LT68a, PM65]. Size [Luc65b, Nic65, WD68c]. Sized [IBH68]. Skeletal [BD69, CB69a, EEKS65, EI67, FK66, Fis67, GP66, Gol68a, GTLE65, Ish68, IBH68, KC68, Rez69, SM69, SFM67, SFM69, SH68b, WS65, WSB68, WS68, WD68a, Zad66]. Skin [DP67, FP65a, KEW67a, Mun65]. Slices [HH67, MHA65]. Slime [CBR65, GG68, NA65, New67b]. Slime-Producing [New67b]. Slow [Fah67, Hes65b, MG68, Pag65]. Small [Ebs67, EB66, Laf65, MW67, PPKA68, RBLH68, SS65d, VF67]. Smallpox [AB65b]. Smooth [Ber68, CB69b, IS69, KR68, KR69a, Lan65b, Lan67a, Mer68, Non68, PH67, Rob69b, RK67c, Sed69, Tha66a]. Smooth-Surfaced [IS69, Sed69]. Snake [Hes65b, HMW66]. Snout [Mun65]. Sodium [Bul69, SK68a, VU68, Zad66]. Soie [Bea68]. Soluble [LT69]. Solution [LS68b, SD69a]. Solutions [CCIC69]. Soma [LR66]. Soma-Somatic [LR66]. Somatic [LR66, Ros65b]. Some [Baj67, CB66, DH66, DdD67, FB66, HTBM65, Ham65, Har65, HL66, MBM67, Pep66, RB66, RB69a, RRG68, SK68a, Sla65, Ted65, VU68, WRJ67, dZG69]. Son [VP69]. Source [Car69b, Mil69]. Sources [DL69]. Space [Coh68, PB69, VCM65]. Spaces [LL66]. Spatial [MI69, MG67]. Special [Ish68, Mat65, PF65, She66]. Specialization [EE69]. Specializations [TL67]. Specialized [GPB68]. Species [KS69, KA65, Phi69, SvA67]. Specific [BF68a, BF68b, BMH68, KC69, Lym67]. Spectrophotometric [Kor67]. Sperm [BF68c, Fra65, MU68, Nag65, Phi66a, Phi69, Rob66, RR69]. Spermatids [MP67b, Moe69b]. Spermatocytes [For65, Hes65c, Moe69b]. Spermatogenesis [CK67, CK68a, DSA65, Tur68]. Spermatozoides [Che67]. Spermiogenesis [AWE67, Jam66, KMK66, Lan69b, Phi66b, Sho67, Vau66]. Sphaerocarpos [Kel69]. Sphere [Vau66]. Spherical [Laf65]. Spheroidal [Lan67b]. Spider [BP69a, HC69b]. Spinach [ADC69, LP66, NM67, PaSN65, PNGM66]. Spinal [BBP65, MS69b, SFM69, Wei68]. Spindle [Baj67, For65, KB65b, WST68]. Spiny [New67b]. Spleen [SH65]. Spontaneous [OJP^{+}68]. Spoon [HR67]. Sporadic [Faw68]. Sporogonic [VR67]. Sporophyte [Kel69]. Squid [MR68, Met69, MI69, SDV68]. Sr [GC66b]. Stability [CCIC69, Kir69]. Stable [HH66, See67]. Stage [HS65b, ZMBB66]. Stages [AHS67, HHS66, HJ68, KC69, Lan65b, LT67, WHL69]. Stain [VC65]. Stained [Bar69, BE65, MU67, MD69, SAW^{+}66, Tan69]. Staining [ADC69, Bar67, CMM66, FP65b, Fri69, Hey66, JEZ69, Les67, Ove68, Pep66, Ree65a, SF69a, SWH66, SG69b]. Stains [WWS66]. Stalked [PO65]. Stalks [SS66a]. Staphylococcus [Sug66]. Starfish [MS69a]. State [Rub69b]. Stationary [WO67b, WF68]. Statistical [MW66]. Steady [Rub69b]. Stein [Ken66]. Stenotele [MPD65]. Stephensi [Sch69]. Steps [SVD68]. Stereological [Lou68, WKS66]. Sterility [Sho67]. Sternal [SS65a]. Steroid [Bla66a, Bla66b]. Stimulated [Pog69, RP66, TMZ68]. Stimulation [FCLT69, GWS65, Ich65, VM66a]. Stomach [MD66b]. Storage [LC68, SS67b, SS68b, TKH69]. Storing [CYD69, New67a]. Strain [BLT69, BK69, KB65b, RTL69, RS67b, See66a, See66b, See66c, See66d, SD68a]. Strains [TVM69]. Strand [JS67]. Strandedness [TW65, WH66]. Streaming [Nac69]. Strength [IKNG68]. Stress [EE69]. Striated [Aro67, BLRG67, Gar68, Gol67, HMS65, RDS^{+}69, RB69b, Ros65a, Ros67b, Ros68, Ste65b, SHR67]. Stroma [Gun65, Mit69]. Structural [AR67, BB69a, BB66, Cog66, DB65, FJH65, HD66, HD67, HB67, Ich65, JB68, KCW67, LT67, MT67, MR68, MB67, Nan68, PB68, Pep66, RK67a, RK67c, SOK66, SK68a, TB65, TM69, TMZ68]. Structure [Afz65, AHS67, Bak65, BP69a, Bon68, BC65, BK68, BK69, BP69b, Bri65b, BL68, BD68, CPR67, CF66a, CMDLB68, Chr65, CB65, CMM66, DDGL66, DFM67, Ell65, Fah67, Faw68, FH65, FB66, FFS67, Fli65, FBS68, Fri65, Fri66, Gal66, GGF68, GAPRC68, GW69, Gou68b, GRH66, GC65c, Gun65, GTLE65, Ham65, Ham66, HSM65, HHS66, Hes65a, Hey66, HF65a, HF65b, Hic65, Jen67, JP68c, Kar69, KMK66, Kel66, Kel67, KZ69a, Kes68, KLW68, KZvW65, KB65b, Lav69, LC65, Loc65, LA68, Mat65, MPD65, MK65a, MdHP66, Moe69a, Mol65, MEK68, MMP65, Nan67, NLS67, New67a, NOP69, Ove65, PO65, PO67a, PJO67, PO67b, PS65a, Phi66a, Pic65, Ram65, RC67a, RLV65, Rin67, Rub69a, Rud65, SOK65, SVR67, SF65, She65, SFM67, Smi66b, SP68, Sri66]. Structure [Ste65a, Ste65c, SE69, SBK67, TESW69, TP68a, TP68b, Urb69, VR67, WSW65, Wol67, WF69, WN65a, WF68, YTST65, ZMBB66, dWA65, vI66].

Structures [All69, Ber69, Buc66, FW67, Hes65c, Lan68, Pag65, RAH69, RDS^{+}69, RB69b, SB69b, TWE68, Wil65b]. Strychnine [Ken66]. Studied [BS69, GB66b, Kar67, NdM66, Sim67]. Studies [AH69a, AH69b, AB68, And67b, ABK66, BHP65, BLD^{+}66, Ber69, BB69b, BRGG65, BRG68, BFDP68, BHDW68, BP68b, BP68c, Car69a, CK67, CK68a, CF66b, CY69, CK68b, CG69, EC65, EV65, FW65, FC67, Fra67, GGM^{+}68, GWT65, GP65, GP66, GMB65, GW69, GC66b, GC66c, GBW69, HB69a, HK65, HA67, Hir65, HBHW68, HB69b, Jam66, JB68, KR69a, KT68a, LCW65, LR66, LN65b, LB68, LT68b, LDRA67, LF65, MMB65, MBB67a, MBB67b, Mat65, MMNT68a, MPO67, MBMB68, MHA65, MYS^{+}67, MMNT68b, Nac68, Nac69, Orr65a, OEC65, RLV65, Rei66, RDS^{+}69, RB69b, Rys67, Sat65, Sat68, SPT65, SHW69, SB69b, Sug66, TMMN68b, THM66, TP67, TB69, TW65, UR66a, UR66b, WSGH69, Wei65, Wei66, Wei67b, Wei67a]. Study [Bar67, BBdD65, BR67, BA69, BBP65, BBPM67, CP67b, CB66, CM65, CB69b, CKKW66, DB65, DD65, DFM65, DS69a, DR66, Fel69a, Fis67, GGK66, GB66a, Jer66, Joh69, KS69, KW65, Kor67, Laf65, LT68a, LT67, MBM67, MW67, Ray67, RC67a, RCE69, She66, SD67, SKHR68, SR67b, SAW^{+}66, SB68b, Tur68, VCM65, VM66b, WSB68, WA65]. Styela [KB65a]. Stylonychia [Amm69]. Subcelluar [HD68]. Subcellular [BEB67, BB67a, Kar65c, LL69a]. Subchromatid [BH69b]. Subfractionation [DN66]. Subfractions [SP66]. Subhuman [SD68b]. Submandibular [KCR67b]. Submicrosomal [DDBE69]. Submitochondrial [SEG67]. Substance [SE67]. Substitution [VCM65]. Substratum [FLV68]. Substructure [Flo69, Par65a, Phi66c, Ter67]. Subsurface [Wei68]. Subsynovial [LC66]. Subtilis [LvI65a, Nan67, Nan68, Nan69, SB65a]. Subunit [ST67, ST68]. Subunits [BE68, GB66b, Met67, Nan67, RK67b]. Succinic [SB65a, SB65b]. Sucrose [KT68a, Ted65]. Sugar [KW65]. Suggestive [Faw68]. Sulfate [KHH68]. Sulfur [FE69]. Sulfur-Containing [FE69]. Supercontraction [HMS65]. Superficiel [VP69]. Superior [NDRFK66]. Supplements [HP66a]. Support [DB68b]. Supporting [Coh68]. Supramedullary [LWS^{+}68]. Supraoptic [EH68]. Supravitally [Tan69]. Surface [BDRN65, EE69, GKS^{+}67, GBW69, GB66b, HFC68, Ito65, Kra67, MD69, RL67, Ram65, SD69b, TL67, WK66]. Surface-Active [GKS^{+}67]. Surfaced [DBN68, IS69, Sed69]. Surfaces [ADC69]. Suspension [Pla69]. Suspensions [Gib65b, HT66]. Sustains [YF69]. Suturalis [For65]. Sweat [Ell65]. Swine [FCLT69]. Swiss [KSB65]. Swordtail [Mat65]. Symmetrical [Bru69]. Sympathetic [BTD66, BS69, DS68, Pic65]. Synapse [Hes65a]. Synaptic [HR67, JEZ69, Las69]. Synaptinemal [RI67]. Synchronization [RSBG66, ZG68]. Synchronized [BHP65, Coo66, RC66, RJM68, Sto68]. Synchronizing [RE69]. Synchronous [HW67]. Synchronously [HEL67]. Synchrony [PS65b]. Syncytia [HC68]. Syncytium [Rob66]. Synovial [BSH68]. SynthÈSe [CA68]. Syntheses [BMBW67, Pre66]. Synthesis [AP69, BHO65, CL66, Coo66, CR66, ETS68, ET69, FCLT69, Gol68a, GW67a, Gou68a, GW66, GW67b, HEL67, HBS69, HH67, HD68, HP68, Jer66, JH65, JLL68, Kar65a, KB67, KL65, Kue69, LPLD69, MSH69, MS69c, MM66b, NL66b, NdM66, OEE65, OE66a, PAM67, PLD69, RP67, RP66, RC66, RM69, RS69, SR67a, SSS68, See66a, See66b, See66c, See66d, SL66, SF69b, SPW66, SCP67, SWP69, SW67, SS67b, SS67a, SS68b, SS69b, SMP65, SE66, Sus65, Tak65, Tak68, VM66a, WHL69, WPP69, Wil65a, vdD67]. Synthesized [SP66]. Synthetase [PM65]. Synthetic [DSA65, KHSR66]. Syrian [BRH65]. System [EE68, EI67, HS67a, Hes65b, HZL66, HD67, Ish68, LDRA67, Met67, Orr65a, OEE65, OE66a, PR69, Ros67a, RKC68, Rub69b, SOK65, SM69, SB65a, SB65b, Sed69, ST65, SKEB67, SLAR69, Ter65, WS65, WF69]. Systole [DS69b].

T [Ish68, EI67, HS67a, Hes65b, PR69, SM69, WS65]. T-System [Ish68]. Taenia [BR67, Non68]. Tail [Nag65]. Tangentialis [HR67]. Tapes [Cob69]. Tarentolae [Sim68]. Taste [BS65]. Tch [SWH66]. Technique [FP65b, GL65, HSC68]. Techniques [ADC69, SB65a, SB65b]. Teleosts [And67a]. Telolecithal [Arn69]. Telophase [MB65]. Temperature [Sus67, WO67a]. Temperatures [MSP66, ML66b, TCL65]. Temporal [Kle69, MG67]. Tentacle [Rud65]. Tentans [DE69, DL69]. Term [BBPM67, HN66, SOZC68]. Terminals [BS69, Hes65b, TR69]. Terrestris [AWE67, Cog66]. Test [KB65a]. Testicular [Chr65]. Testis [DS69a, FC68]. Tetrahymena [All69, Arg65, BHP65, Cam65, CG65, CA68, DS69b, EZ68, Fli65, Fra67, Gib65a, GW69, HEL67, Hof65b, LS68a, LH67, MSP66, Par65b, PR68, RE69, RC69, SPW66, SMP65, Sto68, Sto69, TS65b, Win67, WF68]. Tetramitus [OK67]. Tetranitro [LvI65a, LvI65b, SB65a]. Tetranitro-Blue [LvI65a, LvI65b, SB65a]. Tetraphenylboron [HL66]. Tetrazolium [LvI65a, LvI65b, SB65a]. Tetroxide [AS66, CCIC69, Gun65, HF68, Kor67, SWH66, SHG68]. Tetroxide-Fixed [SWH66, SHG68]. Their [And68, Fli69, GPB68, KK67, Mat65, MO68, MK65a, TL67]. Thermal [MU67]. These [MBM67]. Thick [SM67]. Thickness [SSG69]. Thigh [AC67]. Thiocarbohydrazide [SWH66]. Thiocholine [DK67]. Thiohydantoin [HK65]. Thiohydantoin-Induced [HK65]. Thiolacetic [DK67]. Thoracic [HS68, Lan68]. Thorium [Pad69]. Threadlike [MI69]. Three [DS69a, RR69, Sch69, Tha66a]. Three-Dimensional [RR69]. Throughout [Fli65]. Thylakoids [Fal69b]. Thymic [KSR67, MMP65]. Thymidine [Bry66, Cav66, Fus68, HS65b, Lea68, NAL^{+}67, Par65b, Rez69, SP67a, Sco67, SPW66, SS65c, SMP65, TAP67, VM66b, WD68b]. Thymidine- [WD68b, Rez69, SP67a, Sco67]. Thymus [KT68a]. Thyroglobulin [WHL69]. Thyroid [BW66, GTLE65, PW67, RD69, WSW65]. Thyroidectomy [GTLE65]. Thyroids [WHL69]. Thyrotropic [GH65]. Thyrotropin [WSW65]. Tibial [RC67a, SSW67]. Tight [BE68, NSHW68, TRH66]. Tigrinum [Moe69a]. Time [CP67b, SW67]. Time-Lapse [CP67b]. Timicronucleatum [OBJ^{+}68]. Timing [Wil65a]. Tips [Him67a, New67a]. Tipula [dFM66]. Tissue [BFdD67, BdD67a, BdD67b, Bul69, DFM65, Gar69, GMB65, JL68, JT67, Kah68, Kar65c, LHS^{+}65, LDRA67, LG65, LK67a, LP67, NP67, OVF69, SWH66, SG69b, SW66, UAH68, VCM65, WA65, Yat65, dWA65]. Tissues [Bye66, Fis68, HBS65, MYS^{+}67, PA67, SHG68, TGB69]. Toad [Cho65, DCL69, EP68, HSM65, HB66, Las67b, Ros66]. Tobacco [EC67, Joh69, KZvW65, WRJ67]. Tokophrya [Rud65]. Tomato [GC65b]. Tonofilaments [KW67a]. Tool [ROJ65, Sal67, Sal69]. Torpedo [She65]. Total [AMH69, Eak65, Koe68, Pre66, SE67]. Toxoplasma [VP69]. Tracer [Cho65, Kar67, SKK68, VF69]. Tradescantia [HS65a, Tak65, Tak68]. Traits [AH66a]. Transaminase [LT68a, LT68b]. Transepithelial [DCL69]. Transfer [GPB68, LDRA67, Par67]. Transformation [GL68, GK69, Kah68, OK67, PCS69, SW66, TSW67]. Transformations [HC69a]. Transient [See67]. Transition [RJ69, Vau66]. Transitional [Hic65, Hic66a, Hic66b, KKHE65]. Translation [vdD67]. Translocation [GPB68, dZG69]. Transport [AH66b, Baj67, BD69, JP67a, JP67b, JP68a, JP68b, KWWL66, MG68, PWFP68, SKEB67, SS67b, SS67a, SS68b, SS69b, SS66b, WB69]. Transport* [VU68]. Transporting [DB68a]. Transtubular [GGBO69]. Transverse [Pea65, Zad66]. Treated [ADC69, Ber68, CvD67, DBN68, Jag69, Lan69a, MBM67, MCB68, SvA67, WB69, WARS65, dMN67]. Treatment [AB68, Fus68, GTLE65, KS67, MP68, Pad66, RSL69, SB66]. Tree [SOK66]. Treefrog [Bak65]. Triads [WS68]. Trichium [PVT68]. Trichomonas [AG68]. Tricine [Gar69]. Tricomplex [EVD65]. Triglyceride [CBP67]. Triphosphatase [FP66, GP65, Nag65, SDV68, SF69a, MP67a]. Triphosphate [Gib65b, RB66, SPW66, Tak68]. Triploid [Moe69a]. Tritiated [CP67a, DS68, Ebs67, Fus68, NAL^{+}67, Par65b]. Triticum [BW67, BP68a]. Tritium [Bry66, Cav66, Koe68, Stu65, Tay65b]. Tritium-Labeled [Cav66, Tay65b]. Triton [DC67, HBT69, LPB^{+}68]. Tritrichomonas [Nic68]. Tropomyosin [Faw68]. Trough [Wol68]. Trough-Critical [Wol68]. Trypanosoma [BD68]. Trypanosomatidae [SvA67]. Trypsin [EKG66, SGR67]. Tube [BC65, ECH67]. Tubular [CE67, MM66a, Sed69]. Tubules [GGM^{+}68, GGBO69, HMW66, Ish68, Pea65, Phi66c, Phi69, Ric68, Zad66]. Tumor [BS66b, HBT69, KHSR66, Rys67, See66a, See66b, See66c, See66d]. Tumors [SD68b]. Tunicate [Kes65, KB65a]. Tupaia [SOK66]. Turnover [LPLD69, PLD69, SVWD68]. Turtle [Las69]. Twitch [Ber67a, Hes65b, Pag65]. Two [FK66, PO65, RBLH68, SH68a, SvA67, WD68b]. Type [BRH65, HR67, HS67b, JEZ69, MSL69, Sho66, Sho67]. Types [GP66, IBH69]. Tyrosine [HM67].

Ultrastructural [ADC69, CG66, CK68b, CHE66, DB68a, DS69a, DB66, Fla68b, GGM^{+}68, GK69, GSS69, Hac66, Hac68, HC69a, HC69b, HB69b, Kah68, Kar67, KR69a, LB68, LSL68, LHS^{+}65, MPO67, MBMB68, Par69b, RCE69, RJ69, Ros65a, Ros66, RCV69, SKK68, SPW^{+}67, SKWH68, SRZB66, Tha66a, THNB69, Tur68, WWS66]. Ultrastructurales [Bea68]. Ultrastructure [Aik67, AOS69, Baj67, BD69, BS66a, CFK65a, CFK65b, Car67, DB69, DH67, DFW66, FM69, FLV68, FD65, GC65a, GS67a, HL66, HR67, HF68, Huf69, JEE69, KC68, KO65, dFM66, LK67b, Lou68, MMB65, MF69, MCB68, MW67, Nan69, NDRFK66, PH67, RSL69, RS67a, Riz67, Ros65b, Ros69b, Som65, Sre65, SB68a, SSWG68, SS69c, Tha66b, TS65b, WSL65, WLKS68, Yat65]. Ultrastructures [Moo66]. Ultrastrukturen [Moo66]. Ultrathin [BL67, FP65b, Kol65, LBHT67]. Ultraviolet [For65, LN65a, MCK65, Ste65a]. Unadapted [Hol67]. Unbalanced [CG65]. Unciliated [LK67b]. Uncoupled [BL68]. Undulans [Ken66]. Uniaxial [MS69a]. Unique [SD68b]. Unit [BTH68, BDRN65]. Units [KW67a]. Unsaturated [Kor67]. Unusual [Act66, Fla68b, KB65a, KCR67b, LL65, PO65, RR69, SS65d, TWE68]. Unusually [OVF69, Ros69a]. Upon [DS69a, Fli69, RE69, VM66a, ZOE67]. Uptake [HP69, LC68, MHA65, NL66a, Pad69, Rez69, RM69, Rys67, SP67a, Sed69, Sta65]. Uranyl [FP65b, HF68]. Urchin [And68, Fra65, Har65, Hir65, KS69, Lan66, LA68, MCB68, ST68, TCV69, ZZ67]. Urechis [DMEA67, Gou68a]. Ureter [Hic65]. Uridine [Com66b]. Uridine- [Com66b]. Urinary [Cho65]. Urodele [WD68c]. Use [Aro65, Bul69, CB69a, Gar69, MBMB68, MBM67, Orl68, VC65, VF69, HSC68]. Used [LS68b, SKK68]. Using [FP65b, Joh69, SB65a, SB65b]. Uterine [Ber68, BA69, RK67c]. Uterus [WHT67]. Utilisation [TVP67].

Vacuole [DS69b, SK68c]. Vacuoles [FHC68a, FHC68b, FC68, HFC68, Joh69, MBM67, Mis66]. Variants [BTG69]. Variation [LSL68]. Variations [AR67]. Variety [IBH69]. Various [KA65, Kue69, Lan65b, NL66a, SHMR69]. Vas [FF67, Mer68, Rob69b]. Vascular [DS68, SM65c]. Vasopressin [GGM^{+}68]. Veins [GPB68]. Velocity [Nic65]. Venom [MR68]. Ventral [TP68a, TP68b]. Ventricular [SB68a]. Verified [Nan69]. Verruciformis [LK67b]. Vers [Bea68]. Versus [KHSR66]. Vertebrate [BR69, PH67]. Very [Men69]. Vesicle [HFC68, JEZ69, KK69, New67b]. Vesicles [And69, Fal69a, FF67, Fri69, IS69, MM66a, PB68, SM65a]. Vesiculation [BBFA67]. Vessels [SS65d]. Vestibular [SP68]. Vestibularis [HR67]. Vestigial [Fri68]. Viable [BHB68]. Vicia [Laf65, PS65b, TW65, WD68b]. Viciae [Smi65]. Villi [MG67]. Vinblastine [KHH68, MS69a, Sto68]. Vincristine [KH69]. Vincristine-Induced [KH69]. Virus [AB65b, CHE66, DFM65, ECH67, EC67, HC68, KZvW65, LF65, NAL^{+}67, SFG67, SD68b, SB68b, WRJ67, WR68]. Virus-Antibody [SFG67]. Virus-Induced [HC68, NAL^{+}67, SD68b, WR68]. Viruses [MBMB68, dZG69]. Visceral [CB69b]. Visual [Las69, LF69]. Visualization [WHL69]. Visualized [ADC69, TRH66]. Vital [Fle65]. Vitally [SAW^{+}66]. Vitamin [BP65, DDGL66]. Vitellogenesis [HC69b]. Vitro [BH68, BH69a, Bri65b, BMR67, EC68, EI67, FHC68a, FHC68b, FLV68, FCLT69, GL67, GAPRC68, Hol67, HAKB65, KCR67a, KCR68, LT68a, LF65, MBB67a, MBB67b, NDRFK66, PW67, PCS69, SR67a, SFM67, SCP67, Sil69, SB66, TM67, TMZ68, Tor65, TMK67, WHL69, vdD69, KY66]. Vivo [GAPRC68, Jer66, KHSR66, War68]. Volume [AMH69, PaSN65, PWFP68, SS69c]. Vulgare [BW67, BP68a]. Vulgaris [HD66, LvI65b].

Wall [Bro69, Kel69, MT67, NB65, Ray67]. Walls [Hey66]. Walrus [Gol67]. Water [BB67b, GGBO69, LH65, LT69]. Water-Miscible [LH65]. Water-Soluble [LT69]. Watlingi [Cob69]. Which [AF68, Hol67, RTL69, YF69]. White [FK66]. Whorls [NC69]. Wild [Sho66, Sho67]. Wing [Fri68, RC67a]. Withdrawal [HSM65]. Within [Baj67, BS65, BRH65, Bri65a, BS69, Gar68, Kar65b, LT69, Lin65, SSG69]. Without [Hir65]. Wood [LR66]. Work [Gol68a]. Work-Induced [Gol68a]. Woronin [RA65]. Wound [OR68, RB65, RO68]. Wr [HBT69, LPB^{+}68]. Wr-1339 [HBT69, LPB^{+}68].

X [DC67, FB66, MBB67a, MBB67b, Pic65, EV65, Hes65c, LN65a, PHP67]. X-100 [DC67]. X-Irradiated [FB66, Pic65]. X-Irradiation [MBB67a, MBB67b]. X-Radiation [LN65a]. X-Ray [EV65]. X-Rays [Hes65c]. Xenopus [Ove65]. Xiphophorus [Mat65]. Xylem [CB65]. Xylinum [BHO65].

Yeast [AF68, DNC69, MRH^{+}67, RM66, Moo66]. Yeasts [MB67]. Yellow [WR68]. Yellows [CHE66, ECH67]. Yield [BF69]. Yolk [DR66, KB69]. Young [KB68, MD69].

Z [Kel67, SHR67, Faw68, SHMR69, WSB68]. Z- [SHMR69]. Z-Band [Kel67]. Z-Line [SHR67]. Zellkern [Moo66]. Zett. [LR66]. Zeylanica [She67]. Zinc [NKKW69]. Zona [GWS65]. Zonal [BHdD69, BHB68]. Zootermopsis [SS65a]. Zygote [Foo68]. Zymogen [AS66].

References

Atlas:1965:CGC

[AB65a] M. Atlas and V. P. Bond. The cell generation cycle of the elevenday mouse embryo. Journal of Cell Biology, 26(1):19–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

            (electronic). URL http://jcb.rupress.org/content/26/1/19.

Avakyan:1965:OHS

[AB65b] A. A. Avakyan and A. F. Byckovsky. Ontogenesis of human smallpox virus. Journal of Cell Biology, 24(3):337–??, March 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/24/3/337.

Aikawa:1968:SND

[AB68] Masamichi Aikawa and Richard L. Beaudoin. Studies on nuclear division of a Malarial parasite under pyrimethamine treatment. Journal of Cell Biology, 39(3):749–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/39/3/749.

Ash:1966:SDR

[ABK66] K. Owen Ash, J. E. Bransford, and R. B. Koch. Studies on dispersion of rabbit olfactory cells. Journal of Cell Biology, 29(3):554– ??, June 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/29/
3/554.

Anderson:1967:BFI

[AC67] H. Clarke Anderson and P. R. Coulter. Bone formation induced in mouse thigh by cultured human cells. Journal of Cell Biology, 33(1):165–??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/33/1/165.

Acton:1966:UCP

[Act66] A. B. Acton. An unusual ciliumlike process. Journal of Cell Biology, 29(2):366–??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/29/2/366.

Arntzen:1969:CCM

[ADC69] C. J. Arntzen, R. A. Dilley, and F. L. Crane. A comparison of chloroplast membrane surfaces visualized by freeze-etch and negative staining techniques; and ultrastructural characterization of membrane fractions obtained from digitonin-treated spinach chloroplasts. Journal of Cell Biology, 43(1):16–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/43/1/16.

Avers:1968:OYC

[AF68] Charlotte J. Avers and Micheline Federman. The occurrence in yeast of cytoplasmic granules which resemble microbodies. Journal of Cell Biology, 37(2):555–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/37/2/555.

Afzelius:1965:OSM

[Afz65] Björn A. Afzelius. The occurrence and structure of microbodies. Journal of Cell Biology, 26(3):835–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/26/3/835.

Amos:1968:IMG

[AG68] W. B. Amos and A. V. Grimstone. Intercisternal material in the Golgi body of trichomonas. Journal of Cell Biology, 38(2):466–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
2/466.

Abbott:1966:LPT

[AH66a] Joan Abbott and Howard Holtzer. The loss of phenotypic traits by differentiated cells. Journal of Cell Biology, 28(3):473–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/28/
3/473.

Anderson:1966:ATC

[AH66b] Everett Anderson and William R. Harvey. Active transport by the Cecropia midgut. Journal of Cell Biology, 31(1):107–??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/31/
1/107.

Adams:1969:SHCa

[AH69a] Eleanor C. Adams and Arthur T. Hertig. Studies on the human corpus luteum. Journal of Cell Biology, 41(3):696–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/41/3/696.

Adams:1969:SHCb

[AH69b] Eleanor C. Adams and Arthur T. Hertig. Studies on the human corpus luteum. Journal of Cell Biology, 41(3):716–??, June 1969.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/41/3/716.

Aikawa:1966:FMA

[AHHS66] Masamichi Aikawa, Peter K. Hepler, Clay G. Huff, and Helmuth Sprinz. The feeding mechanism of avian Malarial parasites. Journal of Cell Biology, 28(2):355–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/28/2/355.

Aikawa:1967:FSA

[AHS67] Masamichi Aikawa, Clay G. Huff, and Helmuth Sprinz. Fine structure of the asexual stages of plasmodium elongatum. Journal of Cell Biology, 34(1):229–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/1/229.

Aikawa:1967:UPC

[Aik67] Masamichi Aikawa. Ultrastructure of the pellicular complex of plasmodium fallax. Journal of Cell Biology, 35(1):103–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/35/
1/103.

Aist:1969:MAF

[Ais69] James R. Aist. The mitotic apparatus in fungi, ceratocystis fagacearum and fusarium oxysporum. Journal of Cell Biology, 40 (1):120–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/1/120.

Ashworth:1966:HIO

[ALEW66] C. T. Ashworth, J. S. Leonard, E. H. Eigenbrodt, and F. J. Wrightsman. Hepatic intracellular osmiophilic droplets. Journal of Cell Biology, 31(2):301–??, November 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/31/2/301.

Allen:1968:RCS

[All68] Richard D. Allen. A reinvestigation of cross-sections of cilia. Journal of Cell Biology, 37(3):825–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/37/3/825.

Allen:1969:MBB

[All69] Richard D. Allen. The morphogenesis of basal bodies and accessory structures of the cortex of the ciliated protozoan Tetrahymena Pyriformis. Journal of Cell Biology, 40(3):716–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/40/3/
716.

Abuelo:1969:HC

[AM69] J. G. Abuelo and Dorothy E. Moore. The human chromosome. Journal of Cell Biology, 41(1):73–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/41/1/73.

Amaral:1969:ECV

[AMH69] Leonard Amaral, Louis G. Moriber, and Marion Himes. The effect of cortisone on the volume and total protein content of mouse liver nuclei. Journal of Cell Biology, 42(3):835–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
3/835.

Ammermann:1969:RDB

[Amm69] Dieter Ammermann. Release of DNA breakdown products into the culture medium of stylonychia mytilus exconjugants (protozoa, ciliata) during the destruction of the polytene chromosomes. Journal of Cell Biology, 40(2):576–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/40/2/576.

Anderson:1967:FPE

[And67a] Everett Anderson. The formation of the primary envelope during oocyte differentiation in teleosts. Journal of Cell Biology, 35 (1):193–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/1/193.

Anderson:1967:EMS

[And67b] H. Clarke Anderson. Electron microscopic studies of induced cartilage development and calcification. Journal of Cell Biology, 35(1):81–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/1/81.

Anderson:1968:ODS

[And68] Everett Anderson. Oocyte differentiation in the sea urchin, Arbacia Punctulata, with particular reference to the origin of cortical granules and their participation in the cortical reaction. Journal of Cell Biology, 37(2):514–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/37/2/514.

Anderson:1969:VAC

[And69] H. Clarke Anderson. Vesicles associated with calcification in the matrix of epiphyseal cartilage. Journal of Cell Biology, 41(1):59– ??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/41/
1/59.

Amsterdam:1969:DCU

[AOS69] Avraham Amsterdam, Itzhak Ohad, and Michael Schramm. Dynamic changes in the ultrastructure of the acinar cell of the rat parotid gland during the secretory cycle. Journal of Cell Biology, 41(3):753–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/3/753.

Ashley:1969:EMR

[AP69] Charles A. Ashley and Theodore Peters. Electron microscopic radioautographic detection of sites of protein synthesis and migration in liver. Journal of Cell Biology, 43(2):237–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/43/
2/237.

Allenspach:1967:SVD

[AR67] Allan L. Allenspach and L. E. Roth. Structural variations during mitosis in the chick embryo. Journal of Cell Biology, 33(1):179– ??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
1/179.

Argetsinger:1965:ICB

[Arg65] Joan Argetsinger. The isolation of ciliary basal bodies (kinetosomes) from Tetrahymena Pyriformis. Journal of Cell Biology, 24(1):154–??, January 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/1/154.

Arnold:1969:CFF

[Arn69] John M. Arnold. Cleavage furrow formation in a telolecithal egg (Loligo Pealii). Journal of Cell Biology, 41(3):894–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/41/3/894.

Aronson:1965:UFL

[Aro65] John F. Aronson. The use of fluorescein-labeled heavy meromyosin for the cytological demonstration of actin. Journal of Cell Biology, 26(1):293–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/26/1/293.

Aronson:1966:MPB

[Aro66] John F. Aronson. A melting point for the birefringent component of muscle. Journal of Cell Biology, 30(3):453–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/30/
3/453.

Aronson:1967:PLO

[Aro67] John F. Aronson. Polarized light observations on striated muscle contraction in a mite. Journal of Cell Biology, 32(1):169–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/32/
1/169.

Amsterdam:1966:RRZ

[AS66] Abraham Amsterdam and Michael Schramm. Rapid release of the zymogen granule protein by osmium tetroxide and its retention during fixation by glutaraldehyde. Journal of Cell Biology, 29(2):199–??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/29/2/199.

Arnott:1967:IIG

[AS67] Howard J. Arnott and Kenneth M. Smith. Intracellular inclusions in the gut epithelial cells of piesma cinereum say. Journal of Cell Biology, 34(2):639–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/2/639.

Asbury:1967:SCP

[Asb67] A. K. Asbury. Schwann cell proliferation in developing mouse sciatic nerve. Journal of Cell Biology, 34(3):735–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
3/735.

Ashhurst:1965:MPS

[Ash65] Doreen E. Ashhurst. Mitochondrial particles seen in sections. Journal of Cell Biology, 24(3):497–??, March 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/24/3/497.

Anderson:1967:CDS

[AWE67] W. A. Anderson, A. Weissman, and R. A. Ellis. Cytodifferentiation during spermiogenesis in lumbricus terrestris. Journal of Cell Biology, 32(1):11–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/32/1/11.

Black:1965:AICa

[BA65a] Maurice M. Black and Hudson R. Ansley. Antigen-induced changes in lymphoid cell histones. Journal of Cell Biology, 26(1):201–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/1/201.

Black:1965:AICb

[BA65b] Maurice M. Black and Hudson R. Ansley. Antigen-induced changes in lymphoid cell histones. Journal of Cell Biology, 26 (3):797–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/3/797.

Bullivant:1966:SFF

[BA66] Stanley Bullivant and Adelbert Ames. A simple freeze-fracture replication method for electron microscopy. Journal of Cell Biology, 29(3):435–??, June 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/29/3/435.

Black:1967:AICa

[BA67a] Maurice M. Black and Hudson R. Ansley. Antigen-induced changes in lymphoid cell histones. Journal of Cell Biology, 35 (3):619–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/3/619.

Black:1967:AICb

[BA67b] Maurice M. Black and Hudson R. Ansley. Antigen-induced changes in lymphoid cell histones. Journal of Cell Biology, 35 (3):627–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/3/627.

Brandes:1969:EMC

[BA69] David Brandes and Elsa Anton. An electron microscopic cytochemical study of macrophages during uterine involution. Journal of Cell Biology, 41(2):450–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/41/2/450.

Baccetti:1967:CE

[Bac67] Baccio Baccetti. Collagen of the earthworms. Journal of Cell Biology, 34(3):885–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/34/3/885.

Bajer:1967:NUS

[Baj67] Andrew Bajer. Notes on ultrastructure and some properties of transport within the living mitotic spindle. Journal of Cell Biology, 33(3):713–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/33/3/713.

Baker:1965:FSM

[Bak65] Patricia C. Baker. Fine structure and morphogenic movements in the gastrula of the treefrog, hyla regilla. Journal of Cell Biology, 24(1):95–??, January 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/1/95.

Barnicot:1967:SNM

[Bar67] N. A. Barnicot. A study of newt mitotic chromosomes by negative staining. Journal of Cell Biology, 32(3):585–??, March 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/32/3/585.

Barton:1969:INS

[Bar69] R. Barton. Investigation of negatively stained plant flagellar microtubules by optical diffraction. Journal of Cell Biology, 41(2):637–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/2/637.

Bassot:1966:FMP

[Bas66] J. M. Bassot. Une forme microtubulaire et paracristalline de reticulum endoplasmique dans les photocytes des annelides polynoinæ. (French) [A microtubular and paracrystalline form of endoplasmic reticulum in Annelides Polynoinæ photocytes]. Journal of Cell Biology, 31(1):135–??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/31/1/135.

Bloom:1966:FSL

[BB66] Floyd E. Bloom and Russell J. Barrnett. Fine structural localization of acetylcholinesterase in electroplaque of the electric eel. Journal of Cell Biology, 29(3):475–??, June 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/29/3/475.

Baudhuin:1967:EMEb

[BB67a] Pierre Baudhuin and Jacques Berthet. Electron microscopic examination of subcellular fractions. Journal of Cell Biology, 35 (3):631–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/3/631.

Berthet:1967:RAD

[BB67b] J. Berthet and P. Baudhuin. A remark about the determination of the water content of mitochondria. Journal of Cell Biology, 34(2):701–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/34/2/701.

Bisalputra:1967:CMD

[BB67c] T. Bisalputra and A. A. Bisalputra. Chloroplast and mitochondrial DNA in a Brown alga egregia menziesii. Journal of Cell Biology, 33(3):511–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/33/3/511.

Berendes:1969:SFP

[BB69a] Hans D. Berendes and James B. Boyd. Structural and functional properties of polytene nuclei isolated from salivary glands of Drosophila Hydei. Journal of Cell Biology, 41(2):591–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/41/
2/591.

Bibring:1969:ISP

[BB69b] Thomas Bibring and Jane Baxandall. Immunochemical studies of 22s protein from isolated mitotic apparatus. Journal of Cell Biology, 41(2):577–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/41/2/577.

Boyd:1968:MPN

[BBB68] James B. Boyd, Hans D. Berendes, and Hudson Boyd. Mass preparation of nuclei from the larval salivary glands of Drosophila Hydei. Journal of Cell Biology, 38(2):369–??, August 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/38/2/369.

Baudhuin:1965:CBM

[BBdD65] Pierre Baudhuin, Henri Beaufay, and Christian de Duve. Combined biochemical and morphological study of particulate fractions from rat liver. Journal of Cell Biology, 26(1):219–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/26/
1/219.

Barros:1967:MVF

[BBFA67] C. Barros, J. M. Bedford, L. E. Franklin, and C. R. Austin. Membrane vesiculation as a feature of the mammalian acrosome reaction. Journal of Cell Biology, 34(3):??, September 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/34/3/C1.

Bunge:1965:EMS

[BBP65] Richard P. Bunge, Mary Bartlett Bunge, and Edith R. Peterson. An electron microscope study of cultured rat spinal cord. Journal of Cell Biology, 24(2):163–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/24/2/163.

Bunge:1967:LEM

[BBPM67] Mary Bartlett Bunge, Richard P. Bunge, Edith R. Peterson, and Margaret R. Murray. A light and electron microscope study of long-term organized cultures of rat dorsal root ganglia. Journal of Cell Biology, 32(2):439–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/32/2/439.

Bouck:1965:FSD

[BC65] G. Benjamin Bouck and James Cronshaw. The fine structure of differentiating sieve tube elements. Journal of Cell Biology, 25(1): ??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
1/79.1.

Brzin:1967:CAN

[BD67] Miro Brzin and Wolf-D. Dettbarn. Cholinesterase activity of nodal and internodal regions of myelinated nerve fibers of frog. Journal of Cell Biology, 32(3):577–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/32/3/577.

Burton:1968:FSR

[BD68] Paul R. Burton and Donald G. Dusanic. Fine structure and replication of the kinetoplast of trypanosoma lewisi. Journal of Cell Biology, 39(2):318–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/2/318.

Baskin:1969:CUC

[BD69] R. J. Baskin and D. W. Deamer. Comparative ultrastructure and calcium transport in heart and skeletal muscle microsomes. Journal of Cell Biology, 43(3):610–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/43/3/610.

Bowers:1967:LLTb

[BdD67a] William E. Bowers and Christian de Duve. Lysosomes in lymphoid tissue. Journal of Cell Biology, 32(2):339–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/32/
2/339.

Bowers:1967:LLTc

[BdD67b] William E. Bowers and Christian de Duve. Lysosomes in lymphoid tissue. Journal of Cell Biology, 32(2):349–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/32/
2/349.

Brzin:1965:CAP

[BDRN65] Miro Brzin, Wolf-Dietrich Dettbarn, Philip Rosenberg, and David Nachmansohn. Cholinesterase activity per unit surface area of conducting membranes. Journal of Cell Biology, 26 (2):353–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/2/353.

Benedetti:1965:EMO

[BE65] E. L. Benedetti and P. Emmelot. Electron microscopic observations on negatively stained plasma membranes isolated from rat liver. Journal of Cell Biology, 26(1):299–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/26/1/299.

Benedetti:1968:HAS

[BE68] E. L. Benedetti and P. Emmelot. Hexagonal array of subunits in tight junctions separated from isolated rat liver plasma membranes. Journal of Cell Biology, 38(1):15–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/38/1/15.

Beaulaton:1968:MUC

[Bea68] J. A. Beaulaton. Modifications ultrastructurales des cellules SÉCrÉTrices De la glande prothoracique De vers À soie Au cours des deux derniers Âges larvaires. Journal of Cell Biology, 39 (3):501–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/39/3/501.

Baudhuin:1967:EMEa

[BEB67] Pierre Baudhuin, Philippe Evrard, and Jacques Berthet. Electron microscopic examination of subcellular fractions. Journal of Cell Biology, 32(1):181–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/32/1/181.

Behnke:1967:IMO

[Beh67] O. Behnke. Incomplete microtubules observed in mammalian blood platelets during microtubule polymerization. Journal of Cell Biology, 34(2):697–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/2/697.

Bell:1969:LGF

[Bel69] Mary Bell. Langerhans granules in fetal keratinocytes. Journal of Cell Biology, 41(3):914–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/41/3/914.

Bergman:1967:MNE

[Ber67a] Ronald A. Bergman. Motor nerve endings of twitch muscle fibers in hippocampus hudsonius. Journal of Cell Biology, 32(3):751–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/32/
3/751.

Berlin:1967:LAM

[Ber67b] J. D. Berlin. The localization of acid mucopolysaccharides in the Golgi complex of intestinal goblet cells. Journal of Cell Biology, 32(3):760–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/32/3/760.

Bergman:1968:USM

[Ber68] Ronald A. Bergman. Uterine smooth muscle fibers in castrate and estrogen-treated rats. Journal of Cell Biology, 36(3):639–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
3/639.

Berger:1969:ODS

[Ber69] Jacob E. Berger. Optical diffraction studies of crystalline structures in electron micrographs. Journal of Cell Biology, 43(3): 442–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/43/3/442.

Baglio:1965:RAE

[BF65] Corrado M. Baglio and Emmanuel Farber. Reversal by adenine of the ethionine-induced lipid accumulation in the endoplasmic reticulum of the rat liver. Journal of Cell Biology, 27 (3):591–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/3/591.

Bainton:1966:OGP

[BF66] Dorothy Ford Bainton and Marilyn G. Farquhar. Origin of granules in polymorphonuclear leukocytes. Journal of Cell Biology, 28(2):277–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/28/2/277.

Bainton:1968:DECa

[BF68a] Dorothy Ford Bainton and Marilyn G. Farquhar. Differences in enzyme content of azurophil and specific granules of polymorphonuclear leukocytes. Journal of Cell Biology, 39(2): 286–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/39/2/286.

Bainton:1968:DECb

[BF68b] Dorothy Ford Bainton and Marilyn G. Farquhar. Differences in enzyme content of azurophil and specific granules of polymorphonuclear leukocytes. Journal of Cell Biology, 39(2): 299–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/39/2/299.

Barros:1968:BGM

[BF68c] C. Barros and L. E. Franklin. Behavior of the gamete membranes during sperm entry into the mammalian egg. Journal of Cell Biology, 37(3):??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/37/3/C13.

Berry:1969:HYP

[BF69] M. N. Berry and D. S. Friend. High-yield preparation of isolated rat liver parenchymal cells. Journal of Cell Biology, 43 (3):506–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/43/3/506.

Boone:1969:IPM

[BFB^{+}69] Charles W. Boone, Lincoln E. Ford, Howard E. Bond, Donald C. Stuart, and Dianne Lorenz. Isolation of plasma membrane fragments from HeLa cells. Journal of Cell Biology, 41(2):378–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/41/
2/378.

Bowers:1967:LLTa

[BFdD67] William E. Bowers, John T. Finkenstaedt, and Christian de Duve. Lysosomes in lymphoid tissue. Journal of Cell Biology, 32(2):325–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/32/2/325.

Brega:1968:SMA

[BFDP68] Agnese Brega, Arturo Falaschi, Luigi De Carli, and Mario Pavan. Studies on the mechanism of action of pederine. Journal of Cell Biology, 36(3):485–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/36/3/485.

Bertolini:1967:APA

[BH67] B. Bertolini and G. Hassan. Acid phosphatase associated with the Golgi apparatus in human liver cells. Journal of Cell Biology, 32(1):216–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/32/1/216.

Bischoff:1968:EMI

[BH68] Richard Bischoff and Howard Holtzer. The effect of mitotic inhibitors on myogenesis in vitro. Journal of Cell Biology, 36 (1):111–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/36/1/111.

Bischoff:1969:MPD

[BH69a] Richard Bischoff and Howard Holtzer. Mitosis and the processes of differentiation of myogenic cells in vitro. Journal of Cell Biology, 41(1):188–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/1/188.

Brinkley:1969:ESO

[BH69b] B. R. Brinkley and Ronald M. Humphrey. Evidence for subchromatid organization in marsupial chromosomes. Journal of Cell Biology, 42(3):827–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/42/3/827.

Boone:1968:RMV

[BHB68] Charles W. Boone, George S. Harell, and Howard E. Bond. The resolution of mixtures of viable mammalian cells into homogeneous fractions by zonal centrifugation. Journal of Cell Biology, 36(2):369–??, February 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/36/2/369.

Baggiolini:1969:RGR

[BHdD69] Marco Baggiolini, James G. Hirsch, and Christian de Duve. Resolution of granules from rabbit heterophil leukocytes into distinct populations by zonal sedimentation. Journal of Cell Biology, 40 (2):529–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/2/529.

Brittinger:1968:SL

[BHDW68] Günter Brittinger, Rochelle Hirschhorn, Steven D. Douglas, and Gerald Weissmann. Studies on lysosomes. Journal of Cell Biology, 37(2):394–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/37/2/394.

Brittinger:1969:EPM

[BHHW69] Günter Brittinger, Rochelle Hirschhorn, Kurt Hirschhorn, and Gerald Weissmann. Effect of pokeweed mitogen (pwm) on lymphocyte lysosomes. Journal of Cell Biology, 40(3):843–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/40/3/
843.

Ben-Hayyim:1965:SCA

[BHO65] G. Ben-Hayyim and I. Ohad. Synthesis of cellulose by acetobacter xylinum. Journal of Cell Biology, 25(2):191–??, May 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/25/2/191.

Barber:1965:SNG

[BHP65] Albert A. Barber, Warren W. Harris, and George M. Padilla. Studies of native glycogen isolated from synchronized Tetrahymena Pyriformis (Hsm). Journal of Cell Biology, 27(2):281– ??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/2/281.

Biempica:1966:HHM

[Bie66] Luis Biempica. Human hepatic microbodies with crystalloid cores. Journal of Cell Biology, 29(2):383–??, May 1966. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/29/2/383.

Bowers:1968:FSA

[BK68] Blair Bowers and Edward D. Korn. The fine structure of Acanthamoeba Castellanii. Journal of Cell Biology, 39(1):95–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/39/1/95.

Bowers:1969:FSA

[BK69] Blair Bowers and Edward D. Korn. The fine structure of Acanthamoeba Castellanii (Neff strain). Journal of Cell Biology, 41 (3):786–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/3/786.

Berkowitz:1969:SRM

[BKS69] Daniel M. Berkowitz, Tsuyoshi Kakefuda, and Michael B. Sporn. A simple and rapid method for the isolation of enzymatically active HeLa cell nuclei. Journal of Cell Biology, 42(3): 851–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/3/851.

Bernhard:1967:UFS

[BL67] W. Bernhard and Elizabeth H. Leduc. Ultrathin frozen sections. Journal of Cell Biology, 34(3):757–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/34/3/757.

Bullivant:1968:SCU

[BL68] Stanley Bullivant and Werner R. Loewenstein. Structure of coupled and uncoupled cell junctions. Journal of Cell Biology, 37 (3):621–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/37/3/621.

Blanchette:1966:OSCa

[Bla66a] E. Joan Blanchette. Ovarian steroid cells. Journal of Cell Biology, 31(3):501–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/31/3/501.

Blanchette:1966:OSCb

[Bla66b] E. Joan Blanchette. Ovarian steroid cells. Journal of Cell Biology, 31(3):517–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/31/3/517.

Bauer:1966:SIB

[BLD^{+}66] G. Eric Bauer, Arnold W. Lindall, Padmaker K. Dixit, Gordon Lester, and Arnold Lazarow. Studies on insulin biosynthesis. Journal of Cell Biology, 28(3):413–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/28/3/413.

Bloom:1967:NCF

[Blo67] Gunnar D. Bloom. A nucleus with cytoplasmic features. Journal of Cell Biology, 35(1):266–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/35/1/266.

Brandt:1967:RBM

[BLRG67] Philip W. Brandt, Enrique Lopez, John P. Reuben, and Harry Grundfest. The relationship between myofilament packing density and sarcomere length in frog striated muscle. Journal of Cell Biology, 33(2):255–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/33/2/255.

Bindler:1967:INE

[BLS67] Elliot Bindler, Frank S. Labella, and Madhu Sanwal. Isolated nerve endings (neurosecretosomes) from the posterior pituitary. Journal of Cell Biology, 34(1):185–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/34/1/185.

Bancroft:1969:CGH

[BLT69] Frank C. Bancroft, Lawrence Levine, and Armen H. Tashjian. Control of growth hormone production by a clonal strain of rat pituitary cells. Journal of Cell Biology, 43(3):432–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/43/
3/432.

Blum:1965:OAP

[Blu65] Jacob J. Blum. Observations on the acid phosphatases of euglena gracilis. Journal of Cell Biology, 24(2):223–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/24/2/223.

Bensch:1969:MCM

[BM69] Klaus G. Bensch and Stephen E. Malawista. Microtubular crystals in mammalian cells. Journal of Cell Biology, 40(1):95–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
1/95.

Bloch:1967:SDA

[BMBW67] David P. Bloch, Robert A. MacQuigg, Sheilah D. Brack, and Jung-Rung Wu. The syntheses of deoxyribonucleic acid and histone in the onion root meristem. Journal of Cell Biology, 33(3):451–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/33/3/451.

Blanton:1968:PRC

[BMH68] Patricia L. Blanton, James Martin, and Sol Haberman. Pinocytotic response of circulating erythrocytes to specific blood grouping antibodies. Journal of Cell Biology, 37(3):716–??, June 1968.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/37/3/716.

Brinkley:1967:PES

[BMR67] B. R. Brinkley, Patricia Murphy, and L. Carole Richardson. Procedure for embedding in situ selected cells cultured in vitro. Journal of Cell Biology, 35(1):279–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/35/1/279.

Bonnett:1965:PCA

[BN65a] Howard T. Bonnett and Eldon H. Newcomb. Polyribosomes and cisternal accumulations in root cells of radish. Journal of Cell Biology, 27(2):423–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/27/2/423.

Borle:1965:EPH

[BN65b] Andre B. Borle and William F. Neuman. Effects of parathyroid hormone on HeLa cell cultures. Journal of Cell Biology, 24 (2):316–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/2/316.

Beard:1969:DPM

[BN69] Margaret E. Beard and Alex B. Novikoff. Distribution of peroxisomes (microbodies) in the nephron of the rat. Journal of Cell Biology, 42(2):501–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/42/2/501.

Bonnett:1968:REF

[Bon68] Howard T. Bonnett. The root endodermis: Fine structure and function. Journal of Cell Biology, 37(1):199–??, April 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/37/1/199.

Bonnett:1969:CCD

[Bon69] Howard T. Bonnett. Cortical cell death during lateral root formation. Journal of Cell Biology, 40(1):144–??, January 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/40/1/144.

Boquist:1969:IRP

[Boq69] Lennart Boquist. Intranuclear rods in pancreatic islet β-cells. Journal of Cell Biology, 43(2):377–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/43/2/377.

Borle:1968:CMH

[Bor68] André Bernard Borle. Calcium metabolism in HeLa cells and the effects of parathyroid hormone. Journal of Cell Biology, 36 (3):567–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/36/3/567.

Berns:1969:ALM

[BOR69] M. W. Berns, R. S. Olson, and D. E. Rounds. Argon laser microirradiation of nucleoli. Journal of Cell Biology, 43(3):621–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/43/
3/621.

Bouck:1969:EM

[Bou69] G. Benjamin Bouck. Extracellular microtubules. Journal of Cell Biology, 40(2):446–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/40/2/446.

Bassett:1965:RIL

[BP65] Bruce E. Bassett and Lester Packer. Response of isolated lysosomes to vitamin A. Journal of Cell Biology, 27(2):448– ??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/2/448.

Barr:1966:CMN

[BP66] H. J. Barr and W. Plaut. Comparative morphology of nucleolar DNA in Drosophila. Journal of Cell Biology, 31(3):??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/31/
3/C17.

Bartels:1968:ASD

[BP68a] P. G. Bartels and E. J. Pegelow. The action of sirmate (3,4dichlorobenzyl methylcarbamate) on chloroplast ribosomes of Triticum Vulgare L. seedlings. Journal of Cell Biology, 37(2):??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/37/2/C1.

Bruns:1968:SBCa

[BP68b] Romaine R. Bruns and George E. Palade. Studies on blood capillaries. Journal of Cell Biology, 37(2):244–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/37/2/244.

Bruns:1968:SBCb

[BP68c] Romaine R. Bruns and George E. Palade. Studies on blood capillaries. Journal of Cell Biology, 37(2):277–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/37/2/277.

Bell:1969:CFS

[BP69a] Allen L. Bell and David B. Peakall. Changes in fine structure during silk protein production in the ampullate gland of the spider Araneus Sericatus. Journal of Cell Biology, 42(1):284–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
1/284.

Boyd:1969:FSM

[BP69b] C. A. R. Boyd and D. S. Parsons. The fine structure of the microvilli of isolated brush borders of intestinal epithelial cells. Journal of Cell Biology, 41(2):646–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/41/2/646.

Bennett:1967:SIT

[BR67] M. R. Bennett and D. C. Rogers. A study of the innervation of the Taenia Coli. Journal of Cell Biology, 33(3):573–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/33/3/573.

Brightman:1969:JBI

[BR69] M. W. Brightman and T. S. Reese. Junctions between intimately apposed cell membranes in the vertebrate brain. Journal of Cell Biology, 40(3):648–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/40/3/648.

Brandt:1968:CMP

[BRG68] Philip W. Brandt, John P. Reuben, and Harry Grundfest. Correlated morphological and physiological studies on isolated single muscle fibers. Journal of Cell Biology, 38(1):115–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/38/1/115.

Brandt:1965:CMP

[BRGG65] Philip W. Brandt, John P. Reuben, Lucien Girardier, and Harry Grundfest. Correlated morphological and physiological studies on isolated single muscle fibers. Journal of Cell Biology, 25(3):233– ??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
3/233.

Blinzinger:1965:OPT

[BRH65] K. Blinzinger, N. B. Rewcastle, and H. Hager. Observations on prismatic-type mitochondria within astrocytes of the Syrian hamster brain. Journal of Cell Biology, 25(2):293–??, May 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/25/2/293.

Brightman:1965:DWB

[Bri65a] Milton W. Brightman. The distribution within the brain of ferritin injected into cerebrospinal fluid compartments. Journal of Cell Biology, 26(1):99–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/26/1/99.

Brinkley:1965:FSN

[Bri65b] B. R. Brinkley. The fine structure of the nucleolus in mitotic divisions of Chinese hamster cells in vitro. Journal of Cell Biology, 27 (2):411–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/2/411.

Brown:1969:ORG

[Bro69] R. M. Brown. Observations on the relationship of the Golgi apparatus to wall formation in the marine chrysophycean alga, pleurochrysis scherffelii pringsheim. Journal of Cell Biology, 41 (1):109–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/1/109.

Bruns:1969:SEF

[Bru69] Romaine R. Bruns. A symmetrical, extracellular fibril. Journal of Cell Biology, 42(2):418–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/42/2/418.

Bryant:1966:ITT

[Bry66] B. J. Bryant. The incorporation of tritium from thymidine into proteins of the mouse. Journal of Cell Biology, 29(1):29–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/29/1/29.

Beidler:1965:RCW

[BS65] Lloyd M. Beidler and Ronald L. Smallman. Renewal of cells within taste buds. Journal of Cell Biology, 27(2):263–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/27/
2/263.

Biscoe:1966:UCB

[BS66a] T. J. Biscoe and W. E. Stehbens. Ultrastructure of the carotid body. Journal of Cell Biology, 30(3):563–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/30/3/563.

Bohm:1966:FHN

[BS66b] N. Böhm and W. Sandritter. Feulgen hydrolysis of normal cells and mouse ascites tumor cells. Journal of Cell Biology, 28 (1):1–??, January 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/28/1/1.

Bachmann:1967:ASE

[BS67] L. Bachmann and M. M. Salpeter. Absolute sensitivity of electron microscope radioautography. Journal of Cell Biology, 33(2):299– ??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
2/299.

Budd:1969:DLN

[BS69] G. C. Budd and M. M. Salpeter. The distribution of labeled norepinephrine within sympathetic nerve terminals studied with electron microscope radioautography. Journal of Cell Biology, 41(1):21–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/1/21.

Barland:1968:LHA

[BSH68] Peter Barland, Carol Smith, and David Hamerman. Localization of hyaluronic acid in synovial cells by radioautography. Journal of Cell Biology, 37(1):13–??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/37/1/13.

Borisy:1967:MACa

[BT67a] G. G. Borisy and E. W. Taylor. The mechanism of action of colchicine. Journal of Cell Biology, 34(2):525–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/34/2/525.

Borisy:1967:MACb

[BT67b] G. G. Borisy and E. W. Taylor. The mechanism of action of colchicine. Journal of Cell Biology, 34(2):535–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/34/2/535.

Buschmann:1968:EAL

[BT68] Robert J. Buschmann and A. B. Taylor. Extraction of absorbed lipid (linoleic acid-1-^{14} c) from rat intestinal epithelium during processing for electron microscopy. Journal of Cell Biology, 38(1):252–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/1/252.

Brzin:1966:AFS

[BTD66] Miro Brzin, Virginia M. Tennyson, and Philip E. Duffy. Acetylcholinesterase in frog sympathetic and dorsal root ganglia. Journal of Cell Biology, 31(2):215–??, November 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/31/2/215.

Bottomley:1969:ECC

[BTG69] R. H. Bottomley, A. L. Trainer, and M. J. Griffin. Enzymatic and chromosomal characterization of HeLa variants. Journal of Cell Biology, 41(3):806–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/41/3/806.

Bagnara:1968:DCU

[BTH68] Joseph T. Bagnara, John D. Taylor, and Mac E. Hadley. The dermal chromatophore unit. Journal of Cell Biology, 38(1):67–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/38/1/67.

Bibbiani:1969:QDA

[BTVM69] Cecilia Bibbiani, Roberto Tongiani, and Maria Pia Viola-Magni. I. quantitative determination of the amount of DNA per nucleus by interference microscopy. Journal of Cell Biology, 42(2):444–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
2/444.

Bucciarelli:1966:ICR

[Buc66] Emilio Bucciarelli. Intranuclear cisternae resembling structures of the Golgi complex. Journal of Cell Biology, 30(3): 664–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/30/3/664.

Bulger:1969:UPP

[Bul69] Ruth Ellen Bulger. Use of potassium pyroantimonate in the localization of sodium ions in rat kidney tissue. Journal of Cell Biology, 40(1):79–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/40/1/79.

Bradley:1966:ASF

[BW66] Alan S. Bradley and Steven L. Wissig. The anatomy of secretion in the follicular cell of the thyroid gland. Journal of Cell Biology, 30(2):433–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/30/2/433.

Bartels:1967:PAP

[BW67] Paul G. Bartels and T. Elliot Weier. Particle arrangements in proplastids of triticum vulgare l. seedlings. Journal of Cell Biology, 33(2):243–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/33/2/243.

Byers:1966:RCI

[Bye66] Breck Byers. Ribosome crystallization induced in chick embryo tissues by hypothermia. Journal of Cell Biology, 30 (3):??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/30/3/C1.

Charret:1968:SAM

[CA68] Renée Charret and Jean André. La SynthÈSe De l’adn mitochondrial chez Tetrahymena Pyriformis. Journal of Cell Biology, 39 (2):369–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/39/2/369.

Cameron:1965:MEL

[Cam65] Ivan L. Cameron. Macromolecular events leading to cell division in Tetrahymena Pyriformis after removal and replacement of required pyrimidines. Journal of Cell Biology, 25(2):9–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/25/2/9.

Cancilla:1968:DLG

[Can68] P. A. Cancilla. Demonstration of the langerhans granule by lanthanum. Journal of Cell Biology, 38(1):248–??, July 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/38/1/248.

Carroll:1967:UAD

[Car67] George C. Carroll. The ultrastructure of ascospore delimitation in saccobolus kerverni. Journal of Cell Biology, 33(1):218–??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
1/218.

Carell:1969:SCD

[Car69a] Edgar F. Carell. Studies on chloroplast development and replication in euglena. Journal of Cell Biology, 41(2):431–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/41/2/431.

Caro:1969:CSD

[Car69b] Lucien G. Caro. A common source of difficulty in high-resolution radioautography. Journal of Cell Biology, 41(3):918–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/41/3/
918.

Cave:1966:ITL

[Cav66] Mac Donald Cave. Incorporation of tritium-labeled thymidine and Lysine into chromosomes of cultured human leukocytes. Journal of Cell Biology, 29(2):209–??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/29/2/209.

Cronshaw:1965:FSD

[CB65] James Cronshaw and G. Benjamin Bouck. The fine structure of differentiating xylem elements. Journal of Cell Biology, 24 (3):415–??, March 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/3/415.

Carney:1966:EMS

[CB66] John A. Carney and Arnold L. Brown. An electron microscope study of canine cardiac myosin and some of its aggregates. Journal of Cell Biology, 28(2):375–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/28/2/375.

Canonico:1969:UAO

[CB69a] Peter G. Canonico and John W. C. Bird. The use of acridine orange as a lysosomal marker in rat skeletal muscle. Journal of Cell Biology, 43(2):367–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/43/2/367.

Cobb:1969:SNV

[CB69b] J. L. S. Cobb and Terence Bennett. A study of nexuses in visceral smooth muscle. Journal of Cell Biology, 41(1):287–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/41/
1/287.

Cardell:1967:ITA

[CBP67] Robert R. Cardell, Susan Badenhausen, and Keith R. Porter. Intestinal triglyceride absorption in the rat. Journal of Cell Biology, 34(1):123–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/34/1/123.

Cummins:1965:EAM

[CBR65] J. E. Cummins, E. N. Brewer, and H. P. Rusch. The effect of actidione on mitosis in the slime mold Physarum Polycephalum. Journal of Cell Biology, 27(2):337–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/27/2/337.

Carstensen:1969:SEF

[CCIC69] E. L. Carstensen, A. Coopersmith, M. Ingram, and S. Z. Child. Stability of erythrocytes fixed in osmium tetroxide solutions. Journal of Cell Biology, 42(2):565–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/42/2/565.

Cronshaw:1967:TFC

[CE67] James Cronshaw and Katherine Esau. Tubular and fibrillar components of mature and differentiating sieve elements. Journal of Cell Biology, 34(3):801–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/3/801.

Cronshaw:1968:PPPa

[CE68a] James Cronshaw and Katherine Esau. P protein in the phloem of cucurbita. Journal of Cell Biology, 38(1):25–??, July 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/38/1/25.

Cronshaw:1968:PPPb

[CE68b] James Cronshaw and Katherine Esau. P protein in the phloem of cucurbita. Journal of Cell Biology, 38(2):292–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/38/2/292.

Carmichael:1966:FSO

[CF66a] G. G. Carmichael and H. M. Fullmer. The fine structure of the oxytalan fiber. Journal of Cell Biology, 28(1):33–??, January 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/28/1/33.

Chapman:1966:PSP

[CF66b] D. Chapman and D. J. Fluck. Physical studies of phospholipids. Journal of Cell Biology, 30(1):1–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/30/1/1.

Carlsen:1965:CUGa

[CFK65a] Frits Carlsen, Franklin Fuchs, and Gustav G. Knappeis. Contractility and ultrastructure in glycerol-extracted muscle fibers. Journal of Cell Biology, 27(1):25–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/27/1/25.

Carlsen:1965:CUGb

[CFK65b] Frits Carlsen, Franklin Fuchs, and Gustav G. Knappeis. Contractility and ultrastructure in glycerol-extracted muscle fibers. Journal of Cell Biology, 27(1):35–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/27/1/35.

Cameron:1965:NBC

[CG65] Ivan L. Cameron and E. Ernest Guile. Nucleolar and biochemical changes during unbalanced growth of Tetrahymena Pyriformis. Journal of Cell Biology, 26(3):845–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/26/3/845.

Caplan:1966:BUP

[CG66] Arnold I. Caplan and John W. Greenawalt. Biochemical and ultrastructural properties of osmotically lysed rat-liver mitochondria. Journal of Cell Biology, 31(3):455–??, December 1966. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/31/3/455.

Caplan:1968:EOL

[CG68] Arnold I. Caplan and John W. Greenawalt. The effects of osmotic lysis on the oxidative phosphorylation and compartmentation of rat liver mitochondria. Journal of Cell Biology, 36(1):15–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/36/1/15.

Craig:1969:SOR

[CG69] Nessly Craig and Lester Goldstein. Studies on the origin of ribosomes in Amoeba Proteus. Journal of Cell Biology, 40(3):622–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
3/622.

Chakravarty:1965:GRP

[Cha65] Nirmal Chakravarty. Glycolysis in rat peritoneal mast cells. Journal of Cell Biology, 25(2):123–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/25/2/123.

Coleman:1965:BCN

[CHB65] John R. Coleman, Heinz Herrmann, and Barbara Bess. Biosynthesis of collagen and non-collagen protein during development of the chick cornea. Journal of Cell Biology, 25(1):69–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/25/1/69.

Cronshaw:1966:UFB

[CHE66] J. Cronshaw, L. Hoefert, and K. Esau. Ultrastructural features of beta leaves infected with beet yellows virus. Journal of Cell Biology, 31(3):429–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/31/3/429.

Chevaillier:1967:MEE

[Che67] Philippe Chevaillier. Mise en evidence et etude cytochimique d’une proteine basique extranucleaire dans les spermatozoides des crustaces decapodes. Journal of Cell Biology, 32(3):547–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/32/
3/547.

Choi:1965:EMA

[Cho65] Jae Kwon Choi. Electron microscopy of absorption of tracer materials by toad urinary bladder epithelium. Journal of Cell Biology, 25(2):175–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/2/175.

Christensen:1965:FST

[Chr65] A. Kent Christensen. The fine structure of testicular interstitial cells in guinea pigs. Journal of Cell Biology, 26(3): 911–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/3/911.

Carothers:1967:SSH

[CK67] Zane B. Carothers and Gerald L. Kreitner. Studies of spermatogenesis in the hepaticae. Journal of Cell Biology, 33(1):43–??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
1/43.

Carothers:1968:SSH

[CK68a] Zane B. Carothers and Gerald L. Kreitner. Studies of spermatogenesis in the hepaticae. Journal of Cell Biology, 36(3):603–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
3/603.

Cotran:1968:USP

[CK68b] Ramzi S. Cotran and Morris J. Karnovsky. Ultrastructural studies on the permeability of the mesothelium to horseradish peroxidase. Journal of Cell Biology, 37(1):123–??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/37/1/123.

Coggeshall:1966:MFS

[CKKW66] R. E. Coggeshall, E. R. Kandel, I. Kupfermann, and R. Waziri. A morphological and functional study on a cluster of identifiable neurosecretory cells in the abdominal ganglion of Aplysia Californica. Journal of Cell Biology, 31(2):363–??, November 1966. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/31/2/363.

Coimbra:1966:SGS

[CL66] Antonio Coimbra and C. P. Leblond. Sites of glycogen synthesis in rat liver cells as shown by electron microscope radioautography after administration of glucose-H^{3}. Journal of Cell Biology, 30(1):151–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/30/1/151.

Combs:1965:DPE

[CLB65] J. W. Combs, D. Lagunoff, and E. P. Benditt. Differentiation and proliferation of embryonic mast cells of the rat. Journal of Cell Biology, 25(3):577–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/25/3/577.

Christiansen:1965:SPA

[CM65] Richard G. Christiansen and John M. Marshall. A study of phagocytosis in the ameba chaos chaos. Journal of Cell Biology, 25(3):443–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/3/443.

Costoff:1969:IBP

[CM69] Allen Costoff and W. H. McShan. Isolation and biological properties of secretory granules from rat anterior pituitary glands. Journal of Cell Biology, 43(3):564–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/43/3/564.

Cheng-Minoda:1968:FSE

[CMDLB68] K. Cheng-Minoda, J. Davidowitz, A. Liebowitz, and G. M. Breinin. Fine structure of extraocular muscle in rabbit. Journal of Cell Biology, 39(1):193–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/1/193.

Cunningham:1966:SIP

[CMM66] William P. Cunningham, D. James Morré, and H. H. Mollenhauer. Structure of isolated plant Golgi apparatus revealed by negative staining. Journal of Cell Biology, 28(2):169–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/28/
2/169.

Cobb:1969:MCG

[Cob69] James L. S. Cobb. Modified cilia in the gut of the lamellibranch mollusc tapes watlingi. Journal of Cell Biology, 43(1):192–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/43/
1/192.

Coggeshall:1966:FSA

[Cog66] Richard E. Coggeshall. A fine structural analysis of the epidermis of the earthworm, Lumbricus Terrestris L. Journal of Cell Biology, 28(1):95–??, January 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/28/1/95.

Cohen:1968:NES

[Coh68] Adolph I. Cohen. New evidence supporting the linkage to extracellular space of outer segment saccules of frog cones but not rods. Journal of Cell Biology, 37(2):424–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/37/2/424.

Colonnier:1965:NIR

[Col65] M. Colonnier. On the nature of intranuclear rods. Journal of Cell Biology, 25(3):646–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/25/3/646.

Combs:1966:MRM

[Com66a] J. W. Combs. Maturation of rat mast cells. Journal of Cell Biology, 31(3):563–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/31/3/563.

Comings:1966:UHR

[Com66b] David E. Comings. Uridine-5-H^{3} radioautography of the human sex chromatin body. Journal of Cell Biology, 28(3):437–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/28/
3/437.

Comings:1967:HGA

[Com67] David E. Comings. Histones of genetically active and inactive chromatin. Journal of Cell Biology, 35(3):699–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/35/
3/699.

Cook:1966:SCD

[Coo66] J. R. Cook. The synthesis of cytoplasmic DNA in synchronized euglena. Journal of Cell Biology, 29(2):369–??, May 1966. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/29/2/369.

Camargo:1967:RDD

[CP67a] E. Plessmann Camargo and W. Plaut. The radioautographic detection of DNA with tritiated actinomycin D. Journal of Cell Biology, 35(3):713–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/35/3/713.

Carey:1967:TLS

[CP67b] Francis J. Carey and Olive S. Pettengill. A time-lapse study of effects of anticellular antibody on membrane mobility and phagocytic activity of HeLa cells. Journal of Cell Biology, 33(3):709–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
3/709.

Cooper:1968:IAP

[CP68] George W. Cooper and Darwin J. Prockop. Intracellular accumulation of protocollagen and extrusion of collagen by embryonic cartilage cells. Journal of Cell Biology, 38(3):523–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
3/523.

Clementi:1969:ICa

[CP69a] F. Clementi and G. E. Palade. Intestinal capillaries. Journal of Cell Biology, 41(1):33–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/41/1/33.

Clementi:1969:ICb

[CP69b] Francesco Clementi and George E. Palade. Intestinal capillaries. Journal of Cell Biology, 42(3):706–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/42/3/706.

Cameron:1967:CFS

[CPR67] D. A. Cameron, H. A. Paschall, and R. A. Robinson. Changes in the fine structure of bone cells after the administration of parathyroid extract. Journal of Cell Biology, 33(1):1–??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/33/1/1.

Cummins:1966:LDS

[CR66] J. E. Cummins and H. P. Rusch. Limited DNA synthesis in the absence of protein synthesis in Physarum Polycephalum. Journal of Cell Biology, 31(3):577–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/31/3/577.

Cleary:1967:CCC

[CSJ67] E. G. Cleary, L. B. Sandberg, and D. S. Jackson. The changes in chemical composition during development of the bovine nuchal ligament. Journal of Cell Biology, 33(3):469–??, June 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/33/3/469.

Cassidy:1967:CMI

[CT67] M. M. Cassidy and C. S. Tidball. Cellular mechanism of intestinal permeability alterations produced by chelation depletion. Journal of Cell Biology, 32(3):685–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/32/3/685.

Cohn:1967:CDC

[CvD67] Norman S. Cohn and P. van Duijn. Constancy of DNA content in adrenal medulla nuclei of cold-treated rats. Journal of Cell Biology, 33(2):349–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/33/2/349.

Clark-Walker:1967:BMS

[CWL67] G. D. Clark-Walker and Anthony W. Linnane. The biogenesis of mitochondria in Saccharomyces Cerevisiae. Journal of Cell Biology, 34(1):1–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/34/1/1.

Chen:1969:EMR

[CY69] I-Li Chen and Robert D. Yates. Electron microscopic radioautographic studies of the carotid body following injections of labeled biogenic amine precursors. Journal of Cell Biology, 42(3): 794–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/3/794.

Chen:1969:ERH

[CYD69] I-Li Chen, Robert D. Yates, and Donald Duncan. The effects of reserpine and hypoxia on the amine-storing granules of the hamster carotid body. Journal of Cell Biology, 42(3):804–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
3/804.

Dass:1965:SSR

[DB65] C. M. S. Dass and S. T. Bayley. A structural study of rat liver ribosomes. Journal of Cell Biology, 25(1):9–??, April 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/25/1/9.

Doolin:1966:UOC

[DB66] Paul F. Doolin and Wesley J. Birge. Ultrastructural organization of cilia and basal bodies of the epithelium of the choroid Plexus in the chick embryo. Journal of Cell Biology, 29(2):333– ??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/29/
2/333.

Diamond:1968:FCU

[DB68a] Jared M. Diamond and William H. Bossert. Functional consequences of ultrastructural geometry in “backwards” fluidtransporting epithelia. Journal of Cell Biology, 37(3):694–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/37/
3/694.

Dobelle:1968:CCG

[DB68b] William H. Dobelle and Michael Beer. Chemically cleaved graphite support films for electron microscopy. Journal of Cell Biology, 39(3):733–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/39/3/733.

Deamer:1969:USR

[DB69] D. W. Deamer and R. J. Baskin. Ultrastructure of sarcoplasmic reticulum preparations. Journal of Cell Biology, 42(1):296–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
1/296.

Deter:1967:PLC

[DBdD67] Russell L. Deter, Pierre Baudhuin, and Christian de Duve. Participation of lysosomes in cellular autophagy induced in rat liver by glucagon. Journal of Cell Biology, 35(2):??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/35/2/C11.

Dallner:1968:HRS

[DBN68] Gustav Dallner, Anders Bergstrand, and Robert Nilsson. Heterogeneity of rough-surfaced liver microsomal membranes of adult, phenobarbital-treated, and newborn rats. Journal of Cell Biology, 38(2):257–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/38/2/257.

Verster:1965:MBC

[dBVSH65] F. de Balbian Verster, O. Z. Sellinger, and J. C. Harkin. Morphological and biochemical correlates of cerebral microsomes. Journal of Cell Biology, 25(2):69–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/25/2/69.

Deamer:1967:ATX

[DC67] David W. Deamer and Antony Crofts. Action of triton x-100 on chloroplast membranes. Journal of Cell Biology, 33(2):395–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
2/395.

DiBona:1969:AST

[DCL69] Donald R. DiBona, Mortimer M. Civan, and Alexander Leaf. The anatomic site of the transepithelial permeability barriers of toad bladder. Journal of Cell Biology, 40(1):1–??, January 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/40/1/1.

Davenport:1965:CSC

[DD65] Richard Davenport and Janice C. Davenport. A cytochemical study of cytoplasmic basic proteins in the ascidian oocyte. Journal of Cell Biology, 25(2):319–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/25/2/319.

Dwivedi:1969:ICC

[DDB69] R. S. Dwivedi, S. K. Dutta, and D. P. Bloch. Isolation and characterization of chromatin from neurospora crassa. Journal of Cell Biology, 43(1):51–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/43/1/51.

Dallman:1969:HDE

[DDBE69] Peter R. Dallman, Gustav Dallner, Anders Bergstrand, and Lars Ernster. Heterogeneous distribution of enzymes in submicrosomal membrane fragments. Journal of Cell Biology, 41(2):357–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/41/
2/357.

Deter:1967:IGI

[DdD67] Russell L. Deter and Christian de Duve. Influence of glucagon, an inducer of cellular autophagy, on some physical properties of rat liver lysosomes. Journal of Cell Biology, 33(2):437–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
2/437.

Daniel:1966:AEV

[DDGL66] Mary R. Daniel, J. T. Dingle, Audrey M. Glauert, and J. A. Lucy. The action of excess of vitamin a alcohol on the fine structure of rat dermal fibroblasts. Journal of Cell Biology, 30(3): 465–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/30/3/465.

Daneholt:1969:DBC

[DE69] B. Daneholt and J.-E. Edström. The DNA base composition of individual chromosomes and chromosome segments from chironomus tentans. Journal of Cell Biology, 41(2):620–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/41/2/620.

Dembitzer:1968:DDA

[Dem68] Herbert M. Dembitzer. Digestion and the distribution of acid phosphatase in blepharisma. Journal of Cell Biology, 37(2):329– ??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/37/
2/329.

Dingle:1966:DFA

[DF66] Allan D. Dingle and Chandler Fulton. Development of the flagellar apparatus of naegleria. Journal of Cell Biology, 31(1):43–??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/31/
1/43.

David-Ferreira:1965:EMS

[DFM65] J. F. David-Ferreira and R. A. Manaker. An electron microscope study of the development of a mouse hepatitis virus in tissue culture cells. Journal of Cell Biology, 24(1):57–??, January 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/24/1/57.

Lima-de-Faria:1966:UCM

[dFM66] A. Lima de Faria and M. J. Moses. Ultrastructure and cytochemistry of metabolic DNA in tipula. Journal of Cell Biology, 30(1):177–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/30/1/177.

Dobb:1967:FSS

[DFM67] M. G. Dobb, R. D. B. Fraser, and T. P. Macrae. The fine structure of silk fibroin. Journal of Cell Biology, 32(2):289–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/32/
2/289.

Duvall:1966:USH

[DFW66] Arndt J. Duvall, Åke Flock, and Jan Wersäll. The ultrastructure of the sensory hairs and associated organelles of the cochlear inner hair cell, with reference to directional sensitivity. Journal of Cell Biology, 29(3):497–??, June 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/29/3/497.

Diacumakos:1965:CCN

[DGT65] E. G. Diacumakos, L. Garnjobst, and E. L. Tatum. A cytoplasmic character in neurospora crassa. Journal of Cell Biology, 26(2):427–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/2/427.

Dembitzer:1966:SNC

[DH66] H. M. Dembitzer and H. I. Hirshfield. Some new cytological observations in the heterotrichous ciliate, blepharisma. Journal of Cell Biology, 30(1):201–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/30/1/201.

Doggenweiler:1967:UPN

[DH67] C. F. Doggenweiler and John E. Heuser. Ultrastructure of the prawn nerve sheaths. Journal of Cell Biology, 34(2):407–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
2/407.

Diers:1966:PMO

[Die66] Lothar Diers. On the plastids, mitochondria, and other cell constituents during oögenesis of a plant. Journal of Cell Biology, 28(3):527–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/28/3/527.

Dietert:1967:CCO

[Die67] Scott E. Dietert. The columnar cells occurring in the parietal layer of Bowman’s capsule. Journal of Cell Biology, 35(2): 435–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/2/435.

Dingemans:1969:RBC

[Din69] K. P. Dingemans. The relation between cilia and mitoses in the mouse adenohypophysis. Journal of Cell Biology, 43(2): 361–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/43/2/361.

Davis:1967:EML

[DK67] Richard Davis and George B. Koelle. Electron microscopic localization of acetylcholinesterase and nonspecific cholinesterase at the neuromuscular junction by the gold-thiocholine and goldthiolacetic acid methods. Journal of Cell Biology, 34(1):157–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
1/157.

Djordjevic:1968:DLE

[DK68] Bozidar Djordjevic and Jae Ho Kim. Different lethal effects of mitomycin C and actinomycin D during the division cycle of HeLa cells. Journal of Cell Biology, 38(3):477–??, September 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/38/3/477.

Doyle:1969:SLS

[DL69] D. Doyle and H. Laufer. Sources of larval salivary gland secretion in the dipteran chironomus tentans. Journal of Cell Biology, 40 (1):61–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/1/61.

Das:1967:CBP

[DMEA67] Nirmal K. Das, Julie Micou-Eastwood, and Max Alfert. Cytochemical and biochemical properties of basic proteins of urechis acrosomes. Journal of Cell Biology, 35(2):455–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/35/
2/455.

deMan:1967:LEM

[dMN67] J. C. H. de Man and N. J. A. Noorduyn. Light and electron microscopic radioautography of hepatic cell nucleoli in mice treated with actinomycin D. Journal of Cell Biology, 33(3):489– ??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
3/489.

Dallner:1966:MCE

[DN66] Gustav Dallner and Robert Nilsson. Mechanism of the cation effect in subfractionation of microsomes. Journal of Cell Biology, 31(1):181–??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/31/1/181.

Damsky:1969:MAG

[DNC69] Caroline H. Damsky, William M. Nelson, and Albert Claude. Mitochondria in anaerobically-grown, lipid-limited brewer’s yeast. Journal of Cell Biology, 43(1):174–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/43/1/174.

Dahl:1967:EMN

[DP67] Rolf H. Dahl and James N. Pratley. The effects of magnesium on nucleoside phosphatase activity in frog skin. Journal of Cell Biology, 33(2):411–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/33/2/411.

Droller:1966:EMS

[DR66] Michael J. Droller and Thomas F. Roth. An electron microscope study of yolk formation during oogenesis in lebistes reticulatus guppyi. Journal of Cell Biology, 28(2):209–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/28/2/209.

Dafni:1966:CPP

[DS66] Zipora Dafni and M. Shilo. The cytotoxic principle of the phytoflagellate prymnesium parvum. Journal of Cell Biology, 28(3):461–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/28/3/461.

Devine:1968:LTN

[DS68] C. E. Devine and F. O. Simpson. Localization of tritiated norepinephrine in vascular sympathetic axons of the rat intestine and mesentery by electron microscope radioautography. Journal of Cell Biology, 38(1):184–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/38/1/184.

Dietert:1969:UBS

[DS69a] Scott E. Dietert and Terence J. Scallen. An ultrastructural and biochemical study of the effects of three inhibitors of cholesterol biosynthesis upon murine adrenal gland and testis. Journal of Cell Biology, 40(1):44–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/40/1/44.

Dunham:1969:IPT

[DS69b] Philip B. Dunham and Larry C. Stoner. Indentation of the pellicle of Tetrahymena at the contractile vacuole pore before systole. Journal of Cell Biology, 43(1):184–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/43/1/184.

Das:1965:SAD

[DSA65] Nirmal K. Das, Elsie P. Siegel, and Max Alfert. Synthetic activities during spermatogenesis in the locust. Journal of Cell Biology, 25(2):387–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/2/387.

DiMarco:1965:IEN

[DSDD65] A. Di Marco, R. Silvestrini, S. Di Marco, and T. Dasdia. Inhibiting effect of the new cytotoxic antibiotic daunomycin on nucleic acids and mitotic activity of HeLa cells. Journal of Cell Biology, 27(3):545–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/27/3/545.

Dallner:1966:BERa

[DSP66a] Gustav Dallner, Philip Siekevitz, and George E. Palade. Biogenesis of endoplasmic reticulum membranes. Journal of Cell Biology, 30(1):73–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/30/1/73.

Dallner:1966:BERb

[DSP66b] Gustav Dallner, Philip Siekevitz, and George E. Palade. Biogenesis of endoplasmic reticulum membranes. Journal of Cell Biology, 30(1):97–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/30/1/97.

deFWebster:1965:RIC

[dWA65] Henry deF. Webster and Adelbert Ames. Reversible and irreversible changes in the fine structure of nervous tissue during oxygen and glucose deprivation. Journal of Cell Biology, 26(3): 885–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/3/885.

deZoeten:1969:PII

[dZG69] G. A. de Zoeten and G. Gaard. Possibilities for inter- and intracellular translocation of some icosahedral plant viruses. Journal of Cell Biology, 40(3):814–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/40/3/814.

Eakin:1965:DRC

[Eak65] Richard M. Eakin. Differentiation of rods and cones in total darkness. Journal of Cell Biology, 25(1):162–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/25/1/162.

Estensen:1966:PIN

[EB66] Richard D. Estensen and Renato Baserga. Puromycin-induced necrosis of crypt cells of the small intestine of mouse. Journal of Cell Biology, 30(1):13–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/30/1/13.

Eliasson:1967:RDP

[EBH67] Eva Eliasson, G. Eric Bauer, and Tore Hultin. Reversible degradation of polyribosomes in Chang cells cultured in a glutaminedeficient medium. Journal of Cell Biology, 33(2):287–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/33/2/
287.

Ebstein:1967:TAD

[Ebs67] Bonnie Sue Ebstein. Tritiated actinomycin D as a cytochemical label for small amounts of DNA. Journal of Cell Biology, 35 (3):709–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/3/709.

Eichholz:1965:SOB

[EC65] Alexander Eichholz and Robert K. Crane. Studies on the organization of the brush border in intestinal epithelial cells. Journal of Cell Biology, 26(3):687–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/26/3/687.

Esau:1967:RTM

[EC67] Katherine Esau and James Cronshaw. Relation of tobacco mosaic virus to the host cells. Journal of Cell Biology, 33(3):665–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
3/665.

Ehrenreich:1968:FHP

[EC68] Barbara A. Ehrenreich and Zanvil A. Cohn. Fate of hemoglobin pincytosed by macrophages in vitro. Journal of Cell Biology, 38(1):244–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/1/244.

Echlin:1965:AHA

[Ech65] Patrick Echlin. An apparent helical arrangement of ribosomes in developing pollen mother cells of ipomoea purpurea (l.) Roth. Journal of Cell Biology, 24(1):150–??, January 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/24/1/150.

Esau:1967:RBY

[ECH67] K. Esau, J. Cronshaw, and L. L. Hoefert. Relation of beet yellows virus to the phloem and to movement in the sieve tube. Journal of Cell Biology, 32(1):71–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/32/1/71.

Eisenberg:1968:SDS

[EE68] Brenda Eisenberg and Robert S. Eisenberg. Selective disruption of the sarcotubular system in frog sartorius muscle. Journal of Cell Biology, 39(2):451–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/2/451.

Ernst:1969:DSS

[EE69] Stephen A. Ernst and Richard A. Ellis. The development of surface specialization in the secretory epithelium of the avian salt gland in response to osmotic stress. Journal of Cell Biology, 40(2):305–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/2/305.

Edelman:1965:ISM

[EEKS65] Jean C. Edelman, P. Michael Edelman, Karl M. Knigge, and Irving L. Schwartz. Isolation of skeletal muscle nuclei. Journal of Cell Biology, 27(2):365–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/27/2/365.

Elsdale:1969:MAM

[EF69] Tom Elsdale and Robert Foley. Morphogenetic aspects of multilayering in Petri dish cultures of human fetal lung fibroblasts. Journal of Cell Biology, 41(1):298–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/41/1/298.

Enestrom:1968:RMC

[EH68] Sverker Eneström and Anders Hamberger. Respiration and mitochondrial content in single neurons of the supraoptic nucleus. Journal of Cell Biology, 38(3):483–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/38/3/483.

Ezerman:1967:DSR

[EI67] Elizabeth B. Ezerman and Harunori Ishikawa. Differentiation of the sarcoplasmic reticulum and T system in developing chick skeletal muscle in vitro. Journal of Cell Biology, 35(2): 405–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/2/405.

Emeson:1966:NFA

[EKG66] Eugene E. Emeson, Yutaka Kikkawa, and Boris Gueft. New features of amyloid found after digestion with trypsin. Journal of Cell Biology, 28(3):570–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/28/3/570.

Elkind:1969:CKA

[EKSG69] M. M. Elkind, E. Kano, and H. Sutton-Gilbert. Cell Killing by actinomycin D in relation to the growth cycle of Chinese hamster cells. Journal of Cell Biology, 42(2):366–??, August 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/42/2/366.

Ellis:1965:FSM

[Ell65] Richard A. Ellis. Fine structure of the myoepithelium of the eccrine sweat glands of man. Journal of Cell Biology, 27(3): 551–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/3/551.

Essner:1965:NPA

[ENQ65] Edward Essner, Alex B. Novikoff, and Nelson Quintana. Nucleoside phosphatase activities in rat cardiac muscle. Journal of Cell Biology, 25(2):201–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/25/2/201.

Echandia:1968:MNF

[EP68] E. L. Rodriguez Echandia and R. S. Piezzi. Microtubules in the nerve fibers of the toad bufo arenarum Hensel. Journal of Cell Biology, 39(2):491–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/2/491.

Esterhuizen:1968:ABA

[ESL68] A. C. Esterhuizen, T. L. B. Spriggs, and J. D. Lever. Axon beadings in autonomic cholinergic nerves. Journal of Cell Biology, 38(2):454–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/2/454.

Engel:1966:CPS

[ET66] A. G. Engel and Lois W. Tice. Cytochemistry of phosphatases of the sarcoplasmic reticulum. Journal of Cell Biology, 31(3): 473–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/31/3/473.

Enger:1969:RSC

[ET69] M. D. Enger and R. A. Tobey. RNA synthesis in Chinese hamster cells. Journal of Cell Biology, 42(1):308–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/42/1/308.

Enger:1968:RSC

[ETS68] M. D. Enger, R. A. Tobey, and A. G. Saponara. RNA synthesis in Chinese hamster cells. Journal of Cell Biology, 36(3):583–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
3/583.

Elbers:1965:EMX

[EV65] P. F. Elbers and P. H. J. T. Ververgaert. Electron microscope and X-ray diffraction studies on a homologous series of saturated phosphatidylcholines. Journal of Cell Biology, 25(2):375– ??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
2/375.

Elbers:1965:TFP

[EVD65] P. F. Elbers, P. H. J. T. Ververgaert, and R. Demel. Tricomplex fixation of phospholipids. Journal of Cell Biology, 24(1):23–??, January 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/24/
1/23.

Everingham:1968:AIA

[Eve68a] John W. Everingham. Attachment of intranuclear annulate lamellae to the nuclear envelope. Journal of Cell Biology, 37 (2):540–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/37/2/540.

Everingham:1968:IAL

[Eve68b] John W. Everingham. Intranuclear annulate lamellae in ascidian embryos. Journal of Cell Biology, 37(2):551–??, May 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/37/2/551.

Elliott:1968:GAA

[EZ68] Alfred M. Elliott and Roger G. Zieg. A Golgi apparatus associated with mating in Tetrahymena Pyriformis. Journal of Cell Biology, 36(2):391–??, February 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/36/2/391.

Fabrikant:1968:KCP

[Fab68] Jacob I. Fabrikant. The kinetics of cellular proliferation in regenerating liver. Journal of Cell Biology, 36(3):551–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/36/3/551.

Fahrenbach:1967:FSF

[Fah67] Wolf H. Fahrenbach. The fine structure of fast and slow crustacean muscles. Journal of Cell Biology, 35(1):69–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/35/1/69.

Fahimi:1969:CLP

[Fah69] H. Dariush Fahimi. Cytochemical localization of peroxidatic activity of catalase in rat hepatic microbodies (peroxisomes). Journal of Cell Biology, 43(2):275–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/43/2/275.

Falk:1969:FVP

[Fal69a] Heinz Falk. Fusiform vesicles in plant cells. Journal of Cell Biology, 43(1):167–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/43/1/167.

Falk:1969:RTP

[Fal69b] Heinz Falk. Rough thylakoids: Polysomes attached to chloroplast membranes. Journal of Cell Biology, 42(2):582–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
2/582.

Fawcett:1968:SOC

[Faw68] Don W. Fawcett. The sporadic occurrence in cardiac muscle of anomalous Z bands exhibiting a periodic structure suggestive of tropomyosin. Journal of Cell Biology, 36(1):266–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
1/266.

Flaks:1966:SOF

[FB66] Bojan Flaks and Paul Bresloff. Some observations on the fine structure of the lutein cells of x-irradiated rat ovary. Journal of Cell Biology, 30(2):227–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/30/2/227.

Friedman:1968:FSM

[FBS68] Barry Friedman, Pierre Blais, and Patricia Shaffer. Fine structure of millipore filters. Journal of Cell Biology, 39(1):208–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/39/
1/208.

Fisher:1967:EMS

[FC67] Harold W. Fisher and T. W. Cooper. Electron microscope studies of the microvilli of HeLa cells. Journal of Cell Biology, 34 (2):569–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/34/2/569.

Frank:1968:LAP

[FC68] Arthur L. Frank and A. Kent Christensen. Localization of acid phosphatase in lipofuscin granules and possible autophagic vacuoles in interstitial cells of the guinea pig testis. Journal of Cell Biology, 36(1):1–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/36/1/1.

Florentin:1969:SDS

[FCLT69] R. A. Florentin, B. H. Choi, K. T. Lee, and W. A. Thomas. Stimulation of DNA synthesis and cell division in vitro by serum from cholesterol-Fed swine. Journal of Cell Biology, 41(2):641– ??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/41/
2/641.

Flock:1965:UKS

[FD65] Åke Flock and Arndt J. Duvall. The ultrastructure of the kinocilium of the sensory cells in the inner ear and lateral line organs. Journal of Cell Biology, 25(1):1–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/25/1/1.

Fukuyama:1969:SCP

[FE69] Kimie Fukuyama and William L. Epstein. Sulfur-containing proteins and epidermal keratinization. Journal of Cell Biology, 40 (3):830–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/3/830.

Feldherr:1965:EEO

[Fel65] Carl M. Feldherr. The effect of the electron-opaque pore material on exchanges through the nuclear annuli. Journal of Cell Biology, 25(1):43–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/1/43.

Feldherr:1966:NED

[Fel66] Carl M. Feldherr. Nucleocytoplasmic exchanges during cell division. Journal of Cell Biology, 31(1):199–??, October 1966. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/31/1/199.

Feldherr:1968:NED

[Fel68] Carl M. Feldherr. Nucleocytoplasmic exchanges during early interphase. Journal of Cell Biology, 39(1):49–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/39/1/49.

Feldherr:1969:CSN

[Fel69a] Carl M. Feldherr. A comparative study of nucleocytoplasmic interactions. Journal of Cell Biology, 42(3):841–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
3/841.

Feleppa:1969:HMI

[Fel69b] Ernest Feleppa. Holographic, motion-induced-contrast images. Journal of Cell Biology, 40(3):838–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/40/3/838.

Friend:1967:FCV

[FF67] Daniel S. Friend and Marilyn G. Farquhar. Functions of coated vesicles during protein absorption in the rat vas deferens. Journal of Cell Biology, 35(2):357–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/35/2/357.

Fleischer:1969:ICG

[FFO69] Becca Fleischer, Sidney Fleischer, and Hidehiro Ozawa. Isolation and characterization of Golgi membranes from bovine liver. Journal of Cell Biology, 43(1):59–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/43/1/59.

Fleischer:1967:FSL

[FFS67] Sidney Fleischer, Becca Fleischer, and Walther Stoeckenius. Fine structure of lipid-depleted mitochondria. Journal of Cell Biology, 32(1):193–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/32/1/193.

Fedorko:1965:CSG

[FH65] Martha E. Fedorko and James G. Hirsch. Crystalloid structure in granules of guinea pig basophils and human mast cells. Journal of Cell Biology, 26(3):973–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/26/3/973.

Fedorko:1966:CGF

[FH66] Martha E. Fedorko and James G. Hirsch. Cytoplasmic granule formation in myelocytes. Journal of Cell Biology, 29(2):307–??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/29/
2/307.

Fedorko:1968:AVPa

[FHC68a] Martha E. Fedorko, James G. Hirsch, and Zanvil A. Cohn. Autophagic vacuoles produced in vitro. Journal of Cell Biology, 38(2):377–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/2/377.

Fedorko:1968:AVPb

[FHC68b] Martha E. Fedorko, James G. Hirsch, and Zanvil A. Cohn. Autophagic vacuoles produced in vitro. Journal of Cell Biology, 38(2):392–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/2/392.

Fischman:1967:EMS

[Fis67] Donald A. Fischman. An electron microscope study of myofibril formation in embryonic chick skeletal muscle. Journal of Cell Biology, 32(3):557–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/32/3/557.

Fisher:1968:LER

[Fis68] Donald B. Fisher. Localization of endogenous RNA polymerase activity in frozen sections of plant tissues. Journal of Cell Biology, 39(3):745–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/39/3/745.

Fitz-James:1965:ISL

[FJH65] Philip Fitz-James and Ronald Hancock. The initial structural lesion of penicillin action in bacillus megaterium. Journal of Cell Biology, 26(2):657–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/26/2/657.

Fahimi:1966:CLT

[FK66] H. Dariush Fahimi and Morris J. Karnovsky. Cytochemical localization of two glycolytic dehydrogenases in white skeletal muscle. Journal of Cell Biology, 29(1):113–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/29/1/113.

Flaks:1968:FMG

[Fla68a] Bojan Flaks. Formation of membrane-glycogen arrays in rat hepatoma cells. Journal of Cell Biology, 36(2):410–??, February 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
2/410.

Flaks:1968:UAU

[Fla68b] Bojan Flaks. Unusual aspects of ultrastructural differentiation in rat hepatoma cells. Journal of Cell Biology, 38(1):230–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
1/230.

Flemming:1965:CKC

[Fle65] Walther Flemming. Contributions to the knowledge of the cell and its vital processes. Journal of Cell Biology, 25(1):3–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/25/1/3.

Flickinger:1965:FSN

[Fli65] Charles J. Flickinger. The fine structure of the nuclei of Tetrahymena Pyriformis throughout the cell cycle. Journal of Cell Biology, 27(3):519–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/27/3/519.

Flickinger:1968:EEC

[Fli68] Charles J. Flickinger. The effects of enucleation on the cytoplasmic membranes of Amoeba Proteus. Journal of Cell Biology, 37(2):300–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/37/2/300.

Flickinger:1969:DGC

[Fli69] Charles J. Flickinger. The development of Golgi complexes and their dependence upon the nucleus in amebae. Journal of Cell Biology, 43(2):250–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/43/2/250.

Florendo:1969:RSI

[Flo69] Noel T. Florendo. Ribosome substructure in intact mouse liver cells. Journal of Cell Biology, 41(1):335–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/41/1/335.

Flaxman:1968:UCA

[FLV68] B. Allen Flaxman, Marvin A. Lutzner, and Eugene J. Van Scott. Ultrastructure of cell attachment to substratum in vitro. Journal of Cell Biology, 36(2):406–??, February 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/36/2/406.

Fristrom:1965:PII

[FM65] J. W. Fristrom and H. K. Mitchell. The preparative isolation of imaginal discs from larvae of Drosophila Melanogaster. Journal of Cell Biology, 27(2):445–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/27/2/445.

Fernandez-Madrid:1967:BC

[FM67] F. Fernández-Madrid. Biosynthesis of collagen. Journal of Cell Biology, 33(1):27–??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/33/1/27.

Fawcett:1969:UCM

[FM69] Don W. Fawcett and N. Scott McNutt. The ultrastructure of the cat myocardium. Journal of Cell Biology, 42(1):1–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/42/1/1.

Frederick:1969:CLC

[FN69] Sue Ellen Frederick and Eldon H. Newcomb. Cytochemical localization of catalase in leaf microbodies (peroxisomes). Journal of Cell Biology, 43(2):343–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/43/2/343.

Foor:1968:ZFA

[Foo68] W. Eugene Foor. Zygote formation in ascaris lumbricoides (nematoda). Journal of Cell Biology, 39(1):119–??, October 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/39/1/119.

Forssmann:1969:ECE

[FOP^{+}69] W. G. Forssmann, L. Orci, R. Pictet, A. E. Renold, and C. Rouiller. The endocrine cells in the epithelium of the gastrointestinal mucosa of the rat. Journal of Cell Biology, 40(3):692–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
3/692.

Forer:1965:LRS

[For65] Arthur Forer. Local reduction of spindle fiber birefringence in living nephrotoma suturalis (loew) spermatocytes induced by ultraviolet microbeam irradiation. Journal of Cell Biology, 25(1):95– ??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
1/95.

Farquhar:1965:CJA

[FP65a] Marilyn G. Farquhar and George E. Palade. Cell junctions in amphibian skin. Journal of Cell Biology, 26(1):263–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/26/1/263.

Frasca:1965:RTD

[FP65b] Julio M. Frasca and Verta R. Parks. A routine technique for double-staining ultrathin sections using uranyl and lead salts. Journal of Cell Biology, 25(1):157–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/25/1/157.

Farquhar:1966:ATL

[FP66] Marilyn G. Farquhar and George E. Palade. Adenosine triphosphatase localization in amphibian epidermis. Journal of Cell Biology, 30(2):359–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/30/2/359.

Franklin:1965:MGM

[Fra65] Luther E. Franklin. Morphology of gamete membrane fusion and of sperm entry into oocytes of the sea urchin. Journal of Cell Biology, 25(2):81–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/25/2/81.

Franke:1966:INM

[Fra66] Werner W. Franke. Isolated nuclear membranes. Journal of Cell Biology, 31(3):619–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/31/3/619.

Frankel:1967:SMO

[Fra67] J. Frankel. Studies on the maintenance of oral development in Tetrahymena Pyriformis gl-c. Journal of Cell Biology, 34(3): 841–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/34/3/841.

Friend:1965:FSB

[Fri65] Daniel S. Friend. The fine structure of Brunner’s glands in the mouse. Journal of Cell Biology, 25(3):563–??, June 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/25/3/563.

Friend:1966:FSG

[Fri66] Daniel S. Friend. The fine structure of giardia muris. Journal of Cell Biology, 29(2):317–??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/29/2/317.

Fristrom:1968:CDW

[Fri68] Dianne Fristrom. Cellular degeneration in wing development of the mutant vestigial of Drosophila Melanogaster. Journal of Cell Biology, 39(2):488–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/39/2/488.

Friend:1969:CSM

[Fri69] Daniel S. Friend. Cytochemical staining of multivesicular body and Golgi vesicles. Journal of Cell Biology, 41(1):269–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/41/1/
269.

Franke:1969:NSM

[FS69] Werner W. Franke and Werner Schinko. Nuclear shape in muscle cells. Journal of Cell Biology, 42(1):326–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/42/1/326.

Fujita:1967:QAC

[Fuj67] Setsuya Fujita. Quantitative analysis of cell proliferation and differentiation in the cortex of the postnatal mouse cerebellum. Journal of Cell Biology, 32(2):277–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/32/2/277.

Fussell:1968:CLF

[Fus68] Catharine P. Fussell. Cytoplasmic label following tritiated thymidine treatment of Allium Cepa L. roots. Journal of Cell Biology, 39(2):264–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/39/2/264.

Feder:1965:SNA

[FW65] Ned Feder and Merrill K. Wolf. Studies on nucleic acid metachromasy. Journal of Cell Biology, 27(2):327–??, November 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/27/2/327.

Fischman:1967:FML

[FW67] Donald A. Fischman and George Weinbaum. The formation of multiple layers of membrane-like structures in Escherichia coli B. Journal of Cell Biology, 32(2):524–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/32/2/524.

Gall:1966:MFS

[Gal66] Joseph G. Gall. Microtubule fine structure. Journal of Cell Biology, 31(3):639–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/31/3/639.

Gall:1967:ONP

[Gal67] Joseph G. Gall. Octagonal nuclear pores. Journal of Cell Biology, 32(2):391–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/32/2/391.

Gomez-Acebo:1968:FSD

[GAPRC68] J. Gomez-Acebo, R. Parrilla, and J. L. R-Candela. Fine structure of the A and D cells of the rabbit endocrine pancreas in vivo and incubated in vitro. Journal of Cell Biology, 36(1):33–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
1/33.

Garant:1968:GMC

[Gar68] Philias R. Garant. Glycogen-membrane complexes within mouse striated muscle cells. Journal of Cell Biology, 36(3):648–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
3/648.

Gardner:1969:UTB

[Gar69] Robert S. Gardner. The use of tricine buffer in animal tissue cultures. Journal of Cell Biology, 42(1):320–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/42/1/320.

Glas:1966:QSM

[GB66a] Ulla Glas and Gunter F. Bahr. Quantitative study of mitochondria in rat liver. Journal of Cell Biology, 29(3):507–??, June 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/29/3/507.

Gupta:1966:CRS

[GB66b] Brij L. Gupta and Michael J. Berridge. A coat of repeating subunits on the cytoplasmic surface of the plasma membrane in the rectal papillae of the blowfly, Calliphora Erythrocephala (Meig.), studied in situ by electron microscopy. Journal of Cell Biology, 29(2):376–??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/29/2/376.

Goldfischer:1969:LAP

[GB69a] Sidney Goldfischer and Jay Bernstein. Lipofuscin (aging) pigment granules of the newborn human liver. Journal of Cell Biology, 42(1):253–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/1/253.

Griffin:1969:CCE

[GB69b] M. J. Griffin and R. Ber. Cell cycle events in the hydrocortisone regulation of alkaline phosphatase in HeLa S_{3} cells. Journal of Cell Biology, 40(2):297–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/40/2/297.

Groniowski:1969:EMS

[GBW69] J. Groniowski, W. Biczyskowa, and M. Walski. Electron microscope studies on the surface coat of the nephron. Journal of Cell Biology, 40(3):585–??, March 1969. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/40/3/585.

Gantt:1965:UPC

[GC65a] E. Gantt and S. F. Conti. The ultrastructure of porphyridium cruentum. Journal of Cell Biology, 26(2):365–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/26/2/365.

Gregory:1965:LSI

[GC65b] D. W. Gregory and E. C. Cocking. The large-scale isolation of protoplasts from immature tomato fruit. Journal of Cell Biology, 24(1):143–??, January 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/1/143.

Grimstone:1965:FSF

[GC65c] A. V. Grimstone and L. R. Cleveland. The fine structure and function of the contractile axostyles of certain flagellates. Journal of Cell Biology, 24(3):387–??, March 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/24/3/387.

Gantt:1966:GAC

[GC66a] E. Gantt and S. F. Conti. Granules associated with the chloroplast lamellae of porphyridium cruentum. Journal of Cell Biology, 29(3):423–??, June 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/29/3/423.

Greenawalt:1966:EMS

[GC66b] John W. Greenawalt and Ernesto Carafoli. Electron microscope studies on the active accumulation of Sr^{++} by rat-liver mitochondria. Journal of Cell Biology, 29(1):37–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/29/1/37.

Griffin:1966:SMH

[GC66c] Martin J. Griffin and Rody P. Cox. Studies on the mechanism of hormonal induction of alkaline phosphatase in human cell cultures. Journal of Cell Biology, 29(1):1–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/29/1/1.

Griffin:1967:CMI

[GCG67] M. J. Griffin, R. P. Cox, and N. Grujic. A chemical method for the isolation of HeLa cell nuclei and the nuclear localization of HeLa cell alkaline phosphatase. Journal of Cell Biology, 33(1):200–??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
1/200.

Guttes:1968:RDR

[GG68] Sophie Guttes and Edmund Guttes. Regulation of DNA replication in the nuclei of the slime mold Physarum Polycephalum. Journal of Cell Biology, 37(3):761–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/37/3/761.

Guttes:1969:RNA

[GG69] Edmund Guttes and Sophie Guttes. Replication of nucleolusassociated DNA during “G_{2} phase” in Physarum Polycephalum. Journal of Cell Biology, 43(2):229–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/43/2/229.

Grantham:1969:PTW

[GGBO69] Jared J. Grantham, Charles E. Ganote, Maurice B. Burg, and Jack Orloff. Paths of transtubular water flow in isolated renal collecting tubules. Journal of Cell Biology, 41(2):562–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/41/
2/562.

Gambetti:1968:FSP

[GGF68] Pierluigi Gambetti, Nicholas K. Gonatas, and Louis B. Flexner. The fine structure of puromycin-induced changes in mouse entorhinal cortex. Journal of Cell Biology, 36(2):379–??, February 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
2/379.

Gabrusewycz-Garcia:1966:SNM

[GGK66] Natalia Gabrusewycz-Garcia and Ruth G. Kleinfeld. A study of the nucleolar material in sciara coprophila. Journal of Cell Biology, 29(2):347–??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/29/2/347.

Ganote:1968:USV

[GGM^{+}68] Charles E. Ganote, Jared J. Grantham, Harold L. Moses, Maurice B. Burg, and Jack Orloff. Ultrastructural studies of Vasopressin effect on isolated perfused renal collecting tubules of the rabbit. Journal of Cell Biology, 36(2):355–??, February 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/36/2/355.

Greenspan:1965:ILP

[GH65] Francis S. Greenspan and Judy R. Hargadine. The intracellular localization of pituitary thyrotropic hormone. Journal of Cell Biology, 26(1):177–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/26/1/177.

Greider:1969:IPS

[GHL69] Marie H. Greider, S. L. Howell, and P. E. Lacy. Isolation and properties of secretory granules from rat islets of langerhans. Journal of Cell Biology, 41(1):162–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/41/1/162.

Gibbons:1965:RGC

[Gib65a] I. B. Gibbons. Reactivation of glycerinated cilia from Tetrahymena Pyriformis. Journal of Cell Biology, 25(2):400–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/25/2/
400.

Gibbons:1965:EAT

[Gib65b] I. R. Gibbons. An effect of adenosine triphosphate on the light scattered by suspensions of cilia. Journal of Cell Biology, 26 (3):707–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/3/707.

Gill:1965:IMI

[Gil65] D. M. Gill. An improved method for the isolation of rat liver nuclei by density centrifugation. Journal of Cell Biology, 24 (1):157–??, January 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/1/157.

Girbardt:1965:LEE

[Gir65] Manfred Girbardt. Lebendnachweis Von Einzelelementen Des Endoplasmatischen Retikulums. (German) [Proof of single elements of the endoplasmic reticulum]. Journal of Cell Biology, 27 (2):433–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/2/433.

Gibb:1968:IDC

[GK68] L. Gibb and E. R. M. Kay. Incorporation of DNA by cells of the Ehrlich-lettrÉ ascites carcinoma. Journal of Cell Biology, 38(2):452–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/2/452.

Goyer:1969:UTM

[GK69] R. A. Goyer and R. Krall. Ultrastructural transformation in mitochondria isolated from kidneys of normal and lead-intoxicated rats. Journal of Cell Biology, 41(2):393–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/41/2/393.

Graham:1967:MMO

[GKS^{+}67] R. C. Graham, M. J. Karnovsky, A. W. Shafer, E. A. Glass, and Manfred L. Karnovsky. Metabolic and morphological observations on the effect of surface-active agents on leukocytes. Journal of Cell Biology, 32(3):629–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/32/3/629.

Granick:1965:ILH

[GL65] S. Granick and Richard D. Levere. The intracellular localization of Heme by a fluorescence technique. Journal of Cell Biology, 26(1):167–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/1/167.

Ginsburg:1967:VDM

[GL67] H. Ginsburg and D. Lagunoff. The in vitro differentiation of mast cells. Journal of Cell Biology, 35(3):685–??, December 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/35/3/685.

Ginsburg:1968:ATR

[GL68] Haim Ginsburg and David Lagunoff. Aggregation and transformation of rat lymphocytes on rat embryo monolayers. Journal of Cell Biology, 39(2):392–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/2/392.

Gordon:1965:SID

[GMB65] Gerald B. Gordon, Leonard R. Miller, and Klaus G. Bensch. Studies on the intracellular digestive process in mammalian tissue culture cells. Journal of Cell Biology, 25(2):41–??, May 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/25/2/41.

Gimenez-Martin:1965:NML

[GMGFLS65] G. Giménez-Martín, A. González-Fernández, and J. F. López-Sáez. A new method of labeling cells. Journal of Cell Biology, 26(1):305–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/1/305.

Goldfischer:1967:CLM

[Gol67] Sidney Goldfischer. The cytochemical localization of myoglobin in striated muscle of man and walrus. Journal of Cell Biology, 34(1):398–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/34/1/398.

Goldberg:1968:PSD

[Gol68a] Alfred L. Goldberg. Protein synthesis during work-induced growth of skeletal muscle. Journal of Cell Biology, 36(3):653–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
3/653.

Goldhor:1968:PLR

[Gol68b] S. Goldhor. Protein: Lipid ratios of liver mitochondria during development. Journal of Cell Biology, 37(3):823–??, June 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/37/3/823.

Gould:1968:CFM

[Gou68a] Meredith C. Gould. Completion of the first meiotic division after fertilization in urechis caupo eggs despite inhibition of protein synthesis. Journal of Cell Biology, 38(1):220–??, July 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/38/1/220.

Gouranton:1968:CSM

[Gou68b] Jean Gouranton. Composition, structure, et mode De formation des ConcrÉTions MinÉRales dans l’intestin moyen des homopteres cercopides. Journal of Cell Biology, 37(2):316–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/37/
2/316.

Gauthier:1965:CSA

[GP65] Geraldine F. Gauthier and Helen A. Padykula. Cytochemical studies of adenosine triphosphatase activity in the sarcoplasmic reticulum. Journal of Cell Biology, 27(1):252–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/27/1/252.

Gauthier:1966:CSF

[GP66] Geraldine F. Gauthier and Helen A. Padykula. Cytological studies of fiber types in skeletal muscle. Journal of Cell Biology, 28 (2):333–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/28/2/333.

Goldstein:1967:PNI

[GP67] Lester Goldstein and David M. Prescott. Proteins in nucleocytoplasmic interactions. Journal of Cell Biology, 33(3):637–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
3/637.

Goldstein:1968:PNI

[GP68] Lester Goldstein and David M. Prescott. Proteins in nucleocytoplasmic interactions. Journal of Cell Biology, 36(1):53–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/36/1/53.

Gunning:1968:STC

[GPB68] B. E. S. Gunning, J. S. Pate, and L. G. Briarty. Specialized “transfer cells” in minor veins of leaves and their possible significance in phloem translocation. Journal of Cell Biology, 37(3):??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/37/3/C7.

Green:1965:PCM

[Gre65] Paul B. Green. Pathways of cellular morphogenesis. Journal of Cell Biology, 27(2):343–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/27/2/343.

Greenlee:1966:FSE

[GRH66] Theodore K. Greenlee, Russell Ross, and Jerry L. Hartman. The fine structure of elastic fibers. Journal of Cell Biology, 30(1):59– ??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/30/
1/59.

Gifford:1967:USA

[GS67a] Ernest M. Gifford and Kenneth D. Stewart. Ultrastructure of the shoot apex of chenopodium album and certain other seed plants. Journal of Cell Biology, 33(1):131–??, April 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/33/1/131.

Glick:1967:DIP

[GS67b] J. Leslie Glick and Anthony P. Salim. DNA-induced pigment production in a hamster cell line. Journal of Cell Biology, 33(1):209– ??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
1/209.

Griffin:1969:LMI

[GSS69] Joe L. Griffin, Marvin N. Stein, and Robert E. Stowell. Laser microscope irradiation of Physarum Polycephalum: Dynamic and ultrastructural effects. Journal of Cell Biology, 40(1):108–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
1/108.

Gibbs:1965:PAR

[GSW65] Sarah P. Gibbs, W. R. Sistrom, and Patricia B. Worden. The photosynthetic apparatus of rhodospirillum molischianum. Journal of Cell Biology, 26(2):395–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/26/2/395.

Gustafsson:1965:RBS

[GTLE65] R. Gustafsson, J. R. Tata, O. Lindberg, and L. Ernster. The relationship between the structure and activity of rat skeletal muscle mitochondria after thyroidectomy and thyroid hormone treatment. Journal of Cell Biology, 26(2):555–??, August 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/26/2/555.

Gibbins:1969:MFD

[GTP69] John R. Gibbins, Lewis G. Tilney, and Keith R. Porter. Microtubules in the formation and development of the primary mesenchyme in Arbacia Punctulata. Journal of Cell Biology, 41(1):201–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/1/201.

Gunning:1965:FSC

[Gun65] B. E. S. Gunning. The fine structure of chloroplast stroma following aldehyde osmium-tetroxide fixation. Journal of Cell Biology, 24(1):79–??, January 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/1/79.

Grasso:1966:RBR

[GW66] Joseph A. Grasso and John W. Woodard. The relationship between RNA synthesis and hemoglobin synthesis in amphibian erythropoiesis. Journal of Cell Biology, 31(2):279–??, November 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/31/
2/279.

Gorovsky:1967:HCC

[GW67a] Martin A. Gorovsky and John Woodard. Histone content of chromosomal loci active and inactive in RNA synthesis. Journal of Cell Biology, 33(3):723–??, June 1967. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/33/3/723.

Grasso:1967:DSM

[GW67b] Joseph A. Grasso and John W. Woodard. DNA synthesis and mitosis in erythropoietic cells. Journal of Cell Biology, 33(3):645– ??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
3/645.

Gorovsky:1969:SNS

[GW69] Martin A. Gorovsky and John Woodard. Studies on nuclear structure and function in Tetrahymena Pyriformis. Journal of Cell Biology, 42(3):673–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/42/3/673.

Giacomelli:1965:CAR

[GWS65] Filiberto Giacomelli, Joseph Wiener, and David Spiro. Cytological alterations related to stimulation of the Zona Glomerulosa of the adrenal gland. Journal of Cell Biology, 26(2):499–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/26/
2/499.

Garnjobst:1965:SCC

[GWT65] Laura Garnjobst, J. F. Wilson, and E. L. Tatum. Studies on a cytoplasmic character in neurospora crassa. Journal of Cell Biology, 26(2):413–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/2/413.

Huberman:1966:IMC

[HA66] Joel A. Huberman and Giuseppe Attardi. Isolation of metaphase chromosomes from HeLa cells. Journal of Cell Biology, 31 (1):95–??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/31/1/95.

Hertig:1967:SHO

[HA67] Arthur T. Hertig and Eleanor C. Adams. Studies on the human oocyte and its follicle. Journal of Cell Biology, 34(2):647–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
2/647.

Hancock:1968:NBE

[HA68] Ronald Hancock and Harold Amos. Nuclear binding of exogenous histones by L cells at low pH. Journal of Cell Biology, 36(1):??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
1/C1.

Hill:1969:EAM

[HA69] George C. Hill and Winston A. Anderson. Effects of acriflavine on the mitochondria and kinetoplast of crithidia fasciculata. Journal of Cell Biology, 41(2):547–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/41/2/547.

Hackenbrock:1966:UBM

[Hac66] Charles R. Hackenbrock. Ultrastructural bases for metabolically linked mechanical activity in mitochondria. Journal of Cell Biology, 30(2):269–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/30/2/269.

Hackenbrock:1968:UBM

[Hac68] Charles R. Hackenbrock. Ultrastructural bases for metabolically linked mechanical activity in mitochondria. Journal of Cell Biology, 37(2):345–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/37/2/345.

Hagopian:1966:MAF

[Hag66] Martin Hagopian. The myofilament arrangement in the femoral muscle of the cockroach, leucophaea maderae fabricius. Journal of Cell Biology, 28(3):545–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/28/3/545.

Hsu:1965:NMD

[HAKB65] T. C. Hsu, Frances E. Arrighi, Robert R. Klevecz, and B. R. Brinkley. The nucleoli in mitotic divisions of mammalian cells in vitro. Journal of Cell Biology, 26(2):539–??, August 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/26/2/539.

Hama:1965:SOF

[Ham65] Kiyoshi Hama. Some observations on the fine structure of the lateral line organ of the Japanese sea eel lyncozymba nystromi. Journal of Cell Biology, 24(2):193–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/24/2/193.

Hama:1966:FSS

[Ham66] Kiyoshi Hama. The fine structure of the Schwann cell sheath of the nerve fiber in the shrimp (Penaeus Japonicus). Journal of Cell Biology, 31(3):624–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/31/3/624.

Harris:1965:SOC

[Har65] Patricia Harris. Some observations concerning metakinesis in sea urchin eggs. Journal of Cell Biology, 25(1):73–??, April 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/25/1/73.

Hays:1966:IMT

[HB66] Richard M. Hays and Peter Barland. The isolation of the membrane of the toad bladder epithelial cell. Journal of Cell Biology, 31(1):209–??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/31/1/209.

Hugon:1967:FSL

[HB67] J. Hugon and M. Borgers. Fine structural localization of lysosomal enzymes in the absorbing cells of the duodenal mucosa of the mouse. Journal of Cell Biology, 33(1):212–??, April 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/33/1/212.

Hendelman:1969:RSC

[HB69a] Walter J. Hendelman and Richard P. Bunge. Radioautographic studies of choline incorporation into peripheral nerve myelin. Journal of Cell Biology, 40(1):190–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/40/1/190.

Humphrey:1969:USR

[HB69b] Ronald M. Humphrey and B. R. Brinkley. Ultrastructural studies of radiation-induced chromosome damage. Journal of Cell Biology, 42(3):745–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/42/3/745.

Hirschhorn:1968:SL

[HBHW68] Rochelle Hirschhorn, Günter Brittinger, Kurt Hirschhorn, and Gerald Weissmann. Studies on lysosomes. Journal of Cell Biology, 37(2):412–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/37/2/412.

Hotta:1965:NDC

[HBS65] Yasuo Hotta, Alix Bassel, and Herbert Stern. Nuclear DNA and cytoplasmic DNA from tissues of higher plants. Journal of Cell Biology, 27(3):451–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/27/3/451.

Hendler:1969:RPS

[HBS69] Richard W. Hendler, Amelia H. Burgess, and Raymond Scharff. Respiration and protein synthesis in Escherichia coli membraneenvelope fragments. Journal of Cell Biology, 42(3):715– ??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/3/715.

Horvat:1969:ILE

[HBT69] Agnes Horvat, Jane Baxandall, and Oscar Touster. The isolation of lysosomes from Ehrlich ascites tumor cells following pretreatment of mice with triton wr-1339. Journal of Cell Biology, 42(2):469–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/2/469.

Holmes:1968:RMM

[HC68] Kathryn V. Holmes and Purnell W. Choppin. On the role of microtubules in movement and alignment of nuclei in virus-induced syncytia. Journal of Cell Biology, 39(3):526–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/39/3/526.

Hackenbrock:1969:IIU

[HC69a] Charles R. Hackenbrock and Arnold I. Caplan. Ion-induced ultrastructural transformations in isolated mitochondria. Journal of Cell Biology, 42(1):221–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/42/1/221.

Hinsch:1969:UOV

[HC69b] Gertrude W. Hinsch and M. Virginia Cone. Ultrastructural observations of vitellogenesis in the spider crab, Libinia Emarginata L. Journal of Cell Biology, 40(2):336–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/40/2/336.

Howard:1967:EPI

[HCGP67] Roger B. Howard, A. Kent Christensen, Frederic A. Gibbs, and Leroy A. Pesch. The enzymatic preparation of isolated intact parenchymal cells from rat liver. Journal of Cell Biology, 35 (3):675–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/3/675.

Heuser:1966:FSO

[HD66] John E. Heuser and Carlos F. Doggenweiler. The fine structural organization of nerve fibers, sheaths, and glial cells in the prawn, palaemonetes vulgaris. Journal of Cell Biology, 30(2):381–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/30/
2/381.

Hirano:1967:SAM

[HD67] Asao Hirano and Herbert M. Dembitzer. A structural analysis of the myelin sheath in the central nervous system. Journal of Cell Biology, 34(2):555–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/2/555.

Horwitz:1968:SSS

[HD68] Allen L. Horwitz and Albert Dorfman. Subcelluar sites for synthesis of chondromucoprotein of cartilage. Journal of Cell Biology, 38(2):358–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/38/2/358.

Hardin:1967:SSH

[HEL67] John A. Hardin, Gerald E. Einem, and David T. Lindsay. Simultaneous synthesis of histone and DNA in synchronously dividing Tetrahymena Pyriformis. Journal of Cell Biology, 32(3):709–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/32/
3/709.

Hess:1965:DCS

[Hes65a] Arthur Hess. Developmental changes in the structure of the synapse on the myelinated cell bodies of the chicken ciliary ganglion. Journal of Cell Biology, 25(3):1–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/25/3/1.

Hess:1965:SRS

[Hes65b] Arthur Hess. The sarcoplasmic reticulum, the T system, and the motor terminals of slow and twitch muscle fibers in the garter snake. Journal of Cell Biology, 26(2):467–??, August 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/26/2/467.

Hess:1965:EXR

[Hes65c] Oswald Hess. The effect of X-rays on the functional structures of the Y chromosome in spermatocytes of Drosophila Hydei. Journal of Cell Biology, 25(1):169–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/25/1/169.

Heyn:1966:MSC

[Hey66] A. N. J. Heyn. The microcrystalline structure of cellulose in cell walls of cotton, ramie, and jute fibers as revealed by negative staining of sections. Journal of Cell Biology, 29(2):181–??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/29/
2/181.

Hickman:1965:OSRa

[HF65a] Donald D. Hickman and Albert W. Frenkel. Observations on the structure of rhodospirillum molischianum. Journal of Cell Biology, 25(2):261–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/25/2/261.

Hickman:1965:OSRb

[HF65b] Donald D. Hickman and Albert W. Frenkel. Observations on the structure of rhodospirillum rubrum. Journal of Cell Biology, 25(2):279–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/2/279.

Hirsch:1968:UHL

[HF68] James G. Hirsch and Martha E. Fedorko. Ultrastructure of human leukocytes after simultaneous fixation with glutaraldehyde and osmium tetroxide and “postfixation” in uranyl acetate. Journal of Cell Biology, 38(3):615–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/38/3/615.

Hirsch:1968:VFF

[HFC68] James G. Hirsch, Martha E. Fedorko, and Zanvil A. Cohn. Vesicle fusion and formation at the surface of pinocytic vacuoles in macrophages. Journal of Cell Biology, 38(3):629–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
3/629.

Howell:1969:IPSa

[HFL69] S. L. Howell, C. J. Fink, and P. E. Lacy. Isolation and properties of secretory granules from rat islets of langerhans. Journal of Cell Biology, 41(1):154–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/41/1/154.

Holt:1969:MCD

[HG69] Charles E. Holt and Elizabeth G. Gurney. Minor components of the DNA of Physarum Polycephalum. Journal of Cell Biology, 40 (2):484–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/2/484.

Holt:1966:IOP

[HH66] S. J. Holt and R. Marian Hicks. The importance of osmiophilia in the production of stable azoindoxyl complexes of high contrast for combined enzyme cytochemistry and electron microscopy. Journal of Cell Biology, 29(2):361–??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/29/2/361.

Hokin:1967:RLI

[HH67] Lowell E. Hokin and Dorothy Huebner. Radioautographic localization of the increased synthesis of phosphatidylinositol in response to pancreozymin or acetylcholine in guinea pig pancreas slices. Journal of Cell Biology, 33(3):521–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/33/3/521.

Holmes:1969:CRP

[HHM69] Richard Holmes, Judy Helms, and Gretchen Mercer. Cholesterol requirement of primary diploid human fibroblasts. Journal of Cell Biology, 42(1):262–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/42/1/262.

Hepler:1966:FSE

[HHS66] Peter K. Hepler, Clay G. Huff, and Helmuth Sprinz. The fine structure of the exoerythrocytic stages of plasmodium fallax. Journal of Cell Biology, 30(2):333–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/30/2/333.

Hicks:1965:FST

[Hic65] R. M. Hicks. The fine structure of the transitional epithelium of rat ureter. Journal of Cell Biology, 26(1):25–??, July 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/26/1/25.

Hicks:1966:FGC

[Hic66a] R. M. Hicks. The function of the Golgi complex in transitional epithelium. Journal of Cell Biology, 30(3):623–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/30/3/623.

Hicks:1966:PRT

[Hic66b] R. M. Hicks. The permeability of rat transitional epithelium. Journal of Cell Biology, 28(1):21–??, January 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/28/1/21.

Himes:1967:AHM

[Him67a] Marion Himes. An analysis of heterochromatin in maize root tips. Journal of Cell Biology, 35(1):175–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/35/1/175.

Himes:1967:DPB

[Him67b] Marion Himes. DNA-protein binding in interphase chromosomes. Journal of Cell Biology, 34(1):77–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/34/1/77.

Hiramoto:1965:FSC

[Hir65] Y. Hiramoto. Further studies on cell division without mitotic apparatus in sea urchin eggs. Journal of Cell Biology, 25(1):161– ??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
1/161.

Hirano:1968:COO

[Hir68] Asao Hirano. A confirmation of the oligodendroglial origin of myelin in the adult rat. Journal of Cell Biology, 38(3): 637–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/3/637.

Hepler:1968:MES

[HJ68] Peter K. Hepler and William T. Jackson. Microtubules and early stages of cell-plate formation in the endosperm of haemanthus katherinae Baker. Journal of Cell Biology, 38(2):437–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
2/437.

Herdson:1965:EMS

[HK65] P. B. Herdson and J. P. Kaltenbach. Electron microscope studies on enzyme activity and the isolation of thiohydantoin-induced myelin figures in rat liver. Journal of Cell Biology, 25(3):485–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
3/485.

Howell:1969:BGF

[HKL69] S. L. Howell, M. Kostianovsky, and P. E. Lacy. Beta granule formation in isolated islets of langerhans. Journal of Cell Biology, 42(3):695–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/42/3/695.

Harris:1966:SEE

[HL66] Curtis C. Harris and Charles A. Leone. Some effects of EDTA and tetraphenylboron on the ultrastructure of mitochondria in mouse liver cells. Journal of Cell Biology, 28(2):405–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/28/
2/405.

Hodson:1967:TIR

[HM67] Stuart Hodson and J. Marshall. Tyrosine incorporation into the rabbit retina. Journal of Cell Biology, 35(3):722–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/35/
3/722.

Hoyle:1965:MSS

[HMS65] Graham Hoyle, James H. McAlear, and Allen Selverston. Mechanism of supercontraction in a striated muscle. Journal of Cell Biology, 26(2):621–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/26/2/621.

Hoyle:1966:NIT

[HMW66] Graham Hoyle, Patricia Ann McNeill, and Benjamin Walcott. Nature of invaginating tubules in felderstruktur muscle fibers of the garter snake. Journal of Cell Biology, 30(1):197–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/30/
1/197.

Holtzman:1965:LRS

[HN65] Eric Holtzman and Alex B. Novikoff. Lysosomes in the rat sciatic nerve following crush. Journal of Cell Biology, 27(3):651–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/27/
3/651.

Heneen:1966:PNS

[HN66] Waheeb K. Heneen and Warren W. Nichols. Persistence of nucleoli in short term and long term cell cultures and in direct bone marrow preparations in mammalian materials. Journal of Cell Biology, 31(3):543–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/31/3/543.

Holtzman:1967:LGN

[HNV67] Eric Holtzman, Alex B. Novikoff, and Humberto Villaverde. Lysosomes and gerl in normal and chromatolytic neurons of the rat ganglion nodosum. Journal of Cell Biology, 33(2):419–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
2/419.

Hof:1965:CPK

[Hof65a] Jack Van’t Hof. Cell population kinetics of excised roots of pisum sativum. Journal of Cell Biology, 27(1):179–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/27/1/179.

Hoffman:1965:NAB

[Hof65b] Eugene J. Hoffman. The nucleic acids of basal bodies isolated from Tetrahymena Pyriformis. Journal of Cell Biology, 25(2):217–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/2/217.

Hof:1968:CCP

[Hof68] Jack Van’t Hof. Control of cell progression through the mitotic cycle by carbohydrate provision. Journal of Cell Biology, 37(3):773–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/37/3/773.

Holmes:1967:PHS

[Hol67] Richard Holmes. Preparation from human serum of an alpha-one protein which induces the immediate growth of unadapted cells in vitro. Journal of Cell Biology, 32(2):297–??, February 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/32/2/297.

Holdsworth:1968:PCM

[Hol68] Robert H. Holdsworth. The presence of a crystalline matrix in pyrenoids of the diatom, achnanthes brevipes. Journal of Cell Biology, 37(3):831–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/37/3/831.

Hando:1968:ACH

[HOZ68] T. Hando, D. M. Okada, and L. Zamboni. Atypical cilia in human endometrium. Journal of Cell Biology, 39(2):475–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/39/
2/475.

Healy:1966:CMC

[HP66a] George M. Healy and Raymond C. Parker. Cultivation of mammalian cells in defined media with protein and nonprotein supplements. Journal of Cell Biology, 30(3):539–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/30/3/539.

Healy:1966:ICD

[HP66b] George M. Healy and Raymond C. Parker. An improved chemically defined basal medium (cmrl-1415) for newly explanted mouse embryo cells. Journal of Cell Biology, 30(3): 531–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/30/3/531.

Howard:1968:CDS

[HP68] E. F. Howard and W. Plaut. Chromosomal DNA synthesis in Drosophila Melanogaster. Journal of Cell Biology, 39(2): 415–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/39/2/415.

Holtzman:1969:UPM

[HP69] Eric Holtzman and Edith R. Peterson. Uptake of protein by mammalian neurons. Journal of Cell Biology, 40(3):863–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
3/863.

Hinojosa:1967:UST

[HR67] Raúl Hinojosa and J. David Robertson. Ultrastructure of the spoon type synaptic endings in the nucleus vestibularis tangentialis of the chick. Journal of Cell Biology, 34(2):421–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
2/421.

Hodge:1969:PMR

[HRS69] L. D. Hodge, E. Robbins, and M. D. Scharff. Persistence of messenger RNA through mitosis in HeLa cells. Journal of Cell Biology, 40(2):497–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/40/2/497.

Hof:1965:REG

[HS65a] Jack Van’t Hof and A. H. Sparrow. Radiation effects on the growth rate and cell population kinetics of actively growing and dormant roots of Tradescantia Paludosa. Journal of Cell Biology, 26(1):187–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/1/187.

Hotta:1965:ITK

[HS65b] Yasuo Hotta and Herbert Stern. Inducibility of thymidine kinase by thymidine as a function of interphase stage. Journal of Cell Biology, 25(3):99–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/25/3/99.

Hohman:1966:IDC

[HS66] Wayne Hohman and Harald Schraer. The intracellular distribution of calcium in the mucosa of the avian Shell gland. Journal of Cell Biology, 30(2):317–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/30/2/317.

Hagopian:1967:SRA

[HS67a] Martin Hagopian and David Spiro. The sarcoplasmic reticulum and its association with the T system in an insect. Journal of Cell Biology, 32(3):535–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/32/3/535.

Hug:1967:LTI

[HS67b] George Hug and William K. Schubert. Lysosomes in type ii glycogenosis. Journal of Cell Biology, 35(1):??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/35/1/C1.

Hagopian:1968:FLC

[HS68] Martin Hagopian and David Spiro. The filament lattice of cockroach thoracic muscle. Journal of Cell Biology, 36(3):433–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
3/433.

Hayashi:1968:CCE

[HSC68] Masando Hayashi, Tsuranobu Shirahama, and Alan S. Cohen. Combined cytochemical and electron microscopic demonstration of β-glucuronidase activity in rat liver with the use of a simultaneous coupling azo dye technique. Journal of Cell Biology, 36 (2):289–??, February 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/36/2/289.

Hays:1965:ECW

[HSM65] Richard M. Hays, Bayla Singer, and Sasha Malamed. The effect of calcium withdrawal on the structure and function of the toad bladder. Journal of Cell Biology, 25(3):195–??, June 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/25/3/195.

Harper:1967:EMB

[HSS67] Elvin Harper, Sam Seifter, and Berta Scharrer. Electron microscopic and biochemical characterization of collagen in blattarian insects. Journal of Cell Biology, 33(2):385–??, May 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/33/2/385.

Haydon:1965:MHP

[HT65] Glen B. Haydon and D. Alice Taylor. Microtubules in hamster platelets. Journal of Cell Biology, 26(2):673–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/26/2/673.

Hayek:1966:RAM

[HT66] Dean H. Hayek and Samuel R. Tipton. Respiratory activity and maintenance of cell suspensions of rat liver. Journal of Cell Biology, 29(3):405–??, June 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/29/3/405.

Hadjiolov:1965:ISC

[HTBM65] A. A. Hadjiolov, Z. S. Tencheva, and A. G. Bojadjieva-Mikhailova. Isolation and some characteristics of cell nuclei from brain cortex of adult cat. Journal of Cell Biology, 26(2):383–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/26/
2/383.

Hufnagel:1969:CUP

[Huf69] Linda A. Hufnagel. Cortical ultrastructure of paramecium aurelia. Journal of Cell Biology, 40(3):779–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/40/3/779.

Hoeniger:1966:BBB

[HvIvZ66] Judith F. M. Hoeniger, Woutera van Iterson, and Eva Nijman van Zanten. Basal bodies of bacterial flagella in Proteus mirabilis. Journal of Cell Biology, 31(3):603–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/31/3/603.

Hawley:1967:SMD

[HW67] Emerson S. Hawley and R. P. Wagner. Synchronous mitochondrial division in neurospora crassa. Journal of Cell Biology, 35 (3):489–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/3/489.

Howell:1969:IPSb

[HYL69] S. L. Howell, D. A. Young, and P. E. Lacy. Isolation and properties of secretory granules from rat islets of langerhans. Journal of Cell Biology, 41(1):167–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/41/1/167.

Hirano:1966:MCN

[HZL66] Asao Hirano, H. M. Zimmerman, and Seymour Levine. Myelin in the central nervous system as observed in experimentally induced edema in the rat. Journal of Cell Biology, 31(3):397–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/31/
3/397.

Ishikawa:1968:MIS

[IBH68] H. Ishikawa, R. Bischoff, and H. Holtzer. Mitosis and intermediate-sized filaments in developing skeletal muscle. Journal of Cell Biology, 38(3):538–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/38/3/538.

Ishikawa:1969:FAC

[IBH69] Harunori Ishikawa, Richard Bischoff, and Howard Holtzer. Formation of arrowhead complexes with heavy meromyosin in a variety of cell types. Journal of Cell Biology, 43(2):312–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/43/
2/312.

Ichikawa:1965:FSC

[Ich65] Atsushi Ichikawa. Fine structural changes in response to hormonal stimulation of the perfused canine pancreas. Journal of Cell Biology, 24(3):369–??, March 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/24/3/369.

Ikemoto:1968:EMI

[IKNG68] N. Ikemoto, S. Kitagawa, A. Nakamura, and J. Gergely. Electron microscopic investigations of actomyosin as a function of ionic strength. Journal of Cell Biology, 39(3):620–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/39/
3/620.

Ishikawa:1965:IGP

[IP65] Toyoko Ishikawa and Yen Fen Pei. Intramitochondrial glycogen particles in rat retinal receptor cells. Journal of Cell Biology, 25(2):402–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/2/402.

Ito:1969:PMS

[IS69] Akio Ito and Ryo Sato. Proteolytic microdissection of smoothsurfaced vesicles of liver microsomes. Journal of Cell Biology, 40(1):179–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/1/179.

Ishikawa:1968:FEN

[Ish68] Harunori Ishikawa. Formation of elaborate networks of t-system tubules in cultured skeletal muscle with special reference to the t-system formation. Journal of Cell Biology, 38(1):51–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/38/1/51.

Israel:1968:IRP

[ISS68] H. W. Israel, M. M. Salpeter, and F. C. Steward. The incorporation of radioactive proline into cultured cells. Journal of Cell Biology, 39(3):698–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/3/698.

Ito:1965:ESC

[Ito65] Susumu Ito. The enteric surface coat on cat intestinal microvilli. Journal of Cell Biology, 27(3):475–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/27/3/475.

Jacobson:1968:PIP

[Jac68] Ann B. Jacobson. A procedure for isolation of proplastids from etiolated maize leaves. Journal of Cell Biology, 38(1):238–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
1/238.

Jagiello:1969:MIO

[Jag69] Georgiana M. Jagiello. Meiosis and inhibition of ovulation in mouse eggs treated with actinomycin D. Journal of Cell Biology, 42(2):571–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/2/571.

Jamuar:1966:SSN

[Jam66] Mahendra P. Jamuar. Studies of spermiogenesis in a nematode, nippostrongylus brasiliensis. Journal of Cell Biology, 31 (3):381–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/31/3/381.

Jasper:1967:BML

[Jas67] Donald Jasper. Body muscles of the lamprey. Journal of Cell Biology, 32(1):219–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/32/1/219.

Jasper:1968:SPS

[JB68] D. K. Jasper and J. R. Bronk. Studies on the physiological and structural characteristics of rat intestinal mucosa. Journal of Cell Biology, 38(2):277–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/38/2/277.

Jagendorf-Elfvin:1969:RCU

[JEE69] Myra Jagendorf-Elfvin and Eva Eliasson. Reversible changes in the ultrastructure of Chang liver cell mitochondria following incubation of the cells in a glutamine-deficient medium. Journal of Cell Biology, 41(3):905–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/41/3/905.

Jenkins:1967:FSD

[Jen67] R. A. Jenkins. Fine structure of division in ciliate protozoa. Journal of Cell Biology, 34(2):463–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/34/2/463.

Jensen:1968:CE

[Jen68] William A. Jensen. Cotton embryogenesis. Journal of Cell Biology, 36(2):403–??, February 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/36/2/403.

Jersild:1966:RSG

[Jer66] Ralph A. Jersild. A radioautographic study of glyceride synthesis in vivo during intestinal absorption of fats and labeled glucose. Journal of Cell Biology, 31(3):413–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/31/3/413.

Jaim-Etcheverry:1969:SDT

[JEZ69] Guillermo Jaim-Etcheverry and Luis Maria Zieher. Selective demonstration of a type of synaptic vesicle by phosphotungstic acid staining. Journal of Cell Biology, 42(3):855–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
3/855.

Johnson:1965:RAP

[JH65] Terry C. Johnson and John J. Holland. Ribonucleic acid and protein synthesis in mitotic HeLa cells. Journal of Cell Biology, 27 (3):565–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/3/565.

Jollow:1968:BM

[JKL68] D. Jollow, G. M. Kellerman, and Anthony W. Linnane. The biogenesis of mitochondria. Journal of Cell Biology, 37(2):221– ??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/37/
2/221.

Jamakosmanovic:1968:ICT

[JL68] A. Jamakosmanović and W. R. Loewenstein. Intercellular communication and tissue growth. Journal of Cell Biology, 38(3): 556–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/3/556.

Johnson:1968:IPS

[JLL68] Terry C. Johnson, Michael P. Lerner, and Gerald J. Lancz. Inhibition of protein synthesis in noninfected L cells by partially purified Interferon preparations. Journal of Cell Biology, 36(3):617– ??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
3/617.

Johnson:1969:SNV

[Joh69] J. Morris Johnson. A study of nucleolar vacuoles in cultured tobacco cells using radioautography, actinomycin D, and electron microscopy. Journal of Cell Biology, 43(2):197–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/43/
2/197.

Jamieson:1967:ITSa

[JP67a] James D. Jamieson and George E. Palade. Intracellular transport of secretory proteins in the pancreatic exocrine cell. Journal of Cell Biology, 34(2):577–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/2/577.

Jamieson:1967:ITSb

[JP67b] James D. Jamieson and George E. Palade. Intracellular transport of secretory proteins in the pancreatic exocrine cell. Journal of Cell Biology, 34(2):597–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/2/597.

Jamieson:1968:ITSa

[JP68a] James D. Jamieson and George E. Palade. Intracellular transport of secretory proteins in the pancreatic exocrine cell. Journal of Cell Biology, 39(3):580–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/3/580.

Jamieson:1968:ITSb

[JP68b] James D. Jamieson and George E. Palade. Intracellular transport of secretory proteins in the pancreatic exocrine cell. Journal of Cell Biology, 39(3):589–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/3/589.

Johnson:1968:FSC

[JP68c] Ursula G. Johnson and Keith R. Porter. Fine structure of cell division in chlamydomonas reinhardi. Journal of Cell Biology, 38(2):403–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/2/403.

Johnson:1967:SCM

[JS67] Edward A. Johnson and Joachim R. Sommer. A Strand of cardiac muscle. Journal of Cell Biology, 33(1):103–??, April 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/33/1/103.

Jepsen:1967:EMI

[JT67] Arne Jepsen and Jørgen Theilade. Electron microscopic identification of epithelial cells in tissue culture of oral mucous membrane. Journal of Cell Biology, 32(3):770–??, March 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/32/3/770.

Kornguth:1965:LBP

[KA65] Steven Edward Kornguth and John Walberg Anderson. Localization of a basic protein in the myelin of various species with the aid of fluorescence and electron microscopy. Journal of Cell Biology, 26(1):157–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/1/157.

Kahri:1968:EAD

[Kah68] A. I. Kahri. Effects of actinomycin D and puromycin on the acthinduced ultrastructural transformation of mitochondria of cortical cells of rat adrenals in tissue culture. Journal of Cell Biology, 36(1):181–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/36/1/181.

Kane:1965:MAa

[Kan65] R. E. Kane. The mitotic apparatus. Journal of Cell Biology, 25(1):137–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/1/137.

Kane:1967:MA

[Kan67] R. E. Kane. The mitotic apparatus. Journal of Cell Biology, 32 (2):243–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/32/2/243.

Karasaki:1965:EME

[Kar65a] Shuichi Karasaki. Electron microscopic examination of the sites of nuclear RNA synthesis during amphibian embryogenesis. Journal of Cell Biology, 26(3):937–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/26/3/937.

Karasaki:1965:ICW

[Kar65b] Shuichi Karasaki. Intranuclear crystal within the phagocytes of the ovary of Arbacia Punctulata. Journal of Cell Biology, 25(3):654–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/3/654.

Karlin:1965:AAM

[Kar65c] Arthur Karlin. The association of acetylcholinesterase and membrane in subcellular fractions of the electric tissue of electrophorus. Journal of Cell Biology, 25(2):159–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/25/2/159.

Karnovsky:1967:UBC

[Kar67] Morris J. Karnovsky. The ultrastructural basis of capillary permeability studied with peroxidase as a tracer. Journal of Cell Biology, 35(1):213–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/35/1/213.

Karasaki:1969:FSP

[Kar69] Shuichi Karasaki. The fine structure of proliferating cells in preneoplastic rat livers during azo-dye carcinogenesis. Journal of Cell Biology, 40(2):322–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/40/2/322.

Kessel:1965:UCG

[KB65a] R. G. Kessel and H. W. Beams. An unusual configuration of the Golgi complex in pigment-producing “test” cells of the ovary of the tunicate, styela. Journal of Cell Biology, 25(1):55–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/25/1/55.

Krishan:1965:SMS

[KB65b] Awtar Krishan and Robert C. Buck. Structure of the mitotic spindle in L strain fibroblasts. Journal of Cell Biology, 24(3):433– ??, March 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/24/
3/433.

King:1967:SRM

[KB67] Donald W. King and M. L. Barnhisel. Synthesis of RNA in mammalian cells during mitosis and interphase. Journal of Cell Biology, 33(2):265–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/33/2/265.

Kessel:1968:IMN

[KB68] R. G. Kessel and H. W. Beams. Intranucleolar membranes and nuclear-cytoplasmic exchange in Young crayfish oocytes. Journal of Cell Biology, 39(3):735–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/3/735.

Kessel:1969:ALY

[KB69] R. G. Kessel and H. W. Beams. Annulate lamellae and “yolk nuclei” in oocytes of the dragonfly, libellula pulchella. Journal of Cell Biology, 42(1):185–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/42/1/185.

Kellerman:1969:BM

[KBL69] G. M. Kellerman, D. R. Biggs, and Anthony W. Linnane. Biogenesis of mitochondria. Journal of Cell Biology, 42(2):377–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
2/377.

Knappeis:1968:UML

[KC68] Gustav G. Knappeis and Frits Carlsen. The ultrastructure of the M line in skeletal muscle. Journal of Cell Biology, 38(1):202–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
1/202.

Kraehenbuhl:1969:ESI

[KC69] J. P. Kraehenbuhl and M. A. Campiche. Early stages of intestinal absorption of specific antibodies in the newborn. Journal of Cell Biology, 42(2):345–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/42/2/345.

Kumegawa:1967:EMO

[KCR67a] M. Kumegawa, M. Cattoni, and G. G. Rose. Electron microscopy of oral cells in vitro. Journal of Cell Biology, 34(3): 897–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/34/3/897.

Kumegawa:1967:UDS

[KCR67b] M. Kumegawa, M. Cattoni, and George G. Rose. An unusual droplet in submandibular gland of new Born mice. Journal of Cell Biology, 33(3):720–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/33/3/720.

Kumegawa:1968:EMO

[KCR68] M. Kumegawa, M. Cattoni, and George G. Rose. Electron microscopy of oral cells in vitro. Journal of Cell Biology, 36(3):443– ??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
3/443.

Kistler:1967:DFS

[KCW67] Gonzague S. Kistler, Peter R. B. Caldwell, and Ewald R. Weibel. Development of fine structural damage to alveolar and capillary lining cells in oxygen-poisoned rat lungs. Journal of Cell Biology, 32(3):605–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/32/3/605.

Kelly:1966:FSD

[Kel66] Douglas E. Kelly. Fine structure of desmosomes, hemidesmosomes, and an adepidermal globular layer in developing newt epidermis. Journal of Cell Biology, 28(1):51–??, January 1966. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/28/1/51.

Kelly:1967:MMZ

[Kel67] Douglas E. Kelly. Models of muscle z-band fine structure based on a looping filament configuration. Journal of Cell Biology, 34 (3):827–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/34/3/827.

Kelley:1969:WPS

[Kel69] Carole Kelley. Wall projections in the sporophyte and gametophyte of sphaerocarpos. Journal of Cell Biology, 41(3):910–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/41/
3/910.

Kennedy:1966:ESL

[Ken66] John R. Kennedy. The effect of strychnine and light on pigmentation in blepharisma undulans Stein. Journal of Cell Biology, 28 (2):145–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/28/2/145.

Kessel:1965:ICA

[Kes65] R. G. Kessel. Intranuclear and cytoplasmic annulate lamellae in tunicate oocytes. Journal of Cell Biology, 24(3):471–??, March 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/24/
3/471.

Kessel:1968:FSA

[Kes68] R. G. Kessel. Fine structure of annulate lamellae. Journal of Cell Biology, 36(3):658–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/36/3/658.

Kallman:1967:AFB

[KEW67a] Frances Kallman, Jean Evans, and Norman K. Wessells. Anchor filament bundles in embryonic feather germs and skin. Journal of Cell Biology, 32(1):236–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/32/1/236.

Kallman:1967:NEB

[KEW67b] Frances Kallman, Jean Evans, and Norman K. Wessells. Normal epidermal basal cell behavior in the absence of basement membrane. Journal of Cell Biology, 32(1):231–??, January 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/32/1/231.

Kane:1965:MAb

[KF65] R. E. Kane and Arthur Forer. The mitotic apparatus. Journal of Cell Biology, 25(3):31–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/25/3/31.

Kohler:1969:EIC

[KGO69] Peter O. Kohler, Philip M. Grimley, and Bert W. O’malley. Estrogen-induced cytodifferentiation of the ovalbumin-secreting glands of the chick oviduct. Journal of Cell Biology, 40(1):8–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
1/8.

Krishan:1969:OAH

[KH69] Awtar Krishan and Dora Hsu. Observations on the association of helical polyribosomes and filaments with vincristine-induced crystals in Earle’s L-cell fibroblasts. Journal of Cell Biology, 43 (3):553–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/43/3/553.

Krishan:1968:HGE

[KHH68] Awtar Krishan, Dora Hsu, and Patricia Hutchins. Hypertrophy of granular endoplasmic reticulum and annulate lamellae in Earle’s L cells exposed to vinblastine sulfate. Journal of Cell Biology, 39(1):211–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/39/1/211.

Kuff:1966:VPS

[KHSR66] E. L. Kuff, W. C. Hymer, E. Shelton, and N. E. Roberts. The in vivo protein synthetic activities of free versus membrane-bound ribonucleoprotein in a plasma-cell tumor of the mouse. Journal of Cell Biology, 29(1):63–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/29/1/63.

Kemp:1967:CCG

[KI67] Norman E. Kemp and Nancy L. Istock. Cortical changes in growing oocytes and in fertilized or pricked eggs of rana pipiens. Journal of Cell Biology, 34(1):111–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/1/111.

Kirkpatrick:1969:SBM

[Kir69] Joel B. Kirkpatrick. Stability of brain microtubules in homogenates. Journal of Cell Biology, 42(2):600–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/42/2/600.

Kato:1967:ICN

[KK67] Takahiko Kato and Masanori Kurokawa. Isolation of cell nuclei from the mammalian cerebral cortex and their assortment on a morphological basis. Journal of Cell Biology, 32(3):649–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/32/
3/649.

Kanaseki:1969:VB

[KK69] Toku Kanaseki and Ken Kadota. The “vesicle in a basket”. Journal of Cell Biology, 42(1):202–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/42/1/202.

Kanczak:1965:LFT

[KKHE65] Norbert M. Kanczak, Joseph I. Krall, E. Russell Hayes, and Willard B. Elliott. Lysosomal fractions from transitional epithelium. Journal of Cell Biology, 24(2):259–??, February 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/24/2/259.

Kishimoto:1965:NMC

[KL65] Susumu Kishimoto and Irving Lieberman. Nuclear membranes of cultured mammalian cells in the period preceding DNA synthesis. Journal of Cell Biology, 25(1):103–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/25/1/103.

Klaus:1968:ECM

[Kla68] Sidney N. Klaus. The effect of colchicine on mosaic patterns in cultured cells. Journal of Cell Biology, 36(2):399–??, February 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
2/399.

Klevecz:1969:TOM

[Kle69] Robert R. Klevecz. Temporal order in mammalian cells. Journal of Cell Biology, 43(2):207–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/43/2/207.

Kimberg:1968:CIA

[KLW68] Daniel V. Kimberg, Alden V. Loud, and Joseph Wiener. Cortisone-induced alterations in mitochondrial function and structure. Journal of Cell Biology, 37(1):63–??, April 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/37/1/63.

Kruse:1965:PCM

[KM65] Paul F. Kruse and Ed Miedema. Production and characterization of multiple-layered populations of animal cells. Journal of Cell Biology, 27(2):273–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/27/2/273.

Kaye:1966:FSC

[KMK66] Jerome S. Kaye and Rachel McMaster-Kaye. The fine structure and chemical composition of nuclei during spermiogenesis in the house Cricket. Journal of Cell Biology, 31(1):159–??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/31/
1/159.

Kuhn:1965:UHS

[KO65] Nobuko O. Kuhn and Margaret L. Olivier. Ultrastructure of the hepatic sinusoid of the goat capra hircus. Journal of Cell Biology, 26(3):977–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/26/3/977.

Koenig:1967:OSI

[Koe67] Edward Koenig. Observations on selected isolated retinal elements and an analysis of rod cell RNA of the rabbit. Journal of Cell Biology, 34(1):265–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/1/265.

Koenig:1968:MDT

[Koe68] Edward Koenig. A method for determining total protein of isolated cellular elements and corresponding tritium radioactivity. Journal of Cell Biology, 38(3):562–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/38/3/562.

Koller:1965:MUS

[Kol65] Theodor Koller. Mounting of ultrathin sections with the aid of an electrostatic field. Journal of Cell Biology, 27(2):441–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/27/
2/441.

Korn:1967:CSS

[Kor67] Edward D. Korn. A chromatographic and spectrophotometric study of the products of the reaction of osmium tetroxide with unsaturated lipids. Journal of Cell Biology, 34(2):627–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
2/627.

Kessel:1968:OAP

[KP68] R. G. Kessel and W. R. Panje. Organization and activity in the pre- and postovulatory follicle of necturus maculosus. Journal of Cell Biology, 39(1):1–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/1/1.

Kelly:1968:LMF

[KR68] Robert E. Kelly and Robert V. Rice. Localization of myosin filaments in smooth muscle. Journal of Cell Biology, 37(1):105– ??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/37/
1/105.

Kelly:1969:USC

[KR69a] Robert E. Kelly and Robert V. Rice. Ultrastructural studies on the contractile mechanism of smooth muscle. Journal of Cell Biology, 42(3):683–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/42/3/683.

Kubai:1969:DDG

[KR69b] Donna F. Kubai and Hans Ris. Division in the dinoflagellate gyrodinium cohnii (schiller). Journal of Cell Biology, 40(2):508–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
2/508.

Kraemer:1967:CCS

[Kra67] Paul M. Kraemer. Configuration change of surface sialic acid during mitosis. Journal of Cell Biology, 33(1):197–??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/33/1/197.

Krishan:1969:AND

[KRC69] Awtar Krishan and Ratna Ray-Chaudhuri. Asynchrony of nuclear development in cytochalasin-induced multinucleate cells. Journal of Cell Biology, 43(3):618–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/43/3/618.

Kleinfeld:1966:MKA

[KS66] Ruth G. Kleinfeld and Jesse E. Sisken. Morphological and kinetic aspects of mitotic arrest by and recovery from colcemid. Journal of Cell Biology, 31(3):369–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/31/3/369.

Kato:1967:CPH

[KS67] Hatao Kato and Avery A. Sandberg. Chromosome pulverization in human binucleate cells following colcemid treatment. Journal of Cell Biology, 34(1):35–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/1/35.

Kane:1969:CSI

[KS69] R. E. Kane and R. E. Stephens. A comparative study of the isolation of the cortex and the role of the calcium-insoluble protein in several species of sea urchin egg. Journal of Cell Biology, 41(1):133–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/1/133.

Kislev:1965:NAC

[KSB65] Naomi Kislev, Hewson Swift, and Lawrence Bogorad. Nucleic acids of chloroplasts and mitochondria in Swiss chard. Journal of Cell Biology, 25(2):327–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/25/2/327.

Kramarsky:1967:PEF

[KSR67] Bernhard Kramarsky, Richard Siegler, and Marvin A. Rich. Presence of endothelial fenestrations in thymic capillaries of mice. Journal of Cell Biology, 35(2):464–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/35/2/464.

Kodama:1968:SPC

[KT68a] Robert M. Kodama and Henry Tedeschi. Studies on the permeability of calf thymus nuclei isolated in sucrose. Journal of Cell Biology, 37(3):747–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/37/3/747.

Kornguth:1968:CDH

[KT68b] Steven E. Kornguth and Lawrence G. Tomasi. Changes in the distribution of a histone in dorsal root ganglion neurons during development. Journal of Cell Biology, 38(3):515–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
3/515.

Kuehl:1969:EVI

[Kue69] Leroy Kuehl. Effect of various inhibitors on nuclear protein synthesis in rat liver. Journal of Cell Biology, 41(2):660–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/41/
2/660.

Kinter:1965:ASS

[KW65] William B. Kinter and T. Hastings Wilson. Autoradiographic study of sugar and amino acid absorption by everted sacs of hamster intestine. Journal of Cell Biology, 25(2):19–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/25/2/19.

Kallman:1967:PRU

[KW67a] Frances Kallman and Norman K. Wessells. Periodic repeat units of epithelial cell tonofilaments. Journal of Cell Biology, 32(1): 227–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/32/1/227.

Korn:1967:PLB

[KW67b] Edward D. Korn and Robert A. Weisman. Phagocytosis of latex beads by Acanthamoeba. Journal of Cell Biology, 34(1):219–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
1/219.

Kruse:1969:MNM

[KWM69] Paul F. Kruse, Wilbur Whittle, and Ed Miedema. Mitotic and nonmitotic multiple-layered perfusion cultures. Journal of Cell Biology, 42(1):113–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/42/1/113.

Kaye:1966:FTR

[KWWL66] Gordon I. Kaye, Henry O. Wheeler, Robert T. Whitlock, and Nathan Lane. Fluid transport in the rabbit gallbladder. Journal of Cell Biology, 30(2):237–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/30/2/237.

Kulka:1966:SAE

[KY66] R. G. Kulka and U. Yalovsky. Secretion of α-amylase by the embryonic chick pancreas in vitro. Journal of Cell Biology, 29(2):287–??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/29/2/287.

Kelly:1969:FSM

[KZ69a] A. M. Kelly and S. I. Zacks. The fine structure of motor endplate morphogenesis. Journal of Cell Biology, 42(1):154–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/42/1/154.

Kelly:1969:HRI

[KZ69b] A. M. Kelly and S. I. Zacks. The histogenesis of rat intercostal muscle. Journal of Cell Biology, 42(1):135–??, July 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/42/1/135.

Kolehmainen:1965:SCD

[KZvW65] L. Kolehmainen, H. Zech, and D. von Wettstein. The structure of cells during tobacco mosaic virus reproduction. Journal of Cell Biology, 25(3):77–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/25/3/77.

Longo:1968:FSP

[LA68] Frank J. Longo and Everett Anderson. The fine structure of pronuclear development and fusion in the sea urchin, Arbacia Punctulata. Journal of Cell Biology, 39(2):339–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/39/
2/339.

Lafontaine:1965:LEM

[Laf65] J. G. Lafontaine. A light and electron microscope study of small, spherical nuclear bodies in meristematic cells of Allium Cepa, Vicia Faba, and Raphanus Sativus. Journal of Cell Biology, 26(1):1– ??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/26/
1/1.

Landau:1965:HHP

[Lan65a] Joseph V. Landau. High hydrostatic pressure effects on Amoeba Proteus. Journal of Cell Biology, 24(2):332–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/24/2/332.

Lane:1965:ACD

[Lan65b] Bernard P. Lane. Alterations in the cytologic detail of intestinal smooth muscle cells in various stages of contraction. Journal of Cell Biology, 27(1):199–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/27/1/199.

Landau:1966:HPE

[Lan66] J. V. Landau. High pressure effects on endogenous adenine nucleotide levels in sea urchin eggs prior to first cleavage. Journal of Cell Biology, 28(2):408–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/28/2/408.

Lane:1967:LPA

[Lan67a] Bernard P. Lane. Localization of products of ATP hydrolysis in mammalian smooth muscle cells. Journal of Cell Biology, 34 (3):713–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/34/3/713.

Lane:1967:SRN

[Lan67b] Nancy J. Lane. Spheroidal and ring nucleoli in amphibian oocytes. Journal of Cell Biology, 35(2):421–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/35/2/421.

Lane:1968:DPG

[Lan68] Nancy J. Lane. Distribution of phosphatases in the Golgi region and associated structures of the thoracic ganglionic neurons in the grasshopper, Melanoplus Differentialis. Journal of Cell Biology, 37(1):89–??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/37/1/89.

Lane:1969:IFB

[Lan69a] Nancy J. Lane. Intranuclear fibrillar bodies in actinomycin Dtreated oocytes. Journal of Cell Biology, 40(1):286–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
1/286.

Langreth:1969:SCC

[Lan69b] Susan G. Langreth. Spermiogenesis in cancer crabs. Journal of Cell Biology, 43(3):575–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/43/3/575.

Lasansky:1967:CJO

[Las67a] Arnaldo Lasansky. Cell junctions in ommatidia of limulus. Journal of Cell Biology, 33(2):365–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/33/2/365.

Lasansky:1967:PBH

[Las67b] Arnaldo Lasansky. The pathway between hyaloid blood and retinal neurons in the toad. Journal of Cell Biology, 34(2):617–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
2/617.

Lasansky:1969:BJS

[Las69] Arnaldo Lasansky. Basal junctions at synaptic endings of turtle visual cells. Journal of Cell Biology, 40(2):577–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
2/577.

Lavker:1969:FSM

[Lav69] Robert M. Lavker. Fine structure of mucus granules in rumen epithelium. Journal of Cell Biology, 41(2):657–??, May 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/41/2/657.

Leak:1968:USL

[LB68] L. V. Leak and J. F. Burke. Ultrastructural studies on the lymphatic anchoring filaments. Journal of Cell Biology, 36(1):129–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
1/129.

Leduc:1967:UFS

[LBHT67] Elizabeth H. Leduc, W. Bernhard, S. J. Holt, and J. P. Tranzer. Ultrathin frozen sections. Journal of Cell Biology, 34(3):773– ??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/34/3/773.

Locke:1965:SFP

[LC65] M. Locke and J. V. Collins. The structure and formation of protein granules in the fat body of an insect. Journal of Cell Biology, 26(3):857–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/26/3/857.

Luckenbill:1966:ALD

[LC66] Louise M. Luckenbill and Alan S. Cohen. The association of lipid droplets with cytoplasmic filaments in avian subsynovial adipose cells. Journal of Cell Biology, 31(1):195–??, October 1966. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/31/1/195.

Locke:1968:PUM

[LC68] M. Locke and Janet V. Collins. Protein uptake into multivesicular bodies and storage granules in the fat body of an insect. Journal of Cell Biology, 36(3):453–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/36/3/453.

Lamm:1965:SNA

[LCW65] Michael E. Lamm, Lillian Childers, and Merrill K. Wolf. Studies on nucleic acid metachromasy. Journal of Cell Biology, 27(2): 313–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/2/313.

Lindberg:1967:SME

[LDRA67] O. Lindberg, J. De Pierre, E. Rylander, and B. A. Afzelius. Studies of the mitochondrial energy-transfer system of Brown adipose tissue. Journal of Cell Biology, 34(1):293–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/34/1/293.

Leach:1968:TPG

[Lea68] William M. Leach. The thymidine pool in grasshopper neuroblasts during mitosis. Journal of Cell Biology, 36(1):282–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
1/282.

Lentz:1966:IGG

[Len66] Thomas L. Lentz. Intramitochondrial glycogen granules in digestive cells of hydra. Journal of Cell Biology, 29(1):162–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/29/
1/162.

Lentz:1969:DNJ

[Len69] Thomas L. Lentz. Development of the neuromuscular junction. Journal of Cell Biology, 42(2):431–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/42/2/431.

Lesseps:1967:RPC

[Les67] Roland J. Lesseps. The removal by phospholipase C of a layer of lanthanum-staining material external to the cell membrane in embryonic chick cells. Journal of Cell Biology, 34(1):173–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
1/173.

Love:1965:CCS

[LF65] Robert Love and Mario V. Fernandes. Cytological and cytochemical studies of green monkey kidney cells infected in vitro with simian virus 40. Journal of Cell Biology, 25(3):529–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
3/529.

Lasansky:1969:SOE

[LF69] A. Lasansky and M. G. F. Fuortes. The site of origin of electrical responses in visual cells of the leech, hirudo medicinalis. Journal of Cell Biology, 42(1):241–??, July 1969. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/42/1/241.

Lipetz:1965:IEL

[LG65] Jacques Lipetz and Anthony J. Garro. Ionic effects on lignification and peroxidase in tissue cultures. Journal of Cell Biology, 25(1):109–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/1/109.

Leduc:1965:HMN

[LH65] Elizabeth H. Leduc and S. J. Holt. Hydroxypropyl methacrylate, a new water-miscible embedding medium for electron microscopy. Journal of Cell Biology, 26(1):137–??, July 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/26/1/137.

Levy:1967:HAO

[LH67] Michael R. Levy and Ann E. Hunt. L-α-hydroxy acid oxidase activity in Tetrahymena. Journal of Cell Biology, 34(3): 911–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/34/3/911.

Levy:1965:MCU

[LHS^{+}65] W. Arnoldo Levy, Ivan Herzog, Kunihiko Suzuki, Robert Katzman, and Labe Scheinberg. Method for combined ultrastructural and biochemical analysis of neural tissue. Journal of Cell Biology, 27(1):119–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/1/119.

Lin:1965:MWM

[Lin65] Huai-San Lin. Microcylinders within mitochondrial cristae in the rat pinealocyte. Journal of Cell Biology, 25(3):435–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/25/3/435.

Lin:1967:PCA

[Lin67] Huai-San Lin. A peculiar configuration of agranular reticulum (canaliculate lamellar body) in the rat pinealocyte. Journal of Cell Biology, 33(1):15–??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/33/1/15.

Loewenstein:1967:ICTa

[LK67a] Werner R. Loewenstein and Yoshinobu Kanno. Intercellular communication and tissue growth. Journal of Cell Biology, 33(2):225– ??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
2/225.

Lom:1967:UPV

[LK67b] Jiři Lom and Eugene N. Kozloff. The ultrastructure of phalacrocleptes verruciformis, an unciliated ciliate parasitizing the polychaete schizobranchia insignis. Journal of Cell Biology, 33 (2):355–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/33/2/355.

Leeson:1965:UAR

[LL65] C. Roland Leeson and Thomas S. Leeson. An unusual arrangement of ribosomes in Mesenchymal cells. Journal of Cell Biology, 24(2):324–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/2/324.

Leeson:1966:OLB

[LL66] T. S. Leeson and C. R. Leeson. Osmiophilic lamellated bodies and associated material in lung alveolar spaces. Journal of Cell Biology, 28(3):577–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/28/3/577.

Legato:1969:SLC

[LL69a] Marianne J. Legato and Glenn A. Langer. The subcellular localization of calcium ion in mammalian myocardium. Journal of Cell Biology, 41(2):401–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/41/2/401.

Lord:1969:ONM

[LL69b] A. Lord and J.-G. Lafontaine. The organization of the nucleolus in meristematic plant cells. Journal of Cell Biology, 40(3):633–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
3/633.

Lane:1965:EAD

[LN65a] Nancy J. Lane and Alex B. Novikoff. Effects of arginine deprivation, ultraviolet radiation, and X-radiation on cultured Kb cells. Journal of Cell Biology, 27(3):603–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/27/3/603.

Laufer:1965:DSD

[LN65b] Hans Laufer and Yasukiyo Nakase. Developmental studies of the dipteran salivary gland. Journal of Cell Biology, 25(1):97–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
1/97.

Locke:1965:SSD

[Loc65] Michael Locke. The structure of septate desmosomes. Journal of Cell Biology, 25(1):166–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/25/1/166.

Loud:1968:QSD

[Lou68] Alden V. Loud. A quantitative stereological description of the ultrastructure of normal rat liver parenchymal cells. Journal of Cell Biology, 37(1):27–??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/37/1/27.

Lovlie:1969:IFD

[Løv69] Arne Løvlie. Increase in function during growth of a single cell organelle. Journal of Cell Biology, 43(3):389–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/43/3/389.

Lintilhac:1966:LCS

[LP66] Philip M. Lintilhac and R. B. Park. Localization of chlorophyll in spinach chloroplast lamellae by fluorescence microscopy. Journal of Cell Biology, 28(3):582–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/28/3/582.

Loewenstein:1967:ICTb

[LP67] Werner R. Loewenstein and Richard D. Penn. Intercellular communication and tissue growth. Journal of Cell Biology, 33(2):235– ??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
2/235.

Liau:1969:RPP

[LP69] Ming C. Liau and Robert P. Perry. Ribosome precursor particles in nucleoli. Journal of Cell Biology, 42(1):272–??, July 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/42/1/272.

Leighton:1968:LSS

[LPB^{+}68] Federico Leighton, Brian Poole, Henri Beaufay, Pierre Baudhuin, John W. Coffey, Stanley Fowler, and Christian De Duve. The large-scale separation of peroxisomes, mitochondria, and lysosomes from the livers of rats injected with triton wr-1339. Journal of Cell Biology, 37(2):482–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/37/2/482.

Leighton:1969:STR

[LPLD69] Federico Leighton, Brian Poole, Paul B. Lazarow, and Christian De Duve. The synthesis and turnover of rat liver peroxisomes. Journal of Cell Biology, 41(2):521–??, May 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/41/2/521.

Landolt:1966:EMS

[LR66] Alex M. Landolt and Hans Ris. Electron microscope studies on soma-somatic interneuronal junctions in the corpus pedunculatum of the wood ant (Formica Lugubris Zett.). Journal of Cell Biology, 28(2):391–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/28/2/391.

Lovtrup-Rein:1966:IFR

[LRM66] H. Løvtrup-Rein and B. S. McEwen. Isolation and fractionation of rat brain nuclei. Journal of Cell Biology, 30(2):405–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/30/
2/405.

Labella:1965:INE

[LS65] Frank S. Labella and Madhu Sanwal. Isolation of nerve endings from the posterior pituitary gland. Journal of Cell Biology, 25(3):179–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/3/179.

Lazarus:1968:ARA

[LS68a] L. H. Lazarus and O. H. Scherbaum. Activity of ribonuclease, acid phosphatase, and phosphodiesterase in Tetrahymena Pyriformis during growth. Journal of Cell Biology, 36(2):415–??, February 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
2/415.

Lenard:1968:ACP

[LS68b] John Lenard and S. J. Singer. Alteration of the conformation of proteins in red blood cell membranes and in solution by fixatives used in electron microscopy. Journal of Cell Biology, 37(1):117– ??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/37/
1/117.

Legato:1968:UAP

[LSL68] Marianne J. Legato, David Spiro, and Glenn A. Langer. Ultrastructural alterations produced in mammalian myocardium by variation in perfusate ionic composition. Journal of Cell Biology, 37(1):1–??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/37/1/1.

Lentz:1967:FSS

[LT67] Thomas L. Lentz and J. P. Trinkaus. A fine structural study of cytodifferentiation during cleavage, blastula, and gastrula stages of fundulus heteroclitus. Journal of Cell Biology, 32(1):121–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/32/
1/121.

Lee:1968:BHS

[LT68a] Sin Hang Lee and Richard M. Torack. A biochemical and histochemical study of glutamic oxalacetic transaminase activity of rat hepatic mitochondria fixed in situ and in vitro. Journal of Cell Biology, 39(3):725–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/3/725.

Lee:1968:EMS

[LT68b] Sin Hang Lee and Richard M. Torack. Electron microscope studies of glutamic oxalacetic transaminase in rat liver cell. Journal of Cell Biology, 39(3):716–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/3/716.

Libanati:1969:DWS

[LT69] Cesar M. Libanati and Carlos J. Tandler. The distribution of the water-soluble inorganic orthophosphate ions within the cell: Accumulation in the nucleus. Journal of Cell Biology, 42(3): 754–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/3/754.

Luck:1965:FMN

[Luc65a] David J. L. Luck. Formation of mitochondria in neurospora crassa. Journal of Cell Biology, 24(3):461–??, March 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/24/3/461.

Luck:1965:IPP

[Luc65b] David J. L. Luck. The influence of precursor pool size on mitochondrial composition in neurospora crassa. Journal of Cell Biology, 24(3):445–??, March 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/24/3/445.

Leene:1965:TBTa

[LvI65a] W. Leene and Woutera van Iterson. Tetranitro-blue tetrazolium reduction in bacillus subtilis. Journal of Cell Biology, 27(1): 237–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/1/237.

Leene:1965:TBTb

[LvI65b] W. Leene and Woutera van Iterson. Tetranitro-blue tetrazolium reduction in Proteus vulgaris. Journal of Cell Biology, 27(1): 241–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/1/241.

Lehrer:1968:QHS

[LWS^{+}68] Gerard M. Lehrer, Cipora Weiss, Demetra J. Silides, Cecile Lichtman, Moses Furman, and Robert F. Mathewson. The quantitative histochemistry of supramedullary neurons of puffer fishes. Journal of Cell Biology, 37(2):575–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/37/2/575.

Lyman:1967:SIC

[Lym67] Harvard Lyman. Specific inhibition of chloroplast replication in euglena gracilis by nalidixic acid. Journal of Cell Biology, 35 (3):726–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/3/726.

Maack:1967:CAA

[Maa67] Thomas Maack. Changes in the activity of acid hydrolases during renal reabsorption of lysozyme. Journal of Cell Biology, 35 (1):268–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/1/268.

Manasek:1968:MDC

[Man68] Francis J. Manasek. Mitosis in developing cardiac muscle. Journal of Cell Biology, 37(1):191–??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/37/1/191.

Manasek:1969:AGG

[Man69] Francis J. Manasek. The appearance of granules in the Golgi complex of embryonic cardiac myocytes. Journal of Cell Biology, 43(3):605–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/43/3/605.

Matsumoto:1965:SFS

[Mat65] Jiro Matsumoto. Studies on fine structure and cytochemical properties of erythrophores in swordtail, xiphophorus helleri, with special reference to their pigment granules (pterinosomes). Journal of Cell Biology, 27(3):493–??, December 1965. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/27/3/493.

Matsusaka:1967:AAC

[Mat67] Toshihiko Matsusaka. Atpase activity in the ciliary rootlet of human retinal rods. Journal of Cell Biology, 33(1):203–??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
1/203.

Mole-Bajer:1965:TSC

[MB65] J. Molè-Bajer. Telophase segregation of chromosomes and amitosis. Journal of Cell Biology, 25(1):??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/25/1/79.2.

McClary:1967:SDO

[MB67] Dan O. McClary and Wilbert D. Bowers. Structural differentiation of obligately aerobic and facultatively anaerobic yeasts. Journal of Cell Biology, 32(2):519–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/32/2/519.

Masurovsky:1967:CSOa

[MBB67a] Edmund B. Masurovsky, Mary Bartlett Bunge, and Richard P. Bunge. Cytological studies of organotypic cultures of rat dorsal root ganglia following x-irradiation in vitro. Journal of Cell Biology, 32(2):467–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/32/2/467.

Masurovsky:1967:CSOb

[MBB67b] Edmund B. Masurovsky, Mary Bartlett Bunge, and Richard P. Bunge. Cytological studies of organotypic cultures of rat dorsal root ganglia following x-irradiation in vitro. Journal of Cell Biology, 32(2):497–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/32/2/497.

Mego:1967:UFT

[MBM67] John L. Mego, Francisco Bertini, and J. Donald McQueen. The use of formaldehyde-treated^{131} i-albumin in the study of digestive vacuoles and some properties of these particles from mouse liver. Journal of Cell Biology, 32(3):699–??, March 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/32/3/699.

McCombs:1968:UMF

[MBMB68] Robert M. McCombs, Matilda Benyesh-Melnick, and J. Pierre Brunschwig. The use of millipore filters in ultrastructural studies of cell cultures and viruses. Journal of Cell Biology, 36(1):231–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
1/231.

Merlini:1965:EDA

[MC65] Dino Merlini and Felice Caramia. Effect of dehydroascorbic acid on the islets of langerhans of the rat pancreas. Journal of Cell Biology, 26(1):245–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/26/1/245.

Muscatello:1969:OEE

[MC69] Umberto Muscatello and Ernesto Carafoli. The oxidation of exogenous and endogenous cytochrome C in mitochondria. Journal of Cell Biology, 40(3):602–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/40/3/602.

Metz:1968:EPE

[MCB68] Charles B. Metz, M. Virginia Cone, and Jeanne Bryant. Effect of proteolytic enzymes on the ultrastructure of antibody-treated sea urchin eggs. Journal of Cell Biology, 39(2):481–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/39/
2/481.

McConnell:1965:IRO

[McC65] David G. McConnell. The isolation of retinal outer segment fragments. Journal of Cell Biology, 27(3):459–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/27/3/459.

McCurdy:1969:ELD

[McC69] Harriet M. McCurdy. Enzyme localization during melanogenesis. Journal of Cell Biology, 43(2):220–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/43/2/220.

Montgomery:1965:UMI

[MCK65] P. O’B. Montgomery, James E. Cook, and David Karney. Ultraviolet microbeam irradiation in living cell membranes. Journal of Cell Biology, 26(3):959–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/26/3/959.

Macieira-Coelho:1969:AGB

[MCP69] A. Macieira-Coelho and J. Ponten. Analogy in growth between late passage human embryonic and early passage human adult fibroblasts. Journal of Cell Biology, 43(2):374–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/43/2/374.

Mattern:1966:FCR

[MD66a] Carl F. T. Mattern and Wendell A. Daniel. The flame cell of rotifer. Journal of Cell Biology, 29(3):552–??, June 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/29/3/552.

Mattern:1966:SCR

[MD66b] Carl F. T. Mattern and Wendell A. Daniel. The stomach cell of rotifer. Journal of Cell Biology, 29(3):547–??, June 1966. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/29/3/547.

Marikovsky:1969:EMA

[MD69] Y. Marikovsky and D. Danon. Electron microscope analysis of Young and old red blood cells stained with colloidal iron for surface charge evaluation. Journal of Cell Biology, 43(1):1–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/43/1/1.

Miller:1966:SEL

[MdHP66] Fritz Miller, Etienne de Harven, and George E. Palade. The structure of eosinophil leukocyte granules in rodents and in man. Journal of Cell Biology, 31(2):349–??, November 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/31/2/349.

Mollenhauer:1968:DSE

[MEK68] Hilton H. Mollenhauer, W. Evans, and Clara Kogut. Dictyosome structure in euglena gracilis. Journal of Cell Biology, 37(2):579– ??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/37/
2/579.

Mendelson:1969:EMC

[Men69] Martin Mendelson. Electrical and mechanical characteristics of a very fast lobster muscle. Journal of Cell Biology, 42(2):548–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
2/548.

Merrillees:1968:NEI

[Mer68] Neil C. R. Merrillees. The nervous environment of individual smooth muscle cells of the guinea pig vas deferens. Journal of Cell Biology, 37(3):794–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/37/3/794.

Metuzals:1967:HAS

[Met67] J. Metuzals. Helical arrangement of the subunits of the neurofibrillar bundles isolated from leech nervous system. Journal of Cell Biology, 34(2):690–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/2/690.

Metuzals:1969:CFN

[Met69] J. Metuzals. Configuration of a filamentous network in the axoplasm of the squid (Loligo Pealii L.) giant nerve fiber. Journal of Cell Biology, 43(3):480–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/43/3/480.

McNutt:1969:UCM

[MF69] N. Scott McNutt and Don W. Fawcett. The ultrastructure of the cat myocardium. Journal of Cell Biology, 42(1):46–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/42/1/46.

Mizukami:1966:CR

[MG66] Ikuko Mizukami and Joseph Gall. Centriole replication. Journal of Cell Biology, 29(1):97–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/29/1/97.

Moog:1967:STD

[MG67] Florence Moog and Robert D. Grey. Spatial and temporal differentiation of alkaline phosphatase on the intestinal villi of the mouse. Journal of Cell Biology, 32(2):??, February 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/32/2/C1.

McEwen:1968:FSC

[MG68] Bruce S. McEwen and Bernice Grafstein. Fast and slow components in axonal transport of protein. Journal of Cell Biology, 38 (3):494–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/3/494.

Morgan:1967:LLD

[MH67] Thomas E. Morgan and Gary L. Huber. Loss of lipid during fixation for electron microscopy. Journal of Cell Biology, 32(3):757– ??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/32/
3/757.

Miller:1965:HSU

[MHA65] A. T. Miller, D. M. Hale, and K. D. Alexander. Histochemical studies on the uptake of horseradish peroxidase by rat kidney slices. Journal of Cell Biology, 27(2):305–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/27/2/305.

Metuzals:1969:SPT

[MI69] J. Metuzals and C. S. Izzard. Spatial patterns of threadlike elements in the axoplasm of the giant nerve fiber of the squid (Loligo Pealii L.) as disclosed by differential interference microscopy and by electron microscopy. Journal of Cell Biology, 43 (3):456–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/43/3/456.

Millward:1969:CCP

[Mil69] G. R. Millward. Cellulose contamination: a possible source of error in the interpretation of previous experimental evidence for the α-keratin protofibril. Journal of Cell Biology, 42(1):317–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
1/317.

Mishima:1966:MPV

[Mis66] Yutaka Mishima. Melanosomes in phagocytic vacuoles in langerhans cells. Journal of Cell Biology, 30(2):417–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/30/2/417.

Mitchell:1969:PLH

[Mit69] Carolyn D. Mitchell. Preservation of the lipids of the human erythrocyte stroma during fixation and dehydration for electron microscopy. Journal of Cell Biology, 40(3):869–??, March 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/40/3/869.

Maxwell:1965:FSA

[MK65a] David S. Maxwell and Lawrence Kruger. The fine structure of astrocytes in the cerebral cortex and their response to focal injury produced by heavy ionizing particles. Journal of Cell Biology, 25(2):141–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/2/141.

Morrill:1965:PNA

[MK65b] Gene A. Morrill and Adele B. Kostellow. Phospholipid and nucleic acid gradients in the developing amphibian embryo. Journal of Cell Biology, 25(3):21–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/25/3/21.

Matukas:1968:ECP

[MK68] Victor J. Matukas and George A. Krikos. Evidence for changes in protein polysaccharide associated with the onset of calcification in cartilage. Journal of Cell Biology, 39(1):43–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/39/1/43.

Maisel:1966:AAC

[ML66a] John C. Maisel and Ralph I. Lytle. Affinity of animal cell nucleoli for normal serum. Journal of Cell Biology, 29(3):461–??, June 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/29/
3/461.

Merkow:1966:INA

[ML66b] Leonard Merkow and Joseph Leighton. Increased numbers of annulate lamellae in myocardium of chick embryos incubated at abnormal temperatures. Journal of Cell Biology, 28(1):127–??, January 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/28/
1/127.

Miyamoto:1965:NNB

[MM65] Kaneatsu Miyamoto and Seiichi Matsumoto. The nature of the negri body. Journal of Cell Biology, 27(3):677–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/27/
3/677.

Mollenhauer:1966:TCB

[MM66a] Hilton H. Mollenhauer and D. James Morré. Tubular connections between dictyosomes and forming secretory vesicles in plant Golgi apparatus. Journal of Cell Biology, 29(2):373–??, May 1966. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/29/2/373.

Muckenthaler:1966:DSO

[MM66b] F. A. Muckenthaler and A. P. Mahowald. DNA synthesis in the ooplasm of Drosophila Melanogaster. Journal of Cell Biology, 28 (2):199–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/28/2/199.

Muscatello:1965:DRS

[MMA65] Umberto Muscatello, Alfredo Margreth, and Massimo Aloisi. On the differential response of sarcoplasm and myoplasm to denervation in frog muscle. Journal of Cell Biology, 27(1):1–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/27/1/1.

Maniloff:1965:SUR

[MMB65] Jack Maniloff, Harold J. Morowitz, and Russell J. Barrnett. Studies of the ultrastructure and ribosomal arrangements of the pleuropneumonia-like organism a5969. Journal of Cell Biology, 25(1):139–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/1/139.

Matsuura:1968:SPS

[MMNT68a] Shiro Matsuura, Takashi Morimoto, Sunao Nagata, and Yutaka Tashiro. Studies on the posterior silk gland of the silkworm, bombyx mori. Journal of Cell Biology, 38(3):589–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
3/589.

Morimoto:1968:SPS

[MMNT68b] Takashi Morimoto, Shiro Matsuura, Sunao Nagata, and Yutaka Tashiro. Studies on the posterior silk gland of the silkworm, bombyx mori. Journal of Cell Biology, 38(3):604–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
3/604.

Murray:1965:FSM

[MMP65] Raymond G. Murray, Assia S. Murray, and Anthony Pizzo. The fine structure of mitosis in rat thymic lymphocytes. Journal of Cell Biology, 26(2):601–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/26/2/601.

Matsumoto:1968:MBC

[MO68] Jiro Matsumoto and Masataka Obika. Morphological and biochemical characterization of goldfish erythrophores and their pterinosomes. Journal of Cell Biology, 39(2):233–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/39/
2/233.

Miles:1969:LPP

[MO69] Charles P. Miles and Frank O’Neill. ^{3}H labeling patterns of permanent cell line chromosomes showing pulverization or accentuated secondary constrictions. Journal of Cell Biology, 40 (2):553–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/2/553.

Moens:1969:FSM

[Moe69a] Peter B. Moens. The fine structure of meiotic chromosome pairing in the triploid, Lilium Tigrinum. Journal of Cell Biology, 40 (1):273–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/1/273.

Moens:1969:MCC

[Moe69b] Peter B. Moens. Multiple core complexes in grasshopper spermatocytes and spermatids. Journal of Cell Biology, 40(2):542–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
2/542.

Mollenhauer:1965:ISG

[Mol65] Hilton H. Mollenhauer. An intercisternal structure in the Golgi apparatus. Journal of Cell Biology, 24(3):504–??, March 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/24/3/504.

Moor:1966:UIZ

[Moo66] Hans Moor. Ultrastrukturen im Zellkern der Bäckerhefe. (German) [Ultrastructures in the cell nucleus of baker’s yeast]. Journal of Cell Biology, 29(1):153–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/29/1/153.

Morgan:1968:DEN

[Mor68] Winfield S. Morgan. Ductal excretion of neutral red lysosomes in the mouse pancreas. Journal of Cell Biology, 36(1):261–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
1/261.

Mourad:1965:ECD

[Mou65] Effat Badr Mourad. Evidence for cytoplasmic DNA in root cells of nicotiana. Journal of Cell Biology, 24(2):267–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/24/2/267.

Matoltsy:1965:MCG

[MP65] A. Gedeon Matoltsy and Paul F. Parakkal. Membrane-coating granules of keratinizing epithelia. Journal of Cell Biology, 24 (2):297–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/2/297.

Milhaud:1966:PBH

[MP66] Monique Milhaud and George D. Pappas. Postsynaptic bodies in the habenula and interpeduncular nuclei of the cat. Journal of Cell Biology, 30(2):437–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/30/2/437.

Marchesi:1967:LMK

[MP67a] Vincent T. Marchesi and George E. Palade. The localization of Mg–Na–K-activated adenosine triphosphatase on red cell ghost membranes. Journal of Cell Biology, 35(2):385–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/35/2/
385.

McIntosh:1967:MSD

[MP67b] J. Richard McIntosh and Keith R. Porter. Microtubules in the spermatids of the domestic fowl. Journal of Cell Biology, 35 (1):153–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/1/153.

Milhaud:1968:CFA

[MP68] Monique Milhaud and George D. Pappas. Cilia formation in the adult cat brain after pargyline treatment. Journal of Cell Biology, 37(3):599–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/37/3/599.

Mattern:1965:EMO

[MPD65] Carl F. T. Mattern, Helen D. Park, and Wendell A. Daniel. Electron microscope observations on the structure and discharge of the stenotele of hydra. Journal of Cell Biology, 27(3):621–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/27/
3/621.

Michell:1969:MRP

[MPNK69] Robert H. Michell, Samuel J. Pancake, John Noseworthy, and Manfred L. Karnovsky. Measurement of rates of phagocytosis. Journal of Cell Biology, 40(1):216–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/40/1/216.

Matukas:1967:SUI

[MPO67] Victor J. Matukas, Bernard J. Panner, and J. Lowell Orbison. Studies on ultrastructural identification and distribution of protein-polysaccharide in cartilage matrix. Journal of Cell Biology, 32(2):365–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/32/2/365.

McManus:1965:FDM

[MR65] Sister M. A. McManus and L. E. Roth. Fibrillar differentiation in myxomycete plasmodia. Journal of Cell Biology, 25(2):305– ??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
2/305.

Martin:1968:FSA

[MR68] Rainer Martin and Philip Rosenberg. Fine structural alterations associated with venom action on squid giant nerve fibers. Journal of Cell Biology, 36(2):341–??, February 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/36/2/341.

Marcus:1967:OIC

[MRH^{+}67] Leon Marcus, H. Ris, H. O. Halvorson, R. K. Bretthauer, and R. M. Bock. Occurrence, isolation, and characterization of polyribosomes in yeast. Journal of Cell Biology, 34(2):505–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
2/505.

Malawista:1969:VPU

[MS69a] Stephen E. Malawista and Hidemi Sato. Vinblastine produces uniaxial, birefringent crystals in starfish oocytes. Journal of Cell Biology, 42(2):596–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/42/2/596.

Marquet:1969:CIN

[MS69b] Eugene Marquet and Harold John Sobel. Crystalline inclusions in the nuclear envelope and granular endoplasmic reticulum of the fish spinal cord. Journal of Cell Biology, 41(3):774–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/41/3/774.

McCarl:1969:EAD

[MS69c] Richard L. McCarl and Robert C. Shaler. The effects of actinomycin D on protein synthesis and beating in cultured rat heart cells. Journal of Cell Biology, 40(3):850–??, March 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/40/3/850.

Mastro:1969:DSR

[MSH69] Andrea Mastro, Emma Shelton, and W. C. Hymer. DNA synthesis in the rat anterior pituitary. Journal of Cell Biology, 43 (3):626–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/43/3/626.

Majno:1969:ECI

[MSL69] Guido Majno, Stephen M. Shea, and Monika Leventhal. Endothelial contraction induced by histamine-type mediators. Journal of Cell Biology, 42(3):647–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/42/3/647.

Mackenzie:1966:DGG

[MSP66] T. B. Mackenzie, G. E. Stone, and D. M. Prescott. The durations of G_{1}, S, and G_{2} at different temperatures in Tetrahymena Pyriformis Hsm. Journal of Cell Biology, 31(3):633–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/31/
3/633.

Morgan:1968:IRC

[MSW68] Richard S. Morgan, Henry S. Slayter, and David L. Weller. Isolation of ribosomes from cysts of entamoeba invadens. Journal of Cell Biology, 36(1):45–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/36/1/45.

Mahadevan:1967:LSP

[MT67] P. R. Mahadevan and E. L. Tatum. Localization of structural polymers in the cell wall of neurospora crassa. Journal of Cell Biology, 35(2):295–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/35/2/295.

MacInnes:1967:TDF

[MU67] James W. MacInnes and Robert B. Uretz. Thermal depolarization of fluorescence from polytene chromosomes stained with acridine orange. Journal of Cell Biology, 33(3):597–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/33/3/597.

MacInnes:1968:DOM

[MU68] James W. MacInnes and Robert B. Uretz. DNA organization in the mature sperm of several orthoptera by the method of polarized fluorescence microscopy. Journal of Cell Biology, 38 (2):426–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/2/426.

Munger:1965:IIS

[Mun65] Bryce L. Munger. The intraepidermal innervation of the snout skin of the opossum. Journal of Cell Biology, 26(1):79–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/26/1/79.

Mullins:1966:SEE

[MW66] Raneen Mullins and Reimut Wette. On the statistical expectation and evaluation of centriole orientations in cell profiles. Journal of Cell Biology, 30(3):652–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/30/3/652.

Mukherjee:1967:CSU

[MW67] T. M. Mukherjee and A. Wynn Williams. A comparative study of the ultrastructure of microvilli in the epithelium of small and large intestine of mice. Journal of Cell Biology, 34(2):447–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
2/447.

Movat:1965:PPA

[MWGM65] Henry Z. Movat, William J. Weiser, Michael F. Glynn, and James F. Mustard. Platelet phagocytosis and aggregation. Journal of Cell Biology, 27(3):531–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/27/3/531.

Mintz:1967:MBS

[MYS^{+}67] Helen A. Mintz, David H. Yawn, Brian Safer, Edward Bresnick, Annabel G. Liebelt, Zack R. Blailock, Erwin R. Rabin, and Arnold Schwartz. Morphological and biochemical studies of isolated mitochondria from fetal, neonatal, and adult liver and from neoplastic tissues. Journal of Cell Biology, 34(2):513–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
2/513.

Nakajima:1965:CPB

[NA65] Hiromichi Nakajima and Robert D. Allen. The changing pattern of birefringence in plasmodia of the slime mold, Physarum Polycephalum. Journal of Cell Biology, 25(2):361–??, May 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/25/2/361.

Nachmias:1968:FEM

[Nac68] Vivianne T. Nachmias. Further electron microscope studies on fibrillar organization of the ground cytoplasm of chaos chaos. Journal of Cell Biology, 38(1):40–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/38/1/40.

Nachmias:1969:SS

[Nac69] Vivianne T. Nachmias. Studies on streaming. Journal of Cell Biology, 40(1):160–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/40/1/160.

Nagano:1965:LAT

[Nag65] Toshio Nagano. Localization of adenosine triphosphatase activity in the rat sperm tail as revealed by electron microscopy. Journal of Cell Biology, 25(2):101–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/25/2/101.

Nichols:1967:RTT

[NAL^{+}67] W. W. Nichols, P. Aula, A. Levan, W. Heneen, and E. Norrby. Radioautography with tritiated thymidine in measles and sendai virus-induced chromosome pulverizations. Journal of Cell Biology, 35(1):257–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/35/1/257.

Nanninga:1967:FSO

[Nan67] N. Nanninga. Fine structure observed in 50s ribosomal subunits of bacillus subtilis. Journal of Cell Biology, 33(2):??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/33/2/C1.

Nanninga:1968:SFM

[Nan68] N. Nanninga. Structural features of mesosomes (chondrioids) of bacillus subtilis after freeze-etching. Journal of Cell Biology, 39 (2):251–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/39/2/251.

Nanninga:1969:PUB

[Nan69] N. Nanninga. Preservation of the ultrastructure of bacillus subtilis by chemical fixation as verified by freeze-etching. Journal of Cell Biology, 42(3):733–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/42/3/733.

Newcomb:1965:CMW

[NB65] Eldon H. Newcomb and Howard T. Bonnett. Cytoplasmic microtubule and wall microfibril orientation in root hairs of radish. Journal of Cell Biology, 27(3):575–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/27/3/575.

Nickerson:1969:CWR

[NC69] Peter A. Nickerson and Joseph C. Curtis. Concentric whorls of rough endoplasmic reticulum in adrenocortical cells of the Mongolian gerbil. Journal of Cell Biology, 40(3):859–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/40/3/859.

Niijima:1965:ARMa

[ND65a] Lumiko Niijima and Jean Dan. The acrosome reaction in mytilus edulis. Journal of Cell Biology, 25(2):243–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/25/2/243.

Niijima:1965:ARMb

[ND65b] Lumiko Niijima and Jean Dan. The acrosome reaction in mytilus edulis. Journal of Cell Biology, 25(2):249–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/25/2/249.

Noorduyn:1966:RSR

[NdM66] N. J. A. Noorduyn and J. C. H. de Man. RNA synthesis in rat and mouse hepatic cells as studied with light and electron microscope radioautography. Journal of Cell Biology, 30(3):655–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/30/
3/655.

Nicolescu:1966:EDG

[NDRFK66] P. Nicolescu, M. Dolivo, C. Rouiller, and C. Foroglou-Kerameus. The effect of deprivation of glucose on the ultrastructure and function of the superior cervical ganglion of the rat in vitro. Journal of Cell Biology, 29(2):267–??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/29/2/267.

Newcomb:1967:FSP

[New67a] Eldon H. Newcomb. Fine structure of protein-storing plastids in bean root tips. Journal of Cell Biology, 33(1):143–??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/33/1/143.

Newcomb:1967:SVS

[New67b] Eldon H. Newcomb. A spiny vesicle in slime-producing cells of the bean root. Journal of Cell Biology, 35(2):??, November 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/35/2/C17.

Nicklas:1965:CVD

[Nic65] R. Bruce Nicklas. Chromosome velocity during mitosis as a function of chromosome size and position. Journal of Cell Biology, 25(1):119–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/1/119.

Nichols:1968:MAF

[Nic68] Barbara Nichols. A membranous apparatus in a flagellate, tritrichomonas augusta. Journal of Cell Biology, 37(2):559–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/37/2/
559.

Nicklas:1969:CM

[NK69] R. Bruce Nicklas and Carol A. Koch. Chromosome micromanipulation. Journal of Cell Biology, 43(1):40–??, October 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/43/1/40.

Niebauer:1969:OZI

[NKKW69] Gustav Niebauer, Walter S. Krawczyk, Richard L. Kidd, and George F. Wilgram. Osmium zinc iodide reactive sites in the epidermal langerhans cell. Journal of Cell Biology, 43(1):80–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/43/
1/80.

Neutra:1966:RCU

[NL66a] Marian Neutra and C. P. Leblond. Radioautographic comparison of the uptake of galactose-H^{3} and glucose-H^{3} in the Golgi region of various cells secreting glycoproteins or mucopolysaccharides. Journal of Cell Biology, 30(1):137–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/30/1/137.

Neutra:1966:SCM

[NL66b] Marian Neutra and C. P. Leblond. Synthesis of the carbohydrate of mucus in the Golgi complex as shown by electron microscope radioautography of goblet cells from rats injected with glucose-H^{3}. Journal of Cell Biology, 30(1):119–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/30/1/119.

Napolitano:1967:PML

[NLS67] Leonard Napolitano, Francis Lebaron, and Joseph Scaletti. Preservation of myelin lamellar structure in the absence of lipid.

Journal of Cell Biology, 34(3):817–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/34/3/817.

Nobel:1967:EME

[NM67] Park S. Nobel and Satoru Murakami. Electron microscopic evidence for the location and amount of ion accumulation by spinach chloroplasts. Journal of Cell Biology, 32(1):209–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/32/1/209.

Nonomura:1968:MSM

[Non68] Yoshiaki Nonomura. Myofilaments in smooth muscle of guinea pig’s Taenia Coli. Journal of Cell Biology, 39(3):741–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/39/
3/741.

Nishi:1969:FSP

[NOP69] Katsuhide Nishi, Chikayoshi Oura, and Wazir Pallie. Fine structure of Pacinian corpuscles in the mesentery of the cat. Journal of Cell Biology, 43(3):539–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/43/3/539.

Nash:1965:PDL

[NP65] D. Nash and W. Plaut. On the presence of DNA in larval salivary gland nucleoli in Drosophila Melanogaster. Journal of Cell Biology, 27(3):682–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/27/3/682.

Nakane:1967:ELA

[NP67] Paul K. Nakane and G. Barry Pierce. Enzyme-labeled antibodies for the light and electron microscopic localization of tissue antigens. Journal of Cell Biology, 33(2):307–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/33/2/307.

Novikoff:1966:NPC

[NQVF66] Alex B. Novikoff, Nelson Quintana, Humberto Villaverde, and Regina Forschirm. Nucleoside phosphatase and cholinesterase activities in dorsal root ganglia and peripheral nerve. Journal of Cell Biology, 29(3):525–??, June 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/29/3/525.

Newcomb:1968:ACR

[NSHW68] Eldon H. Newcomb, Martin W. Steer, Peter K. Hepler, and William P. Wergin. An atypical crista resembling a “tight junction” in bean root mitochondria. Journal of Cell Biology, 39 (1):35–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/39/1/35.

Organ:1968:MNA

[OBJ^{+}68] Alan E. Organ, Eugene C. Bovee, Theodore L. Jahn, Duncan Wigg, and James R. Fonseca. The mechanism of the nephridial apparatus of paramecium mul timicronucleatum. Journal of Cell Biology, 37(1):139–??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/37/1/139.

Organ:1969:MNA

[OBJ69] Alan E. Organ, Eugene C. Bovee, and Theodore L. Jahn. The mechanism of the nephridial apparatus of paramecium multimicronucleatum. Journal of Cell Biology, 40(2):389–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
2/389.

Orrenius:1966:EMR

[OE66a] Sten Orrenius and Jan L. E. Ericsson. Enzyme-membrane relationship in phenobarbital induction of synthesis of drugmetabolizing enzyme system and proliferation of endoplasmic membranes. Journal of Cell Biology, 28(2):181–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/28/2/
181.

Orrenius:1966:RLG

[OE66b] Sten Orrenius and Jan L. E. Ericsson. On the relationship of liver glucose-6-phosphatase to the proliferation of endoplasmic reticulum in phenobarbital induction. Journal of Cell Biology, 31 (2):243–??, November 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/31/2/243.

Overton:1965:SOB

[OEC65] Jane Overton, Alexander Eichholz, and Robert K. Crane. Studies on the organization of the brush border in intestinal epithelial cells. Journal of Cell Biology, 26(3):693–??, September 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/26/3/693.

Orrenius:1965:PIS

[OEE65] Sten Orrenius, Jan L. E. Ericsson, and Lars Ernster. Phenobarbitalinduced synthesis of the microsomal drug-metabolizing enzyme system and its relationship to the proliferation of endoplasmic membranes. Journal of Cell Biology, 25(3):627–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/25/3/627.

Orci:1968:GCE

[OJP^{+}68] L. Orci, A. Junod, R. Pictet, A. E. Renold, and C. Rouiller. Granulolysis in a cells of endocrine pancreas in spontaneous and experimental diabetes in animals. Journal of Cell Biology, 38 (2):462–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/2/462.

Outka:1967:AFT

[OK67] Darryll E. Outka and Byron C. Kluss. The ameba-to-flagellate transformation in tetramitus rostratus. Journal of Cell Biology, 35(2):323–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/35/2/323.

Okada:1967:SGM

[Oka67] S. Okada. A simple graphic method of computing the parameters of the life cycle of cultured mammalian cells in the exponential growth phase. Journal of Cell Biology, 34(3):915–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
3/915.

Ollerich:1968:ICE

[Oll68] D. A. Ollerich. An intramitochondrial crystalloid in element III of rat chorioallantoic placenta. Journal of Cell Biology, 37(1):188– ??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/37/
1/188.

Odland:1968:HWR

[OR68] George Odland and Russell Ross. Human wound repair. Journal of Cell Biology, 39(1):135–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/1/135.

Orlic:1968:UFH

[Orl68] Donald Orlic. The use of^{55} Fe in high-resolution radioautography of developing red cells. Journal of Cell Biology, 39(1):201–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/39/
1/201.

Orrenius:1965:FSI

[Orr65a] Sten Orrenius. Further studies on the induction of the drughydroxylating enzyme system of liver microsomes. Journal of Cell Biology, 26(3):725–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/26/3/725.

Orrenius:1965:MDH

[Orr65b] Sten Orrenius. On the mechanism of drug hydroxylation in rat liver microsomes. Journal of Cell Biology, 26(3):713–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/26/
3/713.

Oka:1969:IEPa

[OS69a] Takami Oka and Robert T. Schimke. Interaction of estrogen and progesterone in chick oviduct development. Journal of Cell Biology, 41(3):816–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/41/3/816.

Oka:1969:IEPb

[OS69b] Takami Oka and Robert T. Schimke. Interaction of estrogen and progesterone in chick oviduct development. Journal of Cell Biology, 43(1):123–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/43/1/123.

Ohad:1967:BCMa

[OSP67a] I. Ohad, P. Siekevitz, and G. E. Palade. Biogenesis of chloroplast membranes. Journal of Cell Biology, 35(3):521–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/35/
3/521.

Ohad:1967:BCMb

[OSP67b] I. Ohad, P. Siekevitz, and G. E. Palade. Biogenesis of chloroplast membranes. Journal of Cell Biology, 35(3):553–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/35/
3/553.

Overton:1965:CCF

[Ove65] Jane Overton. Changes in cell fine structure during lens regeneration in Xenopus laevis. Journal of Cell Biology, 24(2):211–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/24/
2/211.

Overton:1966:MMM

[Ove66] Jane Overton. Microtubules and microfibrils in morphogenesis of the scale cells of Ephestia Kühniella. Journal of Cell Biology, 29(2):293–??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/29/2/293.

Overton:1968:LLS

[Ove68] Jane Overton. Localized lanthanum staining of the intestinal brush border. Journal of Cell Biology, 38(2):447–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/38/2/447.

Overton:1969:FIM

[Ove69] Jane Overton. A fibrillar intercellular material between reaggregating embryonic chick cells. Journal of Cell Biology, 40 (1):136–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/1/136.

Ohkawa:1969:UHM

[OVF69] Kin-Ichi Ohkawa, Molly T. Vogt, and Emmanuel Farber. Unusually high mitochondrial alpha glycerophosphate dehydrogenase activity in rat Brown adipose tissue. Journal of Cell Biology, 41(2):441–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/2/441.

Oki:1965:MCI

[OYHM65] Keiichi Oki, Sadaomi Yoshioka, Kiyoharu Hayashi, and Masasuke Masuda. Mitochondrial changes induced by iron absorption in the duodenal absorptive cells of rats. Journal of Cell Biology, 24 (2):328–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/2/328.

Peters:1967:ARD

[PA67] Theodore Peters and Charles A. Ashley. An artefact in radioautography due to binding of free amino acids to tissues by fixatives. Journal of Cell Biology, 33(1):53–??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/33/1/53.

Padawer:1966:ICM

[Pad66] Jacques Padawer. Induction of cellular movements in mast cells by colchicine treatment. Journal of Cell Biology, 29(1):176–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/29/
1/176.

Padawer:1969:UCT

[Pad69] Jacques Padawer. Uptake of colloidal thorium dioxide by mast cells. Journal of Cell Biology, 40(3):747–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/40/3/747.

Page:1965:CFS

[Pag65] Sally G. Page. A comparison of the fine structures of frog slow and twitch muscle fibres. Journal of Cell Biology, 26(2):477–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/26/
2/477.

Pogo:1967:EPR

[PAM67] B. G. T. Pogo, V. G. Allfrey, and A. E. Mirsky. The effect of phytohemagglutinin on ribonucleic acid synthesis and histone acetylation in equine leukocytes. Journal of Cell Biology, 35(2): 477–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/2/477.

Park:1965:SCL

[Par65a] Roderic B. Park. Substructure of chloroplast lamellae. Journal of Cell Biology, 27(1):151–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/27/1/151.

Parsons:1965:MIT

[Par65b] John A. Parsons. Mitochondrial incorporation of tritiated thymidine in Tetrahymena Pyriformis. Journal of Cell Biology, 25 (3):641–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/3/641.

Parakkal:1967:TPK

[Par67] Paul F. Parakkal. Transfer of premelanosomes into the keratinizing cells of albino hair follicle. Journal of Cell Biology, 35 (2):473–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/2/473.

Parakkal:1969:IMC

[Par69a] Paul F. Parakkal. Involvement of macrophages in collagen resorption. Journal of Cell Biology, 41(1):345–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/41/1/345.

Parakkal:1969:UCB

[Par69b] Paul F. Parakkal. Ultrastructural changes of the basal lamina during the hair growth cycle. Journal of Cell Biology, 40 (2):561–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/2/561.

Packer:1965:LIV

[PaSN65] Lester Packer, Paul andré Siegenthaler, and Park S. Nobel. Lightinduced volume changes in spinach chloroplasts. Journal of Cell Biology, 26(2):593–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/26/2/593.

Palade:1968:SMP

[PB68] G. E. Palade and R. R. Bruns. Structural modulations of plasmalemmal vesicles. Journal of Cell Biology, 37(3):633–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/37/
3/633.

Perrelet:1969:EES

[PB69] Alain Perrelet and Fritz Baumann. Evidence for extracellular space in the rhabdome of the honeybee drone eye. Journal of Cell Biology, 40(3):825–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/40/3/825.

Pauly:1969:BTL

[PCS69] J. L. Pauly, G. A. Caron, and R. R. Suskind. Blast transformation of lymphocytes from guinea pigs, rats, and rabbits induced by mercuric chloride in vitro. Journal of Cell Biology, 40(3):847–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
3/847.

Peachey:1965:SRT

[Pea65] Lee D. Peachey. The sarcoplasmic reticulum and transverse tubules of the frog’s sartorius. Journal of Cell Biology, 25(3):209– ??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
3/209.

Penn:1966:ICB

[Pen66] Richard D. Penn. Ionic communication between liver cells. Journal of Cell Biology, 29(1):171–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/29/1/171.

Pepe:1966:SAS

[Pep66] Frank A. Pepe. Some aspects of the structural organization of the myofibril as revealed by antibody-staining methods. Journal of Cell Biology, 28(3):505–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/28/3/505.

Palade:1965:SFD

[PF65] George E. Palade and Marilyn G. Farquhar. A special fibril of the dermis. Journal of Cell Biology, 27(1):215–??, October 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/27/1/215.

Prescott:1968:PNI

[PG68] David Prescott and Lester Goldstein. Proteins in nucleocytoplasmic interactions. Journal of Cell Biology, 39(2):404–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/39/
2/404.

Panner:1967:FUO

[PH67] Bernard J. Panner and Carl R. Honig. Filament ultrastructure and organization in vertebrate smooth muscle. Journal of Cell Biology, 35(2):303–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/35/2/303.

Phillips:1965:OFC

[Phi65] David M. Phillips. An ordered filamentous component in sciara (diptera) salivary gland cell nuclei. Journal of Cell Biology, 26 (2):677–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/2/677.

Phillips:1966:FSS

[Phi66a] David M. Phillips. Fine structure of Sciara Coprophila sperm. Journal of Cell Biology, 30(3):499–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/30/3/499.

Phillips:1966:OSF

[Phi66b] David M. Phillips. Observations on spermiogenesis in the fungus gnat sciara coprophila. Journal of Cell Biology, 30(3):

477–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/30/3/477.

Phillips:1966:SFT

[Phi66c] David M. Phillips. Substructure of flagellar tubules. Journal of Cell Biology, 31(3):635–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/31/3/635.

Phillips:1967:GCF

[Phi67] David M. Phillips. Giant centriole formation in sciara. Journal of Cell Biology, 33(1):73–??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/33/1/73.

Phillips:1969:EPP

[Phi69] David M. Phillips. Exceptions to the prevailing pattern of tubules (9 + 9 + 2) in the sperm flagella of certain insect species. Journal of Cell Biology, 40(1):28–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/40/1/28.

Priest:1967:DOR

[PHP67] Jean H. Priest, Judith E. Heady, and Robert E. Priest. Delayed onset of replication of human X chromosomes. Journal of Cell Biology, 35(2):483–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/35/2/483.

Pick:1965:FSS

[Pic65] Joseph Pick. The fine structure of sympathetic neurons in xirradiated frogs. Journal of Cell Biology, 26(2):335–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/26/
2/335.

Pate:1967:FSCb

[PJO67] Jack L. Pate, John L. Johnson, and Erling J. Ordal. The fine structure of chondrococcus columnaris. Journal of Cell Biology, 35(1):15–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/1/15.

Piper:1969:DSC

[PK69] J. A. Piper and N. W. Klein. The differential sensitivity of cultured chick mesodermal cells to actinomycin D. Journal of Cell Biology, 40(1):209–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/40/1/209.

Plagemann:1969:CMM

[Pla69] Peter G. W. Plagemann. Choline metabolism and membrane formation in rat hepatoma cells grown in suspension culture. Journal of Cell Biology, 42(3):766–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/42/3/766.

Poole:1969:STR

[PLD69] Brian Poole, Federico Leighton, and Christian De Duve. The synthesis and turnover of rat liver peroxisomes. Journal of Cell Biology, 41(2):536–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/41/2/536.

Piddington:1965:CBG

[PM65] Ronald Piddington and A. A. Moscona. Correspondence between glutamine synthetase activity and differentiation in the embryonic retina in situ and in culture. Journal of Cell Biology, 27 (1):247–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/1/247.

Priestley:1968:DMK

[PM68] G. C. Priestley and R. A. Malt. Development of the metanephric kidney. Journal of Cell Biology, 37(3):703–??, June 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/37/3/703.

Priestley:1969:MBR

[PM69] G. C. Priestley and R. A. Malt. Membrane-bound ribosomes in kidney. Journal of Cell Biology, 41(3):886–??, June 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/41/3/886.

Packer:1966:FSC

[PNGM66] Lester Packer, Park S. Nobel, Elizabeth L. Gross, and Howard C. Mel. Fractionation of spinach chloroplasts by flow sedimentationelectrophoresis. Journal of Cell Biology, 28(3):443–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/28/3/
443.

Pate:1965:FST

[PO65] Jack L. Pate and Erling J. Ordal. The fine structure of two unusual stalked bacteria. Journal of Cell Biology, 27(1):133–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/27/
1/133.

Pate:1967:FSCa

[PO67a] Jack L. Pate and Erling J. Ordal. The fine structure of chondrococcus columnaris. Journal of Cell Biology, 35(1):1–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/35/1/1.

Pate:1967:FSCc

[PO67b] Jack L. Pate and Erling J. Ordal. The fine structure of chondrococcus columnaris. Journal of Cell Biology, 35(1):37–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/35/1/37.

Podolsky:1968:DFM

[Pod68] R. J. Podolsky. Deposit formation in muscle fibers following contraction in the presence of lead. Journal of Cell Biology, 39 (1):197–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/39/1/197.

Pogo:1969:EAD

[Pog69] Beatriz G. T. Pogo. The effect of N-acetyl-D-galactosamine and amantadine hydrochloride on equine lymphocytes stimulated with phytohemagglutinin. Journal of Cell Biology, 40(2):571–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
2/571.

Phillips:1969:SNP

[PP69] Stephanie Gordon Phillips and David M. Phillips. Sites of nucleolus production in cultured Chinese hamster cells. Journal of Cell Biology, 40(1):248–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/40/1/248.

Peters:1968:SPN

[PPKA68] Alan Peters, Charmian C. Proskauer, and Ita R. Kaiserman-Abramof. The small pyramidal neuron of the rat cerebral cortex. Journal of Cell Biology, 39(3):604–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/39/3/604.

Parsons:1968:DDA

[PR68] John A. Parsons and Ronald C. Rustad. The distribution of DNA among dividing mitochondria of Tetrahymena Pyriformis. Journal of Cell Biology, 37(3):683–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/37/3/683.

Pasquali-Ronchetti:1969:OSR

[PR69] Ivonne Pasquali-Ronchetti. The organization of the sarcoplasmic reticulum and T system in the femoral muscle of the housefly, Musca Domestica. Journal of Cell Biology, 40(1):269–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
1/269.

Prescott:1966:STM

[Pre66] David M. Prescott. The syntheses of total macronuclear protein, histone, and DNA during the cell cycle in euplotes eurystomus. Journal of Cell Biology, 31(1):1–??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/31/1/1.

Pasquali-Ronchetti:1969:NDM

[PRGC69] Ivonne Pasquali-Ronchetti, John W. Greenawalt, and Ernesto Carafoli. On the nature of the dense matrix granules of normal mitochondria. Journal of Cell Biology, 40(2):565–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
2/565.

Phillips:1965:CFS

[PS65a] David M. Phillips and Hewson Swift. Cytoplasmic fine structure of sciara salivary glands. Journal of Cell Biology, 27(2): 395–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/2/395.

Prensky:1965:MAI

[PS65b] Wolf Prensky and Harold H. Smith. The mechanism of 5aminouracil-induced synchrony of cell division in Vicia Faba root meristems. Journal of Cell Biology, 24(3):401–??, March 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/24/3/401.

Pfuderer:1966:CIP

[PS66] Peter Pfuderer and D. C. Swartzendruber. The configuration of isolated polysomes. Journal of Cell Biology, 30(1):193–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/30/
1/193.

Palay:1968:AHI

[PSPO68] Sanford L. Palay, Constantino Sotelo, Alan Peters, and Paula M. Orkand. The axon hillock and the initial segment. Journal of Cell Biology, 38(1):193–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/38/1/193.

Peters:1967:MFA

[PV67] Alan Peters and James E. Vaughn. Microtubules and filaments in the axons and astrocytes of early postnatal rat optic nerves. Journal of Cell Biology, 32(1):113–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/32/1/113.

Pelc:1969:IDL

[PVM69] S. R. Pelc and Maria Pia Viola-Magni. III. Decrease of labeled DNA in cells of the adrenal medulla after intermittent exposure to cold. Journal of Cell Biology, 42(2):460–??, August 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/42/2/460.

Pieri:1968:ICP

[PVT68] J. Pieri, C. Vaugien, and M. Trouillier. Interpretation cytophotometrique des phenomenes micronucleaires Au cours De la division binaire et des divisions pregamiques chez paramecium trichium. Journal of Cell Biology, 36(3):664–??, March 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/36/3/664.

Pastan:1967:CDF

[PW67] I. Pastan and S. H. Wollman. Colloid droplet formation in dog thyroid in vitro. Journal of Cell Biology, 35(1):262–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/35/
1/262.

Packer:1968:EIM

[PWFP68] L. Packer, J. M. Wrigglesworth, P. A. G. Fortes, and B. C. Pressman. Expansion of the inner membrane compartment and its relation to mitochondrial volume and ion transport. Journal of Cell Biology, 39(2):382–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/2/382.

Reichle:1965:MSW

[RA65] R. E. Reichle and J. V. Alexander. Multiperforate septations, woronin bodies, and septal plugs in fusarium. Journal of Cell Biology, 24(3):489–??, March 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/24/3/489.

Rajalakshmi:1969:ICL

[RAH69] Srinivasan Rajalakshmi, W. Robert Adams, and Robert E. Handschumacher. Isolation and characterization of low density structures from orotic acid-induced fatty livers. Journal of Cell Biology, 41(2):625–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/2/625.

Ramsey:1965:FSS

[Ram65] Helen J. Ramsey. Fine structure of the surface of the cerebral cortex of human brain. Journal of Cell Biology, 26(2):323–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/26/
2/323.

Ramus:1969:DPC

[Ram69] J. Ramus. Dimorphic pit connections in the red alga pseudogloiophloea. Journal of Cell Biology, 41(1):340–??, April 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/41/1/340.

Ray:1967:RSC

[Ray67] Peter M. Ray. Radioautographic study of cell wall deposition in growing plant cells. Journal of Cell Biology, 35(3):659–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/35/
3/659.

Ross:1965:WHC

[RB65] Russell Ross and Earl P. Benditt. Wound healing and collagen formation. Journal of Cell Biology, 27(1):83–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/27/1/83.

Raff:1966:EAT

[RB66] Elizabeth C. Raff and J. J. Blum. The effects of adenosine triphosphate and related compounds on some hydrodynamic properties of glycerinated cilia. Journal of Cell Biology, 31(3): 445–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/31/3/445.

Robinson:1967:LAA

[RB67] Peter M. Robinson and Christopher Bell. The localization of acetylcholinesterase at the autonomic neuromuscular junction. Journal of Cell Biology, 33(1):93–??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/33/1/93.

Raff:1969:SPM

[RB69a] E. C. Raff and J. J. Blum. Some properties of a model assay for ciliary contractility. Journal of Cell Biology, 42(3): 831–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/3/831.

Rogers:1969:QSEb

[RB69b] A. W. Rogers and E. A. Barnard. Quantitative studies on enzymes in structures in striated muscles by labeled inhibitor methods. Journal of Cell Biology, 41(3):686–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/41/3/686.

Ross:1969:EF

[RB69c] Russell Ross and Paul Bornstein. The elastic fiber. Journal of Cell Biology, 40(2):366–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/40/2/366.

Ruhenstroth-Bauer:1968:TPS

[RBLH68] G. Ruhenstroth-Bauer and Christine Lücke-Huhle. Two populations of small lymphocytes. Journal of Cell Biology, 37(1):196–??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/37/
1/196.

Reddy:1969:MEA

[RBS69] Janardan Reddy, Sukhum Bunyaratvej, and Donald Svoboda. Microbodies in experimentally altered cells. Journal of Cell Biology, 42(2):587–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/42/2/587.

Regan:1966:CMA

[RC66] J. D. Regan and E. H. Y. Chu. A convenient method for assay of DNA synthesis in synchronized human cell cultures. Journal of Cell Biology, 28(1):139–??, January 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/28/1/139.

Reger:1967:CSF

[RC67a] J. F. Reger and D. P. Cooper. A comparative study on the fine structure of the basalar muscle of the wing and the tibial extensor muscle of the leg of the lepidopteran achalarus lyciades. Journal of Cell Biology, 33(3):531–??, June 1967. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/33/3/531.

Rosenbaum:1967:FRP

[RC67b] Joel L. Rosenbaum and F. M. Child. Flagellar regeneration in protozoan flagellates. Journal of Cell Biology, 34(1):345–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
1/345.

Reissig:1968:LGC

[RC68] Magdalena Reissig and Anthony V. Colucci. Localization of glycogen in the cestode, hymenolepis diminuta. Journal of Cell Biology, 39(3):754–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/39/3/754.

Rosenbaum:1969:CRT

[RC69] Joel L. Rosenbaum and Kathryn Carlson. Cilia regeneration in Tetrahymena and its inhibition by colchicine. Journal of Cell Biology, 40(2):415–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/40/2/415.

Rifkind:1969:USE

[RCE69] Richard A. Rifkind, David Chui, and Hazel Epler. An ultrastructural study of early morphogenetic events during the establishment of fetal hepatic erythropoiesis. Journal of Cell Biology, 40 (2):343–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/2/343.

Rude:1969:CUI

[RCV69] Sonia Rude, Richard E. Coggeshall, and Lucas S. Van Orden. Chemical and ultrastructural identification of 5hydroxytryptamine in an identified neuron. Journal of Cell Biology, 41(3):832–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/3/832.

Robison:1969:DIC

[RD69] William L. Robison and David Davis. Determination of iodine concentration and distribution in rat thyroid follicles by electronprobe microanalysis. Journal of Cell Biology, 43(1):115–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/43/
1/115.

Rogers:1969:QSEa

[RDS^{+}69] A. W. Rogers, Z. Darzynkiewicz, M. M. Salpeter, K. Ostrowski, and E. A. Barnard. Quantitative studies on enzymes in structures in striated muscles by labeled inhibitor methods. Journal of Cell Biology, 41(3):665–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/41/3/665.

Rooney:1969:EDS

[RE69] David W. Rooney and John J. Eiler. Effects of divisionsynchronizing hypoxic and hyperthermic shocks upon Tetrahymena respiration and intracellular ATP concentration. Journal of Cell Biology, 41(1):145–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/41/1/145.

Reedy:1965:SSE

[Ree65a] Michael K. Reedy. Section staining for electron microscopy. Journal of Cell Biology, 26(1):309–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/26/1/309.

Reese:1965:OCF

[Ree65b] T. S. Reese. Olfactory cilia in the frog. Journal of Cell Biology, 25(2):209–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/2/209.

Reiss:1966:SLM

[Rei66] Oscar K. Reiss. Studies of lung metabolism. Journal of Cell Biology, 30(1):45–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/30/1/45.

Reynolds:1965:LPC

[Rey65] Edward S. Reynolds. Liver parenchymal cell injury. Journal of Cell Biology, 25(3):53–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/25/3/53.

Reznik:1969:TUS

[Rez69] M. Reznik. Thymidine-^{3}H uptake by satellite cells of regenerating skeletal muscle. Journal of Cell Biology, 40(2):568–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
2/568.

Rambourg:1969:DCC

[RHL69] A. Rambourg, W. Hernandez, and C. P. Leblond. Detection of complex carbohydrates in the Golgi apparatus of rat cells. Journal of Cell Biology, 40(2):395–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/40/2/395.

Roth:1967:DDF

[RI67] Thomas F. Roth and Michio Ito. DNA-dependent formation of the synaptinemal complex at meiotic prophase. Journal of Cell Biology, 35(1):247–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/35/1/247.

Rickson:1968:NCT

[Ric68] Fred R. Rickson. Nuclear and cytoplasmic tubules in cortical cells of leaf beltian bodies. Journal of Cell Biology, 38(2):471–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
2/471.

Ringo:1967:FMF

[Rin67] David L. Ringo. Flagellar motion and fine structure of the flagellar apparatus in chlamydomonas. Journal of Cell Biology, 33(3):543–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/33/3/543.

Rizki:1967:USI

[Riz67] T. M. Rizki. Ultrastructure of the secretory inclusions of the salivary gland cell in Drosophila. Journal of Cell Biology, 32 (2):531–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/32/2/531.

Robbins:1969:UCM

[RJ69] Elliott Robbins and Gisela Jentzsch. Ultrastructural changes in the mitotic apparatus at the metaphase-to-anaphase transition. Journal of Cell Biology, 40(3):678–??, March 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/40/3/678.

Robbins:1968:CCS

[RJM68] Elliott Robbins, Gisela Jentzsch, and Anita Micali. The centriole cycle in synchronized HeLa cells. Journal of Cell Biology, 36 (2):329–??, February 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/36/2/329.

Reese:1967:FSL

[RK67a] T. S. Reese and Morris J. Karnovsky. Fine structural localization of a blood-brain barrier to exogenous peroxidase. Journal of Cell Biology, 34(1):207–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/34/1/207.

Revel:1967:HAS

[RK67b] J. P. Revel and M. J. Karnovsky. Hexagonal array of subunits in intercellular junctions of the mouse heart and liver. Journal of Cell Biology, 33(3):??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/33/3/C7.

Ross:1967:FSC

[RK67c] Russell Ross and Seymour J. Klebanoff. Fine structural changes in uterine smooth muscle and fibroblasts in response to estrogen. Journal of Cell Biology, 32(1):155–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/32/1/155.

Rose:1968:CSM

[RKC68] George G. Rose, M. Kumegawa, and M. Cattoni. The circumfusion system for multipurpose culture chambers. Journal of Cell Biology, 39(2):430–??, November 1968. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/2/430.

Rambourg:1967:EMO

[RL67] A. Rambourg and C. P. Leblond. Electron microscope observations on the carbohydrate-rich cell coat present at the surface of cells in the rat. Journal of Cell Biology, 32(1):27–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/32/1/27.

Reimann:1965:SBF

[RLV65] B. E. F. Reimann, Joyce C. Lewin, and B. E. Volcani. Studies on the biochemistry and fine structure of silica Shell formation in diatoms. Journal of Cell Biology, 24(1):39–??, January 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/24/1/39.

Robinow:1966:FAN

[RM66] C. F. Robinow and J. Marak. A fiber apparatus in the nucleus of the yeast cell. Journal of Cell Biology, 29(1):129–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/29/
1/129.

Robbins:1969:OUD

[RM69] Elliott Robbins and Gene A. Morrill. Oxygen uptake during the HeLa cell life cycle and its correlation with macromolecular synthesis. Journal of Cell Biology, 43(3):629–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/43/3/629.

Rosenbaum:1969:FES

[RMR69] Joel L. Rosenbaum, John E. Moulder, and David L. Ringo. Flagellar elongation and shortening in chlamydomonas. Journal of Cell Biology, 41(2):600–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/41/2/600.

Ross:1968:HWR

[RO68] Russell Ross and George Odland. Human wound repair. Journal of Cell Biology, 39(1):152–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/1/152.

Robison:1966:MRM

[Rob66] W. Gerald Robison. Microtubules in relation to the motility of a sperm syncytium in an armored scale insect. Journal of Cell Biology, 29(2):251–??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/29/2/251.

Robb:1969:MID

[Rob69a] James A. Robb. Maintenance of imaginal discs of Drosophila Melanogaster in chemically defined media. Journal of Cell Biology, 41(3):876–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/3/876.

Robinson:1969:CCI

[Rob69b] Peter M. Robinson. A cholinergic component in the innervation of the longitudinal smooth muscle of the guinea pig vas deferens. Journal of Cell Biology, 41(2):462–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/41/2/462.

Rodman:1969:MRI

[Rod69] Toby C. Rodman. Morphology and replication of intranucleolar DNA in polytene nuclei. Journal of Cell Biology, 42(2):575–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
2/575.

Rogers:1968:RPA

[Rog68] M. Elizabeth Rogers. Ribonucleoprotein particles in the amphibian oocyte nucleus. Journal of Cell Biology, 36(3):421–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
3/421.

Rounds:1965:LPT

[ROJ65] Donald E. Rounds, Robert S. Olson, and Fred M. Johnson. The laser as a potential tool for cell research. Journal of Cell Biology, 27(1):191–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/1/191.

Rosenbluth:1965:UOO

[Ros65a] Jack Rosenbluth. Ultrastructural organization of obliquely striated muscle fibers in ascaris lumbricoides. Journal of Cell Biology, 25(3):495–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/3/495.

Rosenbluth:1965:USM

[Ros65b] Jack Rosenbluth. Ultrastructure of somatic muscle cells in ascaris lumbricoides. Journal of Cell Biology, 26(2):579–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/26/
2/579.

Rosenbluth:1966:RMS

[Ros66] Jack Rosenbluth. Redundant myelin sheaths and other ultrastructural features of the toad cerebellum. Journal of Cell Biology, 28(1):73–??, January 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/28/1/73.

Rose:1967:CSM

[Ros67a] George G. Rose. The circumfusion system for multipurpose culture chambers. Journal of Cell Biology, 32(1):89–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/32/1/89.

Rosenbluth:1967:OSM

[Ros67b] Jack Rosenbluth. Obliquely striated muscle. Journal of Cell Biology, 34(1):15–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/34/1/15.

Rosenbluth:1968:OSM

[Ros68] Jack Rosenbluth. Obliquely striated muscle. Journal of Cell Biology, 36(1):245–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/36/1/245.

Rosenbluth:1969:SRU

[Ros69a] Jack Rosenbluth. Sarcoplasmic reticulum of an unusually fastacting crustacean muscle. Journal of Cell Biology, 42(2):534–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
2/534.

Rosenbluth:1969:UDM

[Ros69b] Jack Rosenbluth. Ultrastructure of dyads in muscle fibers of ascaris lumbricoides. Journal of Cell Biology, 42(3):817– ??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/3/817.

Roth:1967:EMM

[Rot67] L. E. Roth. Electron microscopy of mitosis in amebae. Journal of Cell Biology, 34(1):47–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/1/47.

Rasmussen:1966:RSD

[RP66] R. E. Rasmussen and R. B. Painter. Radiation-stimulated DNA synthesis in cultured mammalian cells. Journal of Cell Biology, 29(1):11–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/29/1/11.

Rao:1967:MRS

[RP67] M. V. Narasimha Rao and David M. Prescott. Micronuclear RNA synthesis in paramecium caudatum. Journal of Cell Biology, 33(2):281–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/33/2/281.

Ross:1969:UMP

[RR69] Jean Ross and W. Gerald Robison. Unusual microtubular patterns and three-dimensional movement of mealybug sperm and sperm bundles. Journal of Cell Biology, 40(2):426–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
2/426.

Renaud:1968:SPP

[RRG68] F. L. Renaud, A. J. Rowe, and I. R. Gibbons. Some properties of the protein forming the outer fibers of cilia. Journal of Cell Biology, 36(1):79–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/36/1/79.

Rebhun:1967:UBI

[RS67a] Lionel I. Rebhun and Greta Sander. Ultrastructure and birefringence of the isolated mitotic apparatus of marine eggs. Journal of Cell Biology, 34(3):859–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/3/859.

Robbins:1967:ADG

[RS67b] Elliott Robbins and Matthew D. Scharff. The absence of a detectable G_{1} phase in a cultured strain of Chinese hamster lung cell. Journal of Cell Biology, 34(2):684–??, August 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/34/2/684.

Robbins:1969:SCB

[RS69] Elliott Robbins and Michael Shelanski. Synthesis of a colchicinebinding protein during the HeLa cell life cycle. Journal of Cell Biology, 43(2):371–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/43/2/371.

Rusch:1966:SMF

[RSBG66] H. P. Rusch, W. Sachsenmaier, Kathryn Behrens, and Vera Gruter. Synchronization of mitosis by the fusion of the plasmodia of Physarum Polycephalum. Journal of Cell Biology, 31 (1):204–??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/31/1/204.

Rayns:1969:UDM

[RSL69] D. G. Rayns, F. O. Simpson, and Janet M. Ledingham. Ultrastructure of desmosomes in mammalian intercalated disc; appearances after lanthanum treatment. Journal of Cell Biology, 42(1):322–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/1/322.

Richardson:1969:ECS

[RTL69] U. Ingrid Richardson, Armen H. Tashjian, and Lawrence Levine. Establishment of a clonal strain of hepatoma cells which secrete albumin. Journal of Cell Biology, 40(1):236–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/40/1/236.

Rubin:1969:ECG

[Rub69a] Walter Rubin. Enzyme cytochemistry of gastric parietal cells at a fine structure level. Journal of Cell Biology, 42(1):332–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
1/332.

Rubinow:1969:SMS

[Rub69b] S. I. Rubinow. A simple model of a steady state differentiating cell system. Journal of Cell Biology, 43(1):32–??, October 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/43/1/32.

Rudzinska:1965:FSF

[Rud65] Maria A. Rudzinska. The fine structure and function of the tentacle in Tokophrya Infusionum. Journal of Cell Biology, 25(3):459– ??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
3/459.

Ryser:1967:SPU

[Rys67] Hugues J.-P. Ryser. Studies on protein uptake by isolated tumor cells. Journal of Cell Biology, 32(3):737–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/32/3/737.

Svoboda:1966:RHM

[SA66] Donald J. Svoboda and Daniel L. Azarnoff. Response of hepatic microbodies to a hypolipidemic agent, ethyl chlorophenoxyisobutyrate (cpib). Journal of Cell Biology, 30(2):442–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/30/
2/442.

Salpeter:1967:EMR

[Sal67] Miriam M. Salpeter. Electron microscope radioautography as a quantitative tool in enzyme cytochemistry. Journal of Cell Biology, 32(2):379–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/32/2/379.

Salpeter:1969:EMR

[Sal69] M. M. Salpeter. Electron microscope radioautography as a quantitative tool in enzyme cytochemistry. Journal of Cell Biology, 42(1):122–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/1/122.

Svoboda:1969:MEA

[SAR69] Donald Svoboda, Daniel Azarnoff, and Janardan Reddy. Microbodies in experimentally altered cells. Journal of Cell Biology, 40(3):734–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/3/734.

Satir:1965:SC

[Sat65] Peter Satir. Studies on cilia. Journal of Cell Biology, 26(3): 805–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/3/805.

Satir:1968:SC

[Sat68] Peter Satir. Studies on cilia. Journal of Cell Biology, 39(1):77–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/39/
1/77.

Storb:1966:EMS

[SAW^{+}66] Rainer Storb, Robert L. Amy, Richard K. Wertz, Bernard Fauconnier, and Marcel Bessis. An electron microscope study of vitally stained single cells irradiated with a ruby laser microbeam. Journal of Cell Biology, 31(1):11–??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/31/1/11.

Sedar:1965:DSD

[SB65a] Albert W. Sedar and Ronald M. Burde. The demonstration of the succinic dehydrogenase system in bacillus subtilis using tetranitro-blue tetrazolium combined with techniques of electron microscopy. Journal of Cell Biology, 27(1):53–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/27/1/53.

Sedar:1965:LSD

[SB65b] Albert W. Sedar and Ronald M. Burde. Localization of the succinic dehydrogenase system in Escherichia coli using combined techniques of cytochemistry and electron microscopy. Journal of Cell Biology, 24(2):285–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/24/2/285.

Sommer:1965:CLA

[SB65c] Joachim R. Sommer and Jacob J. Blum. Cytochemical localization of acid phosphatases in euglena gracilis. Journal of Cell Biology, 24(2):235–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/24/2/235.

Stephens:1965:AMF

[SB65d] R. J. Stephens and R. F. Bils. An atypical mitochondrial form in normal rat liver. Journal of Cell Biology, 24(3):500–??, March 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/24/
3/500.

Stubblefield:1966:CFC

[SB66] Elton Stubblefield and B. R. Brinkley. Cilia formation in Chinese hamster fibroblasts in vitro as a response to colcemid treatment. Journal of Cell Biology, 30(3):645–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/30/3/645.

Simard:1967:HSC

[SB67] R. Simard and W. Bernhard. A heat-sensitive cellular function located in the nucleolus. Journal of Cell Biology, 34(1):61–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/34/1/61.

Staley:1968:UFV

[SB68a] Nancy A. Staley and Ellis S. Benson. The ultrastructure of frog ventricular cardiac muscle and its relationship to mechanisms of excitation-contraction coupling. Journal of Cell Biology, 38(1):99–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/1/99.

Sweeny:1968:EMS

[SB68b] P. R. Sweeny and R. Bather. An electron microscopic study of the chorioallantoic membrane following infection with rous sarcoma virus. Journal of Cell Biology, 36(2):299–??, February 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/36/2/299.

Schick:1969:IPR

[SB69a] Lloyd Schick and Larry G. Butler. Inorganic pyrophosphatase of rat liver mitochondria. Journal of Cell Biology, 42(1):235–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/42/
1/235.

Sternlieb:1969:ODS

[SB69b] Irmin Sternlieb and Jacob E. Berger. Optical diffraction studies of crystalline structures in electron micrographs. Journal of Cell Biology, 43(3):448–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/43/3/448.

Susi:1967:FSF

[SBK67] F. R. Susi, W. D. Belt, and J. W. Kelly. Fine structure of fibrillar complexes associated with the basement membrane in human oral mucosa. Journal of Cell Biology, 34(2):686–??, August 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/34/2/686.

Salpeter:1969:REM

[SBS69] M. M. Salpeter, L. Bachmann, and E. E. Salpeter. Resolution in electron microscope radioautography. Journal of Cell Biology, 41(1):1–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/1/1.

Shirahama:1967:HRE

[SC67a] Tsuranobu Shirahama and Alan S. Cohen. High-resolution electron microscopic analysis of the amyloid fibril. Journal of Cell Biology, 33(3):679–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/33/3/679.

Shirahama:1967:RAF

[SC67b] Tsuranobu Shirahama and Alan S. Cohen. Reconstitution of amyloid fibrils from alkaline extracts. Journal of Cell Biology, 35 (2):459–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/2/459.

Schultz:1968:MLA

[SC68] Robert L. Schultz and Norman M. Case. Microtubule loss with acrolein and bicarbonate-containing fixatives. Journal of Cell Biology, 38(3):633–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/38/3/633.

Schneider:1969:ETD

[Sch69] Imogene Schneider. Establishment of three diploid cell lines of anopheles stephensi (diptera: Culicidae). Journal of Cell Biology, 42(2):603–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/2/603.

Scott:1967:TEM

[Sco67] Bronnetta L. Scott. Thymidine-^{3}H electron microscope radioautography of osteogenic cells in the fetal rat. Journal of Cell Biology, 35(1):115–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/35/1/115.

Showacre:1967:NNR

[SCP67] Jane L. Showacre, W. G. Cooper, and D. M. Prescott. Nucleolar and nuclear RNA synthesis during the cell life cycle in monkey and pig kidney cells in vitro. Journal of Cell Biology, 33(2):273– ??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
2/273.

Skutelsky:1967:EMS

[SD67] Ehud Skutelsky and David Danon. An electron microscopic study of nuclear elimination from the late erythroblast. Journal of Cell Biology, 33(3):625–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/33/3/625.

Silverstein:1968:PRR

[SD68a] Samuel C. Silverstein and Samuel Dales. The penetration of reovirus RNA and initiation of its genetic function in l-strain fibroblasts. Journal of Cell Biology, 36(1):197–??, January 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/36/1/197.

Smith:1968:UCM

[SD68b] R. D. Smith and F. Deinhardt. Unique cytoplasmic membranes in rous sarcoma virus-induced tumors of a subhuman primate. Journal of Cell Biology, 37(3):819–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/37/3/819.

Scallen:1969:QRC

[SD69a] Terence J. Scallen and Scott E. Dietert. The quantitative retention of cholesterol in mouse liver prepared for electron microscopy by fixation in a digitonin-containing aldehyde solution. Journal of Cell Biology, 40(3):802–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/40/3/802.

Skutelsky:1969:RSC

[SD69b] E. Skutelsky and D. Danon. Reduction in surface charge as an explanation of the recognition by macrophages of nuclei expelled from normoblasts. Journal of Cell Biology, 43(1):8–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/43/1/8.

Sabatini:1968:ATA

[SDV68] María Teresa Sabatini, Reinaldo Dipolo, and Raimundo Villegas. Adenosine triphosphatase activity in the membranes of the squid nerve fiber. Journal of Cell Biology, 38(1):176–??, July 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/38/1/176.

Schlegel:1967:IRL

[SdZ67] David E. Schlegel and Gustaaf A. de Zoeten. Improved resolution in light microscope radioautography. Journal of Cell Biology, 33(3):728–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/33/3/728.

Studzinski:1966:RBR

[SE66] George P. Studzinski and Kay A. O. Ellem. Relationship between RNA synthesis, cell division, and morphology of mammalian cells. Journal of Cell Biology, 29(3):411–??, June 1966. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/29/3/411.

Slagel:1967:RMC

[SE67] D. E. Slagel and J.-E. Edström. A rapid method for calculation of the total amount of absorbing substance in microextracts. Journal of Cell Biology, 34(1):395–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/1/395.

Stoeckenius:1969:CMS

[SE69] Walther Stoeckenius and Donald M. Engelman. Current models for the structure of biological membranes. Journal of Cell Biology, 42(3):613–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/42/3/613.

Sedar:1969:UPS

[Sed69] Albert W. Sedar. Uptake of peroxidase into the smooth-surfaced tubular system of the gastric acid-secreting cell. Journal of Cell Biology, 43(1):179–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/43/1/179.

Seed:1966:SDRa

[See66a] John Seed. The synthesis of DNA, RNA, and nuclear protein in normal and tumor strain cells. Journal of Cell Biology, 28 (2):233–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/28/2/233.

Seed:1966:SDRb

[See66b] John Seed. The synthesis of DNA, RNA, and nuclear protein in normal and tumor strain cells. Journal of Cell Biology, 28 (2):249–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/28/2/249.

Seed:1966:SDRc

[See66c] John Seed. The synthesis of DNA, RNA, and nuclear protein in normal and tumor strain cells. Journal of Cell Biology, 28 (2):257–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/28/2/257.

Seed:1966:SDRd

[See66d] John Seed. The synthesis of DNA, RNA, and nuclear protein in normal and tumor strain cells. Journal of Cell Biology, 28 (2):263–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/28/2/263.

Seeman:1967:THE

[See67] Philip Seeman. Transient holes in the erythrocyte membrane during hypotonic hemolysis and stable holes in the membrane after lysis by saponin and lysolecithin. Journal of Cell Biology, 32(1):55–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/32/1/55.

Schnaitman:1967:SLM

[SEG67] Carl Schnaitman, V. Gene Erwin, and John W. Greenawalt. The submitochondrial localization of monoamine oxidase. Journal of Cell Biology, 32(3):719–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/32/3/719.

Schaffner:1965:CHS

[SF65] Fenton Schaffner and Philip Felig. Changes in hepatic structure in rats produced by breathing pure oxygen. Journal of Cell Biology, 27(3):505–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/27/3/505.

Smith:1966:LFR

[SF66] Robert E. Smith and Marilyn G. Farquhar. Lysosome function in the regulation of the secretory process in cells of the anterior pituitary gland. Journal of Cell Biology, 31(2):319–??, November 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/31/
2/319.

Scheer:1969:NSA

[SF69a] Ulrich Scheer and Werner W. Franke. Negative staining and adenosine triphosphatase activity of annulate lamellae of newt oocytes. Journal of Cell Biology, 42(2):519–??, August 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/42/2/519.

Shinozuka:1969:RNA

[SF69b] Hisashi Shinozuka and Emmanuel Farber. Reformation of nucleoli after ethione-induced fragmentation in the absence of significant protein synthesis. Journal of Cell Biology, 41(1):280–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/41/
1/280.

Silverman:1967:MIV

[SFG67] Lloyd Silverman, Laurence H. Frommhagen, and David Glick. Measurement of influenza virus-antibody reaction by quantitative electron microscopy. Journal of Cell Biology, 35(1):61–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/35/1/61.

Shimada:1967:FSE

[SFM67] Y. Shimada, D. A. Fischman, and A. A. Moscona. The fine structure of embryonic chick skeletal muscle cells differentiated in vitro. Journal of Cell Biology, 35(2):445–??, November 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/35/2/445.

Shimada:1969:DNM

[SFM69] Yutaka Shimada, D. A. Fischman, and A. A. Moscona. The development of nerve-muscle junctions in monolayer cultures of embryonic spinal cord and skeletal muscle cells. Journal of Cell Biology, 43(2):382–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/43/2/382.

Schnaitman:1968:EPI

[SG68] Carl Schnaitman and John W. Greenawalt. Enzymatic properties of the inner and outer membranes of rat liver mitochondria. Journal of Cell Biology, 38(1):158–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/38/1/158.

Silverman:1969:MPC

[SG69a] Lloyd Silverman and David Glick. Measurement of protein concentration by quantitative electron microscopy. Journal of Cell Biology, 40(3):773–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/40/3/773.

Silverman:1969:RST

[SG69b] Lloyd Silverman and David Glick. The reactivity and staining of tissue proteins with phosphotungstic acid. Journal of Cell Biology, 40(3):761–??, March 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/40/3/761.

Svoboda:1967:MEA

[SGA67] Donald Svoboda, Harold Grady, and Daniel Azarnoff. Microbodies in experimentally altered cells. Journal of Cell Biology, 35(1):127–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/1/127.

Shinozuka:1968:DHC

[SGF68] Hisashi Shinozuka, Peter J. Goldblatt, and Emmanuel Farber. The disorganization of hepatic cell nucleoli induced by ethionine and its reversal by adenine. Journal of Cell Biology, 36(2):313–??, February 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
2/313.

Stromer:1967:MRS

[SGR67] Marvin H. Stromer, D. E. Goll, and L. E. Roth. Morphology of rigor-shortened bovine muscle and the effect of trypsin on preand postrigor myofibrils. Journal of Cell Biology, 34(2):431–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
2/431.

Swartzendruber:1965:EMA

[SH65] D. C. Swartzendruber and M. G. Hanna. Electron microscopic autoradiography of germinal center cells in mouse spleen. Journal of Cell Biology, 25(3):109–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/25/3/109.

Sommer:1967:EGF

[SH67a] J. R. Sommer and W. Hasselbach. The effect of glutaraldehyde and formaldehyde on the calcium pump of the sarcoplasmic reticulum. Journal of Cell Biology, 34(3):902–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/34/3/902.

Sudweeks:1967:CLC

[SH67b] Alan D. Sudweeks and Rolla B. Hill. Control of liver cell replication by albumin need. Journal of Cell Biology, 34(1):404–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
1/404.

Sadowski:1968:ITD

[SH68a] Paul D. Sadowski and Janet Alcock Howden. Isolation of two distinct classes of polysomes from a nuclear fraction of rat liver. Journal of Cell Biology, 37(1):163–??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/37/1/163.

Stanley:1968:MES

[SH68b] D. W. Stanley and H. O. Hultin. Mechanism of emptying of skeletal muscle cell segments. Journal of Cell Biology, 39(1):??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/39/
1/C1.

Sheridan:1965:FSE

[She65] Michael N. Sheridan. The fine structure of the electric organ of torpedo marmorata. Journal of Cell Biology, 24(1):129–??, January 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/24/
1/129.

Sheridan:1966:ESS

[She66] Judson D. Sheridan. Electrophysiological study of special connections between cells in the early chick embryo. Journal of Cell Biology, 31(1):??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/31/1/C1.

Shen:1967:MAN

[She67] Eugene Y. F. Shen. Microspectrophotometric analysis of nuclear DNA in chara zeylanica. Journal of Cell Biology, 35(2): 377–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/2/377.

Sheridan:1968:EEL

[She68] Judson D. Sheridan. Electrophysiological evidence for lowresistance intercellular junctions in the early chick embryo. Journal of Cell Biology, 37(3):650–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/37/3/650.

Spicer:1968:NPP

[SHG68] S. S. Spicer, J. H. Hardin, and W. B. Greene. Nuclear precipitates in pyroantimonate-osmium tetroxide-fixed tissues. Journal of Cell Biology, 39(1):216–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/1/216.

Stromer:1969:EVP

[SHMR69] Marvin H. Stromer, D. J. Hartshorne, Helmut Mueller, and Robert V. Rice. The effect of various protein fractions on Zand M-line reconstitution. Journal of Cell Biology, 40(1):167–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/40/
1/167.

Shoup:1966:DPG

[Sho66] Jane Rearick Shoup. The development of pigment granules in the eyes of wild type and mutant Drosophila Melanogaster. Journal of Cell Biology, 29(2):223–??, May 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/29/2/223.

Shoup:1967:SWT

[Sho67] Jane Rearick Shoup. Spermiogenesis in wild type and in a male sterility mutant of Drosophila Melanogaster. Journal of Cell Biology, 32(3):663–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/32/3/663.

Stromer:1967:RRZ

[SHR67] Marvin H. Stromer, D. J. Hartshorne, and R. V. Rice. Removal and reconstitution of z-line material in a striated muscle. Journal of Cell Biology, 35(3):??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/35/3/C23.

Staubli:1969:CMB

[SHW69] Willy Stäubli, Robert Hess, and Ewald R. Weibel. Correlated morphometric and biochemical studies on the liver cell. Journal of Cell Biology, 42(1):92–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/42/1/92.

Silagi:1969:CPP

[Sil69] Selma Silagi. Control of pigment production in mouse melanoma cells in vitro. Journal of Cell Biology, 43(2):263–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/43/
2/263.

Simard:1967:BAD

[Sim67] René Simard. The binding of actinomycin D-^{3}H to heterochromatin as studied by quantitative high resolution radioautography. Journal of Cell Biology, 35(3):716–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/35/3/716.

Simpson:1968:EAK

[Sim68] Larry Simpson. Effect of acriflavin on the kinetoplast of Leishmania tarentolae. Journal of Cell Biology, 37(3):660–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/37/3/660.

Sommer:1968:CM

[SJ68a] Joachim R. Sommer and Edward A. Johnson. Cardiac muscle. Journal of Cell Biology, 36(3):497–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/36/3/497.

Sommer:1968:PFH

[SJ68b] Joachim R. Sommer and Edward A. Johnson. Purkinje fibers of the heart examined with the peroxidase reaction. Journal of Cell Biology, 37(2):570–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/37/2/570.

Simpson:1967:MDC

[SK67a] Charles F. Simpson and J. M. Kling. The mechanism of denucleation in circulating erythroblasts. Journal of Cell Biology, 35(1):237–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/1/237.

Stirling:1967:HRRa

[SK67b] Charles E. Stirling and William B. Kinter. High-resolution radioautography of galactose-^{3}H accumulation in rings of hamster intestine. Journal of Cell Biology, 35(3):585–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/35/3/585.

Siew:1968:ESC

[SK68a] D. Siew and S. Klein. The effect of sodium chloride on some metabolic and fine structural changes during the greening of etiolated leaves. Journal of Cell Biology, 37(2):590–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/37/2/590.

Simpson:1968:MME

[SK68b] Charles F. Simpson and J. M. Kling. The mechanism of mitochondrial extrusion from phenylhydrazine-induced reticulocytes in the circulating blood. Journal of Cell Biology, 36(1):103–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
1/103.

Stoeckenius:1968:FCP

[SK68c] Walther Stoeckenius and Wolf H. Kunau. Further characterization of particulate fractions from lysed cell envelopes of halobacterium halobium and isolation of gas vacuole membranes. Journal of Cell Biology, 38(2):337–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/38/2/337.

Sottocasa:1967:ETS

[SKEB67] Gian Luigi Sottocasa, Bo Kuylenstierna, Lars Ernster, and Anders Bergstrand. An electron-transport system associated with the outer membrane of liver mitochondria. Journal of Cell Biology, 32(2):415–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/32/2/415.

Slayter:1968:EMS

[SKHR68] Henry Slayter, Yukio Kiho, Cecil E. Hall, and Alexander Rich. An electron microscopic study of large bacterial polyribosomes. Journal of Cell Biology, 37(2):583–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/37/2/583.

Schneeberger-Keeley:1968:UBA

[SKK68] Eveline E. Schneeberger-Keeley and Morris J. Karnovsky. The ultrastructural basis of alveolar-capillary membrane permeability to peroxidase used as a tracer. Journal of Cell Biology, 37(3):781– ??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/37/
3/781.

Simpson:1965:NBP

[SKN65] Charles F. Simpson, J. M. Kling, and F. C. Neal. The nature of bands in parasitized bovine erythrocytes. Journal of Cell Biology, 27(1):225–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/1/225.

Seligman:1968:NUD

[SKWH68] Arnold M. Seligman, Morris J. Karnovsky, Hannah L. Wasserkrug, and Jacob S. Hanker. Nondroplet ultrastructural demonstration of cytochrome oxidase activity with a polymerizing osmiophilic reagent, diaminobenzidine (dab). Journal of Cell Biology, 38(1):1–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/1/1.

Shah:1966:DDR

[SL66] Vinod C. Shah and Harvard Lyman. DNA-dependent RNA synthesis in chloroplasts of euglena gracilis. Journal of Cell Biology, 29(1):174–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/29/1/174.

Slautterback:1965:MCM

[Sla65] David B. Slautterback. Mitochondria in cardiac muscle cells of the canary and some other birds. Journal of Cell Biology, 24 (1):1–??, January 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/1/1.

Stoltze:1969:IMN

[SLAR69] H. J. Stoltze, N. S. T. Lui, O. R. Anderson, and O. A. Roels. The influence of the mode of nutrition on the digestive system of ochromonas malhamensis. Journal of Cell Biology, 43 (3):396–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/43/3/396.

Seraydarian:1965:DGS

[SM65a] K. Seraydarian and W. F. H. M. Mommaerts. Density gradient separation of sarcotubular vesicles and other particulate constituents of rabbit muscle. Journal of Cell Biology, 26(2):641–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/26/
2/641.

Sisken:1965:IKM

[SM65b] Jesse E. Sisken and Luciano Morasca. Intrapopulation kinetics of the mitotic cycle. Journal of Cell Biology, 25(2):179–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
2/179.

Suter:1965:PLA

[SM65c] Elsi R. Suter and Guido Majno. Passage of lipid across vascular endothelium in newborn rats. Journal of Cell Biology, 27 (1):163–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/1/163.

Suzuki:1967:IFB

[SM67] Teruo Suzuki and F. K. Mostofi. Intramitochondrial filamentous bodies in the thick limb of Henle of the rat kidney. Journal of Cell Biology, 33(3):605–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/33/3/605.

Schiaffino:1969:CDS

[SM69] S. Schiaffino and A. Margreth. Coordinated development of the sarcoplasmic reticulum and T system during postnatal differentiation of rat skeletal muscle. Journal of Cell Biology, 41(3):855–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/41/
3/855.

Smith:1965:OFM

[Smi65] David S. Smith. The organization of flight muscle in an aphid, Megoura Viciae (Homoptera). Journal of Cell Biology, 27(2): 379–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/27/2/379.

Smith:1966:OFM

[Smi66a] David S. Smith. The organization of flight muscle fibers in the odonata. Journal of Cell Biology, 28(1):109–??, January 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/28/1/109.

Smith:1966:SIM

[Smi66b] David S. Smith. The structure of intersegmental muscle fibers in an insect, Periplaneta Americana L. Journal of Cell Biology, 29(3):449–??, June 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/29/3/449.

Stone:1965:HTD

[SMP65] G. E. Stone, O. L. Miller, and D. M. Prescott. H^{3}-thymidine derivative pools in relation to macronuclear DNA synthesis in Tetrahymena Pyriformis. Journal of Cell Biology, 25(2):171–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
2/171.

Schoenberg:1967:DFE

[SMW67] Melvin D. Schoenberg, Richard D. Moore, and Austin S. Weisberger. Differentiation and functional expression of potential antibody-producing cells in the presence of chloramphenicol. Journal of Cell Biology, 32(2):401–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/32/2/401.

Samorajski:1965:FSL

[SOK65] T. Samorajski, J. M. Ordy, and J. R. Keefe. The fine structure of lipofuscin age pigment in the nervous system of aged mice. Journal of Cell Biology, 26(3):779–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/26/3/779.

Samorajski:1966:SOR

[SOK66] T. Samorajski, J. M. Ordy, and J. R. Keefe. Structural organization of the retina in the tree shrew (Tupaia Glis). Journal of Cell Biology, 28(3):489–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/28/3/489.

Sommer:1965:UPC

[Som65] Joachim R. Sommer. The ultrastructure of the pellicle complex of euglena gracilis. Journal of Cell Biology, 24(2):253–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/24/
2/253.

Samorajski:1968:ELT

[SOZC68] T. Samorajski, J. M. Ordy, W. Zeman, and H. J. Curtis. Early and long-term effects of deuteron irradiation on cells in the cerebellum of adult mice. Journal of Cell Biology, 38(1):213–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
1/213.

Siekevitz:1966:DNS

[SP66] P. Siekevitz and G. E. Palade. Distribution of newly synthesized amylase in microsomal subfractions of guinea pig pancreas. Journal of Cell Biology, 30(3):519–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/30/3/519.

Scheving:1967:CPR

[SP67a] Lawrence Einar Scheving and John Edward Pauly. Circadian phase relationships of thymidine-^{3}H uptake, labeled nuclei, grain counts, and cell division rate in rat corneal epithelium. Journal of Cell Biology, 32(3):677–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/32/3/677.

Seeman:1967:APL

[SP67b] Philip M. Seeman and George E. Palade. Acid phosphatase localization in rabbit eosinophils. Journal of Cell Biology, 34(3): 745–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/34/3/745.

Sotelo:1968:FSL

[SP68] Constantino Sotelo and Sanford L. Palay. The fine structure of the lateral vestibular nucleus in the rat. Journal of Cell Biology, 36(1):151–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/36/1/151.

Shibko:1965:SRL

[SPT65] S. Shibko, J. Pangborn, and A. L. Tappel. Studies on the release of lysosomal enzymes from kidney lysosomes. Journal of Cell Biology, 25(3):479–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/25/3/479.

Shoup:1966:TTS

[SPW66] G. D. Shoup, D. M. Prescott, and J. R. Wykes. Thymidine triphosphate synthesis in Tetrahymena. Journal of Cell Biology, 31(2):295–??, November 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/31/2/295.

Seligman:1967:UDC

[SPW^{+}67] Arnold M. Seligman, Robert E. Plapinger, Hannah L. Wasserkrug, Chandicharan Deb, and Jacob S. Hanker. Ultrastructural demonstration of cytochrome oxidase activity by the nadi reaction with osmiophilic reagents. Journal of Cell Biology, 34(3): 787–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/34/3/787.

Smith:1965:DEN

[SR65] Catherine A. Smith and Grant L. Rasmussen. Degeneration in the efferent nerve endings in the cochlea after axonal section. Journal of Cell Biology, 26(1):63–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/26/1/63.

Schneider:1967:DRP

[SR67a] Lawrence K. Schneider and William O. Rieke. DNA replication patterns and chromosomal protein synthesis in opossum lymphocytes in vitro. Journal of Cell Biology, 33(3):497–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/33/3/497.

Stoeckenius:1967:MSH

[SR67b] Walther Stoeckenius and Robert Rowen. A morphological study of halobacterium halobium and its lysis in media of low salt concentration. Journal of Cell Biology, 34(1):365–??, July 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/34/1/365.

Sagebiel:1968:SRC

[SR68] Richard W. Sagebiel and Thomas H. Reed. Serial reconstruction of the characteristic granule of the langerhans cell. Journal of Cell Biology, 36(3):595–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/36/3/595.

Sell:1969:LPC

[SR69] David A. Sell and Edward S. Reynolds. Liver parenchymal cell injury. Journal of Cell Biology, 41(3):736–??, June 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/41/3/736.

Srebro:1965:UGB

[Sre65] Zbigniew Srebro. The ultrastructure of gliosomes in the brains of amphibia. Journal of Cell Biology, 26(2):313–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/26/2/313.

Srivastava:1966:FSC

[Sri66] Lalit M. Srivastava. On the fine structure of the cambium of Fraxinus Americana L. Journal of Cell Biology, 31(1):79–??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/31/1/79.

Song:1969:PMR

[SRRK69] C. S. Song, W. Rubin, A. B. Rifkind, and A. Kappas. Plasma membranes of the rat liver. Journal of Cell Biology, 41(1):124– ??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/41/
1/124.

Stein:1966:UCD

[SRZB66] R. J. Stein, W. R. Richter, R. A. Zussman, and G. Brynjolfsson. Ultrastructural characterization of Daphnia heart muscle. Journal of Cell Biology, 29(1):168–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/29/1/168.

Satir:1965:NAM

[SS65a] P. Satir and A. M. Stuart. A new apical microtubule-associated organelle in the sternal gland of zootermopsis nevadensis (hagen), isoptera. Journal of Cell Biology, 24(2):277–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/24/2/277.

Smith:1965:MCE

[SS65b] Una Smith and David S. Smith. A microtubular complex in the epidermal nucleus of an insect, carausius morosus. Journal of Cell Biology, 26(3):961–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/26/3/961.

Steffensen:1965:IHT

[SS65c] Dale M. Steffensen and William F. Sheridan. Incorporation of H^{3}- thymidine into chloroplast DNA of marine algae. Journal of Cell Biology, 25(3):619–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/25/3/619.

Stehbens:1965:UDB

[SS65d] W. E. Stehbens and M. D. Silver. Unusual development of basement membrane about small blood vessels. Journal of Cell Biology, 26(2):669–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/26/2/669.

Schmidt:1966:DCS

[SS66a] Jean M. Schmidt and R. Y. Stanier. The development of cellular stalks in bacteria. Journal of Cell Biology, 28(3):423–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/28/
3/423.

Stevens:1966:RTN

[SS66b] Barbara J. Stevens and Hewson Swift. RNA transport from nucleus to cytoplasm in chironomus salivary glands. Journal of Cell Biology, 31(1):55–??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/31/1/55.

Stein:1967:LSIb

[SS67a] Olga Stein and Yechezkiel Stein. Lipid synthesis, intracellular transport, and secretion. Journal of Cell Biology, 34(1):251–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
1/251.

Stein:1967:LSIa

[SS67b] Olga Stein and Yechezkiel Stein. Lipid synthesis, intracellular transport, storage, and secretion. Journal of Cell Biology, 33(2):319–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/33/2/319.

Sadowski:1968:EMB

[SS68a] Paul D. Sadowski and Jan W. Steiner. Electron microscopic and biochemical characteristics of nuclei and nucleoli isolated from rat liver. Journal of Cell Biology, 37(1):147–??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/37/1/147.

Stein:1968:LSI

[SS68b] Olga Stein and Yechezkiel Stein. Lipid synthesis, intracellular transport, and storage. Journal of Cell Biology, 36(1):63–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/36/1/63.

Stuart:1968:MFA

[SS68c] A. M. Stuart and P. Satir. Morphological and functional aspects of an insect epidermal gland. Journal of Cell Biology, 36(3):527– ??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/36/
3/527.

Shortman:1969:SDC

[SS69a] Ken Shortman and Kathrin Seligman. The separation of different cell classes from lymphoid organs. Journal of Cell Biology, 42 (3):783–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/3/783.

Stein:1969:LSI

[SS69b] Olga Stein and Yechezkiel Stein. Lecithin synthesis, intracellular transport, and secretion in rat liver. Journal of Cell Biology, 40 (2):461–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/2/461.

Stoner:1969:OIA

[SS69c] Clinton D. Stoner and Howard D. Sirak. Osmotically-induced alterations in volume and ultrastructure of mitochondria isolated from rat liver and bovine heart. Journal of Cell Biology, 43 (3):521–??, December 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/43/3/521.

Silverman:1969:MTW

[SSG69] Lloyd Silverman, Berit Schreiner, and David Glick. Measurement of thickness within sections by quantitative electron microscopy. Journal of Cell Biology, 40(3):768–??, March 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/40/3/768.

Stein:1969:MCC

[SSGF69] Olga Stein, Yechezkiel Stein, Dewitt S. Goodman, and Noel H. Fidge. The metabolism of chylomicron cholesteryl ester in rat liver. Journal of Cell Biology, 43(3):410–??, December 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/43/3/410.

Scott:1968:SCD

[SSS68] N. Steele Scott, Vinod C. Shah, and Robert M. Smillie. Synthesis of chloroplast DNA in isolated chloroplasts. Journal of Cell Biology, 38(1):151–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/38/1/151.

Schafer:1967:AAD

[SSSR67] Irwin A. Schafer, Lloyd Silverman, Julia C. Sullivan, and William Van B. Robertson. Ascorbic acid deficiency in cultured human fibroblasts. Journal of Cell Biology, 34(1):83–??, July 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/34/1/83.

Schenk:1967:CRT

[SSW67] Robert K. Schenk, David Spiro, and Joseph Wiener. Cartilage resorption in the tibial epiphyseal plate of growing rats. Journal of Cell Biology, 34(1):275–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/34/1/275.

Stocking:1968:UCI

[SSWG68] C. R. Stocking, L. K. Shumway, T. E. Weier, and D. Greenwood. Ultrastructure of chloroplasts isolated by nonaqueous extraction. Journal of Cell Biology, 36(1):270–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/36/1/270.

Smith:1965:EML

[ST65] David S. Smith and J. E. Treherne. The electron microscopic localization of cholinesterase activity in the central nervous system of an insect, Periplaneta Americana L. Journal of Cell Biology, 26(2):445–??, August 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/26/2/445.

Shelanski:1967:IPS

[ST67] M. L. Shelanski and E. W. Taylor. Isolation of a protein subunit from microtubules. Journal of Cell Biology, 34(2):549–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
2/549.

Shelanski:1968:PPS

[ST68] Michael L. Shelanski and Edwin W. Taylor. Properties of the protein subunit of central-pair and outer-doublet microtubules of sea urchin flagella. Journal of Cell Biology, 38(2):304–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
2/304.

Stay:1965:PUO

[Sta65] Barbara Stay. Protein uptake in the oocytes of the Cecropia moth. Journal of Cell Biology, 26(1):49–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/26/1/49.

Stephens:1965:AMC

[Ste65a] R. E. Stephens. Analysis of muscle contraction by ultraviolet microbeam disruption of sarcomere structure. Journal of Cell Biology, 25(2):129–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/25/2/129.

Stephens:1965:ACI

[Ste65b] R. E. Stephens. Anomalous contraction of invertebrate striated muscle. Journal of Cell Biology, 27(3):639–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/27/3/639.

Stevens:1965:FSN

[Ste65c] Barbara J. Stevens. The fine structure of the nucleolus during mitosis in the grasshopper neuroblast cell. Journal of Cell Biology, 24(3):349–??, March 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/3/349.

Stephens:1967:MA

[Ste67] R. E. Stephens. The mitotic apparatus. Journal of Cell Biology, 32(2):255–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/32/2/255.

Stirling:1967:HRRb

[Sti67] Charles E. Stirling. High-resolution radioautography of phlorizin-3H in rings of hamster intestine. Journal of Cell Biology, 35 (3):605–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/3/605.

Stine:1968:EAD

[Sti68] G. J. Stine. Enzyme activities during the asexual cycle of neurospora crassa. Journal of Cell Biology, 37(1):81–??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/37/1/81.

Stone:1968:SCD

[Sto68] G. E. Stone. Synchronized cell division in Tetrahymena Pyriformis following inhibition with vinblastine. Journal of Cell Biology, 39(3):556–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/39/3/556.

Stone:1969:CHD

[Sto69] G. E. Stone. Change in Histone/ DNA ratio during population growth in Tetrahymena. Journal of Cell Biology, 42(3): 837–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/3/837.

Stubblefield:1965:QTA

[Stu65] Elton Stubblefield. Quantitative tritium autoradiography of mammalian chromosomes. Journal of Cell Biology, 25(3):137–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
3/137.

Suganuma:1966:FSP

[Sug66] Atsushi Suganuma. Further studies on the plasma membrane of staphylococcus aureus. Journal of Cell Biology, 30(1):208–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/30/
1/208.

Suskind:1965:ACE

[Sus65] R. Gerald Suskind. Autoradiographic and cytochemical evidence for synthesis of a Lysine-containing ribonucleoprotein in nucleoli inhibited by actinomycin D. Journal of Cell Biology, 24 (2):309–??, February 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/2/309.

Suskind:1967:ETA

[Sus67] R. Gerald Suskind. Effect of temperature of aldehyde fixation on the radioautographic localization of ribonucleoprotein in nucleoli of HeLa cells. Journal of Cell Biology, 34(3):721–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
3/721.

Steinert:1967:LKD

[SvA67] M. Steinert and Suzanne van Assel. The loss of kinetoplastic DNA in two species of trypanosomatidae treated with acriflavine. Journal of Cell Biology, 34(2):489–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/34/2/489.

Soeiro:1968:EPI

[SVD68] R. Soeiro, M. H. Vaughan, and J. E. Darnell. The effect of puromycin on intranuclear steps in ribosome biosynthesis. Journal of Cell Biology, 36(1):91–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/36/1/91.

Schaefer:1967:FSM

[SVR67] Carl W. Schaefer, Jerome P. Vanderberg, and Johannes Rhodin. The fine structure of mosquito midgut muscle. Journal of Cell Biology, 34(3):905–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/34/3/905.

Soeiro:1968:TND

[SVWD68] R. Soeiro, M. H. Vaughan, J. R. Warner, and J. E. Darnell. The turnover of nuclear DNA-like RNA in HeLa cells. Journal of Cell Biology, 39(1):112–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/1/112.

Sutton:1966:TMT

[SW66] Jerry S. Sutton and Leon Weiss. Transformation of monocytes in tissue culture into macrophages, epithelioid cells, and multinucleated giant cells. Journal of Cell Biology, 28(2):303–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/28/
2/303.

Sisken:1967:TSC

[SW67] Jesse E. Sisken and Elaina Wilkes. The time of synthesis and the conservation of mitosis-related proteins in cultured human amnion cells. Journal of Cell Biology, 34(1):97–??, July 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/34/1/97.

Sisken:1967:IMA

[SWDK67] Jesse E. Sisken, Elaina Wilkes, Grace M. Donnelly, and Tsuyoshi Kakefuda. The isolation of the mitotic apparatus from mammalian cells in culture. Journal of Cell Biology, 32(1):212–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/32/
1/212.

Seligman:1966:NSM

[SWH66] Arnold M. Seligman, Hannah L. Wasserkrug, and Jacob S. Hanker. A new staining method (oto) for enhancing contrast of lipid-containing membranes and droplets in osmium tetroxidefixed tissue with osmiophilic thiocarbohydrazide (tch). Journal of Cell Biology, 30(2):424–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/30/2/424.

Siev:1969:SIN

[SWP69] Moshe Siev, Robert Weinberg, and Sheldon Penman. The selective interruption of nucleolar RNA synthesis in HeLa cells by cordycepin. Journal of Cell Biology, 41(2):510–??, May 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/41/2/510.

Saunders:1968:LAL

[SWR68] David R. Saunders, Janet Wilson, and Cyrus E. Rubin. Loss of absorbed lipid during fixation and dehydration of jejunal mucosa. Journal of Cell Biology, 37(1):183–??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/37/1/183.

Stein:1968:ALP

[SWS68] Yechezkiel Stein, Christopher Widnell, and Olga Stein. Acylation of lysophosphatides by plasma membrane fractions of rat liver. Journal of Cell Biology, 39(1):185–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/39/1/185.

Szollosi:1965:ENP

[Szo65] Daniel Szollosi. Extrusion of nucleoli from pronuclei of the rat. Journal of Cell Biology, 25(3):545–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/25/3/545.

Takats:1965:NRN

[Tak65] S. T. Takats. Non-random nuclear orientation during DNA synthesis in Tradescantia pollen grains. Journal of Cell Biology, 25(1):151–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/1/151.

Takats:1968:RLD

[Tak68] S. T. Takats. Radioautographic localization of deoxythymidine triphosphate in Tradescantia pollen grains during DNA synthesis. Journal of Cell Biology, 38(3):509–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/38/3/509.

Tanaka:1969:ERL

[Tan69] Yasukazu Tanaka. Effect of the ruby laser microbeam on mitochondria of Kb cells supravitally stained by pinacyanol. Journal of Cell Biology, 41(2):424–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/41/2/424.

Tobey:1967:ETD

[TAP67] R. A. Tobey, E. C. Anderson, and D. F. Petersen. The effect of thymidine on the duration of G_{1} in Chinese hamster cells. Journal of Cell Biology, 35(1):53–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/35/1/53.

Taylor:1965:MCI

[Tay65a] Edwin W. Taylor. The mechanism of colchicine inhibition of mitosis. Journal of Cell Biology, 25(1):145–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/25/1/145.

Taylor:1965:DTL

[Tay65b] J. Herbert Taylor. Distribution of tritium-labeled DNA among chromosomes during meiosis. Journal of Cell Biology, 25(2):57– ??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
2/57.

Taylor:1966:MMC

[Tay66] A. C. Taylor. Microtubules in the microspikes and cortical cytoplasm of isolated cells. Journal of Cell Biology, 28(2):155–??, February 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/28/
2/155.

Tice:1965:DRF

[TB65] Lois Withrow Tice and Russell J. Barrnett. Diazophthalocyanins as reagents for fine structural cytochemistry. Journal of Cell Biology, 25(1):23–??, April 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/25/1/23.

Tilney:1969:SMH

[TB69] Lewis G. Tilney and Breck Byers. Studies on the microtubules in heliozoa. Journal of Cell Biology, 43(1):148–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/43/1/148.

Thorell:1965:MCS

[TCL65] B. Thorell, B. Chance, and V. Legallais. Microspectrophotometry of cytochromes in the single cell at room and liquid nitrogen temperatures. Journal of Cell Biology, 26(3):741–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/26/
3/741.

Thaler:1969:INS

[TCV69] M. M. Thaler, M. C. L. Cox, and C. A. Villee. Isolation of nuclei from sea urchin eggs and embryos. Journal of Cell Biology, 42 (3):846–??, September 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/3/846.

Tooze:1965:CAE

[TD65] J. Tooze and H. G. Davies. Cytolysomes in amphibian erythrocytes. Journal of Cell Biology, 24(1):146–??, January 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/24/1/146.

Tice:1966:CPS

[TE66] Lois W. Tice and A. G. Engel. Cytochemistry of phosphatases of the sarcoplasmic reticulum. Journal of Cell Biology, 31(3): 489–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/31/3/489.

Tedeschi:1965:SOP

[Ted65] Henry Tedeschi. Some observations on the permeability of mitochondria to sucrose. Journal of Cell Biology, 25(2):229–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
2/229.

Terzakis:1965:NCS

[Ter65] John A. Terzakis. The nucleolar channel system of human endometrium. Journal of Cell Biology, 27(2):293–??, November 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/27/
2/293.

Terzakis:1967:SEB

[Ter67] John A. Terzakis. Substructure in an epithelial basal lamina (basement membrane). Journal of Cell Biology, 35(1):273–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/35/
1/273.

Tandler:1969:RMH

[TESW69] Bernard Tandler, Robert A. Erlandson, Archie L. Smith, and Ernest L. Wynder. Riboflavin and mouse hepatic cell structure and function. Journal of Cell Biology, 41(2):477–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/41/2/477.

Thomas:1968:MEM

[TG68] Richard S. Thomas and John W. Greenawalt. Microincineration, electron microscopy, and electron diffraction of calcium phosphate-loaded mitochondria. Journal of Cell Biology, 39 (1):55–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/39/1/55.

Tilney:1969:MFD

[TG69] Lewis G. Tilney and John R. Gibbins. Microtubules in the formation and development of the primary mesenchyme in Arbacia Punctulata. Journal of Cell Biology, 41(1):227–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/41/1/227.

Trench:1969:CFO

[TGB69] Robert K. Trench, Richard W. Greene, and Barbara G. Bystrom. Chloroplasts as functional organelles in animal tissues. Journal of Cell Biology, 42(2):404–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/42/2/404.

Thaemert:1966:UIN

[Tha66a] J. C. Thaemert. Ultrastructural interrelationships of nerve processes and smooth muscle cells in three dimensions. Journal of Cell Biology, 28(1):37–??, January 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/28/1/37.

Thaemert:1966:UCM

[Tha66b] J. C. Thaemert. Ultrastructure of cardiac muscle and nerve contiguities. Journal of Cell Biology, 29(1):156–??, April 1966. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/29/1/156.

Tilney:1966:SMH

[THM66] Lewis G. Tilney, Yukio Hiramoto, and Douglas Marsland. Studies on the microtubules in heliozoa. Journal of Cell Biology, 29(1):77–??, April 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/29/1/77.

Thomson:1969:UBC

[THNB69] John F. Thomson, Duane A. Habeck, Sharron L. Nance, and Karen L. Beetham. Ultrastructural and biochemical changes in Brown fat in cold-exposed rats. Journal of Cell Biology, 41 (1):312–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/1/312.

Tesar:1969:HSG

[TKH69] Joseph T. Tesar, Harold Koenig, and Charles Hughes. Hormone storage granules in the beef anterior pituitary. Journal of Cell Biology, 40(1):225–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/40/1/225.

Trinkaus:1967:SSF

[TL67] J. P. Trinkaus and Thomas L. Lentz. Surface specializations of fundulus cells and their relation to cell movements during gastrulation. Journal of Cell Biology, 32(1):139–??, January 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/32/1/139.

Tokumitsu:1967:CPI

[TM67] Takashi Tokumitsu and Karl Maramorosch. Cytoplasmic protrusions in insect cells during mitosis in vitro. Journal of Cell Biology, 34(2):677–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/34/2/677.

Tilney:1969:FSA

[TM69] Lewis G. Tilney and Douglas Marsland. A fine structural analysis of cleavage induction and furrowing in the eggs of Arbacia Punctulata. Journal of Cell Biology, 42(1):170–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/42/1/170.

Tsukada:1966:NRL

[TMF66] Hideyuki Tsukada, Yohichi Mochizuki, and Seiki Fujiwara. The nucleoids of rat liver cell microbodies. Journal of Cell Biology, 28(3):449–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/28/3/449.

Tulchin:1967:DSG

[TMK67] N. Tulchin, G. M. Mateyko, and M. J. Kopac. Drosophila salivary glands in vitro. Journal of Cell Biology, 34(3):891–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
3/891.

Tsukada:1968:MDM

[TMK68] Hideyuki Tsukada, Yohichi Mochizuki, and Tadako Konishi. Morphogenesis and development of microbodies of hepatocytes of rats during pre- and postnatal growth. Journal of Cell Biology, 37(2):231–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/37/2/231.

Tashiro:1968:ENM

[TMMN68a] Yutaka Tashiro, Shiro Matsuura, Takashi Morimoto, and Sunao Nagata. Extrusion of nuclear materials into cytoplasm in the posterior silk gland cells of silkworm, bombyx mori. Journal of Cell Biology, 36(3):??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/36/3/C5.

Tashiro:1968:SPS

[TMMN68b] Yutaka Tashiro, Takashi Morimoto, Shiro Matsuura, and Sunao Nagata. Studies on the posterior silk gland of the silkworm, bombyx mori. Journal of Cell Biology, 38(3):574–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
3/574.

Tokuyasu:1968:FSA

[TMZ68] K. Tokuyasu, S. C. Madden, and L. J. Zeldis. Fine structural alterations of interphase nuclei of lymphocytes stimulated to growth activity in vitro. Journal of Cell Biology, 39(3): 630–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/39/3/630.

Tooze:1965:IEM

[Too65] J. Tooze. An investigation by electron microscopy of the nucleoside phosphatase activity of amphibian and mammalian erythrocytes. Journal of Cell Biology, 26(1):209–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/26/1/209.

Tormey:1965:ALF

[Tor65] John Mcd. Tormey. Artifactual localization of ferritin in the ciliary epithelium in vitro. Journal of Cell Biology, 25(2):1–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/25/2/1.

Tilney:1967:SMH

[TP67] Lewis G. Tilney and Keith R. Porter. Studies on the microtubules in heliozoa. Journal of Cell Biology, 34(1):327–??, July 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/34/1/327.

Toselli:1968:FSVa

[TP68a] Paul A. Toselli and Frank A. Pepe. The fine structure of the ventral intersegmental abdominal muscles of the insect rhodnius prolixus during the molting cycle. Journal of Cell Biology, 37(2):445– ??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/37/
2/445.

Toselli:1968:FSVb

[TP68b] Paul A. Toselli and Frank A. Pepe. The fine structure of the ventral intersegmental abdominal muscles of the insect rhodnius prolixus during the molting cycle. Journal of Cell Biology, 37(2):462– ??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/37/
2/462.

Tandler:1969:ONT

[TR69] Bernard Tandler and Leonard L. Ross. Observations of nerve terminals in human labial salivary glands. Journal of Cell Biology, 42(1):339–??, July 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/1/339.

Trelstad:1966:TJB

[TRH66] Robert L. Trelstad, Jean-Paul Revel, and Elizabeth D. Hay. Tight junctions between cells in the early chick embryo as visualized with the electron microscope. Journal of Cell Biology, 31(1):??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/31/
1/C6.

Trotter:1965:EOL

[Tro65] Nancy L. Trotter. Electron-opaque, lipid-containing bodies in mouse liver at early intervals after partial hepatectomy and Sham operation. Journal of Cell Biology, 25(3):41–??, June 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/25/3/41.

Trotter:1967:EOB

[Tro67] Nancy L. Trotter. Electron-opaque bodies and fat droplets in mouse liver after fasting or glucose injection. Journal of Cell Biology, 34(3):703–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/34/3/703.

Tice:1965:LMA

[TS65a] Lois W. Tice and David S. Smith. The localization of myofibrillar ATP ase activity in the flight muscles of the blowfly, Calliphora Erythrocephala. Journal of Cell Biology, 25(3):121–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/25/
3/121.

Tokuyasu:1965:UMP

[TS65b] Kiyoteru Tokuyasu and Otto H. Scherbaum. Ultrastructure of mucocysts and pellicle of Tetrahymena Pyriformis. Journal of Cell Biology, 27(1):67–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/27/1/67.

Tandler:1969:NOI

[TS69] Carlos J. Tandler and Alberto J. Solari. Nucleolar orthophosphate ions. Journal of Cell Biology, 41(1):91–??, April 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/41/1/91.

Thines-Sempoux:1969:ACN

[TSACBB69] D. Thines-Sempoux, A. Amar-Costesec, H. Beaufay, and J. Berthet. The association of cholesterol, 5’-nucleotidase, and alkaline phosphodiesterase I with a distinct group of microsomal particles. Journal of Cell Biology, 43(1):189–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/43/1/189.

Terzakis:1967:TPG

[TSW67] John A. Terzakis, Helmuth Sprinz, and Ronald A. Ward. The transformation of the plasmodium gallinaceum oocyst in aedes aegypti mosquitoes. Journal of Cell Biology, 34(1):311–??, July 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/34/1/
311.

Troughton:1969:PGC

[TT69] W. David Troughton and Jerry S. Trier. Paneth and goblet cell renewal in mouse duodenal crypts. Journal of Cell Biology, 41(1):251–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/1/251.

Turner:1968:USP

[Tur68] F. R. Turner. An ultrastructural study of plant spermatogenesis. Journal of Cell Biology, 37(2):370–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/37/2/370.

Tongiani:1969:IDA

[TVM69] Roberto Tongiani and Maria Pia Viola-Magni. Ii. differences in adrenal medulla nuclear DNA content among rats of different strains following intermittent exposure to cold. Journal of Cell Biology, 42(2):452–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/42/2/452.

Tixier-Vidal:1967:EQP

[TVP67] A. Tixier-Vidal and R. Picart. Etude quantitative Par radioautographie Au microscope electronique De l’utilisation De la dlleucine-^{3} h Par les cellules De l’hypophyse du Canard en culture organotypiqe. Journal of Cell Biology, 35(3):501–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/35/
3/501.

Trosko:1965:SVF

[TW65] James E. Trosko and Sheldon Wolff. Strandedness of Vicia Faba chromosomes as revealed by enzyme digestion studies. Journal of Cell Biology, 26(1):125–??, July 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/26/1/125.

Tandler:1968:UFS

[TWE68] Bernard Tandler, David L. Williamson, and Lee Ehrman. Unusual filamentous structures in the paragonia of male Drosophila. Journal of Cell Biology, 38(2):329–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/38/2/329.

Uno:1968:EAE

[UAH68] H. Uno, K. Adachi, and F. Hu. Enzyme activities of the epithelial outgrowth of the hair follicles in tissue culture. Journal of Cell Biology, 38(3):640–??, September 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/38/3/640.

Utsunomiya:1966:SFIa

[UR66a] Takeshi Utsunomiya and Jay S. Roth. Studies on the function of intracellular ribonucleases. Journal of Cell Biology, 29(3):387– ??, June 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/29/
3/387.

Utsunomiya:1966:SFIb

[UR66b] Takeshi Utsunomiya and Jay S. Roth. Studies on the function of intracellular ribonucleases. Journal of Cell Biology, 29(3):395– ??, June 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/29/
3/395.

Urban:1969:FSP

[Urb69] Paul Urban. The fine structure of pronuclear fusion in the coenocytic marine alga bryopsis hypnoides lamouroux. Journal of Cell Biology, 42(2):606–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/42/2/606.

Ulsamer:1969:LAC

[USK69] Andrew G. Ulsamer, Frank Rees Smith, and Edward D. Korn. Lipids of Acanthamoeba Castellanii. Journal of Cell Biology, 43 (1):105–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/43/1/105.

Vaes:1968:MBR

[Vae68] Gilbert Vaes. On the mechanisms of bone resorption. Journal of Cell Biology, 39(3):676–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/3/676.

Vaughn:1966:RSC

[Vau66] Jack C. Vaughn. The relationship of the “ sphere chromatophile “ to the fate of displaced histones following histone transition in rat spermiogenesis. Journal of Cell Biology, 31(2):257–??, November 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/31/
2/257.

vonBorstel:1966:INN

[vBPB66] R. C. von Borstel, D. M. Prescott, and F. J. Bollum. Incorporation of nucleotides into nuclei of fixed cells by DNA polymerase. Journal of Cell Biology, 29(1):21–??, April 1966. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/29/1/21.

Venable:1965:SLC

[VC65] John H. Venable and Richard Coggeshall. A simplified lead citrate stain for use in electron microscopy. Journal of Cell Biology, 25(2):407–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/2/407.

VanHarreveld:1965:SES

[VCM65] A. Van Harreveld, Jane Crowell, and S. K. Malhotra. A study of extracellular space in central nervous tissue by freezesubstitution. Journal of Cell Biology, 25(1):117–??, April 1965.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/25/1/117.

vonderDecken:1967:ERP

[vdD67] Alexandra von der Decken. Evidence for regulation of protein synthesis at the translation level in response to dietary alterations. Journal of Cell Biology, 33(3):657–??, June 1967. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/33/3/657.

vonderDecke:1969:AVR

[vdD69] Alexandra von der Decke. Activation in vitro of rat liver polyribosomes. Journal of Cell Biology, 43(1):138–??, October 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/43/1/138.

Verbin:1967:ECC

[VF67] Robert S. Verbin and Emmanuel Farber. Effect of cycloheximide on the cell cycle of the crypts of the small intestine of the rat. Journal of Cell Biology, 35(3):649–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/35/3/649.

Venkatachalam:1969:UBL

[VF69] M. A. Venkatachalam and H. Dariush Fahimi. The use of beef liver catalase as a protein tracer for electron microscopy. Journal of Cell Biology, 42(2):480–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/42/2/480.

Vasington:1968:OLR

[VG68] Frank D. Vasington and John W. Greenawalt. Osmotically lysed rat liver mitocondria. Journal of Cell Biology, 39(3):661–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/39/
3/661.

Vincent:1965:MDB

[VHR65] Jacques Vincent, Stanislas Haumont, and Joseph Roels. Microscopic determination of bone phosphorus by quantitative autoradiography of neutron-activated sections. Journal of Cell Biology, 24(1):31–??, January 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/1/31.

vanIterson:1966:FSR

[vI66] Woutera van Iterson. The fine structure of the ribonucleoprotein in bacterial cytoplasm. Journal of Cell Biology, 28(3):563–??, March 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/28/
3/563.

vanIterson:1966:BBB

[vIHvZ66] Woutera van Iterson, Judith F. M. Hoeniger, and Eva Nijman van Zanten. Basal bodies of bacterial flagella in Proteus mirabilis. Journal of Cell Biology, 31(3):585–??, December 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/31/3/585.

vanIterson:1967:MB

[vIHvZ67] Woutera van Iterson, Judith F. M. Hoeniger, and Eva Nijman van Zanten. A “microtubule” in a bacterium. Journal of Cell Biology, 32(1):1–??, January 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/32/1/1.

Viola-Magni:1965:CDC

[VM65] Maria Pia Viola-Magni. Changes in the DNA content of adrenal medulla nuclei of rats intermittently exposed to cold. Journal of Cell Biology, 25(3):415–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/25/3/415.

Viola-Magni:1966:ADL

[VM66a] Maria Pia Viola-Magni. An analysis of DNA loss and synthesis in the rat adrenal medulla nuclei upon cold stimulation. Journal of Cell Biology, 30(2):213–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/30/2/213.

Viola-Magni:1966:RSH

[VM66b] Maria Pia Viola-Magni. A radioautographic study with H^{3}- thymidine on adrenal medulla nuclei of rats intermittently exposed to cold. Journal of Cell Biology, 28(1):9–??, January 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/28/1/9.

Vivier:1969:CMS

[VP69] Emile Vivier and André Petitprez. Le complexe membranaire superficiel et son evolution lors De l’elaboration des individusfils chez toxoplasma gondii. Journal of Cell Biology, 43(2): 329–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/43/2/329.

Vanderberg:1967:DNC

[VR67] Jerome Vanderberg and Johannes Rhodin. Differentiation of nuclear and cytoplasmic fine structure during sporogonic development of plasmodium berghei. Journal of Cell Biology, 32(3):??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/32/
3/C7.

Vatter:1968:EL

[VRN^{+}68] A. E. Vatter, O. K. Reiss, Joyce K. Newman, Karin Lindquist, and Elly Groeneboer. Enzymes of the lung. Journal of Cell Biology, 38(1):80–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/38/1/80.

Voute:1968:SMA

[VU68] Cornelis L. Voûte and Hans H. Ussing. Some morphological aspects of active sodium transport*. Journal of Cell Biology, 36 (3):625–??, March 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/36/3/625.

Weiss:1965:EMS

[WA65] Leon Weiss and Alan C. Aisenberg. An electron microscope study of lymphatic tissue in runt disease. Journal of Cell Biology, 25(3):149–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/3/149.

Warren:1968:ECM

[War68] Robert H. Warren. The effect of colchicine on myogenesis in vivo in rana pipiens and rhodnius prolixus (hemiptera). Journal of Cell Biology, 39(3):544–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/3/544.

Warshaw:1969:CEM

[War69] Joseph B. Warshaw. Cellular energy metabolism during fetal development. Journal of Cell Biology, 41(2):651–??, May 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/41/2/651.

Wiesner:1965:DRA

[WARS65] R. Wiesner, G. Acs, E. Reich, and A. Shafiq. Degradation of ribonucleic acid in mouse fibroblasts treated with actinomycin. Journal of Cell Biology, 27(1):47–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/27/1/47.

Whetsell:1969:RAG

[WB69] William O. Whetsell and Richard P. Bunge. Reversible alterations in the Golgi complex of cultured neurons treated with an inhibitor of active Na and K transport. Journal of Cell Biology, 42(2):490–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/42/2/490.

Wirtschafter:1967:HCD

[WCJS67] Z. T. Wirtschafter, E. G. Cleary, D. S. Jackson, and L. B. Sandberg. Histological changes during the development of the bovine nuchal ligament. Journal of Cell Biology, 33(3):481–??, June 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/33/
3/481.

Wanson:1968:RSM

[WD68a] Jean-Claude Wanson and Pierre Drochmans. Rabbit skeletal muscle glycogen. Journal of Cell Biology, 38(1):130–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/38/1/
130.

Webster:1968:ETH

[WD68b] P. L. Webster and D. Davidson. Evidence from thymidine-^{3} Hlabeled meristems of Vicia Faba of two cell populations. Journal of Cell Biology, 39(2):332–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/2/332.

Wolstenholme:1968:SDB

[WD68c] David R. Wolstenholme and Igor B. Dawid. A size difference between mitochondrial DNA molecules of urodele and anuran amphibia. Journal of Cell Biology, 39(1):222–??, October 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/39/1/222.

Weiss:1965:SCD

[Wei65] Leonard Weiss. Studies on cell deformability. Journal of Cell Biology, 26(3):735–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/26/3/735.

Weiss:1966:SCD

[Wei66] Leonard Weiss. Studies on cell deformability. Journal of Cell Biology, 30(1):39–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/30/1/39.

Weiss:1967:SCDb

[Wei67a] Leonard Weiss. Studies of cell deformability. Journal of Cell Biology, 35(2):347–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/35/2/347.

Weiss:1967:SCDa

[Wei67b] Leonard Weiss. Studies on cell deformability. Journal of Cell Biology, 33(2):341–??, May 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/33/2/341.

Weis:1968:CSC

[Wei68] Peddrick Weis. Confronting subsurface cisternae in chick embryo spinal ganglia. Journal of Cell Biology, 39(2):485–??, November 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/39/
2/485.

Went:1966:IMA

[Wen66] Hans A. Went. An indirect method to assay for mitotic centers in sand dollar (Dendraster Excentricus) eggs. Journal of Cell Biology, 30(3):555–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/30/3/555.

Wunderlich:1968:SME

[WF68] Frank Wunderlich and Werner W. Franke. Structure of macronuclear envelopes of Tetrahymena Pyriformis in the stationary phase of growth. Journal of Cell Biology, 38(2):458–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
2/458.

Wolken:1969:ESO

[WF69] Jerome J. Wolken and Robert G. Florida. The eye structure and optical system of the crustacean copepod, copilia. Journal of Cell Biology, 40(1):279–??, January 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/40/1/279.

Wolken:1965:CEC

[WG65] J. J. Wolken and G. J. Gallik. The compound eye of a crustacean, leptodora kindtii. Journal of Cell Biology, 26(3):968–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/26/
3/968.

Warren:1968:MAC

[WG68a] L. Warren and M. C. Glick. Membranes of animal cells. Journal of Cell Biology, 37(3):729–??, June 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/37/3/729.

Wissig:1968:MMA

[WG68b] S. L. Wissig and D. O. Graney. Membrane modifications in the apical endocytic complex of ileal epithelial cells. Journal of Cell Biology, 39(3):564–??, December 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/39/3/564.

Wolfe:1966:SMC

[WH66] Stephen L. Wolfe and Godfrey M. Hewitt. The strandedness of meiotic chromosomes from oncopeltus. Journal of Cell Biology, 31(1):31–??, October 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/31/1/31.

Wallace:1968:BM

[WHL68] P. G. Wallace, M. Huang, and Anthony W. Linnane. The biogenesis of mitochondria. Journal of Cell Biology, 37(2):207–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/37/
2/207.

Whur:1969:RVI

[WHL69] P. Whur, Annette Herscovics, and C. P. Leblond. Radioautographic visualization of the incorporation of galactose-^{3}H and mannose-^{3}H by rat thyroids in vitro in relation to the stages of thyroglobulin synthesis. Journal of Cell Biology, 43(2): 289–??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/43/2/289.

Widnell:1967:IEA

[WHT67] C. C. Widnell, T. H. Hamilton, and J. R. Tata. The isolation of enzymically active nuclei from the rat heart and uterus. Journal of Cell Biology, 32(3):766–??, March 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/32/3/766.

Williamson:1965:TDA

[Wil65a] D. H. Williamson. The timing of deoxyribonucleic acid synthesis in the cell cycle of Saccharomyces Cerevisiae. Journal of Cell Biology, 25(3):517–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/25/3/517.

Wills:1965:CSM

[Wil65b] E. J. Wills. Crystalline structures in the mitochondria of normal human liver parenchymal cells. Journal of Cell Biology, 24(3):511–??, March 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/24/3/511.

Williams:1967:MIB

[Wil67] M. A. Williams. More intramitochondrial bodies. Journal of Cell Biology, 35(3):730–??, December 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/35/3/730.

Wilson:1969:ABM

[Wil69] H. J. Wilson. Arms and bridges on microtubules in the mitotic apparatus. Journal of Cell Biology, 40(3):854–??, March 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/40/3/854.

Winicur:1967:REC

[Win67] Sandra Winicur. Reactivation of ethanol-calcium-isolated cilia from Tetrahymena Pyriformis. Journal of Cell Biology, 35(1):??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/35/
1/C7.

Wallach:1966:CSA

[WK66] Donald F. Hoelzl Wallach and Virendra B. Kamat. The contribution of sialic acid to the surface charge of fragments of plasma membrane and endoplasmic reticulum. Journal of Cell Biology, 30(3):660–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/30/3/660.

Wetzel:1969:PLB

[WK69] Mary G. Wetzel and Edward D. Korn. Phagocytosis of latex beads by Acanthamoeba Castellanii (Neff). Journal of Cell Biology, 43(1):90–??, October 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/43/1/90.

Wallach:1966:PDB

[WKG66] Donald F. Hoelzl Wallach, Virendra B. Kamat, and Mitchell H. Gail. Physicochemical differences between fragments of plasma membrane and endoplasmic reticulum. Journal of Cell Biology, 30(3):601–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/30/3/601.

Weibel:1966:PSM

[WKS66] Ewald R. Weibel, Gonzague S. Kistler, and Walter F. Scherle. Practical stereological methods for morphometric cytology. Journal of Cell Biology, 30(1):23–??, July 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/30/1/23.

Wiener:1968:QDC

[WLKS68] Joseph Wiener, Alden V. Loud, Daniel V. Kimberg, and David Spiro. A quantitative description of cortisone-induced alterations in the ultrastructure of rat liver parenchymal cells. Journal of Cell Biology, 37(1):47–??, April 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/37/1/47.

Wooding:1965:FSD

[WN65a] F. B. P. Wooding and D. H. Northcote. The fine structure and development of the companion cell of the phloem of Acer Pseudoplatanus. Journal of Cell Biology, 24(1):117–??, January 1965. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/24/1/117.

Woods:1965:CAR

[WN65b] John F. Woods and George Nichols. Collagenolytic activity in rat bone cells. Journal of Cell Biology, 26(3):747–??, September 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/26/
3/747.

Watanabe:1967:ETG

[WO67a] I. Watanabe and S. Okada. Effects of temperature on growth rate of cultured mammalian cells (l5178y). Journal of Cell Biology, 32(2):309–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/32/2/309.

Watanabe:1967:SPC

[WO67b] I. Watanabe and S. Okada. Stationary phase of cultured mammalian cells (l5178y). Journal of Cell Biology, 35(2): 285–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/35/2/285.

Wolfe:1965:IM

[Wol65] Stephen L. Wolfe. Isolated microtubules. Journal of Cell Biology, 25(2):408–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/2/408.

Wolff:1967:FSL

[Wol67] Klaus Wolff. The fine structure of the langerhans cell granule. Journal of Cell Biology, 35(2):468–??, November 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/35/2/468.

Wolfe:1968:EPD

[Wol68] Stephen L. Wolfe. The effect of prefixation on the diameter of chromosome fibers isolated by the Langmuir trough-critical point method. Journal of Cell Biology, 37(3):610–??, June 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/37/3/610.

Wolken:1969:MPP

[Wol69] Jerome J. Wolken. Microspectrophotometry and the photoreceptor of Phycomyces I . Journal of Cell Biology, 43(2):354– ??, November 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/43/2/354.

Willems:1969:RRS

[WPP69] Margherita Willems, Maria Penman, and Sheldon Penman. The regulation of RNA synthesis and processing in the nucleolus during inhibition of protein synthesis. Journal of Cell Biology, 41 (1):177–??, April 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/41/1/177.

Weintraub:1968:ICB

[WR68] M. Weintraub and H. W. J. Ragetli. Intracellular characterization of bean yellow mosaic virus-induced inclusions by differential enzyme digestion. Journal of Cell Biology, 38(2):316–??, August 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
2/316.

Weintraub:1967:SCA

[WRJ67] M. Weintraub, H. W. Ragetli, and V. T. John. Some conditions affecting the intracellular arrangement and concentration of tobacco mosaic virus particles in local lesions. Journal of Cell Biology, 35(1):183–??, October 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/35/1/183.

Walker:1965:CSS

[WS65] Sheppard M. Walker and G. Randolph Schrodt. Continuity of the T system with the sarcolemma in rat skeletal muscle fibers. Journal of Cell Biology, 27(3):671–??, December 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/27/3/671.

Walker:1968:TSM

[WS68] Sheppard M. Walker and G. Randolph Schrodt. Triads in skeletal muscle fibers of 19-day fetal rats. Journal of Cell Biology, 37(2):564–??, May 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/37/2/564.

Walker:1968:EMS

[WSB68] Sheppard M. Walker, G. Randolph Schrodt, and Maxine Bingham. Electron microscope study of the sarcoplasmic reticulum at the Z line level in skeletal muscle fibers of fetal and newborn rats. Journal of Cell Biology, 39(2):469–??, November 1968. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/39/2/469.

Weibel:1969:CMB

[WSGH69] Ewald R. Weibel, Willy Stäubli, Hans Rudolf Gnägi, and Felix A. Hess. Correlated morphometric and biochemical studies on the liver cell. Journal of Cell Biology, 42(1):68–??, July 1969. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/42/1/68.

Wiener:1965:UPN

[WSL65] Joseph Wiener, David Spiro, and Werner R. Loewenstein. Ultrastructure and permeability of nuclear membranes. Journal of Cell Biology, 27(1):107–??, October 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/27/1/107.

Wisniewski:1968:EMS

[WST68] H. Wisniewski, M. L. Shelanski, and R. D. Terry. Effects of mitotic spindle inhibitors on neurotubules and neurofilaments in anterior Horn cells. Journal of Cell Biology, 38(1):224–??, July 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/38/
1/224.

Wetzel:1965:CFS

[WSW65] Bruce K. Wetzel, Samuel S. Spicer, and Seymour H. Wollman. Changes in fine structure and acid phosphatase localization in rat thyroid cells following thyrotropin administration. Journal of Cell Biology, 25(3):593–??, June 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/25/3/593.

Wetzel:1966:ULA

[WWS66] Mary G. Wetzel, Bruce K. Wetzel, and Samuel S. Spicer. Ultrastructural localization of acid mucosubstances in the mouse colon with iron-containing stains. Journal of Cell Biology, 30 (2):299–??, August 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/30/2/299.

Yatsu:1965:UCT

[Yat65] Lawrence Y. Yatsu. The ultrastructure of cotyledonary tissue from gossypium hirsutum l. seeds. Journal of Cell Biology, 25(2):193–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/2/193.

Yamauchi:1967:NMN

[YB67] Akio Yamauchi and Geoffrey Burnstock. Nerve-myoepithelium and nerve-glandular epithelium contacts in the lacrimal gland of the sheep. Journal of Cell Biology, 34(3):917–??, September 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
3/917.

Young:1969:PRP

[YB69] Richard W. Young and Dean Bok. Participation of the retinal pigment epithelium in the rod outer segment renewal process.

Journal of Cell Biology, 42(2):392–??, August 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/42/2/392.

Young:1968:RPR

[YD68] Richard W. Young and Bernard Droz. The renewal of protein in retinal rods and cones. Journal of Cell Biology, 39(1):169–??, October 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/39/
1/169.

Yeh:1969:DFW

[YF69] Jen Yeh and Harold W. Fisher. A diffusible factor which sustains contact inhibition of replication. Journal of Cell Biology, 40(2):382–??, February 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/40/2/382.

Young:1967:RPC

[You67] Richard W. Young. The renewal of photoreceptor cell outer segments. Journal of Cell Biology, 33(1):61–??, April 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/33/1/61.

Yamamoto:1965:FSO

[YTST65] Toshiyuki Yamamoto, Kyoji Tasaki, Yoshio Sugawara, and Akira Tonosaki. Fine structure of the octopus retina. Journal of Cell Biology, 25(2):345–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/25/2/345.

Zadunaisky:1966:LST

[Zad66] Jose A. Zadunaisky. The location of sodium in the transverse tubules of skeletal muscle. Journal of Cell Biology, 31 (2):??, November 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/31/2/C11.

Zalokar:1969:ICS

[Zal69] Marko Zalokar. Intracellular centrifugal separation of organelles in Phycomyces. Journal of Cell Biology, 41(2):494–??, May 1969. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/41/2/494.

Zamboni:1968:IBS

[ZG68] Luciano Zamboni and Bernard Gondos. Intercellular bridges and synchronization of germ cell differentiation during oogenesis in the rabbit. Journal of Cell Biology, 36(1):276–??, January 1968. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/36/1/276.

Zamboni:1966:FSH

[ZMBB66] Luciano Zamboni, Daniel R. Mishell, James H. Bell, and Manuel Baca. Fine structure of the human ovum in the pronuclear stage. Journal of Cell Biology, 30(3):579–??, September 1966. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/30/3/579.

Zeidenberg:1967:ILG

[ZOE67] Phillip Zeidenberg, Sten Orrenius, and Lars Ernster. Increase in levels of glucuronylating enzymes and associated rise in activities of mitochondrial oxidative enzymes upon phenobarbital administration in the rat. Journal of Cell Biology, 32(2):528–??, February 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/32/
2/528.

Zimmerman:1965:EHP

[ZR65] Arthur M. Zimmerman and Ronald C. Rustad. Effects of high pressure on pinocytosis in Amoeba Proteus. Journal of Cell Biology, 25(2):397–??, May 1965. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/25/2/397.

Zimmerman:1967:ACS

[ZZ67] Arthur M. Zimmerman and Selma Zimmerman. Action of colcemid in sea urchin eggs. Journal of Cell Biology, 34(2):483–??, August 1967. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/34/
2/483.